A Complete Bibliography of Publications in Acta Informatica

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

22 November 2024 Version 1.54

Title word cross-reference

#SAT [1268].

(n, k) [370]. (N − 1) [1203]. + [1117, 903]. 0 [349, 852]. 1 [32, 939, 293, 778, 516, 916, 607, 563, 548, 946, 231, 852, 578]. 1, 1 [488]. 2 [1234, 1246, 737, 714, 715, 313, 788, 92, 166, 202]. 2.5n [535]. 3 [1234, 313, 92, 202]. 5′ → 3′ [1631, 1586]. 7 [1141]. ^{∗} [466]. ^{0} [1621]. α − β [235]. α_{2} [1355]. B [515]. B^{∗} [442]. B^{+} [680]. mod5 [1141]. P [1090]. QH [1090]. D [1209, 401, 136, 1078, 1017, 1185]. DTIMEF(F) =?DSPACEF(F) [1014]. ∀ [832]. K [305, 401, 130, 1445, 467, 1078, 618, 1624, 252, 1343, 1187, 851, 1293, 157, 326, 338, 972, 961, 1043, 280, 359, 852]. K_{s,t} [1201]. L [820, 875]. λ [1664]. LR [318]. LU [659]. M [1021, 41, 1658, 561]. M^{X} [1521]. L [1625]. MTL_{0,∞} [1406]. μ [731, 885]. N [1203, 467, 140, 12, 1659, 852]. n = 2^{k} − 1 [1600]. ¬ [832]. ν_{2} [1355]. O(1) [669]. O(n) [136]. O(n^{2}) [1486]. O(V^{5/3}E^{2/3}) [321]. P [1183, 1184, 1226, 1379, 1146]. P = NP [399, 399]. P_{5} [1276]. π [1580, 1206, 1001, 1210, 1299, 912, 1198]. q [1442]. R(1) [1390]. S [892]. Sometime = always + recursionalways [596]. T [1426]. t < n/3 [1486]. T_{3}^{n} [535]. × [1130]. ε [377]. wp [678]. x_{1}^{i}x_{2}^{i} ⋯ x_{m}^{i} = y_{1}^{i}y_{2}^{i} ⋯ y_{n}^{i} [957]. Y [801, 806].

* [891, 889, 880].

-anonymization [1343]. -Arcs [140]. -ary [1442, 1658, 561]. -automata [1390]. -avoidable [1246]. -Based [516]. -Bounded [370]. -Calculus [1299, 912, 1664, 1206, 1001, 731, 1580, 1210, 1198]. -Charts [136]. -Column [852]. -comma [1445]. -conflictable [851]. -decomposition [659]. -Dimensional [788, 1187, 1203]. -edge-deficient [1624]. -Expressions [885]. -free [1201, 1276]. -Head [92]. -invariant [892]. -languages [1017]. -limited [1293]. -Matrices [852]. -Minimal [41]. -minimale [41]. -nets [1021]. -Parser [563]. -primitive [1379, 1185]. -Problems [399]. -productions [377]. -products [1355]. -projections [467, 852]. -PS-codes [1659]. -pure [1146]. -reduction [1625]. -Structures [714, 737]. -Structures. [715]. -Systems [820, 875, 1426]. -Tree [801, 806, 515]. -trees [903, 305, 442, 680]. -Valent [618]. -vectors [467]. -words [1017]. -writer [916].

/G/1 [772]. /M/1 [1521].

0L [1058].

1 [1322, 458, 865, 712, 387, 1521, 772]. 1988 [709]. 1D [1600].

2 [539, 556]. 2-3 [1666]. 2-infix-outfix [1659]. 2014 [1488]. 2015 [1502, 1522]. 2017 [1578]. 21st [1494, 1464]. 23 [671]. 26th [1471]. 2detLIN [1586].

4NF [1034].

60th [1424]. 68 [336, 199, 105].

87k [671].

90 [913].

A* [1327]. A. [841]. Abelian [1246, 595]. Abstract [66, 416, 452, 756, 887, 821, 811, 1373, 800, 298, 1518, 1001, 1461, 1347, 217, 464]. Abstraction [1057, 907, 1097, 1290]. Abstractions [1222, 1251, 1558, 1458]. AC [1621]. Acceptable [559]. Accepted [272, 806]. Accepting [1382, 1068, 1325]. Acception [215]. Acceptors [370, 967]. Access [93, 99, 462, 187, 782]. accommodating [1152]. According [177]. account [1475, 1468]. Achievable [868]. Acknowledgement [909, 989, 1065, 1127, 1150, 1176, 1200, 1248, 1272]. Acta [671, 1133]. Action [993, 1455, 1317, 1145, 1285, 1102]. Actions [1225, 1212, 1076]. Active [309, 1226, 1183, 1594]. Active-Operations [309]. actor [1480]. actor-based [1480]. Actual [379]. Acyclic [968]. Adaptive [1274, 108, 1332]. Addenda [74, 162]. Addendum [761]. Adder [48]. Adding [1140, 1117]. Addition [48, 684]. Additionsdauer [48]. Additive [695]. Additivity [952]. Address [918]. addressing [1381]. adequate [1103]. Admissibility [647, 1296]. admissible [1473]. Admitting [361]. Affine [847, 1668, 123]. Affix [480, 168]. Affix-Grammars [168]. After [536]. Age [723, 1670]. agents [1041]. aggregate [1170]. Aggregation [1587, 682, 1674]. Aggregation-based [1587]. Ahead [897]. Aid [249, 96]. Alan [749]. Algebra [923, 785, 747, 940, 855, 251, 1282, 1021, 1304, 1425, 1512, 1102, 1193, 1640, 1081]. Algebraic [970, 869, 927, 742, 218, 705, 1223, 257, 478, 1285, 549, 167, 835, 955, 804, 411, 1529, 637, 1203, 217, 1054, 1239, 1051, 1354, 641, 306]. algebraically [1040]. Algebras [766, 962, 1374, 1326, 1605, 1098, 1142]. algébriques [498, 257, 306, 167]. ALGOL [199, 418, 105, 336]. ALGOL-like [418]. Algorithm [469, 100, 508, 134, 694, 784, 778, 793, 321, 77, 76, 37, 256, 788, 630, 340, 716, 612, 528, 308, 312, 702, 532, 826, 606, 136, 255, 569, 1298, 1344, 1264, 1588, 1474, 1636, 1358, 1304, 1250, 1666, 1230, 1393, 1552, 1327, 1663, 1462, 1607, 1342, 423, 1037, 1403, 1266, 1540, 1601, 1033]. Algorithme [312, 76]. Algorithmic [647, 286, 182, 105, 1520, 1372]. algorithmically [1603]. Algorithms [687, 1494, 1510, 23, 565, 728, 639, 87, 425, 115, 241, 413, 434, 374, 812, 872, 408, 522, 884, 703, 354, 25, 526, 424, 769, 707, 108, 959, 2, 403, 44, 1489, 1072, 1551, 1181, 1035, 1165, 1675, 1191, 1270, 1086, 1098, 1470, 1047, 1137, 1652, 299, 660]. Alias [913]. Aliasing [356]. All-Round [788]. Allen [1452]. allocate [1333]. Allocation [479, 61, 315, 118, 189, 600, 127, 1189, 1393, 1044]. Almost [740, 702]. Alphabet [191, 1058, 777]. alphabets [997]. Alternating [854, 819, 1642, 791, 1388, 1370, 1602, 1633, 1263]. Alternation [322, 331]. Alternative [429, 130]. ambients [1247, 1214]. Ambiguity [91, 685, 138, 1212]. Ambiguous [325, 1588]. Among [698, 663, 819, 123]. Amongst [937]. Amortized [1022]. Amounts [297]. Analyse [76, 84, 279]. Analysis [233, 234, 335, 1494, 941, 305, 334, 979, 592, 421, 1078, 499, 610, 1224, 458, 374, 126, 304, 286, 630, 84, 977, 504, 160, 628, 756, 85, 522, 8, 194, 600, 1444, 222, 524, 428, 795, 817, 205, 375, 944, 29, 103, 163, 606, 968, 897, 1227, 1243, 1571, 1306, 1115, 1667, 892, 1505, 1111, 1480, 991, 1210, 1066, 1401, 1424, 913, 1005, 1245, 1521, 76, 279]. Analytic [61, 686, 1056]. Analyzer [511]. Analyzing [443, 1534]. anarchy [1336]. Ancestor [634]. Andersson [826]. Andrei [709]. angelic [1077]. annihilators [1371]. anonymization [1343]. antichains [1532]. antiport [1184]. Apex [651, 856, 963]. appearance [1632]. Application [592, 618, 153, 9, 121, 192, 879, 1021, 1309, 1063, 1512, 1410, 1099]. Applications [664, 287, 497, 547, 441, 448, 5, 22, 520, 988, 258, 1323, 1350, 1660, 1367]. Applied [375]. Applying [1282, 242, 102]. Approach [726, 130, 834, 915, 215, 878, 838, 686, 164, 494, 70, 381, 482, 317, 331, 762, 606, 186, 142, 938, 1584, 1130, 1161, 1057, 1237, 1354, 1005, 1113]. Approaches [869]. approximability [1115, 1181]. Approximate [353, 334, 1495, 1465]. Approximating [772, 1661, 1009]. Approximation [925, 1175, 676, 716, 712, 528, 616, 569, 1498]. Approximations [387, 1174, 865, 277]. April [709]. Äquivalente [226]. Arbitrary [766, 555, 527, 414]. Arbitration [393]. Architecture [696, 1153, 927]. Architectures [602]. Arcs [140]. Area [504]. Argument [557]. Arguments [283]. Arithmetic [402, 348, 67, 879]. Arithmetical [1566]. ARQ [1401, 1444]. Array [691]. Arrays [302, 847, 159, 178, 1126]. Arrivals [431, 277]. ary [1442, 1658, 561]. Ascending [367]. Asian [1250]. aspects [1674]. assembly [1430]. assertional [1108]. Assertions [346, 596]. assignment [1215]. associated [1604]. association [1398]. Associative [127, 1533]. Associativity [67]. Assume [1473]. Assume-admissible [1473]. Assumption [915]. Assumption-Commitment [915]. asymmetric [1215]. Asymptotic [1024]. Asymptotics [870, 799]. Asynchronous [368, 302, 857, 812, 1531, 1004, 1122, 1435, 1561, 1486]. asynchrony [1424]. Atomic [916, 1072]. Atomicity [368, 884, 1108, 1255]. atoms [1577]. ATP [814]. Attribute [382, 506, 376, 651, 725, 780, 579, 446, 645, 867, 667, 453, 396, 638, 543, 882, 1487]. attribute-based [1487]. Attributed [611, 700, 300]. Attributes [233, 234, 628, 543]. ATVA [1522]. Author [658]. autodense [1289, 1328]. Automat [18]. Automata [905, 120, 302, 507, 432, 588, 708, 965, 857, 79, 640, 60, 80, 239, 240, 801, 806, 1535, 874, 1320, 461, 92, 497, 854, 926, 47, 297, 219, 388, 13, 810, 46, 1630, 28, 685, 463, 670, 895, 1457, 1388, 1163, 1587, 1594, 1592, 1309, 1000, 1378, 1319, 1485, 1310, 1505, 1355, 1512, 1184, 1589, 1602, 1633, 1627, 1596, 1316, 1595, 1631, 1337, 1418, 1614, 1099, 1331, 1653, 1010, 1105, 1434, 1217, 1411, 1190, 1390, 1586, 1443, 1599, 1456, 1644, 1573, 1351, 1628, 1671, 1611, 1570, 1469, 303, 1610]. Automata-based [1320]. Automated [1557, 1571, 1615]. Automaten [79, 60, 303, 28, 80, 47]. Automates [13]. Automatic [90, 97, 292, 1113, 29]. Automatically [117, 579]. Automatically-Generated [579]. Automating [738, 494]. Automaton [551, 18, 1308, 1650]. Automorphismes [13]. Automorphisms [13]. Availability [591, 1020]. Available [136, 1091]. Average [1503, 996, 263, 561, 991, 515]. Average-case [996, 991]. Average-energy [1503]. averages [1357]. Avoidable [1341, 1361, 1246, 777]. Avoiding [1387, 1442]. Aware [1159]. Axiomatic [503, 51, 74, 164, 70, 373, 344, 133, 310]. Axiomatics [118]. Axiomatising [490, 1105]. axiomatizability [1326, 1006]. axiomatization [1178, 1576]. Axiomatizing [940]. Axioms [945, 1029, 196, 1426]. ayant [566].

B [23, 193, 701, 903, 679]. B-Trees [23, 193, 701, 679]. Backtracking [175, 337]. backward [1410, 1365]. Bag [973]. Bag-Containment [973]. bakery [1344]. balance [1092]. Balanced [432, 830, 194, 669, 433, 1204]. Balancing [1658, 776, 655]. bar [1378]. Base [312, 261]. Based [807, 503, 336, 731, 962, 662, 930, 516, 304, 948, 871, 629, 975, 371, 401, 447, 981, 310, 173, 968, 358, 1298, 1581, 1579, 1587, 1612, 1542, 1465, 1169, 1186, 1487, 1320, 1480, 1675, 1210, 270, 1635, 1019, 786, 1573, 1005, 1645, 1113, 1197, 1540, 1363, 1671, 1417]. Bases [1116, 426]. Basic [714]. Basis [882, 90]. Batch [151, 277]. Batch-Interactive [151]. Bauer [1432]. Bayesian [1477]. Beautiful [283]. Behandlung [364]. Behavior [210, 701, 391]. behavioral [1051]. behaviors [1359]. Behaviour [680, 552, 129, 149, 800, 1539, 1235]. Behavioural [869, 1026]. Behaviours [385, 1310]. below [1565]. Bendix [738]. Berechnung [11, 26, 35]. Bernoulli [348]. Bernstein [1029]. Best [380, 1174]. Between [948, 379, 960, 439, 638, 959, 845, 1014, 1561, 775, 1484, 913, 1567, 1093, 174, 1259, 879]. beyond [842]. Bi [1016, 1162, 1364]. Bi-catenation [1016]. bi-criterion [1162]. bi-ideals [1364]. Bicenters [498]. Bicentres [498]. bicluster [1635]. Bifaithful [417]. bifidèles [417]. bifix [1387]. Big [1507, 1456]. Bijection [307]. Bin [540, 208, 694, 1131, 1483]. Bin-Packing [540]. Binary [469, 427, 740, 23, 824, 626, 969, 77, 496, 98, 267, 24, 493, 497, 263, 4, 194, 584, 527, 109, 251, 933, 959, 826, 262, 333, 1387, 1341, 1361, 990, 1636, 1246, 1438, 775, 1058, 1182, 1635, 1486, 777, 1221]. bio [1483]. bio-inspired [1483]. biochemical [1490]. BIP [1581, 1579]. Bipartite [964, 354]. bipolar [1447]. birthday [1424, 1601]. bisimilarity [1423, 1278, 1604, 1616, 1142]. Bisimulation [912, 1569, 1576, 1479, 1516, 1640]. Bisimulations [743]. Bitopological [1605]. Black [383, 1538, 1666, 1114]. Black-White [383]. Blackboard [896]. Blankinship [308]. Blocking [687, 334, 177]. Blocks [828, 1037]. boolean [1383, 1301, 244, 351, 677, 225, 545, 87, 100, 467, 275, 323, 41, 885, 1280, 245, 281, 1635, 1257, 101, 449, 291, 341]. Boolean-reasoning-based [1635]. Boole’sche [41, 101]. Bordered [571, 1481]. borderenergetic [1662]. bornés [257]. bottom [1318]. bottom-up [1318]. Bound [100, 535, 24, 462, 291, 1240, 1501, 1073]. Boundary [640, 574, 1086]. Bounded [1520, 272, 648, 886, 542, 370, 856, 918, 854, 257, 219, 584, 763, 823, 938, 1344, 1233, 1505, 1585, 1143, 1352]. Bounded-Iteration [763]. Boundedly [851, 159]. boundedness [1580]. Bounds [244, 266, 728, 425, 540, 759, 77, 691, 681, 67, 602, 527, 25, 799, 868, 449, 720, 963, 677, 1044, 1342, 1644, 1142]. boxed [1214]. boxes [1085, 1110, 892]. BPA [1036]. BPP [1036, 1604]. Bracket [53]. Bracket-Closed [53]. Branch [100]. Branch-and-Bound [100]. Branching [476, 610, 758, 1654, 1576, 1570, 1610]. branching-time [1654, 1570, 1610]. Breakdown [693]. Brinch [30]. BRNGLR [1266]. Broadcast [83, 1180]. Bruijn [1677]. Bucket [979, 1056, 1124, 1179]. Buddy [315, 624, 546, 1189]. Buffer [812, 420]. building [1070]. bunches [1083]. BURS [975]. Byzantine [1486]. c [1096, 1224]. C. [132]. C/C [1224]. CA [1601]. Cache [812]. caching [1202, 1230]. CADP [1435]. cake [677]. calculate [1049]. Calculation [721, 11, 26, 35]. Calculi [943, 371, 1306]. Calculus [653, 509, 862, 381, 1299, 912, 808, 1025, 1664, 1206, 1207, 1027, 1001, 731, 1509, 1166, 1208, 1079, 260, 1580, 1210, 1324, 750, 1198]. Calls [577, 176]. Canonical [605, 1156, 9]. Cantor [1029]. Cantor-Bernstein [1029]. Capacity [290, 235]. capture [1090]. Careful [1628]. Carlo [595]. Case [620, 294, 624, 194, 546, 1122, 996, 991, 669, 1]. cases [1452]. categorical [988]. Categories [879]. Category [1097, 930]. Category-Based [930]. Category-theoretic [1097]. catenation [1016]. Catenative [45]. Causal [978, 1212]. causalities [1508]. Causality [923, 1041, 871, 1001, 1421]. CCS [844, 662, 583, 1582, 1373, 724, 1428]. CD [1390]. CDAWGs [1676]. Cells [543, 1061]. Cellular [507, 463, 1000, 1595]. Central [76, 178]. centrale [76]. centralisées [279]. Centralized [279]. Cerný [1319]. Certain [141, 474, 342, 455]. Certaines [342, 474]. Certificate [985]. Certificates [860]. CFFD [1565]. CFL [697]. Chain [409, 548, 1650, 1484]. Chain-Free [548]. Chains [799, 1518, 1608]. changes [1651]. Channel [521]. channels [1227, 1539, 1044, 1198]. CHAP [1353]. Characterisation [840]. Characteristic [1575, 1479, 166]. Characteristics [572, 79]. Characterization [976, 382, 495, 497, 366, 361, 339, 251, 809, 737, 1340, 1051]. Characterizations [1445, 806, 741, 1369]. Characterize [453]. characterizing [1655, 1514, 1190]. Charakteristik [166]. Charting [1292]. Charts [136]. check [1112]. Checking [1013, 661, 731, 950, 796, 884, 1466, 426, 1497, 1536, 1431, 1385, 1294, 1251, 1527, 1496, 1491, 1321, 1543, 1373, 1302]. Checkpointing [335]. Checkpoints [723]. CHOCS [808]. Choice [491, 392, 1520]. choice-free [1520]. choices [1546, 1077]. Choosing [75]. Chromatic [933]. ciliates [1286]. circuit [677]. Circuits [7, 595, 753, 677, 1263, 1182]. Circular [500, 1394]. Circularity [506, 396]. Circus [1345]. Clarifications [268]. Class [807, 179, 793, 229, 278, 110, 162, 754, 197, 795, 1062, 895, 963, 1600, 1073]. Classes [71, 801, 819, 967, 453, 809, 186, 228, 1090, 1659, 995, 474]. Classically [858]. Classification [703, 1452]. Classifications [1229]. Clause [284]. Clean [372, 254]. CLEAR [510]. Clint [132]. cliques [1201]. clocked [1060]. Closed [42, 53, 550, 1116, 1081]. Closer [111, 1672]. Closure [474, 574, 823, 1303, 1369, 185, 1386, 1280]. Clustering [575, 1343]. Co [697, 34, 825]. Co-CFL [697]. Co-Nested [825]. Co-routines [34]. coalgebraically [1421]. Cocomparability [964]. Code [421, 874, 636, 975, 404, 394, 1311, 1442, 1655, 1241, 1209, 1456]. Coders [807]. Codes [140, 1445, 1514, 1169, 1334, 1185, 1659]. coding [1014]. Codings [86, 78]. coffee [1563]. Cohérence [312, 1478]. Coincidental [1138, 1149]. coinduction [1475, 1365]. Collatz [994, 1071]. collection [1091]. colored [1296, 1677]. Coloring [955, 1395, 1201, 1276, 1500]. colors [1500]. coloured [1638]. Column [852]. combination [1313]. Combinational [753]. Combinator [629]. Combinator-Based [629]. Combinatorial [698, 576, 1171]. Combined [151, 1606]. Combining [1420, 1433]. comma [1445]. Command [116]. Comments [188, 30, 58, 43, 95, 243]. Commit [688]. Commitment [915]. Common [393, 634, 44, 423, 1367]. Communicating [639, 509, 350, 558, 822, 795, 644, 1592, 1082]. Communication [521, 759, 974, 854, 791, 1413, 1173, 1050, 1561, 1158, 1044]. Communication-Bounded [854]. communities [1061]. Commutative [544]. Commutativity [67]. Compact [494]. Comparative [869, 632, 1677]. Comparing [1036, 1122, 871]. Comparison [30, 801, 579, 15, 43, 768, 1335, 1350, 1148]. Comparisons [873]. compatibility [1535, 1434]. Competitive [983]. Compilation [284, 629]. Compiler [838, 206, 955, 364, 364]. Compilers [274]. Compiling [1223, 310, 1196]. Complement [1602, 946, 1633]. Complementary [63]. Complementation [236]. Complements [228]. Complete [514, 542, 466, 969, 119, 682, 171, 813, 329, 371, 288, 343, 706, 1452, 1356, 1155, 1188, 473, 1576, 1051]. Completeness [7, 389, 615, 746, 1292]. Completion [649, 1039]. Complexity [266, 351, 411, 198, 204, 593, 514, 203, 179, 499, 535, 759, 286, 62, 390, 544, 183, 819, 379, 718, 245, 890, 928, 121, 192, 191, 68, 650, 688, 644, 147, 211, 809, 36, 435, 443, 576, 291, 341, 1172, 677, 1416, 1069, 1297, 1271, 1357, 1593, 1622, 1517, 1300, 1585, 996, 1090, 1181, 1447, 1125, 746, 889, 1642, 1167, 1362, 1342, 1142, 1506, 1193, 430, 437, 1648, 998]. component [1531, 1581, 1579, 1535]. component-based [1581, 1579]. components [1533, 1353]. composed [1302]. Composite [81]. Composition [915, 1318, 645, 1161, 1441, 1585, 1512, 891, 1533]. Compositional [1385, 896, 1224, 1435, 904, 1480, 1437, 1113]. Compositionality [1420, 1433]. Comprehending [1424]. comprehensive [1197]. Compression [485]. compressions [1658]. computability [1000]. computable [1093]. Computation [1013, 495, 511, 951, 453, 274, 1625]. Computational [687, 1069, 545, 953, 859, 707, 970]. computationally [1188]. Computations [827, 1225, 708, 940, 175, 883, 1661, 1004, 677, 1374, 1466]. Compute [502, 284, 309, 455]. Computer [152, 83, 126, 590, 144, 591, 150, 614, 504, 189, 212, 153, 151, 96, 277]. Computers [40, 454]. Computing [353, 986, 552, 27, 136, 729, 142, 185, 1002, 1192, 1044]. Concatenation [982]. Concepts [15, 1027, 1197, 30, 43]. Conceptual [995, 988]. concerning [1011, 430, 437]. CONCUR [1471]. Concurrency [1471, 193, 915, 966, 440, 689, 717, 312, 779, 621, 1420, 1433, 1436, 1612, 1205, 1530, 1284, 889, 1567]. Concurrent [743, 1461, 188, 715, 313, 434, 950, 900, 597, 311, 110, 162, 933, 187, 243, 443, 265, 755, 1021, 1370, 1028, 1435, 1145, 1212, 1113, 1656, 1076]. condensed [1235]. Condition [7, 662, 1517]. condition-response [1517].

Condition/Event [662]. Conditional [834, 813, 128]. Conditions [541, 1241]. cones [417]. Conference [1471, 1494]. Configurable [1611]. conflict [1015]. conflict-free [1015]. conflictable [851]. conflicts [1279]. Confluence [936]. Confluent [652, 754]. Congruence [1564, 754, 1166]. Congruences [840, 13, 1565, 13]. Conjecture [994, 73, 1319, 1397]. Conjunctive [1388, 877, 1499, 973, 1261, 1112]. Connected [691, 602]. connectedness [1117]. Connection [845, 289]. Connective [832]. Connective-Free [832]. Connectivity [1349, 1203]. Conquer [502]. Conradi [243]. Consensus [1265, 919, 1486, 1540]. Conservation [223]. Conservative [56]. Considerations [403, 1162]. Considered [181, 260]. considering [1492]. consistency [1180, 1032]. Consistent [1004, 63, 1574]. Constant [648, 463, 1069, 1538, 1596, 1022, 1486, 1644]. Constant-time-maintainable [648]. Constants [958]. Constrained [1649, 1608]. constraint [235, 1450]. Constraints [622, 530, 954, 635, 755, 1416, 1140, 1271, 1583, 1020, 1657]. Construct [763]. Constructibility [603, 1429]. Constructing [469, 249, 253, 157, 573]. Construction [1494, 747, 946, 974, 619, 106, 447, 292, 475, 128, 1437, 1084]. Constructions [916]. Constructive [878, 1436]. constructors [1109]. Contained [225, 156, 1339]. container [1367]. Containing [222]. Containment [973, 1107, 1499, 1112]. Contention [606]. Contentions [517]. Context [498, 409, 997, 59, 848, 780, 856, 982, 419, 704, 84, 390, 926, 366, 325, 145, 139, 733, 890, 928, 327, 221, 324, 72, 301, 33, 46, 64, 114, 103, 410, 211, 264, 1164, 1482, 377, 1280, 1138, 1149, 1293, 1346, 1362, 1190, 750]. Context-Free [498, 780, 856, 982, 419, 84, 390, 926, 366, 928, 72, 301, 33, 410, 264, 997, 848, 325, 1164, 1482, 377, 1280]. Context-Sensitive [59, 704, 46, 64, 103, 211, 1190]. Context-Sensitivity [114]. Contextfree [266, 1408]. Contexts [359, 1508]. Contextual [1439, 935, 901, 813]. Continuation [174]. Continued [11]. Continuity [557]. Continuous [438, 549, 698, 1557, 1451, 1643, 1352]. continuous-time [1451, 1352]. Contour [502]. Contrôle [312]. contracted [1548]. Contraction [883, 1416]. contribution [260]. Contributions [657]. Control [948, 440, 151, 312, 205, 178, 779, 727, 1311, 1603, 1541, 1134, 889, 104, 1352, 1244]. controllability [1469, 1470]. Controllable [583, 1449, 1643]. Controlled [366, 1411, 238, 19, 1528]. controller [1406]. controllers [1557, 1437, 1556]. controlling [1209]. Conventions [632]. convergence [1254]. conversion [1657]. Converting [1012]. Convex [513, 964, 354]. convexity [1672]. cooperation [1269]. Cooperative [655, 1132, 1401, 1444]. Coordinating [110, 162]. Coordination [494, 392, 1290]. cope [1219]. Copy [371]. core [1509]. cores [1453]. Coroutine [451]. Coroutines [295]. Correct [200, 548, 38, 1354, 265]. Correcting [292]. Correction [1581, 1618, 784, 293, 1609, 1634, 760, 97, 1610, 1342]. Corrections [660]. Correctness [1308, 347, 661, 415, 295, 206, 1223, 525, 70, 1098, 108, 310, 196, 567, 255, 1026, 21]. Correspondence [829, 459]. Corresponding [72]. Corrigenda [74, 162, 192]. corrigendum [1172]. Cost [353, 24, 974, 165, 727, 1046, 1474, 1191, 1455, 1035]. cost-effectiveness [1046]. costlier [1504]. Costs [427, 919, 414, 209, 1181, 1160, 1504]. COSY [270]. Counter [550, 388, 191, 1184]. Counterexamples [954]. counterparts [1164]. Counting [676, 1495, 1642, 1084, 1506]. Coupled [1568, 645, 189]. Cover [528]. Coverage [101]. Covering [501, 410, 1201, 1375]. Coverings [554]. Covers [471, 324, 1268, 1460]. CPP [1096]. CPP/GE/c/L [1096]. CPS [1324]. CPS-calculus [1324]. Craig [1458]. CRCW [937]. Creation [785, 1398]. Crick [1485, 1589, 1631, 1586]. Criterion [112, 1108, 1255, 1162]. Critical [305, 590, 58, 128]. crochet [53]. cross [1387, 775]. cross-bifix-free [1387]. Crossing [960]. Crossing-Over [960]. Crypt [637]. Crypt-equivalent [637]. CSMA [304]. CSP [1618, 1613, 831, 604, 567]. CTL [1527, 1543]. CTMDPs [1352]. cube [1199]. cubic [1266]. customers [1096]. cut [1536]. cut-points [1536]. Cycle [958, 1671]. Cyclic [431, 201, 126, 668, 683, 699, 1462]. Cylinder [258]. Cylindre [258].

D [336]. D0L [1240, 1340, 1501]. dans [261, 279, 566]. Darlington [299, 660]. Darstellungen [364]. Data [233, 234, 23, 487, 294, 445, 893, 500, 269, 210, 406, 666, 49, 93, 99, 452, 81, 682, 977, 395, 249, 160, 756, 1054, 821, 165, 170, 570, 739, 279, 344, 721, 672, 426, 428, 401, 375, 187, 1067, 209, 992, 800, 89, 598, 186, 298, 441, 448, 5, 22, 917, 274, 642, 846, 1457, 1140, 1128, 1507, 1251, 1347, 1392, 1617, 217, 21, 1331, 519, 1277, 553, 786, 1314, 1221, 1222, 464, 988]. Data-Flow [977]. Data-Intensive [917]. Database [771, 487, 661, 236, 261, 320, 360, 648, 654, 361, 987, 529, 954, 312, 414, 441, 448, 635, 947, 779, 727, 1308, 1038, 1007, 1460, 1197, 1034]. Databases [440, 603, 690, 843, 866, 1048, 1271, 996, 995, 1106, 760, 761]. Dataflow [904, 630, 1098]. datalog [1090, 1237, 837]. datatypes [1566]. Datenstrukturen [93, 99]. d’attente [566]. Daubechies [965]. David [135]. DB [156]. DB/DC [156]. DC [156]. deadlines [1053]. Deadlock [517, 177, 633, 1363]. Deadlock-Freedom [517]. Deadlocking [674]. deal [450]. deallocation [1189]. debugging [1041]. December [709]. Decentralized [1645]. Decidability [588, 112, 950, 520, 1294, 1537, 1644, 1351, 1629]. Decidable [204, 405, 538, 352, 389, 1545]. decide [1296]. Deciding [1278, 560, 1451, 1193]. Decision [1253, 692, 1011, 547, 754, 959, 1489, 1275, 1657]. declaratively [1612]. Decomposable [129]. Decomposing [828, 1037]. Decomposition [82, 594, 205, 744, 96, 1305, 1318, 1313, 659]. decompositions [737]. Decreasing [417, 1204]. décroissants [417]. Dedicated [222, 1432, 1424]. Deduction [357]. Deductive [866]. Deep [1217, 1315]. deficient [1624]. Defined [379]. Defining [71, 328]. Definiteness [722]. Definition [252, 378, 51, 74, 373, 1508]. Definitions [1223, 996]. deformation [1273]. Degree [856, 685, 832, 823, 1416, 1639]. Degrees [114, 1577]. Delay [752, 751, 490, 918, 1401, 1603]. delay-resilient [1603]. Delays [304, 1603]. Delegation [755]. Deleted [987, 1007]. Deleting [400]. deletion [1528, 1666, 1447]. Delivery [861]. delta [1385, 1509]. delta-oriented [1385, 1509]. Demand [54, 212, 234]. Demonic [1077]. Denotational [1436, 518, 428, 1018]. Dense [1229, 1452]. Density [325]. Dependability [849]. Dependence [457, 913]. Dependencies [771, 236, 952, 570, 471, 866, 456, 1140, 1465, 1128, 1038, 1292, 1102]. Dependency [843]. Dependent [723, 663]. Depletable [1539]. Depth [1620, 305, 538, 125, 341, 275, 323]. Depth-First [305, 125, 1620]. Deques [789]. derandomization [1035]. Derivation [627, 72, 265]. Derivational [183]. Derivations [653, 733]. derivatives [1597]. derived [1047]. Deriving [1111, 424, 102]. Descending [1518]. descent [1257]. Describing [552, 37, 711]. Description [91]. descriptional [1596, 1648]. Descriptively [329]. Design [336, 838, 206, 482, 609, 2, 173, 227, 1539, 1057, 1030, 1031, 1034]. Designed [336]. Detecting [898]. Detection [438]. Determinancy [27]. Determinism [331, 1168]. Deterministic [184, 588, 857, 797, 461, 153, 505, 482, 609, 191, 967, 1390, 1443, 586, 420, 895, 273, 1529, 1519, 1588, 1316, 1337, 1187, 1607, 1586, 553, 1628, 86]. deux [439]. Development [705, 249, 804, 641, 1136, 659]. Developmental [45]. device [1061]. Devices [27, 447, 1650, 104]. DFA [1574]. DFAs [1342]. Diagrams [959, 1103]. diameter [1672]. Different [629, 86, 78, 959]. Diffusion [277]. digital [1557, 1232, 1104]. Dijkstra [1133, 1027, 413]. Dijkstra-Scholten [1027]. Dimension [759, 544, 899]. Dimensional [425, 788, 401, 1405, 1279, 1203, 1187, 1601]. dioids [1529]. Diophantine [308]. Direct [483, 946, 178]. Directed [366, 859, 170, 968, 1620, 1173, 1373]. Direction [19]. Directory [1151]. Disambiguating [931]. discounting [1402]. Discovering [1656]. Discrete [348, 966, 309, 762, 365, 1611]. Discrete-Time [348]. discretization [1558]. discretization-free [1558]. Disjoint [125]. disjointedness [1107]. Disjunctive [1379, 1339]. Disjunctivity [1481]. Disk [76, 812, 55, 1202]. Dispatcher [106]. Disque [76]. Distance [951, 1665]. Distance-edge-monitoring [1665]. Distances [1232, 353]. distinct [1674]. distinctness [1155]. Distinguishing [1380]. Distributed [479, 261, 915, 662, 1408, 1265, 978, 591, 529, 810, 312, 688, 644, 655, 643, 656, 1413, 1525, 1660, 1050, 1132, 1393, 1192, 1159, 1663, 1044, 1084, 1136]. Distribution [810, 365, 956, 1392]. distribution-sensitive [1392]. Distributions [348]. Distributive [1547]. Diverse [42]. diverses [42]. Divide [502]. Divide-and-Conquer [502]. DKL [1607]. DLOGTIME [1263]. DLOGTIME-uniform [1263]. DNA [1002, 1157]. do [1069, 260]. Does [1381]. DOL [220]. Domain [978, 862, 1458]. domain-specific [1458]. Domains [711, 1518]. domination [1672]. Données [312, 261, 279]. Dot [1652]. double [816, 1533]. double-ended [816]. double-sided [1533]. Down [640, 8, 526, 1128]. Drawing [533, 435, 1172]. Drum [248, 512]. DSL [445]. Dual [540, 571, 1262, 1344, 1444, 1288]. dual-hop [1444]. Duality [726, 1605]. duel [1675]. duel-and-sweep [1675]. d’un [76, 312]. d’une [279, 312]. Duplicate [681]. Duplicate-Key [681]. duplicated [1669]. durch [364]. Dynamic [315, 572, 881, 1469, 701, 282, 401, 337, 961, 655, 1477, 634, 262, 635, 823, 1536, 1508, 1509, 1517, 1582, 1044, 1470]. Dynamic-Sized [881]. dynamical [1245]. dynamically [1449]. Dynamics [654, 1539].

eager [1552]. Earley [1317]. Early [1317]. Easily [613]. Easy [985]. ECO [1171]. Edge [125, 1416, 1624, 1500, 1665]. Edge-Disjoint [125]. Edges [353, 1043]. Editor [1075, 1419]. Editorial [1651, 1670, 1242, 1404, 1075]. Edsger [1133]. Effect [407, 165, 318]. Effective [747, 778]. Effectively [534]. effectiveness [1046]. Effects [164]. Efficiency [805, 626, 355, 454, 1539, 1122, 1401]. Efficient [460, 815, 32, 294, 1406, 979, 293, 287, 1487, 563, 630, 768, 1449, 408, 1343, 736, 1086, 884, 716, 354, 564, 1470, 424, 707, 959, 128, 1358, 1666, 1607, 786, 1403, 1459]. Efficiently [1483, 1070]. Eigenschaften [79]. Einbettungssatz [204]. Electing [599]. electronic [1670]. Element [1155, 1514, 1185]. Elemental [621]. Elementary [798, 1066]. Elements [12, 652]. Eliminate [500]. Eliminates [598]. Eliminating [195, 769, 318, 359]. Elimination [44, 1413]. ELL [563, 252, 231]. ELR [252]. embedded [1164]. Embedding [204, 1323, 618, 1213]. Emptiness [370]. Emptiness-Problem [370]. Emulator [969]. eNCE [1086]. Encodements [496]. encoding [1664, 1214, 1671]. encoding-based [1671]. Encodings [209]. ended [816]. endliche [18, 28]. endlichen [60]. endlicher [60, 80, 47]. Endogenous [151]. energy [1503, 1401]. enforcement [1654, 1032]. Enhanced [1202]. entre [439]. Entreeings [815]. Entropy [458, 267, 379, 585, 668, 712, 772, 683, 699]. entscheidbare [204]. Enumerable [345, 741]. Enumeration [670]. Environment [402, 107, 974, 451, 735, 910]. environments [1180, 1459]. enzymatic [1377]. EOL [172, 290, 104, 389]. Epsilon [460]. Equal [345]. equals [802]. Equation [308]. Equational [1360, 930, 949, 1006, 1064]. Equations [847, 929, 957]. Equilibria [1336, 1606, 1375]. Equilibrium [592, 458, 1365]. equipment [1350]. Equivalence [351, 1536, 545, 577, 82, 983, 220, 588, 587, 174, 895, 1413, 596, 1287, 1170, 1240, 1501, 1342, 1467, 1351, 1302, 1301, 1259, 1076, 226]. Equivalences [608, 871, 1216, 1386]. Equivalent [278, 637]. erase [1293]. erkennbaren [79]. Erratum [1361, 360, 671, 683, 699, 1007, 1179, 1149]. Error [939, 452, 293, 778, 231, 292, 1535, 1327, 938]. Errors [784, 97]. Ershov [709]. escalation [1365]. estimating [1357]. estimation [1495]. ETOL [141]. EU [1455]. EU-COST [1455]. Euclid [216]. eukaryotic [1584]. Eulerian [767]. Evaluate [263]. Evaluating [424]. Evaluation [382, 87, 100, 623, 484, 645, 849, 62, 67, 312, 638, 667]. Evaluators [579]. Even [1087]. Event [845, 662, 1512, 522, 762]. Every [777]. evolution [1496]. Evolutionary [1061, 1236, 1144, 1188, 1384, 1325]. evolutions [1160]. evolving [1061]. Exact [144, 1663, 1250]. Example [298]. examples [1038]. Exception [735]. Exclusion [818, 931, 1298, 1389]. Execution [523, 403, 882, 1025, 970, 1641, 996, 1449, 1468]. Exhaustive [1171, 1264, 233]. Existence [639, 1537]. Expansion [757]. Expansions [792, 11, 1024]. Expected [680, 830, 695]. Experience [796]. Experimentally [248]. Explaining [412]. Explanation [745]. explicit [1269]. Exploiting [1296, 1415]. Exploration [1624]. Exponential [431, 334, 530, 150, 566]. Exponentially [1204]. exponentielles [566]. EXPRESS [1578]. EXPRESS/SOS [1578]. Expressed [242]. Expressibility [802, 1036, 999]. expression [1637, 1070]. Expressions [402, 87, 885, 67, 136, 1588, 1400, 1303, 1378, 1129, 1648]. Expressive [751, 631, 570, 763, 1216]. Expressiveness [1579, 1452, 1581, 1052, 1380, 1581]. ExpSpace [1545]. ExpSpace-hard [1545]. EXSPACE [179]. Extended [180, 259, 3, 1318, 839, 145, 170, 338, 895, 1457]. Extendible [484, 159]. Extending [1152, 1090]. Extensibility [648]. Extension [52, 1605, 1138, 1149, 1080]. Extensional [608]. Extensions [589, 916, 1228, 268, 1326]. External [666, 935]. Extracting [1453]. Extraction [681]. Extremal [353].

F [760]. face [1135]. Facility [335]. factor [1442]. Factorisation [1511]. factorization [1397]. factors [1069, 1063, 1110]. Failure [878, 1603, 972, 1417]. failure-based [1417]. failure-taming [1603]. Failures [591, 723, 599, 706, 656, 674]. Fair [993, 509, 1247, 1074, 713, 833, 1304, 1565, 1136, 1428]. Fairness [1225, 450]. Families [42, 850, 171, 1190, 1662, 53]. Familles [53, 42]. Family [150, 682, 883, 550, 574, 1182]. fan [1182]. fan-out [1182]. Fast [1519, 87, 1189, 766, 134, 784, 778, 340, 44, 1035, 11, 166]. Faster [569, 1087, 1601]. Fault [479, 986, 827, 686, 190, 696, 1192]. Fault-Tolerance [686]. Fault-Tolerant [479, 986, 827]. faulty [1312]. FDs [1259]. Features [381, 1584]. Feedback [189, 753, 964]. Feedback-Coupled [189]. fermées [42, 53]. fermeture [474]. Festschrift [1590, 1619]. few [1003]. fibers [1286]. fibration [1097]. Fields [60, 166, 20]. FIFO [633]. File [675, 407, 786]. Files [566]. filters [1384]. Final [1310]. Find [639, 308]. Finding [764, 1376, 354, 908, 1043, 1358, 1415, 1629]. Fine [818, 977]. Fine-Grained [818, 977]. fingerprints [1543]. Finis [13]. Finite [514, 840, 906, 965, 405, 60, 80, 239, 240, 247, 490, 461, 92, 854, 47, 18, 297, 1614, 219, 251, 652, 369, 754, 795, 13, 1352, 28, 147, 685, 436, 724, 732, 803, 1326, 1305, 1587, 1594, 1592, 1589, 1100, 1631, 1418, 1010, 1411, 1586, 1351, 1302, 1628]. finite-state [1302]. Finitely [418, 794, 1364]. Finiteness [844]. First [305, 1369, 344, 125, 328, 1654, 1620, 1568, 1514, 1203, 1042, 1642, 1402, 1380, 1151]. First-Order [328, 1369, 1654, 1203, 1642, 1402, 1380]. Fitting [1605]. Fixed [544, 407, 762, 1436, 1500]. Fixed-Length [407]. fixed-point [1436]. Fixpoint [557]. fixpoints [1262]. flat [1657]. flexible [1546, 1468]. Flow [693, 345, 233, 234, 893, 622, 115, 321, 977, 444, 160, 787, 739, 56, 428, 375, 274, 1134, 1210, 1005]. Flow-Time [693, 622, 787]. Flowgraphs [394]. Flows [56]. Flußdiagramme [226]. FM15 [1515]. FO [1117]. foggy [1523]. Fooling [551]. forbidding [1382]. Forced [670, 1348]. forest [1457]. Forests [850]. Forgetting [926]. Fork [623]. Form [838, 324, 280, 1014, 1107, 1121]. Formal [66, 1118, 1413, 368, 203, 215, 376, 378, 63, 597, 110, 162, 975, 197, 981, 804, 879, 265, 1557, 1584, 1011, 1493, 1468, 1555, 641, 1197, 1477, 659]. Formalism [170]. Formalisms [541]. Formelübersetzung [158]. Formes [397]. Forms [59, 113, 246, 213, 139, 172, 290, 389, 33, 397]. Formula [976, 594, 158, 568, 646, 449, 832, 1113]. formula-based [1113]. Formulae [323, 275, 1038]. Formulas [825, 537, 45, 920, 1012]. Forte [342]. Fortran [913]. Forward [764, 1410, 1219]. forwards [1670]. foundation [1529, 93, 99]. Foundations [748, 1349, 1034]. Fourier [309]. Fractal [953]. Fractions [11, 1566]. fragment [1406]. Fragmentation [315]. Fragmented [600, 529]. Framework [977, 428, 1615, 1211, 988]. Frameworks [160, 739, 811, 1581, 1579]. Free [498, 491, 409, 780, 856, 982, 419, 957, 548, 918, 84, 390, 926, 366, 145, 733, 928, 324, 560, 72, 301, 33, 921, 410, 832, 264, 1164, 1520, 1387, 997, 1201, 1276, 1482, 848, 374, 1558, 377, 1157, 1256, 325, 1280, 1167, 1486, 1348, 1015]. Freedom [517, 1363]. French [258, 397, 498, 42, 53, 261, 76, 257, 306, 417, 279, 439, 474, 13, 167, 312, 342, 566]. Friedrich [1432]. Fringe [830, 1658]. Fringe-Balanced [830]. frontier [1292]. FSSP [1601, 1600]. full [1586]. fully [1001]. Function [820, 1223, 1463, 1297, 1152, 781]. Functional [771, 552, 210, 559, 952, 471, 866, 456, 1130, 1507]. Functionally [222]. Functions [351, 545, 100, 594, 17, 41, 797, 757, 858, 859, 629, 524, 455, 147, 449, 341, 1000, 1441, 245, 132]. Fundamental [152]. Fundamentals [966, 1134]. Funktionen [41]. Further [421, 512]. Fusion [1025]. Fuzzy [1527].

G [772, 223, 458, 865, 585, 712, 387, 223, 1096, 458, 585, 712, 387]. G-queue [1096]. G/G/1 [712, 387]. G/G/1/N [585]. G/G/m [223]. G/M/1 [458]. gain [1307, 1044]. Game [1469, 224, 1368, 1513, 1375]. games [1503, 1370, 1474, 1572, 1606, 1491, 1550, 1352, 1609]. GandALF [1502]. Gap [379]. gauches [258]. Gauntlet [135]. GE [1096]. gene [1496]. General [431, 592, 320, 360, 348, 1409, 149, 465, 673, 630, 183, 525, 668, 164, 481, 308, 157, 744, 1452, 1264, 1475, 892, 1437, 1211, 683, 699, 1112]. generalised [1535]. Generalization [902, 1059, 992]. Generalized [593, 1508, 269, 276, 626, 493, 442, 1121, 956, 492, 273, 601, 1087, 1220, 1636, 1128, 1029, 1663, 1293]. generalizing [1182]. Generally [600]. Generated [856, 579, 349, 718, 1364]. générateurs [306]. Generating [470, 648, 885, 780, 374, 991, 584, 1, 1061]. Generation [421, 488, 875, 215, 636, 25, 564, 975, 808, 1171, 1264, 1220, 1089, 1573, 1073]. Generative [486, 290]. Generator [563, 231, 1650]. Generators [306, 290]. Genericity [1206]. genomics [1677]. Geometric [625, 1245]. Geometrical [847]. Geometry [953, 140, 573]. German [198, 204, 7, 48, 79, 41, 93, 99, 60, 80, 39, 84, 399, 47, 18, 65, 158, 68, 101, 314, 288, 343, 303, 226, 28, 11, 166, 29, 211, 364, 36, 26, 35, 20]. gestion [261, 76, 279]. GI [1521]. GI/M/1 [1521]. Given [198, 534, 1442, 1639]. Glitch [745]. Global [2, 543]. GLR [1168, 1266]. Go [465]. Gobang [288]. Golomb [1415]. Good [213]. Goto [357, 1414]. goto-graphs [1414]. Gotos [250]. GR [1551]. Grained [818, 977]. grammaires [397]. Grammar [419, 113, 246, 139, 967, 822, 586, 543, 1082, 1132, 397]. Grammarforms [146]. Grammarhomomorphisms [146]. Grammars [460, 259, 409, 203, 488, 59, 506, 112, 179, 555, 640, 848, 376, 651, 780, 856, 982, 446, 516, 793, 229, 645, 607, 872, 252, 946, 84, 390, 366, 300, 867, 349, 486, 339, 253, 9, 851, 145, 890, 928, 238, 327, 564, 453, 324, 369, 863, 72, 480, 338, 396, 586, 638, 19, 301, 33, 131, 280, 410, 211, 931, 882, 168, 319, 264, 1388, 1164, 1118, 1231, 1593, 1111, 377, 1293, 1346, 1362, 1089, 1257]. Grammatiken [211]. Granularity [914, 505, 1309]. Graph [958, 618, 651, 856, 982, 229, 256, 692, 408, 366, 718, 383, 331, 650, 908, 574, 955, 963, 1218, 1276, 1376, 1086, 1211, 1047, 1527]. Graphen [303, 314]. Graphics [107]. Graphs [565, 353, 594, 533, 618, 982, 39, 398, 361, 964, 354, 170, 56, 814, 314, 303, 102, 131, 289, 902, 968, 569, 823, 1620, 1548, 1178, 1286, 1233, 1414, 1201, 1276, 1517, 1662, 1174, 925, 1623, 1505, 1029, 1624, 1673, 1132, 1372, 1614, 932, 1182, 1677, 1033]. Gray [1264, 1442, 1123]. Greed [770]. Greedy [508]. Greibach [73]. grid [786]. grids [1213]. Gries [577, 135, 604]. ground [746]. Group [419, 595, 1100, 560, 921, 1173, 1114, 1158]. Groups [198, 204, 555, 1627, 1462]. growth [1414]. Grundlagen [93, 99]. Gruppen [198, 204, 204]. GSM [807, 697]. Guarantee [508]. Guaranteeing [31, 1363]. guarding [1033]. guided [1591]. Guiding [1458]. Gurevich [1137].

H [760]. Habermann [841]. Hadzilacos [1389]. hairpin [1256]. hairpin-free [1256]. Hamiltonicity [1672]. Hamming [1125]. Handling [452, 748, 735]. Hansen [30]. Hard [985, 575, 1545]. Hardness [1302, 1534]. Hardware [199, 1354]. Hashing [484, 689]. Hausdorff [899]. Having [353]. HDTOL [708]. Head [92, 1418]. Heap [400, 1457, 816]. Heaps [287, 532, 1281]. Height [740, 499, 830, 584, 1283, 1104, 1644]. Heights [610]. Hennessy [1575]. Heuristic [975]. Heuristics [609, 817, 861, 1551, 1635]. Hex [343]. Hiding [717]. Hierarchical [1525, 10, 849, 753, 575, 919, 1635, 317, 744]. Hierarchical-Tree [575]. Hiérarchie [474]. Hierarchies [119, 568, 1418, 245, 464]. Hierarchy [247, 388, 967, 341, 963, 990, 1392, 1586, 1062, 474]. High [665, 110, 162, 1021, 1572, 1609, 782]. High-Level [110, 162, 1021, 1572, 1609]. high-performance [782]. Higher [466, 38, 835, 808, 1479, 1288, 1380]. Higher-Order [835, 1479, 1288, 1380]. Highly [776]. Higman [204, 204, 1059]. History [295, 362]. Hitting [799]. Hoare [132, 503, 466, 311, 371]. Hoare-like [371]. Hoare-Logic [466].

Home [491]. Homogeneous [594, 439]. homogènes [439]. Homomorphism [124]. Homomorphisms [71, 86, 78, 1376]. hop [1444]. Hopcroft [37]. Horn [351]. hot [1186]. Hotz [555, 419]. HPP [1226]. HSDPA [1350]. Huang [1636]. Huffman [1055, 380]. Hybrid [1188, 814, 1242, 1060, 1180, 1325]. Hybridization [1243]. Hyper [932]. Hypercube [902]. Hyperedge [692, 718, 1439]. Hypergraph [780, 1415]. hyperproperties [1554].

I/O [1308]. Ianov [225]. IC0901 [1455]. Ideal [1147, 1358]. ideals [1364]. Identical [765]. Identification [879, 1125]. Identities [328]. Identity [1542, 1364]. Identity-based [1542]. Idioms [404]. IEH [902]. II [204, 99, 204, 234, 715, 99, 240, 277, 69, 78, 331, 108, 1031, 720, 35, 448, 22, 437]. illustrated [1029]. ILP [1629]. ILP-instance [1629]. ILP-instances [1629]. Image [875]. Images [898, 1591]. impact [1348]. Implementation [429, 979, 134, 742, 451, 716, 407, 128, 1660, 90]. Implementations [627, 1026, 553, 641]. Implementing [965]. implicants [1116]. implicates [1116]. implication [1301]. Implicit [625]. Impossibility [800, 910]. Improved [244, 792, 1633, 1258, 341, 1189, 1676]. Improvement [130, 812]. Improves [117]. Improving [626, 426]. Impurity [1106]. In-Place [979]. In-Tree [530]. Inclusion [771, 1088, 1140]. Incomplete [952, 934]. Incompletely [554]. increase [1585]. increasing [1053, 1307]. Incremental [460, 562, 728, 237, 705, 638, 882, 1414]. incrementally [1558]. Indecision [1603]. indefinite [1641]. Indépendance [342]. Independence [98, 862, 342]. Independent [771, 143, 186, 947]. Index [1632, 684, 658, 782]. index-sequential [782]. Indexes [14]. indirect [1381]. Induction [283, 1635]. Inductive [649, 882, 1493]. Inefficiency [1375]. inevitability [596]. inexact [1270]. Inference [911, 176, 1287, 1292]. Inferring [456]. infima [1577]. Infinitary [347, 1254, 1386, 831]. Infinite [857, 950, 330, 549, 1278, 1233, 997, 1275, 1374, 1440]. Infinite-State [950, 1278]. infix [1659]. Inflectional [29]. Influence [190]. Inform [671]. Informatica [1133]. informatics [1432]. Information [592, 207, 865, 1223, 221, 387, 274, 1525, 1499, 1391, 1097, 1050, 1274, 1111, 1270, 1210, 889, 1646]. Inherently [120, 325]. inheritance [842]. Inherited [859]. Inhibition [20]. Inhibitionsfelder [20]. inhibitors [1120]. Initial [534, 855, 251]. Inner [394]. Innermost [936]. Innermost-Confluence [936]. Input [348, 447, 576, 802]. Insertion [429, 313, 1528]. insertion-deletion [1528]. Insertion-Only [429]. Insertions [680, 293]. inspired [1483, 1061]. Instability [730]. instance [1629]. instances [1629]. Instruction [1366, 1598]. INTEGER [364, 374, 773, 937, 880, 1657, 364]. INTEGER-Darstellungen [364]. INTEGER-Representation [364]. integers [1234, 1638]. Integrating [636]. integration [1391, 1197]. Integrity [661, 748, 426]. Intelligent [107]. Intensive [917]. Interacting [88]. Interaction [943]. Interactive [384, 107, 151, 1555]. intercodes [1445]. Interconnection [883]. Interdependence [960]. Interface [1535, 1630, 1430, 1434, 1570, 1610]. Interfaces [447, 1533]. Interference [604]. intermittent [596]. Internal [561]. International [1471, 1494, 1464]. Internet [974, 1177]. Interpolation [286, 62, 1507, 1458, 786]. interpolation-based [786]. Interpretation [446, 887, 811]. Interpreted [418]. Interprocedural [944]. Interruptions [581, 536]. Intersection [513, 595, 1303, 1396]. Intersections [257, 257]. Interval [1081, 1452, 1465, 1277, 1466, 1081]. interval-based [1465]. introduction [1064]. invariance [1440]. Invariant [1668, 390, 524, 596, 892]. Invariants [891, 1651, 1575, 1537]. inverse [258]. inverses [258, 781]. Investigations [555]. involution [1438]. Involving [587]. IoT [1650]. irregular [1181]. Iso [502]. Iso-Rectangles [502]. issue [1471, 1494, 1464, 1488, 1502, 1455, 1522, 1424, 1404, 1578, 1448, 1515, 1420, 1433]. issues [1520, 998]. Iterated [184, 697, 665, 1541, 1491]. Iteration [700, 259, 763, 296, 1374]. Iterative [372, 302, 906, 57, 278, 630, 178, 1474, 1126, 1656]. ITNC [948].

J [749]. J. [660]. Java [1290]. Java-like [1290]. Job [271, 31, 171, 177, 1091]. John [299]. Join [840, 623, 603, 456]. Jörn [1619]. Journeys [580]. Judgmental [1396]. Jumping [40, 1631]. Jumps [132, 17, 164]. just [1571].

Kemeny [1674]. Kernel [932, 205]. Kernel-Control [205]. Kernels [106]. Kettenbruchentwicklungen [11]. Key [681, 873, 1314]. Keys [81, 489]. Kinded [911]. kinds [1412]. Klaus [332, 1619]. Klaus-Jörn [1619]. Knapsack [462, 616]. Knuth [738]. Kolmogorov [593]. Komplexität [198, 204, 68, 198]. kontextfreier [84]. kontextsensitiven [211]. Körpern [166, 60]. Korth [760]. Kripke [1607, 1403]. Kruskal [1059]. Kutrib [1590].

L [1432, 1096]. Label [366]. labeled [737]. Labelled [1095]. labels [1183]. Laguerre [140]. LALR [488, 607, 1089, 564]. Lamplighter [1627]. Landscape [935]. Langages [417, 397, 498, 42, 53, 257, 167]. Lange [1619]. Language [66, 445, 336, 466, 58, 465, 232, 378, 51, 74, 206, 95, 821, 216, 932, 463, 173, 227, 328, 6, 105, 755, 397, 1061, 1623, 1209, 1501, 270, 1591]. Languages [122, 345, 498, 726, 42, 53, 272, 470, 215, 49, 71, 141, 935, 856, 806, 949, 692, 63, 444, 926, 366, 948, 325, 38, 257, 550, 741, 718, 417, 97, 1229, 381, 587, 191, 197, 72, 167, 810, 684, 981, 574, 46, 64, 963, 864, 1069, 1118, 1021, 1218, 1309, 1621, 1303, 997, 1379, 1290, 1583, 1017, 1289, 1328, 1371, 1260, 1011, 1058, 1074, 1059, 1316, 1157, 1256, 1086, 1280, 1339, 1138, 1149, 1190, 1443, 1016, 1146]. Laplacian [1662]. LAR [182]. Large [14, 793, 883, 748, 710, 441, 448]. Lattice [726, 531, 846, 1250]. Lattices [840, 998, 1282]. Law [223]. Laws [152, 672, 1182, 1547]. layer [677]. Layered [752]. Layout [969]. Lazy [888, 1230, 1552]. LBA [1231, 191, 228]. LBA-Problem [191]. LCS [612]. Leader [599]. Leaders [639]. learners [1347]. Learning [52, 938, 1493, 998, 1162, 1607]. learns [1342]. Least [24, 1367]. Left [619, 9, 327, 273, 258]. left-inverse [258]. Lemma [296, 264]. lemmas [1622, 1263]. Length [115, 876, 407, 561, 1596, 777, 1600]. Lengths [803, 772]. lens [1672]. Letter [1075, 587, 191, 1419, 1669, 1351, 777]. letter-duplicated [1669]. letters [1141]. Level [49, 665, 110, 162, 238, 337, 319, 1021, 1572, 1609]. Levin [604]. Levin-Gries [604]. Lexical [897]. lexicographically [991]. Life [590, 1432]. Life-Critical [590]. Lifetime [628]. Lifting [1326]. Like [296, 1085, 1290, 418, 371]. Limit [1182]. limited [1010, 1293]. Line [425, 694, 779, 1201, 1276, 1039, 1091]. linéaires [306]. Linear [259, 79, 935, 80, 1585, 757, 1676, 47, 612, 306, 600, 230, 344, 308, 424, 702, 1621, 1667, 1593, 1553, 1213, 1541, 1117, 1086, 1280, 1462, 1167, 1237, 1506, 1570, 1610, 1023, 1504]. Linear-bounded [1585]. Linear-size [1676]. linear-time [1462, 1506, 1570, 1610]. lineare [80, 60]. linearen [79, 47]. linearizability [1524]. Linearly [1233, 239, 240, 229, 60]. lines [1385]. Link [599]. Linked [492]. links [1336]. List [522, 1008, 255, 492, 879, 1187, 1023]. List-Arithmetic [879]. Lists [395, 876, 344, 879]. Liveness [1304, 1571, 1427, 1526, 1219, 1545]. LL [939, 293, 778, 516, 961, 280, 1249, 1252]. Load [776, 151, 655, 656]. Loading [190]. Local [504, 899, 1092]. Locality [962, 871]. localize [1348]. Locally [45]. Locations [181]. Locking [361, 914]. Log [924, 813]. LOG-Specifications [813]. Log-Structured [924]. Logic [347, 501, 1013, 476, 503, 647, 356, 466, 930, 583, 894, 614, 858, 182, 311, 657, 169, 478, 426, 282, 329, 450, 837, 1095, 1368, 1311, 1423, 1451, 1116, 1605, 1553, 1275, 1615, 1604, 1575, 1642, 1099, 1467, 1454, 1005, 1506, 1139]. logic-based [1005]. Logical [317, 447, 809, 90]. logics [1452, 1130, 1369, 1330, 1402, 1504]. Long [617]. Longest [413, 367, 1669, 423]. Look [111, 367, 897, 1672]. Look-Ahead [897]. Lookahead [369]. Looking [1523]. Looks [945]. Loop [374, 1329]. Loop-free [374]. Loops [330, 394, 1040, 1668, 1285]. Losslessness [603]. Lösung [399]. Lower [266, 728, 425, 535, 691, 462, 602, 449, 720, 291, 677, 1142, 1073]. lower-bounds-to-circuit-size [677]. LR [32, 130, 548, 946, 349, 851, 157, 326, 338, 359, 1087, 1414, 1327, 460, 237, 253, 145, 195, 480]. LR-Parsing [145]. LRU [1335]. LRU-K [1335]. LSM [924]. LSM-Tree [924]. LTL [1294, 1550, 1453, 1492, 1504]. LTS [1429]. Lucid [284]. lumpability [1484, 1616]. Lycklama [1389].

M [772, 1322, 458, 1521, 1310, 1322, 458, 865, 1291, 387, 223]. M-automata [1310]. M-nets [1291]. M. [132]. M/G/1 [458, 865, 387]. M/M/1 [1322]. Machine [693, 66, 1311, 421, 232, 787, 1020, 147, 861, 677, 1053, 1039, 1191, 998, 1162, 1643, 1375]. Machines [272, 40, 819, 65, 462, 505, 795, 554, 791, 1463, 1155, 1461, 1165, 1598, 1375, 1091]. Macro [259, 1148]. Main [710, 214]. maintainable [648]. Maintaining [1180, 719, 720, 492]. Maintenance [14, 23, 787]. Management [487, 661, 261, 76, 736, 1277, 279, 221]. Managing [1419]. manipulating [1457]. Many [317, 1393, 765]. Many-sorted [317]. Map [847]. Mapping [914]. mappings [1323]. Mark [760]. Marking [256, 255]. Markov [1489, 1096, 285, 1484, 1352, 799, 1608, 1521]. Markovian [1350]. Marktoberdorf [1133]. Martin [577, 1590]. Maschinen [65]. MAT [1347]. Matching [708, 605, 908, 477, 1637, 1675]. Matchings [354]. Mathematical [283, 126]. Matrices [852, 1544, 934]. Matrix [244, 872, 165, 1258]. matter [1069, 1381]. Max [267, 816, 1175]. Max-Entropy [267]. Maximal [828, 321, 573, 1037, 1492]. Maximizing [208]. Maximum [458, 585, 668, 683, 699, 712, 354, 908, 772, 1553, 1307]. Mealy [302]. Mean [622, 115, 1489, 1523, 972]. mean-payoff [1489]. mean-time [972]. meanings [1066]. Measure [502]. Measures [951, 197, 457, 1106, 515]. Measuring [1608]. Mechanical [705, 161]. Mechanism [231, 1097]. Median [48]. Meet [840, 1504]. Mellin [870]. membership [1597, 1653]. Membrane [1120, 1356]. Membranes [1226, 1183]. Memo [736]. Memo-Table [736]. Mémoire [76]. memoriam [709]. Memories [3, 505]. Memory [518, 82, 137, 315, 1045, 640, 118, 76, 812, 719, 710, 1630, 720, 910, 333]. Merge [924, 55]. Merge-Tree [924]. Mergeable [287, 816]. Merged [1235]. Mergesort [870, 1024, 1019]. Merging [589, 12, 532, 710]. Merlin [580]. Merlin-Randell [580]. Mesh [691, 602]. Mesh-Connected [691, 602]. meshes [1008]. Message [1154, 1151, 915, 894, 1645]. message-based [1645]. messages [1486]. Meta [238, 1254]. Meta-Controlled [238]. meta-theory [1254]. Metalanguage [122]. Metalgorithm [108]. Metatheorems [692]. Method [233, 234, 881, 201, 813, 282, 157, 280, 670, 1343, 1197]. Methodology [582, 671]. Methods [336, 82, 315, 121, 192, 292, 173, 1243, 1584, 596]. Metric [980, 1467, 1454, 621, 1451, 1093]. microprocessors [927]. Milner [1575]. min [1474, 816, 1175]. min-cost [1474]. min-max [816, 1175]. mincing [677]. Minimal [41, 1055, 949, 974, 101, 554, 1574, 1073, 41]. minimale [41, 41]. Minimally [1103]. Minimalpolynome [101]. minimization [1587, 1351]. minimize [1540]. Minimizing [622, 115, 832, 727, 1039]. Minimum [752, 1639, 853, 964, 471, 359, 1043]. Minimum-Delay [752]. Mining [1465]. misconceptions [1027]. mittlere [48]. Mixed [511, 581, 992]. mixins [1207]. MIZAR [539, 556]. MM [1096]. mobile [1356, 1207, 1041, 1209, 1180]. mobility [1483]. Modal [1425, 894, 1431, 1423, 1130, 1409, 1604, 1396]. Model [248, 1013, 445, 368, 61, 82, 137, 572, 552, 731, 129, 784, 950, 1455, 1496, 150, 796, 1491, 399, 212, 859, 151, 701, 739, 64, 956, 5, 22, 1497, 1308, 1140, 1294, 1542, 1251, 1321, 1284, 1631, 1543, 1555, 1460, 1373, 1415]. Model-checking [1491]. Modeled [126]. Modeling [675, 917, 988]. Modell [399]. Modelling [1153, 966, 1288, 156]. Models [1455, 148, 144, 686, 836, 549, 707, 937, 744, 1536, 1638, 970, 1413, 1275, 1111, 277, 927, 1386, 1124, 1177, 1567, 553, 1244, 1506, 1179]. Modes [629]. Modifiable [393]. Modifications [638]. Modified [413]. Modular [1026, 796, 674]. Modularisation [930]. Modularity [867, 999]. modulated [1096]. modules [1186]. MOMI [1207]. Monadic [538]. Monge [934]. Monitoring [748, 1510, 1665]. Monitors [128]. Monoid [957, 560, 1256]. Monoïdes [167]. Monoids [733, 928, 167]. Monotone [545, 594, 535, 160, 291, 1003, 275, 323, 1487, 1547]. Monotonic [100, 1489]. Monte [595]. Morphemes [29]. Morphic [898]. Morphisms [587, 586, 1101]. most [1437, 1043]. Motion [663]. Motzkin [1220]. Motzkins [1123]. Movement [638]. moves [1569]. Moving [1430, 1670]. MR [671]. MRE [744]. Multi [1378, 775, 468, 914, 127, 401, 337, 1351, 155, 1269, 1202, 1096, 1405, 1395, 1318, 1558, 1418]. multi-coloring [1395]. Multi-Dimensional [401, 1405]. multi-disk [1202]. Multi-Granularity [914]. multi-head [1418]. Multi-letter [1351]. Multi-Level [337]. Multi-Queue [155, 468]. Multi-relations [775]. multi-resolution [1558]. multi-server [1096]. multi-sets [775]. multi-threading [1269]. Multi-tilde-bar [1378]. multicast [1159, 1401]. Multiclass [212, 868]. Multiconnection [475]. Multidimensional [194]. Multihead [461, 551, 388, 46]. Multiparty [1478]. Multipass [85]. Multiple [500, 896, 276, 784, 271, 668, 598, 1227, 683, 699, 1444, 1601, 1521]. Multiplication [244, 766, 134, 166]. Multiplier [613]. Multiplikation [166]. Multiplikations [57]. Multiprocessing [402]. Multiprocessor [986, 222, 729, 1307]. Multiprogrammed [189]. Multiprogramming [154, 148]. MultiSet [1664]. Multisets [853]. Multitape [708]. Multitype [148]. Multivalued [236, 456, 1072, 1292, 1486]. Multiversion [779]. Multiway [792, 666, 489, 316]. Multiwriter [916]. mutex [1304]. Mutual [818, 1356, 1298, 1660, 1389]. Mutually [820]. My [1562].

N [585]. Naive [799]. Name [756, 1398]. narrowing [1518]. Nash [1365]. Natural [1591, 1605, 1166]. Nature [944]. Near [427]. Near-Optimal [427]. Nearest [788, 634]. Nearest-Neighbor [788]. Nearly [469, 109, 568]. Necessary [541, 455]. Needed [263, 309]. Negation [878]. negative [1096, 802]. Negotiation [1530]. Neighbor [788]. Neighborhood [1660, 1467]. Nested [1532, 839, 734, 825, 914, 760, 690, 786, 761]. Net [940, 872, 948, 1529, 1536, 1638, 1520, 1156, 1235]. Net-Based [948]. Nets [905, 941, 893, 743, 886, 922, 758, 904, 901, 477, 942, 457, 724, 520, 674, 1529, 1296, 1021, 1094, 1425, 1427, 633, 1604, 1267, 1545, 1291, 1383, 1349, 999, 1398, 1049, 1656, 1359, 1015]. Network [351, 438, 276, 535, 149, 150, 65, 212, 568, 147, 696, 744, 291, 1539, 519]. Networks [687, 764, 893, 495, 639, 752, 1144, 1384, 83, 796, 271, 355, 504, 153, 475, 1236, 1227, 1519, 1450, 1188, 1173, 1496, 1449, 1401, 1325, 1559, 1470, 1477, 1665]. Nicely [533, 435, 1172]. Nico [841]. Nilpotency [1595]. NLC [574, 963]. no [671]. Node [366, 56]. Nodes [963, 1236, 1182]. noisy [1198]. nominal [1482]. Non [1220, 323, 844, 757, 27, 206, 604, 553, 895, 1326, 1529, 1115, 1412, 1141, 1595, 566, 1168, 1160]. non-approximability [1115]. Non-Determinancy [27]. non-determinism [1168]. Non-deterministic [553, 1529]. non-finite [1326]. Non-Interference [604]. Non-Linear [757]. Non-monotone [323]. non-primitive [1412]. Non-Procedural [206]. Non-Real-Time [895]. non-repetitive [1141]. Non-Standard [844]. Non-Stationary [566]. non-stationnaires [566]. non-uniform [1595, 1160]. Nonatomic [1344]. nondeterminacy [1262, 1288, 1245]. Nondeterminism [518, 297, 831, 331, 1143, 1010]. Nondeterministic [347, 551, 819, 670, 1423, 1042, 1399, 1067, 78]. Nonfinite [1006]. Nonlinear [94, 1243]. Nonplanar [120]. Nonstandard [677]. Nonterminal [1362, 963]. Nonterminals [71, 86, 78, 1293]. NOR [495]. Normal [59, 838, 324, 1107, 1121]. Normalisierungsproblems [211]. Normalization [211]. normed [1036]. Note [468, 325, 703, 190, 46, 114, 228, 1397, 1328, 1338, 1334, 1146]. Notions [415, 714, 1076]. Nouns [29]. Novel [902]. NP [1356, 542, 473, 682, 575]. NP-Complete [682, 1356, 473]. NP-Hard [575]. Null [760, 690, 761]. nulled [1168]. Number [208, 57, 551, 263, 309, 455, 963, 1236, 1594, 1350, 1650, 1500, 1204, 1307]. Numbers [528, 767, 990]. numeral [1617]. Numerical [403, 1377].

O [1308]. obituary [1432]. Object [842, 1109, 907, 755, 1140, 1306, 1111, 995, 1591, 1018]. Object-Oriented [907, 755, 1140, 1306, 1111]. objectives [1405, 1606]. Objects [549, 1171]. OBSCURE [821]. Observable [1427]. Observation [939]. Observational [742, 1324]. Obstacles [663]. Obtain [100]. occasion [1424]. od [260]. Off [1571, 275]. Off-the-shelf [1571]. Offs [224]. Ogden [296]. Ogden-Like [296]. OL [213, 86, 78]. OL-Systems [86, 78]. Omega [519]. On-Line [425, 694, 779, 1201, 1276, 1039, 1091]. One [611, 1072, 376, 461, 550, 1346, 191, 1155, 1000, 1185, 1362, 1601]. One-Counter [550, 191]. one-dimensional [1601]. One-Letter [191]. One-pass [611]. One-sided [1346, 1362]. one-tape [1155]. One-Visit [376]. One-Way [461, 1000]. One-write [1072]. Ones [916, 278, 1012]. Online [1395, 1199, 1500, 1165, 1191, 1270, 1151]. Only [429, 293]. Open [50, 69]. Operating [106]. Operation [251]. Operational [845, 152, 1593, 1622, 662, 452, 411, 1143]. Operations [193, 884, 165, 309, 960, 789, 342, 866, 679, 258, 342, 258, 990, 1371, 1204, 880]. Operator [447, 1564, 1117]. Operators [751, 490, 455, 1366, 1386, 1543, 1429]. Opportunistic [769]. Optimal [469, 427, 740, 613, 620, 495, 100, 986, 235, 16, 154, 969, 502, 1165, 140, 681, 12, 1270, 619, 158, 1192, 1044, 109, 547, 568, 1460, 55, 826, 873, 489, 1037, 1352, 1033]. Optimal-Time [613]. Optimality [440]. Optimally [263, 696, 1556]. optimierende [364]. optimisation [279]. optimistic [1434]. Optimization [877, 888, 1450, 725, 287, 636, 279, 205, 698, 404, 394, 1391]. Optimizations [579, 1216]. optimize [1237]. Optimizing [919, 318, 364, 274]. Optimum [723, 24, 4, 471, 547, 414, 316]. options [1250]. OPTRAN [667]. OR-parallel [970]. Oracle [272]. orbit [1462].

Order [980, 759, 242, 855, 344, 652, 835, 789, 808, 492, 328, 642, 621, 1654, 1369, 633, 1203, 1321, 1675, 1042, 1479, 1642, 1402, 1288, 1067, 1380]. order-preserving [1675]. order-sensitive [1067]. Order-Statistic [789]. Ordered [14, 300, 549, 1457, 1443, 1599]. Ordering [738, 10]. Orders [840, 864, 1282, 1442, 1213]. organisation [927]. Organization [14, 441, 448]. Oriented [907, 558, 755, 1311, 1385, 1140, 1306, 1509, 1111]. Orthogonality [207]. other [1481, 1429]. Outermost [936]. Outermost-Confluence [936]. outfix [1659]. outline [1095]. Output [504, 802]. Overflow [442]. overhang [1157]. overhang-free [1157]. overhead [1540]. Overloading [911]. overview [1057].

P [1121, 1377]. Packed [208]. Packet [83]. Packing [425, 540, 208, 1119, 694, 214, 1199, 1279, 1483]. PAFAS [1122]. Page [54, 190]. Paged [3]. pagination [76]. Paging [248, 129, 54, 76, 212, 512]. pair [1258]. papers [1471, 1464]. paradigm [1675]. Parallel [346, 622, 915, 48, 1082, 691, 1035, 776, 883, 67, 339, 494, 505, 454, 133, 822, 616, 762, 773, 108, 937, 463, 403, 790, 968, 1249, 1252, 880, 861, 142, 970, 1592, 1505, 1336, 1057, 1675, 1067, 1023, 1033, 659, 925]. Paralleladdierwerks [48]. Parallelism [518, 831]. Parameter [538, 38, 629, 1490, 1671]. Parameterisation [804]. Parameterization [598]. Parameterized [1416, 1674, 1559, 171, 1504, 1626]. Parameters [422, 566]. paramètres [566]. Parametric [911, 1431]. Parametrized [1440]. Paramodulated [289]. paraphrasing [1111]. parents [1603]. Parikh [1544]. Parity [1513, 1666, 1550]. parity-seeking [1666]. Parsable [229, 967]. Parse [931, 1070]. Parser [460, 488, 563, 231, 1317]. Parsers [32, 130, 66, 237, 548, 349, 253, 564, 157, 195, 326, 292, 318, 359, 1519, 1089, 1168]. Parsing [402, 516, 793, 607, 145, 369, 480, 338, 961, 273, 168, 319, 138, 1087, 1637, 1588, 1257, 1266, 1249, 1252]. Part [488, 715, 253, 494, 564, 719, 720, 1089, 1420, 1433, 233, 234, 714, 93, 99, 144, 1030, 1031, 441, 448, 22]. Partage [439]. Partial [792, 980, 416, 714, 715, 41, 614, 525, 858, 47, 194, 131, 598, 1341, 1361, 737, 1441, 1321, 1270, 1597, 1646, 1628, 1600]. Partial-Predicate [614]. Partiell [47]. Partiell-linearen [47]. partielle [41]. parties [1561]. partition [1661, 1359]. partitioned [1107]. Partitioning [1003, 508]. Partitions [783, 374, 719]. partly [1091]. Pascal [268, 254, 58, 51, 95, 913, 539, 556, 6, 74]. pass [611]. Passes [382, 700]. Passing [915, 894, 629]. Past [1167, 1330]. Path [305, 705, 876, 561, 1489, 1395, 1528, 1117]. path-controlled [1528]. Pathlistings [375]. paths [1571, 1220, 1500, 270]. Pattern [94, 486, 898, 477, 1218, 1303, 1583, 1675, 777, 1367]. Patterns [824, 242, 55, 617, 1341, 1361, 1246]. Paxos [1540]. payoff [1489, 1474, 1523]. PCP [1394]. Pebble [224, 1148, 1441]. Pebbles [383]. pedagogically [1666]. Peephole [636]. perfectly [1641]. Performance [152, 624, 3, 508, 983, 623, 1551, 484, 77, 150, 812, 222, 546, 527, 868, 403, 968, 1350, 1057, 782, 515, 312]. Performances [469, 312]. performing [659].

Periodic [335, 1085, 1309, 1595]. Periodic-like [1085]. periodicity [1063, 1085]. Periods [1228]. Perlis [749]. Permanent [177]. Permitted [536]. permitting [1382, 1044]. Permutation [595, 908, 1296]. Permutationen [36]. Permutations [379, 36]. permuted [934]. Permuting [519]. Perpetual [657]. Persistence [352]. persistent [1305, 1422]. Perspective [512, 1308, 1274, 1673, 1570, 1610]. pessimistic [1434]. Petri [1609, 872, 1529, 1536, 1638, 1296, 941, 893, 743, 1021, 1094, 1520, 886, 1425, 922, 1427, 892, 758, 1572, 904, 1267, 1545, 1156, 1235, 1398, 1049, 942, 457, 520, 674, 1359, 1015]. Petrovich [709]. Phase [276, 707]. Phenomenon [745]. Physical [441, 448]. pi [1166, 1208, 1209]. pi-calculus [1166]. Picture [1273, 1443]. Pieces [208]. Pipelines [1153, 523]. Place [979]. Plain [808]. Planar [565, 314, 303, 1620, 925]. planaren [314, 303]. Plane [533]. Planning [1468, 663]. plants [1557]. Point [951, 762, 1436, 1564, 1093]. pointer [913]. pointers [913]. Points [414, 1536, 1115, 1595]. points-to [1115]. polarizations [1183]. Policies [127]. Policy [189]. poly [1117]. Pólya [1177]. Polyhedra [513]. polyhedral [1537]. Polymorphism [472]. Polynomen [166]. Polynomial [124, 1356, 1649, 1093]. Polynomials [766, 101, 166]. Polytime [1013]. Polyvariant [511]. population [1476]. Portable [227]. Position [619]. Positive [399, 802, 1629]. positiven [399]. Possibilities [914]. Possibility [910]. Possible [829]. Post [829, 459]. Postorder [767]. Power [409, 751, 875, 376, 780, 486, 339, 570, 763, 296, 1384, 1340, 1596, 1238, 1315, 1377]. powers [1234, 1544]. Practical [666, 516, 485, 1550, 157, 298]. PRAMs [1263, 505, 1046]. Pre [840]. Pre-Orders [840]. Precedence [530, 91, 246, 85, 327, 578]. Precise [1490]. Predicate [465, 673, 614, 481, 713, 724, 358, 1027, 1161]. Predicate-Transformer [673]. Predicate/Transition [724]. predicates [1487, 1555, 1067]. Predicative [582, 671]. predicting [1584]. Prediction [391]. Predictions [3]. preemptions [1307]. Preemptive [1191, 1147, 1165, 1270, 1643]. Preface [1549, 1502, 1590, 1619, 1522, 1560, 1578, 1515]. Preference [1135]. prefetching [1202]. Prefix [1371, 548, 883, 1338]. Prefix-Correct [548]. Prefix-primitive [1371]. Prefixes [485, 1156]. Preorder [805, 983]. preorders [1632]. Prepaging [391]. Prescribed [863]. Preservation [843]. Preserving [1314, 1675, 1565]. Presortedness [783]. Prevention [177]. price [1336]. Primal [562]. Primary [686]. Primitive [855, 1379, 1412, 1530, 1338, 1371, 1185]. Primitives [180, 623, 110, 162, 373, 106]. Principle [98, 772, 1195]. Principles [336, 173, 336]. Priorities [941]. Prioritized [304]. Priority [597, 712, 151, 177, 536, 765, 1015, 816]. prisoner [1646]. private [1641]. Probabilistic [1013, 540, 370, 499, 144, 277, 1516, 1569, 1423, 1278, 1495, 1534, 1277, 1077, 1285, 1139, 1459]. Problem [818, 514, 370, 220, 922, 413, 618, 714, 83, 321, 605, 705, 399, 544, 595, 444, 462, 753, 716, 580, 612, 952, 191, 528, 68, 810, 367, 392, 459, 211, 895, 168, 1601, 228, 569, 897, 1088, 1356, 1483, 1071, 1231, 1125, 1240, 1501, 1597, 1162, 1663, 1462, 1646, 1364, 1073, 1033, 1175]. Problematic [381]. Problems [542, 370, 759, 860, 119, 287, 473, 692, 682, 788, 399, 408, 575, 418, 354, 829, 121, 754, 101, 616, 944, 698, 520, 576, 214, 1416, 1282, 1253, 1626, 1548, 1014, 925, 1165, 1011, 1270, 1669, 1301, 430, 437, 192]. Procedural [206]. Procedure [577, 176, 422, 38, 1128, 1275, 1112, 1657, 279]. Procedure-Type [422]. Procedures [678, 466, 278, 50, 69, 481, 261, 1287, 430, 437]. Process [923, 905, 785, 962, 1081, 285, 774, 329, 1102, 310, 1326, 1294, 1304, 1142, 1540, 1521, 1640]. Processes [385, 805, 639, 844, 940, 610, 10, 758, 57, 583, 796, 900, 894, 143, 110, 657, 350, 88, 558, 644, 420, 643, 808, 765, 1423, 1489, 1036, 1028, 1393, 1235, 1479, 270, 1177, 1440, 1373, 1477, 1656, 1380, 162]. processeurs [439]. Processing [971, 1366, 1053, 1643, 1470]. Processor [16, 276, 691, 127]. Processors [622, 222, 439, 403, 790, 1236, 1144, 1147, 1188, 1384, 1325]. PROCOL [755]. Product [483, 1385, 1016]. Production [102]. Productions [460, 195, 318, 963, 377]. products [1355]. Prof. [1432]. PROFIT [1035]. PROFIT/COST [1035]. Program [611, 700, 132, 653, 501, 647, 17, 34, 129, 1527, 363, 768, 218, 242, 249, 391, 604, 317, 190, 424, 205, 944, 226, 833, 1160, 882, 96, 90, 123, 1647, 1005, 1030, 1031, 659]. Program-Analysis [944]. Program-Substitution [647]. Programm [26, 35]. Programmable [18]. Programme [7]. Programmed [1231, 932, 19, 264]. programmierbare [18]. Programming [122, 930, 559, 49, 384, 58, 465, 51, 74, 63, 541, 38, 95, 97, 216, 381, 478, 1555, 981, 64, 227, 6, 411, 1069, 1021, 1509, 260]. Programs [346, 259, 347, 503, 500, 372, 511, 7, 254, 200, 117, 1224, 422, 201, 978, 627, 311, 418, 907, 494, 169, 70, 482, 609, 116, 133, 539, 556, 804, 163, 436, 567, 26, 35, 724, 443, 968, 837, 96, 265, 1457, 1536, 1638, 1495, 596, 1641, 1057, 1320, 1534, 1399, 1077, 1237, 641, 1031, 1139, 1329, 1196, 1312]. Project [1181, 603]. Project-Join [603]. projection [1275]. Projections [852, 467, 1221]. Prolog [970, 887]. promoters [1120]. promoters/inhibitors [1120]. Proof [250, 577, 678, 201, 583, 21, 330, 914, 350, 735, 216, 169, 481, 813, 133, 643, 706, 1308, 1128, 1394]. Proofs [501, 1223, 537, 604, 920, 1297, 1413, 1026, 1478]. Properties [115, 405, 1623, 483, 538, 284, 183, 936, 739, 88, 474, 450, 574, 539, 556, 1654, 1510, 1118, 1571, 1369, 596, 1392, 1438, 615, 1481, 1339, 1167, 1466, 990]. Proportional [1616]. proposition [1029]. Propositional [731, 282, 1275, 1038, 1467]. Protocol [304, 361, 1308, 1159]. Protocols [688, 755, 1476, 1559, 972]. Prototyping [981]. provably [1557]. Providing [1158]. Proving [132, 347, 728, 834, 295, 17, 34, 200, 844, 141, 796, 242, 116, 88, 539, 556, 161, 677, 596, 1287]. PS [1659]. pseudo [1481]. pseudo-bordered [1481]. Pseudopolynomial [1474]. PSO [1497]. PSPACE [288, 343]. PSPACE-Complete [288, 343]. PSPACE-vollständig [288, 343]. Publication [1151]. Publisher [1154, 1151]. pumping [1622]. pure [1334, 1146]. Purely [526]. Purpose [320, 360]. Push [640]. Push-Down [640]. Pushdown [225, 588, 665, 1644, 895, 1388, 1253, 1316, 1217, 1390, 1315]. pushdown-CD-systems [1390].

Q [1339]. QoS [1159]. QoS-Aware [1159]. Quad [81, 194, 401]. Quad- [401]. quadratic [1126]. Quadrature [108]. Quadtree [828, 1037]. qualitative [1450, 1606]. Quantification [219]. quantitative [1405, 1409, 1606]. Quantum [536, 1329, 1589, 1399, 1663, 1351, 1608, 1196, 1312]. Quasi [864]. Quasi-Orders [864]. Queries [877, 973, 1261, 1203, 1170, 1342, 1367]. Query [888, 1391, 1499, 1357, 1342, 1112]. Querying [1048, 987, 529, 837, 1007]. Queue [348, 126, 149, 585, 712, 512, 581, 536, 365, 155, 772, 1298, 1096, 816, 1322, 1350, 865, 1596, 468, 1521]. Queueing [223, 334, 592, 458, 149, 153, 566, 744]. Queues [687, 592, 271, 355, 668, 789, 868, 683, 699]. Queuing [150, 387]. Quicksort [163]. Quorum [919]. Quotient [550].

R [132]. R. [336]. RAM [1641]. Ramsey [528]. Randell [580]. Random [680, 664, 876, 462, 1177, 28, 956, 1301, 202, 1232, 1220, 1283, 1078, 1650, 1132, 1124, 1179, 1182, 1346, 1362]. Randomization [1332]. Randomized [505]. Range [294, 719, 720, 1286, 1078, 1187]. Rangierkomplexität [36]. ranging [1090]. Rank [898, 1674]. ranking [1297, 1333, 1008]. ranking-function [1297]. rankings [1135]. Rate [190]. rates [1053]. ratio [1661]. Rational [385, 42, 948, 339, 478, 307, 684, 296, 921, 1400, 1340, 1059]. rationality [1365]. rationelle [42]. re [667]. Reachability [1667, 1505, 1331, 644, 1474, 1517, 1352, 1608]. reaction [1508]. Reactive [1553, 1472, 917, 1640, 1550, 1634]. Reactivity [836]. Read [884]. Readers [188, 243]. Real [507, 378, 455, 895, 1261, 1510, 1079, 1480, 1145, 1212, 1030, 1031]. Real-Time [507, 378, 1510, 1079, 1480, 1212, 1030, 1031]. Realisierbarkeit [60, 80]. Realisierungen [47]. Realistic [981, 1516]. Realizability [80, 1553, 1492]. Realizable [60, 239, 240]. Realizations [47, 1347]. Rearrangement [337]. Reasonable [31]. Reasoning [1040, 1464, 552, 1038, 1042, 182, 1139, 1312, 1618, 1613, 1580, 1285, 1635, 1448]. Rebalancing [933, 1022, 1092]. Rebeca [1321]. Receptive [774]. Recognition [266, 404, 463, 934, 1086]. Recognizability [1218]. Recognizable [747, 949, 219, 810, 794]. Recognized [79]. Recommended [268]. reconfiguration [1548]. Reconstructed [858]. reconstruction [1652]. Record [407, 1632]. Recording [640]. Records [664, 1004]. Recovery [429, 293, 231, 279, 1192, 1540, 261]. Rectangles [502, 1119]. Recurrence [870, 847, 424, 929]. recurrences [1024]. Recursion [1094, 673, 855, 1262, 1023]. Recursive [346, 225, 269, 820, 175, 538, 1267, 278, 1223, 629, 855, 1257, 956, 892, 996, 1287]. Recursively [345, 741]. Red [1538, 1666, 1114]. Red-black [1538, 1114]. Reduced [958]. reducibility [746]. Reducing [57, 652, 1168]. Reduction [446, 285, 942, 1286, 1625, 1321, 1373]. Reductions [698, 359, 1513]. Redundancy [939]. Redundant [604]. reengineering [1422]. Reexecution [887]. Referees [909, 989, 1065, 1127, 1150, 1176, 1200, 1272, 1248]. reference [1084]. references [913]. Referential [722]. Refined [1348]. Refinement [993, 1431, 200, 1230, 1212, 721, 672, 798, 833, 1076, 846, 1025, 1345, 1517, 1079, 1054, 1145, 992]. Refinements [653, 896, 1068, 1251]. Regenerative [271, 355]. Region [194]. Regions [128]. Register [259, 421]. Registers [263, 1072]. regret [1472, 1634]. Regular [470, 488, 1482, 1617, 1289, 583, 796, 486, 253, 9, 587, 230, 847, 564, 369, 864, 1164, 1637, 1588, 1621, 1303, 1072, 1330, 1623, 1070, 1408, 1059, 1316, 1089, 1129, 1629, 1657, 1648]. Regularity [571]. Regulated [984]. regulatory [1496]. Reidar [243]. Rejection [687]. Related [370, 374, 354, 547, 450, 707, 1379, 1328, 1551, 615, 1669, 1659]. Relating [936, 1380]. Relation [415, 40, 597, 948, 410, 1282, 1231, 1014, 185, 1467]. Relational [49, 862, 966, 1284, 760, 603, 482, 609, 426, 954, 690, 455, 436, 917, 996, 1460, 761, 1193, 1034]. Relations [385, 236, 246, 839, 819, 948, 219, 952, 529, 424, 1452, 1153, 1438, 1668, 775, 1107, 1484, 999, 750, 1259]. Relationship [471]. relationships [123]. Relative [728]. Relatively [466]. Relaxed [1092, 1114, 816, 1281, 1022]. relay [1444]. relays [1444]. Reliability [590, 812, 363]. RelView [1282]. remainder [1660]. Remarks [132, 299, 336, 390, 281, 243]. Removing [232]. Reoptimization [1626]. Reorganization [414]. Repair [939, 778, 1615, 1647, 1327]. Répartie [312]. réparties [261]. repetitive [1141]. Replaceability [545]. Replacement [692, 718, 352, 1439]. Replication [1517, 1308, 1166, 1]. Reply [43]. Report [199, 105]. representable [737, 1673]. Representation [714, 199, 398, 1239, 165, 364, 186, 576, 1507, 1235, 1448]. Representations [1464, 625, 949, 753, 21, 1052]. Representative [572]. Representing [1234, 395, 724, 1456]. reprise [261, 279]. requirements [1045, 903, 1030]. Residue [520]. resilient [1603, 1556, 972]. Resolution [646, 920, 606, 138, 1558, 1453]. Resource [764, 622, 517, 770, 189, 127, 1393]. respect [1043]. Respecting [840]. Response [271, 355, 407, 868, 365, 1517]. responsiveness [1417]. restarting [1337, 1443]. restarts [1091]. Restricted [505, 929, 1182]. Restriction [557, 551]. Result [618]. Results [144, 512, 910, 1326, 1071, 1258, 1331, 802]. Retention [451]. Retracts [649]. retrial [1322, 865]. Retrieval [81, 1646, 1591]. RETRPROV [945]. Reusability [867]. Reversal [551, 1505]. reversal-bounded [1505]. reversals [1160]. Reversibility [1599, 1582, 1484]. Reversible [1592, 1546, 1485, 1463, 1598]. review [1333]. Revised [105]. Revisited [804, 1620, 1395, 1131, 1625, 1265, 1255, 1195, 1454, 641]. Revisiting [1423, 1421, 1666]. Rewrite [834, 531, 1294, 1353]. Rewriting [514, 71, 984, 704, 936, 932, 813, 754, 301, 1170, 1254, 1386, 615, 1211, 1121, 1064, 1664]. Rich [1455]. Rich-Model [1455]. Richer [1434]. Right [488, 640, 619, 253, 9, 564, 273, 1593, 1089, 1168]. Right-Boundary [640]. Ring [599]. rings [1600]. Rob [1563, 1562]. Robin [348, 365]. Robinson [702]. Robust [406, 1407, 1333, 786]. Role [161]. Root [400, 434]. rotations [1022]. Roth [760]. Round [348, 788, 365]. Round-Robin [348, 365].

Routines [326, 34]. Routing [565, 1159]. RRWW [1599]. RRWW-automata [1599]. Rule [678, 236, 267, 481]. rulers [1415]. Rules [250, 647, 577, 409, 176, 330, 216, 371, 1226, 942, 1287]. Run [776, 884]. Run-Time [776]. Running [1091, 1186]. Runs [710]. runtime [1654, 1645].

s [1561]. S. [73]. Safe [1221, 1557, 1290, 1214, 1209]. Safety [1558, 1526, 1671]. Safraless [1492]. Same [780]. Samelson [332]. sample [1574]. SAT [1057, 1226]. Satellite [521]. Satisfiability [734, 825, 1451]. satisfying [1271]. Saturated [417]. saturés [417]. Saves [827]. SB [782]. SB-tree [782]. scalability [972]. Scattered [704, 890, 1293, 1138, 1149]. SCCS [751]. Schaltkreise [7]. schedulability [1480]. Schedule [115]. Scheduled [787]. Schedulers [320, 360, 1516]. Schedules [752, 1147, 1358]. Scheduling [693, 530, 16, 827, 770, 468, 31, 787, 386, 171, 177, 439, 155, 861, 1436, 1053, 1039, 1181, 1165, 1191, 1270, 1307, 1162, 1643, 1020]. Schema [75, 1499, 1197]. Schemas [538, 473]. Schemata [93, 93]. Schematic [809]. Schemes [771, 225, 648, 985, 616, 159, 226, 947, 1175, 1215, 1192, 925]. Schnelle [11, 166]. Scholten [1027]. Schorr [256, 255]. Science [614]. Search [427, 740, 626, 313, 286, 267, 24, 4, 194, 669, 561, 109, 933, 975, 526, 817, 337, 2, 826, 873, 489, 125, 262, 316, 333, 433, 1232, 1620, 1078, 1104, 1658, 1187, 1092, 1182]. Searches [194]. Searching [294, 625, 434, 484, 98]. searchlight [1033]. Second [855, 808, 642, 1655, 242]. second-Order [242]. Secondary [719, 720, 333]. Secret [1641]. Secret-shared [1641]. secure [1641, 1487, 1398]. seeking [1666]. Selected [1471, 1464, 1301]. Selection [874, 404, 790, 1261, 1444, 1056]. Selective [301]. Self [1164, 886, 1173, 730, 156, 1332, 1660, 1559]. Self-embedded [1164]. Self-Stabilization [730, 1332]. Self-Stabilizing [886, 1173, 1660, 1559]. SEMANOL [122]. Semantic [336, 661, 834, 284, 326, 711, 843, 173, 1034, 96, 889, 1237]. Semantics [122, 845, 980, 445, 503, 893, 518, 678, 1306, 472, 962, 844, 662, 452, 904, 465, 632, 673, 63, 597, 1066, 859, 110, 162, 657, 478, 481, 717, 558, 835, 174, 310, 510, 706, 674, 621, 357, 1436, 1021, 1001, 1143, 1383, 1018, 1363, 1417, 1640]. Semaphore [180]. Semi [1295, 1203, 1165, 1270]. semi-algebraic [1203]. semi-online [1165, 1270]. Semi-synchronous [1295]. Semigroup [534]. Semigroups [239, 240, 544, 794]. sensing [1586]. Sensitive [59, 704, 46, 64, 103, 211, 1392, 1190, 1067]. Sensitivity [114]. Sentential [33]. Separation [958, 383, 1226]. Separatoren [314]. Separators [314]. Sequence [920, 1003, 1366, 1048, 1583, 1639, 1240]. Sequences [783, 453, 1645, 1359]. Sequential [639, 40, 10, 797, 350, 282, 420, 944, 554, 258, 1310, 1079, 1077, 782, 1137, 1139, 84]. sequentialization [1413]. Sequentielle [84]. séquentielles [258]. Sequnces [28]. Serial [1675]. Serializability [601]. Series [747, 339, 296, 1505, 1347, 1340, 1033]. series-parallel [1033]. Server [431, 581, 868, 1096]. Servers [668, 1350, 683, 699]. Service [335, 530, 276, 348, 153, 536, 1363]. services [1158]. Session [1612, 1478, 1208, 1580, 1479]. Session-based [1612]. sessions [1546]. Set [625, 508, 714, 715, 502, 471, 478, 510, 136, 737, 1598, 1359, 1629]. Set-Partitioning [508]. Sets [593, 41, 951, 985, 748, 964, 307, 230, 520, 467, 1116, 1203, 775, 1481, 1093, 1221, 1665]. settings [1215]. Several [299, 241, 660]. shape [1652]. Shared [518, 915, 1630, 910, 1641]. Shared-Memory [518]. Sharing [82, 888, 276, 107, 656]. Sharp [678]. shelf [1571]. Short [860, 537, 920]. shortest [1489]. shrinkage [1414]. Shuffle [960, 1006, 1016]. Side [164]. sided [1346, 1362, 1533]. Sieve [562, 412]. signal [1615]. Signature [1486, 1487]. Signature-free [1486]. signatures [1542]. Silberschatz [760]. silent [1569, 1045]. similar [423]. similarity [1568]. Simon [1663]. Simple [557, 958, 1650, 778, 916, 27, 361, 1174, 1449, 1, 1470, 1652, 1142]. Simplest [829]. simplification [1389]. Simplified [473, 1676]. simulating [1184]. Simulation [285, 271, 355, 504, 522, 762, 937, 1370, 1403]. Simulations [505, 1219]. simultanée [76]. Simultaneous [495, 76, 819, 1025]. Singers [1601]. Single [693, 421, 1053, 149, 787, 581, 868, 318, 1202, 1039, 1162, 1643]. single- [1202]. single-machine [1162]. Single-Queue [149]. Single-Server [868]. Site [686]. Situational [381]. Situational-Calculus [381]. six [777]. Size [594, 275, 323, 54, 568, 449, 489, 677, 1585, 1676, 1597, 1124, 1179]. Size-depth [275, 323]. Sized [881, 694]. Skip [876, 1187]. Slack [386]. slender [1058]. slenderness [1011]. Slices [768]. Slicings [573]. SLR [607, 578]. Small [593, 1325, 880]. smallest [1600]. SMT [1495, 1573]. SMT-based [1573]. Software [589, 974, 1385, 1153]. Solution [818, 399, 308, 1155, 1548]. Solutions [705, 829, 1600]. solve [1282, 1474]. Solved [82]. Solves [513]. Solving [1356, 1572, 1268, 1226, 1483, 1548, 1033, 1609]. Some [180, 1071, 625, 1412, 188, 405, 413, 283, 71, 759, 1438, 408, 1339, 587, 281, 127, 754, 707, 459, 243, 515, 1369, 1371, 1659, 430, 437]. Song [1601]. sophisticated [1652]. Sort [979]. Sorted [395, 710, 317]. Sortierproblemen [68]. Sorting [299, 241, 691, 660, 853, 602, 703, 68, 773, 937, 1274, 1056, 880]. SOS [1578]. Sound [1613, 466, 813, 371, 1666, 1618]. Space [244, 266, 272, 542, 626, 179, 513, 681, 408, 853, 612, 600, 224, 318, 727, 938, 1426]. Space-Bounded [272, 938]. Spaces [210, 715, 1125]. Spanning [1526, 125, 1043]. Sparse [165]. Sparsification [1498]. spatial [786]. Special [1471, 1494, 1464, 1063, 1488, 1455, 1424, 754, 1448, 1420, 1433, 1502, 1522, 1404, 1578, 1515]. specialization [1062]. specific [1458]. Specification [726, 445, 1161, 654, 583, 907, 821, 855, 549, 317, 116, 558, 800, 298, 729, 358, 1409, 217, 270, 1424, 1018]. Specification-Oriented [558]. Specifications [869, 915, 452, 742, 344, 813, 981, 804, 637, 1306, 1553, 1408, 1042, 1550, 1547, 992, 641, 1030, 1031]. Specified [554, 1623]. specifunction [1194]. Specify [978]. Specifying [1050]. spectrum [1569, 1526]. Speech [438]. Speeding [506, 823]. speeds [1601]. Speedup [976, 1126]. Speicheroptimale [158]. Spelling [784]. splicing [1382]. Split [493, 1636]. Splits [675]. Splitting [77, 1219, 790]. splittings [1323]. Sprachen [84]. Sprouts [711, 1177]. square [1199]. Stability [83, 1565]. stability-preserving [1565]. Stabilization [730, 1045, 1332]. Stabilizing [886, 1173, 1660, 1559]. Stable [681, 712, 1558]. Stably [853]. Stack [388, 174]. Stack-Counter [388]. Stacks [725, 631, 789, 1315]. Stage [126, 475]. Staggered [304]. Standard [445, 844, 199, 1542, 1022]. Starting [1175]. State [1426, 405, 715, 950, 247, 1447, 795, 436, 1587, 1461, 1278, 1290, 1418, 1614, 1440, 1302]. statechart [1103]. Stateless [1497, 1337, 1390, 1599]. Statement [465]. Statements [524, 357]. States [491, 1594, 1351]. Static [941, 572, 1582, 214, 1536]. Station [668, 683, 699]. stationarity [1299]. Stationary [566]. stationnaires [566]. Statistic [789]. Statistical [355]. status [1639]. Steiner [340, 716, 569, 1073]. Step [1446, 1383, 706, 1655, 1514, 1160]. Stepwise [1136, 1349]. STG [1313]. sticker [1074, 1002]. Sticky [1157]. Sticky-free [1157]. still [1504]. Stochastic [1368, 1132, 1490, 1477]. Storage [61, 315, 725, 579, 75, 442, 159, 543, 214, 903, 158]. Store [764, 600]. Store-and-Forward [764]. Stoß [379]. Strassen [134]. Strategies [736, 171, 1525, 1202, 1552, 1313]. Strategy [1405, 776]. Stratified [526, 433, 1090]. Stream [717, 971, 621]. streams [1098]. stretching [1131]. Strictness [1223, 1348]. String [366, 25, 754, 1657, 1664]. String-Rewriting [754]. string/integer [1657]. strings [1, 423, 1574]. strip [1119]. Strong [1142, 1336, 342]. Strongest [524]. Strongly [1319, 265]. Structural [806, 1143, 1545, 1166, 1299]. Structure [845, 23, 888, 666, 81, 395, 249, 65, 933, 2, 443, 411, 1311, 1054, 782, 786]. Structured [990, 170, 924, 1260, 1477]. Structures [675, 294, 269, 207, 49, 714, 93, 99, 966, 170, 401, 89, 598, 441, 448, 5, 22, 1264, 737, 1392, 1617, 1284, 1331, 1607, 1067, 1403]. Structures. [715]. Study [869, 664, 153, 1581, 1579, 1508, 1372]. Sturmian [1063]. stuttering [1195]. style [1266]. Sub [931]. Sub-Parse [931]. Subclass [877]. Subclasses [542]. subexponential [1498, 1250]. subexponential-time [1250]. Subexpressions [44]. Subrecursive [198, 204]. subregular [1384]. Subrekursive [198, 204]. Subroutines [50, 69]. Subsequence [1583, 367, 1669, 423]. subsequences [1003]. Subset [616, 1661, 1052, 1216]. subset-sum [1052]. Subsets [921]. Substitution [184, 647, 113, 748, 238, 891]. substring [1583]. Subtypes [835]. Subtyping [1208, 842, 1051, 1396]. subword [1642]. Succinct [576, 1677]. Sufficient [746, 541, 615]. Suffix [939, 1676]. Suffixes [485]. suitable [423]. Sum [616, 1661, 1052]. Sums [281, 291]. Supercube [696]. Supersets [769]. Supertype [907]. Supnick [934]. Support [487]. suprema [1577]. Surjective [852, 467, 1376]. Survey [324, 971, 1291, 1652]. Survivability [635]. swapping [1186]. Sweep [513, 788, 1675]. switch [1372]. switched [1558]. Switching [83, 36]. Symblicit [1489]. Symbol [470]. Symbolic [1527, 1456, 1244, 1410, 1552, 1573]. Symmetric [23, 639, 1559]. symmetries [1049]. Symmetry [1321]. symport [1184]. symport/antiport [1184]. Symposium [1464]. Synchronic [457].

Synchronisationsgraphen [39]. Synchronizable [588]. Synchronization [623, 143, 373, 822, 420, 443, 1601, 1628, 39]. Synchronized [385, 854, 1129, 1080, 1648]. Synchronizing [30, 15, 43]. Synchronous [1269, 1561, 1436, 1295]. synchrony [1158]. SYNT [1549, 1488]. Syntactic [747, 890, 928, 1256, 167]. syntactiques [167]. Syntax [91, 8, 859, 97]. Syntax-Directed [859]. Syntax-Errors [97]. Synthesis [299, 1422, 1488, 241, 922, 1554, 660, 703, 89, 1557, 1525, 1520, 1473, 1406, 1490, 1405, 1553, 1551, 1408, 1558, 781, 1472, 1634, 1552, 1493, 1550, 1559, 1404, 1555, 1354, 1671, 1492]. Synthesized [859]. synthesizes [1103]. Synthesizing [1407, 1556, 1459, 1603]. System [945, 82, 137, 152, 592, 1225, 117, 126, 591, 957, 212, 407, 279, 106, 643, 706, 790, 1413, 1584, 1189, 1028, 1615, 270, 995, 360]. Systematic [524, 1343]. système [279]. systèmes [261, 474]. Systems [993, 1494, 479, 223, 514, 661, 368, 491, 834, 261, 334, 986, 470, 571, 320, 552, 896, 624, 16, 509, 220, 820, 875, 531, 608, 662, 559, 71, 715, 93, 99, 458, 950, 590, 54, 144, 849, 776, 597, 189, 936, 151, 9, 222, 546, 352, 52, 847, 652, 814, 86, 78, 798, 1242, 754, 1226, 474, 822, 684, 450, 566, 45, 387, 833, 656, 917, 156, 765, 357, 1531, 1183, 1356, 1382, 1243, 411, 1508, 1431, 1305, 1422, 1426, 1407, 1592, 1120, 1294, 1278, 1202, 1490, 1667, 1097, 1122, 1061, 1101, 1511, 1050, 1617, 1409, 1512, 1082]. systems [1528, 1184, 1435, 1132, 277, 1558, 1480, 615, 1002, 1060, 1292, 1550, 1121, 1212, 1285, 1537, 1486, 1390, 104, 1067, 1645, 1113, 1302, 1030, 1031, 1396, 1244, 1080, 1377, 1136, 1245, 1076, 1459]. Systolic [432, 801, 806, 627, 705, 497, 1098].

t [772]. tâches [439]. Table [736]. Tableau [731]. Tableau-Based [731]. Tabular [441, 448]. Tabulation [881]. TACAS [1494]. tag [474]. tag-systèmes [474]. Tag-Systems [474]. taming [1603]. Tandem [592]. Tape [191, 1155]. Task [479, 523, 439]. Tasking [735]. Tasks [530, 468, 155, 1333]. TCSP [845]. teacher [1103]. Team [1604]. Teams [863]. Technique [429, 442, 350, 279, 133, 424, 358, 578]. Techniques [261, 625, 285, 1321]. Teil [93, 99]. template [1518]. Temporal [476, 1464, 748, 713, 450, 837, 1451, 1130, 1465, 1450, 1116, 1330, 1553, 1275, 1615, 1624, 1449, 1099, 1167, 1470, 1454, 1448, 1453, 1506, 1504]. temps [566]. Tennent [336]. Term [780, 936, 615]. Term-Generating [780]. Terminal [139, 1082, 1182, 1033]. Terminals [107]. Termination [834, 372, 254, 632, 111, 418, 1399, 169, 713, 1297]. Test [571, 1073]. Testable [613]. Testing [973, 415, 1345, 77, 363, 900, 218, 1653, 527, 396, 96, 1203, 1565]. Tests [506]. Text [848, 485, 423]. Their [415, 483, 379, 524, 192, 209, 1463, 1231, 1378, 40, 588, 951, 1441, 746, 121]. them [1603]. Theorem [204, 547, 388, 161, 1305, 1126, 1137]. Theorems [247, 800, 1059, 1348]. Theoretic [592, 331, 574, 510, 387, 1097, 1274, 865]. Theoretical [748, 408, 726, 1335]. Theorie [93, 99, 47]. Theories [605, 63, 1409, 1430]. Theory [1471, 66, 664, 93, 99, 384, 94, 363, 872, 249, 774, 631, 47, 440, 735, 52, 344, 975, 912, 601, 642, 411, 1529, 1309, 1006, 1535, 1493, 1254, 1647, 1324, 1299, 1083, 1032, 1197]. thinness [1011]. threading [1269]. threads [1253]. Three [1185, 475, 457, 1416, 1655, 1514, 1561]. Three-element [1185, 1514]. Three-Stage [475]. three-word [1655]. Thrifty [523]. Throughput [1227, 521, 1643]. Thrown [135]. Thue [652]. tier [1281]. tilde [1378]. TIME [1464, 693, 244, 613, 476, 622, 82, 552, 507, 48, 115, 626, 348, 1175, 107, 691, 378, 776, 681, 619, 462, 787, 669, 386, 224, 799, 536, 566, 868, 365, 942, 463, 211, 663, 895, 727, 1654, 1356, 1510, 1451, 1309, 1000, 648, 1250, 1538, 1300, 1039, 1079, 1480, 1462, 1145, 1212, 1486, 1562, 1352, 1093, 972, 1030, 1031, 1506, 1570, 1610]. TIME-2014 [1464]. time-bounded [1352]. Time-Sharing [107]. Time-Space [224, 727]. TIME’13 [1448]. Timed [1163, 1013, 1099, 814, 1068, 1370, 1512, 1331, 1653, 1105, 1671, 1611, 1469]. timelines [1468]. timeouts [1640]. Times [530, 271, 355, 31, 861, 1053, 277, 1643]. Timing [923]. TMs [1263]. tock} [1618, 1613]. tock}-CSP [1618, 1613]. tokens [1344]. Tolerance [686]. Tolerant [479, 986, 827, 696, 1192]. Tomita [1266]. Tomita-style [1266]. tool [1451]. Toolkit [1455]. Tools [1494, 847]. Top [8, 526, 1128]. Top-Down [8, 526, 1128]. Topologies [883]. Total [347, 525, 70, 196, 567, 1474, 1039, 80]. total-payoff [1474]. Trace [905, 810, 1417, 1363]. Trace- [1417]. trace-based [1363]. Traces [857, 1482, 1009, 1446]. Trade [224, 275]. trade-off [275]. Trade-Offs [224]. Tradeoff [323, 1297]. Trading [1183]. Traffic [790, 1584]. Train [580]. Trajectory [1169]. Trajectory-based [1169]. transactional [1180]. Transactions [914, 1271]. Transducer [1577, 803]. Transducers [665, 247, 732, 1048, 1148, 1318, 1300, 1441, 1585, 1410, 1100]. Transductions [906, 587, 42, 1239, 1295]. Transfer [984]. Transform [965, 309]. Transformation [700, 881, 424, 959, 1028, 1238, 1314]. Transformational [121, 192, 1030, 1031]. Transformationen [226]. Transformations [611, 242, 226, 1360]. Transformer [465, 673, 481, 1161]. Transformers [713, 358, 1161]. Transforming [666, 280]. Transforms [870]. transient [1398]. Transition [608, 798, 450, 833, 724, 917, 1095, 1431, 1305, 1519, 1511, 1409, 1512, 1537, 1359]. transitions [1163]. Transitive [823, 1319, 1174, 185]. Translatable [53, 9]. translatables [53]. Translating [1567]. Translation [1638, 158]. Translational [1263]. Translations [405]. Transmission [38, 1082]. Transparency [722]. Transposition [441, 448]. Traversals [500, 598]. Treatment [364]. Tree [469, 1013, 100, 530, 1101, 665, 801, 806, 874, 949, 496, 473, 98, 267, 883, 497, 263, 522, 716, 575, 669, 251, 924, 638, 685, 826, 634, 1088, 1360, 1347, 1070, 1148, 1318, 1300, 1585, 1410, 1320, 1316, 1238, 1099, 1019, 782, 1043, 1367, 515, 1073]. tree-based [1019]. Tree-systems [1101]. tree-transformation [1238]. tree-walking [1300, 1585]. Trees [611, 700, 427, 740, 792, 23, 193, 620, 305, 747, 432, 626, 666, 610, 664, 830, 969, 313, 81, 287, 24, 493, 695, 619, 4, 340, 442, 701, 194, 584, 561, 380, 478, 109, 251, 547, 221, 933, 526, 401, 719, 817, 646, 956, 477, 720, 873, 435, 489, 125, 931, 262, 316, 767, 333, 202, 679, 433, 1232, 1172, 680, 1499, 990, 1283, 1078, 235, 1636, 1104, 1538, 1500, 1055, 1658, 1666, 1639, 1059, 1204, 1092, 1114, 1124, 1179, 1182, 903]. Trellis [507, 801]. Trial [938]. Tricky [537, 920]. Trie [484, 689]. Tries [499, 1676]. Triesort [499]. true [1566, 1145]. Truly [548]. truncated [1444]. TSO [1497, 1524]. TSO-to-TSO [1524]. Tuples [987, 879, 1007]. Turing [1463, 1155, 819, 1598, 147]. Turn [31]. Turn-Around [31]. Two [425, 470, 415, 16, 30, 1281, 1441, 1279, 126, 247, 15, 43, 551, 468, 668, 238, 191, 439, 1047, 800, 155, 873, 319, 850, 1147, 1347, 1336, 374, 1602, 1633, 1165, 683, 699, 777, 403, 1033]. Two-Dimensional [425, 1279]. two-letter [777]. Two-Level [238, 319]. Two-Processor [16]. Two-Station [668, 683, 699]. Two-Symbol [470]. two-terminal [1033]. Two-tier [1281]. Two-Way [247, 551, 191, 873, 1441, 1602, 1633]. typability [1193]. Type [445, 466, 911, 1186, 422, 756, 1210, 1064, 344, 1385, 1290, 1347, 1109, 1521]. Type-based [1186, 1210]. typechecking [1300]. Typed [858, 1214, 1012]. Types [416, 406, 531, 1205, 452, 271, 821, 251, 549, 581, 835, 959, 800, 298, 642, 1478, 1208, 217, 1580, 1211, 553, 1396, 1222, 464]. typing [1348].

Überdeckungsprobleme [101]. Ultimate [220]. ultimately [1309]. Ultralinear [339]. unary [1418]. Unavoidable [824, 617]. Unbalanced [792, 666]. Unbounded [831, 1331, 1077, 1262]. uncertainty [1135, 1449, 1470]. Unconditional [984]. unconstrained [1646]. unconstructibility [1429]. Undecidability [605, 929, 1258, 1394, 1648]. undecidable [1524]. Undefinedness [501]. undirected [1614]. Unfoldability [722]. unfolding [1313]. unfoldings [1156]. Unification [605, 525, 702, 161]. Unified [215, 62, 739]. Uniform [512, 717, 929, 1220, 1014, 1165, 1263, 1595, 1375, 1160]. uniformization [1239]. Unifying [915]. Uniquely [967]. uniqueness [1140]. Unit [353, 195, 359]. UNITY [992]. univalent [1110]. Universal [302, 454, 1236, 1325]. Universality [541, 1463, 1002]. Universe [444]. Unnecessary [181]. Unrestricted [607, 444, 638]. unrevealed [1641]. Unsafe [679]. unsatisfiable [1453]. unspecified [1531]. untyped [1012]. up-to [1475]. Update [669, 866, 1538]. updates [1320, 1114]. Updating [987, 526, 1007]. Upper [24]. Upsequence [413]. urgent [1163]. urn [1124, 1179]. Use [362, 640, 987, 524, 274, 1007, 430, 437]. Useful [264]. user [1350]. users [1227]. Using [336, 500, 1548, 82, 283, 293, 978, 812, 1580, 966, 67, 907, 914, 308, 424, 955, 710, 873, 359, 1359, 772, 1356, 1536, 1499, 1507, 1251, 1111, 1435, 1558, 1327, 889, 1022, 1098, 1268, 1113, 880]. Utilization [442].

vacations [1322, 1521]. Valent [618]. Validated [248]. Validation [150, 1638]. value [1495]. Valuedness [732]. Values [760, 690, 803, 761]. Variable [915, 694, 916, 547, 489]. Variable-Sized [694]. Variablenmengen [41]. Variables [295, 362, 41, 393, 455, 123, 1611].

Variance [285, 903, 1104]. Variants [459]. Variations [86, 78, 1056]. Various [364]. varying [1350]. Vector [1215, 948, 352, 684, 520]. vectors [467]. Verification [1531, 1457, 422, 1476, 654, 483, 1241, 1060, 907, 494, 317, 89, 1435, 1320, 1230, 90, 1470, 1349, 1645, 1113, 1477, 1611]. verify [978]. Verifying [1389, 1167]. VeriSoft [1224]. Vernetzungsstruktur [65]. verschiedener [364]. Version [727]. Versus [71, 704, 92, 948, 331, 1081, 1583, 1336, 1203, 1582]. Vertex [964, 528, 1376, 1268]. vertex-surjective [1376]. Very [424]. vesicle [1584]. via [1661, 1069, 1469, 1493, 1550, 1124, 1179, 1453]. View [1261, 62, 1564, 1158]. Virtual [82, 137, 753]. Visibly [1400]. Visit [376]. visiting [1123]. vital [1043]. VLSI [613, 432, 969]. VLSI-Multiplier [613]. Vogler [1424]. vollständig [288, 343]. Vollständigkeitssatz [7]. Volumes [658]. vorgeschriebener [198].

W [1133]. Waite [256, 255]. waiting [277]. walking [1300, 1585]. walks [1301]. Walter [1424]. Watson [1485, 1589, 1631, 1586]. Wavelet [965]. Way [247, 461, 551, 191, 453, 873, 1068, 1000, 1441, 1602, 1633, 1293]. WCS [103]. WCS-Analysis [103]. Weak [169, 1356]. weakest [1194]. Weakly [64, 947]. Weight [695, 652]. Weight-Reducing [652]. Weighted [1283, 624, 906, 965, 1541, 1643, 262, 1360, 1357, 1100, 1576, 1402, 1033]. Weights [527]. Well [1260, 864]. Well-structured [1260]. WFA [953]. Wheel [412]. Whether [560]. which [1141, 117]. While [524, 1285]. while-loops [1285]. White [383]. Wijngaarden [112, 179, 516, 793]. Window [828, 1037]. windows [1523]. wire [1621]. wireless [1227, 1401]. within [169]. Without [295, 557, 393, 633, 1472, 1634, 831]. Wong [1636]. Word [514, 1673, 544, 733, 928, 898, 1655, 1141, 957, 1597, 1185]. Word-representable [1673]. Words [438, 1110, 79, 1228, 1382, 1544, 1068, 1442, 1387, 1341, 1361, 1063, 1085, 1379, 1412, 1017, 1338, 1481, 991]. workflow [1349, 1656]. workflows [1134]. working [1322, 1521]. Workload [572]. Worse [578]. Worst [620, 294, 624, 194, 669, 546, 1122]. Worst-Case [294, 194, 669, 1122]. Worte [79]. Write [884, 1072]. writer [916]. Writers [188, 243]. WS1S [1532]. WSNs [1444].

XML [1118, 1314, 1259]. XSLT [1238, 1216].

years [1568]. Yo [793]. Yo-Yo [793].

Z [775]. Zeitkomplexität [211]. Zeno [1068]. zonotopes [1649]. Zufallsfolgen [28]. Zugriffssysteme [93, 99, 99]. Zur [47, 68].

References

Mihalache:19xx:GSR

[1] V. Mihalache, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Generating strings by replication: a simple case. Acta Informatica, ?? (??):??, ???? 19xx. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). To appear (in 1999 or 2000??).

Smith:19xx:SDG

[2] Douglas R. Smith. Structure and design of global search algorithms. Acta Informatica, ??(??):??, ???? 19xx. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). To appear (in 1999 or 2000??).

Coffman:1971:PPE

[3] Edward G. Coffman and Brian Randell. Performance predictions for extended paged memories. Acta Informatica, 1(1):1–13, January 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Knuth:1971:OBS

[4] Donald E. Knuth. Optimum binary search trees. Acta Informatica, 1 (1):14–25, January 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Turski:1971:MDS

[5] Wladyslaw M. Turski. A model for data structures and its applications. I. Acta Informatica, 1(1):26–34, January 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wirth:1971:PLP

[6] Niklaus Wirth. The programming language Pascal. Acta Informatica, 1(1):35–63, January 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See comments [58, 95].

Claus:1971:VPS

[7] Volker Claus. Ein Vollständigkeitssatz für Programme und Schaltkreise. (German) [A completeness condition for programs and circuits]. Acta Informatica, 1(1):64–78, January 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Knuth:1971:TSA

[8] Donald E. Knuth. Top-down syntax analysis. Acta Informatica, 1(??): 79–110, ???? 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Langmaack:1971:ARC

[9] Hans Langmaack. Application of regular canonical systems to grammars translatable from left to right. Acta Informatica, 1(??):111–114, ???? 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1971:HOS

[10] Edsger W. Dijkstra. Hierarchical ordering of sequential processes. Acta Informatica, 1(2):115–138, October 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Reprinted in Operating Systems Techniques, C. A. R. Hoare and R. H. Perrot, Eds., Academic Press, 1972, pp. 72–93. This paper introduces the classical synchronization problem of Dining Philosophers.

Schonhage:1971:SBK

[11] A. Schönhage. Schnelle Berechnung von Kettenbruchentwicklungen. (German) [Fast calculation of expansions of continued fractions]. Acta Informatica, 1(??):139–144, ???? 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hwang:1971:OME

[12] F. K. Hwang and Shen Lin. Optimal merging of 2 elements with n elements. Acta Informatica, 1(??):145–158, ???? 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Perrin:1971:CAA

[13] Dominique Perrin and J.-F. Perrot. Congruences et automorphismes des automates finis. (French) [Congruences and automorphisms of finite automata]. Acta Informatica, 1(??):159–172, ???? 1971. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bayer:1972:OML

[14] Rudolf Bayer and Edward M. McCreight. Organization and maintenance of large ordered indexes. Acta Informatica, 1(3):173–189, February 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Also published in/as: ACM SIGFIDET 1970, pp.107–141.

Hansen:1972:CTS

[15] Per Brinch Hansen. A comparison of two synchronizing concepts. Acta Informatica, 1(3):190–199, February 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See comments [30] and reply [43].

Coffman:1972:OST

[16] Edward G. Coffman and R. L. Graham. Optimal scheduling for twoprocessor systems. Acta Informatica, 1(3):200–213, February 1972. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Clint:1972:PPJ

[17] M. Clint and C. A. R. Hoare. Program proving: Jumps and functions. Acta Informatica, 1(??):214–224, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See remarks [132].

Kaufholz:1972:PEA

[18] Gerd Kaufholz. Der programmierbare endliche Automat. (German) [The programmable finite automaton]. Acta Informatica, 1(??):225–241, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rozenberg:1972:DCP

[19] Grzegorz Rozenberg. Direction controlled programmed grammars. Acta Informatica, 1(??):242–252, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Walter:1972:IGI

[20] Hermann Walter. Inhibitionsfelder. (German) [Inhibition fields]. Acta Informatica, 1(??):253–269, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoare:1972:PCD

[21] C. Anthony R. Hoare. Proof of correctness of data representations. Acta Informatica, 1(4):271–281, November 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Turski:1972:MDS

[22] Wladyslaw M. Turski. A model for data structures and its applications. (Part II). Acta Informatica, 1(??):282–289, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bayer:1972:SBB

[23] Rudolf Bayer. Symmetric binary B-trees: Data structure and maintenance algorithms. Acta Informatica, 1(4):290–306, November 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hu:1972:LUB

[24] T. C. Hu and K. C. Tan. Least upper bound on the cost of optimum binary search trees. Acta Informatica, 1(??):307–310, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

McKellar:1972:BAS

[25] A. C. McKellar and C. K. Wong. Bounds on algorithms for string generation. Acta Informatica, 1(??):311–319, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Strassen:1972:BPG

[26] Volker Strassen. Berechnung und Programm. I. (German) [Calculation and programs. I]. Acta Informatica, 1(??):320–335, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hartmanis:1972:NDS

[27] Juris Hartmanis. On non-determinancy in simple computing devices. Acta Informatica, 1(??):336–344, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schnorr:1972:EAZ

[28] Claus-Peter Schnorr and H. Stimm. Endliche Automaten und Zufallsfolgen. (German) [finite automata and random sequnces]. Acta Informatica, 1(??):345–359, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schott:1972:AAI

[29] G. Schott. Automatic analysis of inflectional morphemes in German nouns. Acta Informatica, 1(??):360–374, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Courtois:1972:CCT

[30] P. J. Courtois, F. Heymans, and David Lorge Parnas. Comments on A Comparison of Two Synchronizing Concepts by Per Brinch Hansen. Acta Informatica, 1(??):375–376, ???? 1972. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [15, 43].

Lauesen:1973:JSG

[31] Søren Lauesen. Job scheduling guaranteeing reasonable turn-around times. Acta Informatica, 2(??):1–11, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anderson:1973:ELP

[32] T. Anderson, J. Eve, and James J. Horning. Efficient LR(1) parsers. Acta Informatica, 2(1):12–39, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Salomaa:1973:SFC

[33] Arto K. Salomaa. On sentential forms of context-free grammars. Acta Informatica, 2(??):40–49, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Clint:1973:PPC

[34] M. Clint. Program proving: Co-routines. Acta Informatica, 2(??):50– 63, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Strassen:1973:BPI

[35] Volker Strassen. Berechnung und Programm. II. (German) [Calculation and programs. II]. Acta Informatica, 2(??):64–79, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Stoss:1973:RPG

[36] H.-J. Stoß. Rangierkomplexität von Permutationen. (German) [The switching complexity of permutations]. Acta Informatica, 2(??):80–96, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See remarks [379].

Gries:1973:DAH

[37] David Gries. Describing an algorithm by Hopcroft. Acta Informatica, 2(??):97–109, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Langmaack:1973:CPP

[38] Hans Langmaack. On correct procedure parameter transmission in higher programming languages. Acta Informatica, 2(??):110–142, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Genrich:1973:SGS

[39] H. J. Genrich and Kurt Lautenbach. Synchronisationsgraphen. (German) [Synchronization graphs]. Acta Informatica, 2(2):143–161, August 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1973:SJM

[40] Karel Culík and Michael A. Arbib. Sequential and jumping machines and their relation to computers. Acta Informatica, 2(??):162–171, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrich:1973:MMV

[41] Hans-Dieter Ehrich. Minimale und m-minimale Variablenmengen für partielle Boole’sche Funktionen. (German) [Minimal and m-minimal sets of variables for partial Boolean functions]. Acta Informatica, 2(??):172– 179, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Boasson:1973:DFL

[42] Luc Boasson and Maurice Nivat. Sur diverses familles de langages fermées par transductions rationelle. (French) [On diverse families of languages closed by rational transductions]. Acta Informatica, 2(??):180–188, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hansen:1973:RCC

[43] Per Brinch Hansen. A reply to comments on A Comparison of Two Synchronizing Concepts. Acta Informatica, 2(??):189–190, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [15, 30].

Ullman:1973:FAE

[44] Jeffrey D. Ullman. Fast algorithms for the elimination of common subexpressions. Acta Informatica, 2(3):191–213, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rozenberg:1973:DSL

[45] Grzegorz Rozenberg and Aristid Lindenmayer. Developmental systems with locally catenative formulas. Acta Informatica, 2(??):214–248, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Savitch:1973:NMA

[46] Walter J. Savitch. A note on multihead automata and context-sensitive languages. Acta Informatica, 2(??):249–252, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kandzia:1973:TPL

[47] Peter Kandzia. Zur Theorie der Partiell-linearen Realisierungen endlicher Automaten. (German) [On the theory of partial linear realizations of finite automata]. Acta Informatica, 2(??):253–282, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Claus:1973:MAP

[48] Volker Claus. Die mittlere Additionsdauer eines Paralleladdierwerks. (German) [The median addition time of a parallel adder]. Acta Informatica, 2(??):283–291, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Earley:1973:RLD

[49] Jay Earley. Relational level data structures for programming languages. Acta Informatica, 2(4):293–309, December 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Langmaack:1973:POS

[50] Hans Langmaack. On procedures as open subroutines. I. Acta Informatica, 2(??):311–333, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoare:1973:ADP

[51] C. A. R. Hoare and Niklaus Wirth. An axiomatic definition of the programming language Pascal. Acta Informatica, 2(4):335–355, December 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See addenda and corrigenda [74].

Menzel:1973:ETL

[52] W. Menzel. An extension of the theory of learning systems. Acta Informatica, 2(??):357–381, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Boasson:1973:FLT

[53] Luc Boasson, J. P. Crestin, and Maurice Nivat. Familles de langages translatables et fermées par crochet. (French) [Families of translatable bracket-closed languages]. Acta Informatica, 2(??):383–393, ???? 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1973:PSD

[54] Erol Gelenbe, Paolo Tiberio, and J. C. A. Boekhorst. Page size in demand paging systems. Acta Informatica, 3(1):1–23, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schlumberger:1973:ODM

[55] Maurice Schlumberger and Jean Vuillemin. Optimal disk merge patterns. Acta Informatica, 3(1):25–35, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nahapetian:1973:NFG

[56] A. Nahapetian. Node flows in graphs with conservative flow. Acta Informatica, 3(1):37–41, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fox:1973:RNM

[57] B. L. Fox. Reducing the number of multiplikations in iterative processes. Acta Informatica, 3(1):43–45, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Habermann:1973:CCP

[58] A. Nico Habermann. Critical comments on the programming language Pascal. Acta Informatica, 3(1):47–57, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [6, 95].

Cremers:1973:NFC

[59] Armin B. Cremers. Normal forms for context-sensitive grammars. Acta Informatica, 3(1):59–73, December 31, 1973. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Eichner:1973:LRE

[60] Lutz Eichner. Lineare Realisierbarkeit endlicher Automaten über endlichen Körpern. (German) [Linearly realizable finite automata over finite fields]. Acta Informatica, 3(1):75–100, December 31, 1973. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Betteridge:1974:ASA

[61] Terry Betteridge. An analytic storage allocation model. Acta Informatica, 3(2):101–122, May 7, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Horowitz:1974:UVC

[62] Ellis Horowitz. A unified view of the complexity of evaluation and interpolation. Acta Informatica, 3(2):123–133, May 7, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoare:1974:CCF

[63] C. A. R. Hoare and Peter E. Lauer. Consistent and complementary formal theories of the semantics of programming languages. Acta Informatica, 3(2):135–153, June 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schuler:1974:WCS

[64] P. F. Schuler. Weakly context-sensitive languages as model for programming languages. Acta Informatica, 3(2):155–170, May 7, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kaufholz:1974:VMG

[65] Gerd Kaufholz. Über die Vernetzungsstruktur von Maschinen. (German) [On the network structure of machines]. Acta Informatica, 3(2):171–186, May 7, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Benson:1974:AMT

[66] David B. Benson. An abstract machine theory for formal language parsers. Acta Informatica, 3(2):187–202, May 7, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kuck:1974:BPE

[67] David J. Kuck and Yoichi Muraoka. Bounds on the parallel evaluation of arithmetic expressions using associativity and commutativity. Acta Informatica, 3(3):203–216, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paul:1974:KSG

[68] Wolfgang J. Paul and H.-J. Stoß. Zur Komplexität von Sortierproblemen. (German) [On the complexity of the sorting problem]. Acta Informatica, 3(3):217–225, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Langmaack:1974:POS

[69] Hans Langmaack. On procedures as open subroutines. II. Acta Informatica, 3(3):227–241, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Manna:1974:AAT

[70] Zohar Manna and Amir Pnueli. Axiomatic approach to total correctness of programs. Acta Informatica, 3(3):243–263, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1974:NVH

[71] Andrzej Ehrenfeucht and Grzegorz Rozenberg. Nonterminals versus homomorphisms in defining languages for some classes of rewriting systems. Acta Informatica, 3(3):265–283, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Penttonen:1974:DLC

[72] Martti Penttonen. On derivation languages corresponding to contextfree grammars. Acta Informatica, 3(3):285–291, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berstel:1974:CG

[73] Jean Berstel. Sur une conjecture de S. Greibach. Acta Informatica, 3 (3):293–295, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoare:1974:ACA

[74] C. A. R. Hoare and N. Wirth. Addenda and corrigenda to An Axiomatic Definition of the Programming Language Pascal. Acta Informatica, 3(3): 296, July 22, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [51].

Gotlieb:1974:CSS

[75] C. C. Gotlieb and Frank Wm. Tompa. Choosing a storage schema. Acta Informatica, 3(4):297–319, October 4, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1974:AAG

[76] Erol Gelenbe, Jaques Lenfant, and Dominique Potier. Analyse d’un algorithme de gestion simultanée Mémoire centrale—disque de pagination. (French) [Analysis of a simultaneous management algorithm for central memory and paging disk]. Acta Informatica, 3(4):321–345, October 4, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Garey:1974:PBS

[77] M. R. Garey and R. L. Graham. Performance bounds on the splitting algorithm for binary testing. Acta Informatica, 3(4):347–355, October 4, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nielsen:1974:NHCb

[78] Mogens Nielsen, Grzegorz Rozenberg, Arto Salomaa, and Sven Skyum. Nonterminals, homomorphisms and codings in different variations of OLsystems. II. Nondeterministic systems. Acta Informatica, 3(4):357–364, October 4, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ecker:1974:ELA

[79] K. Ecker and H. Ratschek. Eigenschaften der von linearen Automaten erkennbaren Worte. (German) [Characteristics of words recognized by linear automata]. Acta Informatica, 3(4):365–383, October 4, 1974. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Eichner:1974:TLR

[80] Lutz Eichner. Total lineare Realisierbarkeit endlicher Automaten. (German) [Total linear realizability of finite automata]. Acta Informatica, 3 (4):385–397, October 4, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Finkel:1974:QTD

[81] Raphael A. Finkel and Jon Louis Bentley. Quad trees: a data structure for retrieval on composite keys. Acta Informatica, 4(1):1–9, March 1, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Brandwajn:1974:MTS

[82] A. Brandwajn. A model of a time sharing virtual memory system solved using equivalence and decomposition methods. Acta Informatica, 4(1): 11–47, March 1, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fayolle:1974:SPB

[83] Guy Fayolle, Erol Gelenbe, Jacques Labetoulle, and D. Bastin. The stability problem of broadcast packet switching computer networks. Acta Informatica, 4(1):49–53, March 1, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hotz:1974:SAK

[84] Günter Hotz. Sequentielle Analyse kontextfreier Sprachen. (German) [Sequential analysis of context-free grammars]. Acta Informatica, 4(1): 55–75, March 1, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Khabbaz:1974:MPA

[85] Nabil A. Khabbaz. Multipass precedence analysis. Acta Informatica, 4 (1):77–85, March 1, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nielsen:1974:NHCa

[86] Mogens Nielsen, Grzegorz Rozenberg, Arto Salomaa, and Sven Skyum. Nonterminals, homomorphisms and codings in different variations of OLsystems. I. Deterministic systems. Acta Informatica, 4(1):87–106, March 1, 1974. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Breitbart:1975:AFE

[87] Yuri Breitbart and Allen Reiter. Algorithms for fast evaluation of Boolean expressions. Acta Informatica, 4(2):107–116, June 1975. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Newton:1975:PPI

[88] Glen Newton. Proving properties of interacting processes. Acta Informatica, 4(2):117–126, June 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Spitzen:1975:VSD

[89] Jay M. Spitzen and Ben Wegbreit. The verification and synthesis of data structures. Acta Informatica, 4(2):127–144, June 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Igarashi:1975:APVa

[90] Shigeru Igarashi, Ralph L. London, and David C. Luckham. Automatic program verification. I. A logical basis and its implementation. Acta Informatica, 4(2):145–182, June 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Earley:1975:APS

[91] Jay Earley. Ambiguity and precedence in syntax description. Acta Informatica, 4(2):183–192, June 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ibarra:1975:HVH

[92] Oscar H. Ibarra and Chul E. Kim. On 3-head versus 2-head finite automata. Acta Informatica, 4(2):193–200, June 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrich:1975:GTDa

[93] Hans-Dieter Ehrich. Grundlagen einer Theorie der Datenstrukturen und Zugriffssysteme. Teil I: Datenstrukturen und Schemata. (German) [Foundation of a theory of data structures and access systems. Part I. data structures and schemata]. Acta Informatica, 4(3):201–211, September 31, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gimpel:1975:NPT

[94] James F. Gimpel. Nonlinear pattern theory. Acta Informatica, 4(3):213– 229, September 31, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lecarme:1975:MCP

[95] Olivier Lecarme and Pierre Desjardins. More comments on the programming language Pascal. Acta Informatica, 4(3):231–243, September 31, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [6, 58]. An earlier version appeared in ACM SIGPLAN Notices, October, 1975.

Zislis:1975:SDC

[96] Paul M. Zislis. Semantic decomposition of computer programs: An aid to program testing. Acta Informatica, 4(3):243–269, September 31, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Levy:1975:ACS

[97] J.-P. Lévy. Automatic correction of syntax-errors in programming languages. Acta Informatica, 4(3):271–292, September 31, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Guibas:1975:PIB

[98] Leonidas J. Guibas. A principle of independence for binary tree searching. Acta Informatica, 4(4):293–298, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrich:1975:GTDb

[99] Hans-Dieter Ehrich. Grundlagen einer Theorie der Datenstrukturen und Zugriffssysteme. Teil II: Zugriffssysteme. (German) [Foundation of a theory of data structures and access systems. Part II. access systems]. Acta Informatica, 4(4):299–310, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Breitbart:1975:BBA

[100] Yuri Breitbart and Allen Reiter. A branch-and-bound algorithm to obtain an optimal evaluation tree for monotonic Boolean functions. Acta Informatica, 4(4):311–319, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paul:1975:BMU

[101] Wolfgang J. Paul. Boolesche Minimalpolynome und Überdeckungsprobleme. (German) [Boolean minimal polynomials and coverage problems]. Acta Informatica, 4(4):321–336, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rosen:1975:DGG

[102] Barry K. Rosen. Deriving graphs from graphs by applying a production. Acta Informatica, 4(4):337–357, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schuler:1975:WAC

[103] P. F. Schuler. WCS-analysis of the context-sensitive. Acta Informatica, 4(4):359–371, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nielsen:1975:ESC

[104] Mogens Nielsen. EOL systems with control devices. Acta Informatica, 4(4):373–386, ???? 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vanWijngaarden:1975:RRA

[105] Adriaan van Wijngaarden, B. J. Mailloux, J. E. L. Peck, C. H. A. Koster, M. Sintzoff, C. H. Lindsey, L. G. L. T. Meertens, and R. G. Fisker. Revised report on the algorithmic language ALGOL 68. Acta Informatica, 5(1–3):1–236, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nehmer:1975:DPC

[106] J. Nehmer. Dispatcher primitives for the construction of operating system kernels. Acta Informatica, 5(4):237–255, December 19, 1975. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Giloi:1975:IGI

[107] Wolfgang K. Giloi, J. Encarnação, and S. Savitt. Interactive graphics on intelligent terminals in a time-sharing environment. Acta Informatica, 5(4):257–271, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rice:1975:PAA

[108] John R. Rice. Parallel algorithms for adaptive quadrature II metalgorithm correctness. Acta Informatica, 5(4):273–285, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mehlhorn:1975:NOB

[109] Kurt Mehlhorn. Nearly optimal binary search trees. Acta Informatica, 5(4):287–295, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lauer:1975:FSC

[110] P. E. Lauer and Roy H. Campbell. Formal semantics of a class of highlevel primitives for coordinating concurrent processes. Acta Informatica, 5(4):297–332, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See addenda and corrigenda [162].

Katz:1975:CLT

[111] Shmuel M. Katz and Zohar Manna. A closer look at termination. Acta Informatica, 5(4):333–352, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Deussen:1975:DCV

[112] Peter Deussen. A decidability criterion for van Wijngaarden grammars. Acta Informatica, 5(4):353–375, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ginsburg:1975:SGF

[113] Seymour Ginsburg and Edwin H. Spanier. Substitution of grammar forms. Acta Informatica, 5(4):377–386, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schuler:1975:NDC

[114] P. F. Schuler. A note on degrees of context-sensitivity. Acta Informatica, 5(4):387–394, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coffman:1976:AMM

[115] E. G. Coffman, Jr. and Ravi Sethi. Algorithms minimizing mean flow time: Schedule length properties. Acta Informatica, 6(1):1–14, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). ctr127.

Neuhold:1976:SPC

[116] Erich J. Neuhold and T. Weller. Specification and proving of command programs. Acta Informatica, 6(1):15–40, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Darlington:1976:SWA

[117] John L. Darlington and Rod M. Burstall. A system which automatically improves programs. Acta Informatica, 6(1):41–60, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ershov:1976:AMA

[118] A. P. Ershov. Axiomatics for memory allocation. Acta Informatica, 6 (1):61–75, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Galil:1976:HCP

[119] Zvi Galil. Hierarchies of complete problems. Acta Informatica, 6(1): 77–88, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Book:1976:INA

[120] Ronald V. Book and Ashok K. Chandra. Inherently nonplanar automata. Acta Informatica, 6(1):89–94, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Monien:1976:TMT

[121] Burkhard Monien. Transformational methods and their application to complexity problems. Acta Informatica, 6(1):95–108, March 31, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See corrigenda [192].

Anderson:1976:SMP

[122] E. R. Anderson, F. C. Belz, and E. K. Blum. SEMANOL(73). A metalanguage for programming the semantics of programming languages. Acta Informatica, 6(2):109–131, July 2, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Karr:1976:ARA

[123] Michael Karr. On affine relationships among variables of a program. Acta Informatica, 6(2):133–151, July 2, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Moenck:1976:APH

[124] Robert T. Moenck. Another polynomial homomorphism. Acta Informatica, 6(2):153–169, July 2, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tarjan:1976:EDS

[125] Robert Endre Tarjan. Edge-disjoint spanning trees and depth-first search. Acta Informatica, 6(2):171–185, July 2, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Franta:1976:MAC

[126] W. R. Franta. The mathematical analysis of the computer system modeled as a two stage cyclic queue. Acta Informatica, 6(2):187–209, July 2, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nutt:1976:SRA

[127] Gary J. Nutt. Some resource allocation policies in a multi associative processor. Acta Informatica, 6(3):211–225, August 10, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schmid:1976:EIC

[128] Hans Albrecht Schmid. On the efficient implementation of conditional critical regions and the construction of monitors. Acta Informatica, 6(3): 227–249, August 10, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Courtois:1976:DMP

[129] P.-J. Courtois and H. Vantilborgh. A decomposable model of program paging behaviour. Acta Informatica, 6(3):251–275, August 10, 1976. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Backhouse:1976:AAI

[130] Roland C. Backhouse. An alternative approach to the improvement of LR(k) parsers. Acta Informatica, 6(3):277–296, August 10, 1976. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schneider:1976:GPG

[131] Hans Jürgen Schneider and Hartmut Ehrig. Grammars on partial graphs. Acta Informatica, 6(3):297–316, August 10, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ashcroft:1976:RPP

[132] E. A. Ashcroft, M. Clint, and C. A. R. Hoare. Remarks on Program Proving: Jumps and Functions by M. Clint and C. A. R. Hoare. Acta Informatica, 6(3):317–318, August 10, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [17].

Owicki:1976:APT

[133] Susan Speer Owicki and David Gries. An axiomatic proof technique for parallel programs I. Acta Informatica, 6(4):319–340, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cohen:1976:ISF

[134] Jacques Cohen and Martin Roth. On the implementation of Strassen’s fast multiplication algorithm. Acta Informatica, 6(4):341–355, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1976:GTD

[135] Edsger W. Dijkstra. On a gauntlet thrown by David Gries. Acta Informatica, 6(4):357–359, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Taniguchi:1976:ACS

[136] Kenichi Taniguchi and Tadeo Kasami. An O(n) algorithm for computing the set of available expressions of D-charts. Acta Informatica, 6(4):361– 364, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Brandwajn:1976:MVM

[137] A. Brandwajn. A model of a virtual memory system. Acta Informatica, 6(4):365–386, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wharton:1976:RAP

[138] R. M. Wharton. Resolution of ambiguity in parsing. Acta Informatica, 6(4):387–395, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Maurer:1976:GFT

[139] Hermann A. Maurer and D. Wood. On grammar forms with terminal context. Acta Informatica, 6(4):397–402, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Heise:1976:OCA

[140] Werner Heise. Optimal codes, n-arcs and Laguerre geometry. Acta Informatica, 6(4):403–406, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1976:PCL

[141] Andrzej Ehrenfeucht and Grzegorz Rozenberg. On proving that certain languages are not ETOL. Acta Informatica, 6(4):407–415, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Zuczek:1976:NAP

[142] R. Zuczek. A new approach to parallel computing. Acta Informatica, 7 (1):1–13, October 15, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lamport:1976:SIP

[143] Leslie Lamport. The synchronization of independent processes. Acta Informatica, 7(1):15–34, October 15, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1976:PMC

[144] Erol Gelenbe and Richard R. Muntz. Probabilistic models of computer systems—Part I (exact results). Acta Informatica, 7(1):35–60, October 15, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Madsen:1976:LPE

[145] Ole Lehrmann Madsen and Bent Bruun Kristensen. LR-parsing of extended context free grammars. Acta Informatica, 7(1):61–73, October 15, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Walter:1976:GG

[146] H. K.-G. Walter. Grammarforms and grammarhomomorphisms. Acta Informatica, 7(1):75–93, October 15, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schnorr:1976:NCT

[147] Claus-Peter Schnorr. The network complexity and the Turing machine complexity of finite functions. Acta Informatica, 7(1):95–107, October 15, 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gaver:1976:MMM

[148] Donald P. Gaver and George Humfeld. Multitype multiprogramming models. Acta Informatica, 7(2):111–121, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1976:BSQ

[149] Erol Gelenbe and Guy Pujolle. The behaviour of a single-queue in a general queueing network. Acta Informatica, 7(2):123–136, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Giammo:1976:VCP

[150] Thomas Giammo. Validation of a computer performance model of the exponential queuing network family. Acta Informatica, 7(2):137–152, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Landwehr:1976:EPM

[151] Carl E. Landwehr. An endogenous priority model for load control in combined batch-interactive computer systems. Acta Informatica, 7(2): 153–166, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Buzen:1976:FOL

[152] Jeffrey P. Buzen. Fundamental operational laws of computer system performance. Acta Informatica, 7(2):167–182, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Labetoulle:1976:SQN

[153] Jacques Labetoulle and Guy Pujolle. A study of queueing networks with deterministic service and application to computer networks. Acta Informatica, 7(2):183–195, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Denning:1976:OM

[154] Peter J. Denning, Kevin C. Kahn, Jacques Leroudier, Dominique Potier, and Rajan Suri. Optimal multiprogramming. Acta Informatica, 7(2): 197–216, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Spirn:1976:MQS

[155] Jeffrey R. Spirn. Multi-queue scheduling of two tasks. Acta Informatica, 7(2):217–226, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Welch:1976:SCM

[156] Peter D. Welch. On the self contained modelling of DB/DC systems. Acta Informatica, 7(2):227–247, ???? 1976. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pager:1977:PGM

[157] David Pager. A practical general method for constructing LR(k) parsers. Acta Informatica, 7(3):249–268, January 20, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kern:1977:SFG

[158] Werner Kern. Speicheroptimale Formelübersetzung. (German) [Storage optimal formula translation]. Acta Informatica, 7(3):269–287, January 20, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rosenberg:1977:SSB

[159] Arnold L. Rosenberg and Larry J. Stockmeyer. Storage schemes for boundedly extendible arrays. Acta Informatica, 7(3):289–303, January 20, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kam:1977:MDF

[160] John B. Kam and Jeffrey D. Ullman. Monotone data flow analysis frameworks. Acta Informatica, 7(3):305–317, January 20, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Data Flow Analysis.

Shostak:1977:RUM

[161] Robert E. Shostak. On the role of unification in mechanical theorem proving. Acta Informatica, 7(3):319–323, January 20, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lauer:1977:ACF

[162] P. E. Lauer and Roy H. Campbell. Addenda and corrigenda: Formal Semantics of a Class of High-Level Primitives for Coordinating Concurrent Processes. Acta Informatica, 7(3):325, January 20, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [110].

Sedgewick:1977:AQP

[163] Robert Sedgewick. The analysis of Quicksort programs. Acta Informatica, 7(4):327–355, February 1, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kowaltowski:1977:AAS

[164] Tomasz Kowaltowski. Axiomatic approach to side effects and general jumps. Acta Informatica, 7(4):357–360, February 1, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

MacVeigh:1977:EDR

[165] Donal T. MacVeigh, S. J. Effect of data representation on cost of sparse matrix operations. Acta Informatica, 7(4):361–394, February 1, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schonhage:1977:SMP

[166] A. A. Schönhage. Schnelle Multiplikation von Polynomen über Körpern der Charakteristik 2. (German) [Fast multiplication of polynomials over fields of characteristic 2]. Acta Informatica, 7(4):395–398, February 1, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Perrot:1977:MSL

[167] J.-F. Perrot. Monoïdes syntactiques des langages algébriques. (French) [Syntactic monoids of algebraic languages]. Acta Informatica, 7(4):399– 413, February 1, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Watt:1977:PPA

[168] David A. Watt. The parsing problem for affix-grammars. Acta Informatica, 8(1):1–20, March 29, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Luckham:1977:PTW

[169] David C. Luckham and Norihisa Suzuki. Proof of termination within a weak logic of programs. Acta Informatica, 8(1):21–36, March 29, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Majster:1977:EDG

[170] Mila E. Majster. Extended directed graphs, a formalism for structured data and data structures. Acta Informatica, 8(1):37–59, March 29, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mitrani:1977:CPF

[171] Isi Mitrani and J. H. Hine. Complete parameterized families of job scheduling strategies. Acta Informatica, 8(1):61–73, March 29, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Maurer:1977:EF

[172] Hermann A. Maurer, Arto Salomaa, and Derick Wood. EOL forms. Acta Informatica, 8(1):75–96, March 29, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tennent:1977:LDM

[173] Robert D. Tennent. Language design methods based on semantic principles. Acta Informatica, 8(2):97–112, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See remarks [336].

Russell:1977:EBC

[174] Bruce Russell. On an equivalence between continuation and stack semantics. Acta Informatica, 8(2):113–123, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Francez:1977:BRC

[175] Nissim Francez, Boris Klebansky, and Amir Pnueli. Backtracking in recursive computations. Acta Informatica, 8(2):125–144, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ernst:1977:RIP

[176] George W. Ernst. Rules of inference for procedure calls. Acta Informatica, 8(2):145–152, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Munch-Andersen:1977:SAJ

[177] Bo Munch-Andersen and Torben U. Zahle. Scheduling according to job priority with prevention of deadlock and permanent blocking. Acta Informatica, 8(2):153–175, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Seiferas:1977:IAD

[178] Joel I. Seiferas. Iterative arrays with direct central control. Acta Informatica, 8(2):177–192, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Deussen:1977:VWG

[179] Peter Deussen and Kurt Mehlhorn. Van Wijngaarden grammars and space complexity class EXSPACE. Acta Informatica, 8(2):193–199, May 24, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Agerwala:1977:SES

[180] Tilak Agerwala. Some extended semaphore primitives. Acta Informatica, 8(3):201–220, August 9, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Donahue:1977:LCU

[181] James E. Donahue. Locations considered unnecessary. Acta Informatica, 8(3):221–242, August 9, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kroger:1977:LLA

[182] Fred Kröger. LAR: a logic of algorithmic reasoning. Acta Informatica, 8 (3):243–266, August 9, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Igarashi:1977:GPD

[183] Yoshihide Igarashi. General properties of derivational complexity. Acta Informatica, 8(3):267–283, August 9, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Asveld:1977:IDS

[184] Peter R. J. Asveld and Joost Engelfriet. Iterated deterministic substitution. Acta Informatica, 8(3):285–302, August 9, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Eve:1977:CTC

[185] J. Eve and R. Kurki-Suonio. On computing the transitive closure of a relation. Acta Informatica, 8(4):303–314, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tennent:1977:NAR

[186] Robert D. Tennent. On a new approach to representation independent data classes. Acta Informatica, 8(4):315–324, November 25, 1977. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ramsperger:1977:CAD

[187] N. Ramsperger. Concurrent access to data. Acta Informatica, 8(4):325– 334, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Conradi:1977:SCC

[188] Reidar Conradi. Some comments on Concurrent Readers and Writers. Acta Informatica, 8(4):335–340, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See remarks [243].

Kameda:1977:FCR

[189] Hisao Kameda and Calvin C. Gotlieb. A feedback-coupled resource allocation policy for multiprogrammed computer systems. Acta Informatica, 8(4):341–357, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Parent:1977:NIP

[190] Michel Parent and Dominique Potier. A note on the influence of program loading on the page fault rate. Acta Informatica, 8(4):359–370, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Monien:1977:LPD

[191] Burkhard Monien. The LBA-problem and the deterministic tape complexity of two-way one-counter languages over a one-letter alphabet. Acta Informatica, 8(4):371–382, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Monien:1977:CTM

[192] Burkhard Monien. Corrigenda: Transformational Methods and Their Application to Complexity Problems. Acta Informatica, 8(4):383–384, November 25, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [121].

Bayer:1977:COB

[193] Rudolf Bayer and Mario Schkolnick. Concurrency of operations on B-trees. Acta Informatica, 9(1):1–21, December 13, 1977. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Reprinted in M. Stonebraker, Readings in Database Sys., Morgan Kaufmann, San Mateo, CA, 1988. Also published in/as: IBM, San Jose, Res. Lab, Res. R. RJ 1791, May, 1976.

Lee:1978:WCA

[194] D. T. Lee and C. K. Wong. Worst-case analysis for region and partial region searches in multidimensional binary search trees and balanced quad trees. Acta Informatica, 9(2):23–29, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pager:1978:EUP

[195] David Pager. Eliminating unit productions from LR parsers. Acta Informatica, 9(2):31–59, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sokolowski:1978:ATC

[196] Stefan Sokolowski. Axioms for total correctness. Acta Informatica, 9 (2):61–71, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paredaens:1978:CMF

[197] Jan Paredaens and R. Vyncke. A class of measures on formal languages. Acta Informatica, 9(2):73–86, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Avenhaus:1978:SKGa

[198] Jürgen Avenhaus and Klaus Madlener. Subrekursive Komplexität bei Gruppen: I. Gruppen mit vorgeschriebener Komplexität. (German) [Subrecursive complexity on groups. I. groups of given complexity]. Acta Informatica, 9(2):87–104, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hansen:1978:RSH

[199] Wilfried J. Hansen and Hendrik Boom. The report on the standard hardware representation for ALGOL 68. Acta Informatica, 9(2):105– 119, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Correll:1978:PPC

[200] Claus H. Correll. Proving programs correct through refinement. Acta Informatica, 9(2):121–132, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Francez:1978:PMC

[201] Nissim Francez and Amir Pnueli. A proof method for cyclic programs. Acta Informatica, 9(2):133–157, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Yao:1978:RT

[202] Andrew Chi-Chih Yao. On random 2-3 trees. Acta Informatica, 9(2):159– 170, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Book:1978:CFG

[203] Ronald V. Book. On the complexity of formal grammars. Acta Informatica, 9(2):171–181, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Avenhaus:1978:SKGb

[204] Jürgen Avenhaus and Klaus Madlener. Subrekursive Komplexität bei Gruppen: II. Der Einbettungssatz von Higman für entscheidbare Gruppen. (German) [Subrecursive complexity on groups. II. Higman’s embedding theorem for decidable groups]. Acta Informatica, 9(2):183–193, April 3, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pratt:1978:PAO

[205] Terrence W. Pratt. Program analysis and optimization through kernelcontrol decomposition. Acta Informatica, 9(3):195–216, May 5, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoffmann:1978:DCC

[206] Christoph M. Hoffmann. Design and correctness of a compiler for a non-procedural language. Acta Informatica, 9(3):217–241, May 5, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cremers:1978:OIS

[207] Armin B. Cremers and Thomas N. Hibbard. Orthogonality of information structures. Acta Informatica, 9(3):243–261, May 5, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coffman:1978:PMN

[208] Edward G. Coffman, Jr., Joseph Y.-T. Leung, and D. W. Ting. Bin packing: Maximizing the number of pieces packed. Acta Informatica, 9 (3):263–271, May 5, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rosenberg:1978:DET

[209] Arnold L. Rosenberg. Data encodings and their costs. Acta Informatica, 9(3):273–292, May 5, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cremers:1978:FBD

[210] Armin B. Cremers and Thomas N. Hibbard. Functional behavior in data spaces. Acta Informatica, 9(4):293–307, June 21, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Stadel:1978:ZNK

[211] Manfred Stadel. Die Zeitkomplexität des Normalisierungsproblems bei kontextsensitiven Grammatiken. (German) [The time complexity of the normalization problem of context-sensitive grammars]. Acta Informatica, 9(4):309–329, June 21, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Krzesinski:1978:MNM

[212] Anthony E. Krzesinski and Peter Teunissen. A multiclass network model of a demand paging computer system. Acta Informatica, 9(4):331–343, June 21, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hule:1978:GOF

[213] H. Hule, Hermann A. Maurer, and Thomas Ottmann. Good OL forms. Acta Informatica, 9(4):345–353, June 21, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Warren:1978:SMS

[214] Henry S. Warren, Jr. Static main storage packing problems. Acta Informatica, 9(4):355–376, June 21, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Deussen:1978:UAG

[215] Peter Deussen. A unified approach to the generation and the acception of formal languages. Acta Informatica, 9(4):377–390, June 21, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

London:1978:PRP

[216] Ralph L. London, John V. Guttag, James J. Horning, Butler W. Lampson, J. G. Mitchell, and Gerald J. Popek. Proof rules for the programming language Euclid. Acta Informatica, 10(1):1–26, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Guttag:1978:ASA

[217] John V. Guttag and James J. Horning. The algebraic specification of abstract data types. Acta Informatica, 10(1):27–52, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Howden:1978:APT

[218] William E. Howden. Algebraic program testing. Acta Informatica, 10 (1):53–66, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Laurinolli:1978:BQR

[219] Teuvo Laurinolli. Bounded quantification and relations recognizable by finite automata. Acta Informatica, 10(1):67–78, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1978:UEP

[220] Karel Culik II. The ultimate equivalence problem for DOL systems. Acta Informatica, 10(1):79–84, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Montangero:1978:IMC

[221] Carlo Montangero, Giuliano Pacini, Maria Simi, and Franco Turini. Information management in context trees. Acta Informatica, 10(1):85–94, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Liu:1978:PAM

[222] Jane W. S. Liu and C. L. Liu. Performance analysis of multiprocessor systems containing functionally dedicated processors. Acta Informatica, 10(1):95–104, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bartsch:1978:CLG

[223] B. Bartsch and G. Bolch. A conservation law for G/G/m queueing systems. Acta Informatica, 10(1):105–109, August 4, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paul:1978:TST

[224] Wolfgang J. Paul and Robert Endre Tarjan. Time-space trade-offs in a pebble game. Acta Informatica, 10(2):111–115, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ben-Ari:1978:IPS

[225] Mordechai Ben-Ari. Ianov pushdown schemes are contained in Boolean recursive schemes. Acta Informatica, 10(2):117–125, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sabelfeld:1978:ATF

[226] V. K. Sabel′fel′d. Äquivalente Transformationen für Flußdiagramme. (German) [Equivalence transformations of program schemes]. Acta Informatica, 10(2):127–155, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wallis:1978:DPP

[227] Peter J. L. Wallis. The design of a portable programming language. Acta Informatica, 10(2):157–167, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wotschke:1978:NCC

[228] Detlef Wotschke and Celia Wrathall. A note on classes of complements and the LBA problem. Acta Informatica, 10(2):169–173, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Franck:1978:CLP

[229] Reinhold Franck. A class of linearly parsable graph grammars. Acta Informatica, 10(2):175–201, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Maxson:1978:LRS

[230] Carlton J. Maxson. Linear regular sets. Acta Informatica, 10(2):203–208, September 7, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lewi:1978:EPG

[231] J. Lewi, Karel De Vlaminck, J. Huens, and M. Huybrechts. The ELL(1) parser generator and the error recovery mechanism. Acta Informatica, 10 (3):209–228, October 2, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hehner:1978:RML

[232] Eric C. R. Hehner. On removing the machine from the language. Acta Informatica, 10(3):229–243, October 2, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Babich:1978:MADa

[233] Wayne A. Babich and Mehdi Jazayeri. The method of attributes for data flow analysis: Part I. Exhaustive analysis. Acta Informatica, 10(3): 245–264, October 2, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Babich:1978:MADb

[234] Wayne A. Babich and Mehdi Jazayeri. The method of attributes for data flow analysis: Part II. Demand analysis. Acta Informatica, 10(3): 265–272, October 2, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Choy:1978:OTC

[235] D. M. Choy and C. K. Wong. Optimal α − β trees with capacity constraint. Acta Informatica, 10(3):273–296, October 2, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Biskup:1978:CRM

[236] Joachim Biskup. On the complementation rule for multivalued dependencies in database relations. Acta Informatica, 10(3):297–305, October 2, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Celentano:1978:ILP

[237] Augusto Celentano. Incremental LR parsers. Acta Informatica, 10(4): 307–321, November 27, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Meersman:1978:TLM

[238] Robert Meersman and Grzegorz Rozenberg. Two-level meta-controlled substitution grammars. Acta Informatica, 10(4):323–339, November 27, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Eichner:1978:SLRa

[239] Lutz Eichner. The semigroups of linearly realizable finite automata I. Acta Informatica, 10(4):341–367, November 27, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Eichner:1978:SLRb

[240] Lutz Eichner. The semigroups of linearly realizable finite automata II. Acta Informatica, 10(4):369–390, November 27, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Darlington:1978:SSS

[241] John Darlington. A synthesis of several sorting algorithms. Acta Informatica, 11(1):1–30, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See remarks [299] and corrections [660].

Huet:1978:PAP

[242] Gérard P. Huet and Bernard Lang. Proving and applying program transformations expressed with second-order patterns. Acta Informatica, 11(1):31–55, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Silberschatz:1978:RSC

[243] Abraham Silberschatz and Brian Johnson. Remarks on Some Comments on Concurrent Readers and Writers by Reidar Conradi. Acta Informatica, 11(1):57–60, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [188].

Adleman:1978:ITS

[244] Leonard M. Adleman, Kellogg S. Booth, Franco P. Preparata, and Walter L. Ruzzo. Improved time and space bounds for Boolean matrix multiplication. Acta Informatica, 11(1):61–70, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

McColl:1978:CHB

[245] William F. McColl. Complexity hierarchies for Boolean functions. Acta Informatica, 11(1):71–77, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ginsburg:1978:PRG

[246] Seymour Ginsburg and Derick Wood. Precedence relations in grammar forms. Acta Informatica, 11(1):79–88, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Greibach:1978:HTT

[247] S. A. Greibach. Hierarchy theorems for two-way finite state transducers. Acta Informatica, 11(1):89–101, December 18, 1978. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Adams:1979:EVM

[248] C. Adams, Erol Gelenbe, and J. Vicard. An experimentally validated model of the paging drum. Acta Informatica, 11(2):103–117, January 3, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Jones:1979:CTD

[249] Cliff B. Jones. Constructing a theory of a data structure as an aid to program development. Acta Informatica, 11(2):119–137, January 3, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Arbib:1979:PRG

[250] Michael A. Arbib and Suad Alagić. Proof rules for gotos. Acta Informatica, 11(2):139–148, January 3, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Merzenich:1979:BOT

[251] Wolfgang Merzenich. A binary operation on trees and an initial algebra characterization for finite tree types. Acta Informatica, 11(2):149–168, January 3, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Heilbrunner:1979:DEE

[252] Stephan Heilbrunner. On the definition of ELR(k) and ELL(k) grammars. Acta Informatica, 11(2):169–176, January 3, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

LaLonde:1979:CLP

[253] Wilf R. LaLonde. Constructing LR parsers for regular right part grammars. Acta Informatica, 11(2):177–193, January 3, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coleman:1979:CTP

[254] D. Coleman and J. W. Hughes. The clean termination of Pascal programs. Acta Informatica, 11(3):195–210, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Topor:1979:CSW

[255] Rodney W. Topor. The correctness of the Schorr–Waite list marking algorithm. Acta Informatica, 11(3):211–221, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gries:1979:SWG

[256] David Gries. The Schorr–Waite graph marking algorithm. Acta Informatica, 11(3):223–232, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Latteux:1979:ILA

[257] Michel Latteux. Intersections de langages algébriques bornés. (French) [Intersections of bounded algebraic languages]. Acta Informatica, 11(3): 233–240, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Autebert:1979:OCA

[258] Jean-Michel Autebert. Opérations de cylindre et applications séquentielles gauches inverses. (French) [cylinder operations and sequential left-inverse applications]. Acta Informatica, 11(3):241–258, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Asveld:1979:ELM

[259] Peter R. J. Asveld and Joost Engelfriet. Extended linear macro grammars, iteration grammars, and register programs. Acta Informatica, 11 (3):259–285, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hehner:1979:DCO

[260] Eric C. R. Hehner. do considered od: a contribution to the programming calculus. Acta Informatica, 11(3):287–304, March 22, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bouchet:1979:PRS

[261] P. Bouchet. Procédures de reprise dans les systèmes de gestion de base de données réparties. (French) [Recovery techniques for distributed database management systems]. Acta Informatica, 11(4):305–340, May 30, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Unterauer:1979:DWB

[262] Karl Unterauer. Dynamic weighted binary search trees. Acta Informatica, 11(4):341–362, May 30, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kemp:1979:ANR

[263] R. Kemp. The average number of registers needed to evaluate a binary tree optimally. Acta Informatica, 11(4):363–372, May 30, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vanLeeuwen:1979:ULC

[264] Jan van Leeuwen. A useful lemma for context-free programmed grammars. Acta Informatica, 11(4):373–386, May 30, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vanLamsweerde:1979:FDS

[265] Axel van Lamsweerde and Michel Sintzoff. Formal derivation of strongly correct concurrent programs. Acta Informatica, 12(1):1–31, June 25, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Alt:1979:LBS

[266] Helmut Alt. Lower bounds on space complexity for contextfree recognition. Acta Informatica, 12(1):33–61, June 25, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Horibe:1979:MER

[267] Yasuichi Horibe and Tibor O. H. Nemetz. On the max-entropy rule for a binary search tree. Acta Informatica, 12(1):63–72, June 25, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Steensgaard-Madsen:1979:PCR

[268] Jørgen Steensgaard-Madsen. Pascal—clarifications and recommended extensions. Acta Informatica, 12(1):73–94, June 25, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Burton:1979:GRD

[269] Warren Burton. Generalized recursive data structures. Acta Informatica, 12(2):95–108, July 24, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lauer:1979:CSS

[270] P. E. Lauer, P. R. Torrigiani, and M. W. Shields. COSY — a system specification language based on paths and processes. Acta Informatica, 12(2):109–158, July 24, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iglehart:1979:RSR

[271] Donald L. Iglehart and Gerald S. Shedler. Regenerative simulation of response times in networks of queues with multiple job types. Acta Informatica, 12(2):159–175, July 24, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Book:1979:LAS

[272] Ronald V. Book. On languages accepted by space-bounded oracle machines. Acta Informatica, 12(2):177–185, July 24, 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Turnbull:1979:GDL

[273] C. J. M. Turnbull and E. S. Lee. Generalized deterministic left to right parsing. Acta Informatica, 12(3):187–207, October 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wilhelm:1979:CUD

[274] Reinhard Wilhelm. Computation and use of data flow information in optimizing compilers. Acta Informatica, 12(3):209–225, October 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Commentz-Walter:1979:SDT

[275] Beate Commentz-Walter. Size-depth trade-off in monotone Boolean formulae. Acta Informatica, 12(3):227–243, October 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cohen:1979:MPS

[276] J. W. Cohen. The multiple phase service network with generalized processor sharing. Acta Informatica, 12(3):245–284, October 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1979:PMC

[277] Erol Gelenbe. Probabilistic models of computer systems. II. Diffusion approximations, waiting times and batch arrivals. Acta Informatica, 12 (4):285–303, November 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hikita:1979:CRP

[278] Teruo Hikita. On a class of recursive procedures and equivalent iterative ones. Acta Informatica, 12(4):305–320, November 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mikou:1979:AOP

[279] N. Mikou and S. Tucci. Analyse et optimisation d’une procédure de reprise dans un système de gestion de données centralisées. (French) [Analysis and optimization of a recovery technique in a management system of centralized data]. Acta Informatica, 12(4):321–338, November 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Soisalon-Soininen:1979:MTG

[280] Eljas Soisalon-Soininen and Esko Ukkonen. A method for transforming grammars into LL(k) form. Acta Informatica, 12(4):339–369, November 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mehlhorn:1979:SRB

[281] Kurt Mehlhorn. Some remarks on Boolean sums. Acta Informatica, 12 (4):371–375, November 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nishimura:1979:SMP

[282] Hirokazu Nishimura. Sequential method in propositional dynamic logic. Acta Informatica, 12(4):377–400, November 1979. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1980:SBA

[283] Edsger W. Dijkstra. Some beautiful arguments using mathematical induction. Acta Informatica, 13(1):1–8, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoffmann:1980:SPL

[284] Christoph M. Hoffmann. Semantic properties of Lucid’s compute clause and its compilation. Acta Informatica, 13(1):9–20, January 1980. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Heidelberger:1980:VRT

[285] Philip Heidelberger. Variance reduction techniques for the simulation of Markov process. Acta Informatica, 13(1):21–37, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gonnet:1980:ACA

[286] Gaston H. Gonnet, Lawrence D. Rogers, and J. Alan George. An algorithmic and complexity analysis of interpolation search. Acta Informatica, 13(1):39–52, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Galil:1980:AEM

[287] Zvi Galil. Applications of efficient mergeable heaps for optimization problems on trees. Acta Informatica, 13(1):53–58, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reisch:1980:GPV

[288] Stefan Reisch. Gobang ist PSPACE-vollständig. (German) [Gobang is PSPACE-complete]. Acta Informatica, 13(1):59–66, January 1980. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Siekmann:1980:PCG

[289] Jörg H. Siekmann and Graham Wrightson. Paramodulated connection graphs. Acta Informatica, 13(1):67–86, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Maurer:1980:GGC

[290] Hermann A. Maurer, Arto K. Salomaa, and Derick Wood. On generators and generative capacity of EOL forms. Acta Informatica, 13(1):87–107, January 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wegener:1980:NLB

[291] Ingo Wegener. A new lower bound on the monotone network complexity of Boolean sums. Acta Informatica, 13(2):109–114, February 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rohrich:1980:MAC

[292] Johannes Röhrich. Methods for the automatic construction of error correcting parsers. Acta Informatica, 13(2):115–139, February 1980. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fischer:1980:ELE

[293] Charles N. Fischer, Donn R. Milton, and Sam B. Quiring. Efficient LL(1) error correction and recovery using only insertions. Acta Informatica, 13 (2):141–154, February 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bentley:1980:EWC

[294] Jon Louis Bentley and Hermann A. Maurer. Efficient worst-case data structures for range searching. Acta Informatica, 13(2):155–168, February 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Clarke:1980:PCC

[295] Edmund Melson Clarke. Proving correctness of coroutines without history variables. Acta Informatica, 13(2):169–188, February 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reutenauer:1980:OLI

[296] Christophe Reutenauer. An Ogden-like iteration lemma for rational power series. Acta Informatica, 13(2):189–197, February 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kintala:1980:ANF

[297] Chandra M. R. Kintala and Detlef Wotschke. Amounts of nondeterminism in finite automata. Acta Informatica, 13(2):199–204, February 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tompa:1980:PES

[298] Frank Wm. Tompa. A practical example of the specification of abstract data types. Acta Informatica, 13(3):205–224, March 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Barstow:1980:RSS

[299] David R. Barstow and John Darlington. Remarks on A Synthesis of Several Sorting Algorithms by John Darlington. Acta Informatica, 13 (3):225–227, March 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [241, 660].

Kastens:1980:OAG

[300] Uwe Kastens. Ordered attributed grammars. Acta Informatica, 13(3): 229–256, March 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: Bericht 7/78, Institut für Informatik II, University Karlsruhe (1978).

Rozenberg:1980:CFG

[301] Grzegorz Rozenberg and D. Wood. Context-free grammars with selective rewriting. Acta Informatica, 13(3):257–268, March 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Buning:1980:UAI

[302] Hans Kleine Büning and Lutz Priese. Universal asynchronous iterative arrays of Mealy automata. Acta Informatica, 13(3):269–285, March 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rollik:1980:APG

[303] H. A. Rollik. Automaten in planaren graphen. (German) [Automata in planar graphs]. Acta Informatica, 13(3):287–298, March 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Franta:1980:APC

[304] William R. Franta and Mark Benedict Bilodeau. Analysis of a prioritized CSMA protocol based on staggered delays. Acta Informatica, 13(4):299– 324, May 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berztiss:1980:DFT

[305] A. T. Berztiss. Depth-first K-trees and critical path analysis. Acta Informatica, 13(4):325–346, May 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Latteux:1980:GAL

[306] Michel Latteux. Sur les générateurs algébriques et linéaires. (French) [Algebraic and linear generators]. Acta Informatica, 13(4):347–363, May 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Maurer:1980:RBR

[307] Hermann A. Maurer and Maurice Nivat. Rational bijection of rational sets. Acta Informatica, 13(4):365–378, May 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Morito:1980:UBA

[308] Susumu Morito and Harvey M. Salkin. Using the Blankinship algorithm to find the general solution of a linear Diophantine equation. Acta Informatica, 13(4):379–382, May 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mescheder:1980:NAO

[309] Bernhard Mescheder. On the number of active-operations needed to compute the discrete Fourier transform. Acta Informatica, 13(4):383– 408, May 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Russell:1980:CCP

[310] Bruce Russell. Correctness of the compiling process based on axiomatic semantics. Acta Informatica, 14(1):1–20, June 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lamport:1980:HLC

[311] Leslie Lamport. The “Hoare logic” of concurrent programs. Acta Informatica, 14(1):21–37, June 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Plateau:1980:EPA

[312] Brigitte Plateau. Evaluation des performances d’un algorithme de contrôle de la cohérence d’une base de données Répartie. (French) [Performance evaluation of a concurrency control algorithm for a distributed database]. Acta Informatica, 14(1):39–62, June 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ellis:1980:CSI

[313] Carla Schlatter Ellis. Concurrent search and insertion in 2-3 trees. Acta Informatica, 14(1):63–86, June 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Philipp:1980:SPG

[314] Reiner Philipp and Ernst-Jürgen Prauß. Über Separatoren in planaren Graphen. (German) [separators in planar graphs]. Acta Informatica, 14 (1):87–106, June 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bromley:1980:MFB

[315] Allan G. Bromley. Memory fragmentation in buddy methods for dynamic storage allocation. Acta Informatica, 14(2):107–117, August 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vaishnavi:1980:OMS

[316] Vijay K. Vaishnavi, Hans-Peter Kriegel, and D. Wood. Optimum multiway search trees. Acta Informatica, 14(2):119–133, August 1980. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nakajima:1980:HPS

[317] Reiji Nakajima, Michio Honda, and Hayao Nakahara. Hierarchical program specification and verification — a many-sorted logical approach. Acta Informatica, 14(2):135–155, August 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Soisalon-Soininen:1980:SOE

[318] Eljas Soisalon-Soininen. On the space optimizing effect of eliminating single productions from LR parsers. Acta Informatica, 14(2):157–174, August 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wegner:1980:PTL

[319] Lutz Michael Wegner. On parsing two-level grammars. Acta Informatica, 14(2):175–193, August 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Casanova:1980:GPS

[320] Marco A. Casanova and Philip A. Bernstein. General purpose schedulers for database systems. Acta Informatica, 14(3):195–220, September 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See erratum [360].

Galil:1980:AMF

[321] Zvi Galil. An O(V^{5/3}E^{2/3}) algorithm for the maximal flow problem. Acta Informatica, 14(3):221–242, September 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Preliminary version in Proc. 19th FOCS, pages 231–245, 1978.

Paul:1980:A

[322] Wolfgang J. Paul, Ernst J. Prauß, and Rüdiger Reischuk. On alternation. Acta Informatica, 14(3):243–255, September 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Commentz-Walter:1980:SDT

[323] Beate Commentz-Walter and Jürgen Sattler. Size-depth tradeoff in nonmonotone Boolean formulae. Acta Informatica, 14(3):257–269, September 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nijholt:1980:SNF

[324] Anton Nijholt. A survey of normal form covers for context free grammars. Acta Informatica, 14(3):271–294, September 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kemp:1980:NDI

[325] R. Kemp. A note on the density of inherently ambiguous context-free languages. Acta Informatica, 14(3):295–298, September 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Purdom:1980:SRL

[326] Paul Walton Purdom, Jr. and Cynthia A. Brown. Semantic routines and LR(k) parsers. Acta Informatica, 14(4):299–315, October 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Moll:1980:LCP

[327] Karl-Rudolf R. Moll. Left context precedence grammars. Acta Informatica, 14(4):317–335, October 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wand:1980:FOI

[328] Mitchell Wand. First-order identities as a defining language. Acta Informatica, 14(4):337–357, October 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Originally Report 29, Computer Science Department, Indiana University, 1977.

Nishimura:1980:DCP

[329] Hirokazu Nishimura. Descriptively complete process logic. Acta Informatica, 14(4):359–369, October 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kroger:1980:IPR

[330] Fred Kröger. Infinite proof rules for loops. Acta Informatica, 14(4):371– 389, October 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paul:1980:AIG

[331] Wolfgang J. Paul and Rüdiger Reischuk. On alternation II. A graph theoretic approach to determinism versus nondeterminism. Acta Informatica, 14(4):391–403, October 1980. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1981:KS

[332] Friedrich L. Bauer, Andrei P. Ershov, M. Paul, and Alan J. Perlis. Klaus Samelson. Acta Informatica, 15(1):1–2, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wright:1981:BST

[333] William E. Wright. Binary search trees in secondary memory. Acta Informatica, 15(1):3–17, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Boxma:1981:AAE

[334] Onno J. Boxma and Alan G. Konheim. Approximate analysis of exponential queueing systems with blocking. Acta Informatica, 15(1):19–66, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baccelli:1981:ASF

[335] François Baccelli. Analysis of a service facility with periodic checkpointing. Acta Informatica, 15(1):67–81, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berry:1981:RRD

[336] Daniel M. Berry. Remarks on R. D. Tennent’s Language Design Methods Based on Semantic Principles: Algol 68, a Language Designed Using Semantic Principles. Acta Informatica, 15(1):83–98, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [173].

Purdom:1981:BML

[337] Paul Walton Purdom, Jr., Cynthia A. Brown, and Edward L. Robertson. Backtracking with multi-level dynamic search rearrangement. Acta Informatica, 15(2):99–113, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Purdom:1981:PEL

[338] Paul Walton Purdom, Jr. and Cynthia A. Brown. Parsing extended LR(k) grammars. Acta Informatica, 15(2):115–127, December 1981. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kuich:1981:CPU

[339] Werner Kuich. The characterization of parallel ultralinear grammars by rational power series. Acta Informatica, 15(2):129–139, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kou:1981:FAS

[340] L. Kou, George Markowsky, and Leonard Berman. A fast algorithm for Steiner trees. Acta Informatica, 15(2):141–145, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wegener:1981:ICH

[341] Ingo Wegener. An improved complexity hierarchy on the depth of Boolean functions. Acta Informatica, 15(2):147–152, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rodriguez:1981:IFC

[342] F. Rodriguez. Indépendance forte de certaines opérations. (French) [Strong independence of certain operations]. Acta Informatica, 15(2): 153–166, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reisch:1981:HPV

[343] Stefan Reisch. Hex ist PSPACE-vollständig. (German) [Hex is PSPACEcomplete]. Acta Informatica, 15(2):167–191, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Moore:1981:ADT

[344] Daniel J. Moore and Bruce Russell. Axiomatic data type specifications: a first order theory of linear lists. Acta Informatica, 15(3):193–207, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Araki:1981:FLE

[345] Toshiro Araki and Nobuki Tokura. Flow languages equal recursively enumerable languages. Acta Informatica, 15(3):209–217, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Apt:1981:RAP

[346] Krzysztof R. Apt. Recursive assertions and parallel programs. Acta Informatica, 15(3):219–232, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Back:1981:PTC

[347] R. J. R. Back. Proving total correctness of nondeterministic programs in infinitary logic. Acta Informatica, 15(3):233–249, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Daduna:1981:DTR

[348] Hans Daduna and R. Schassberger. A discrete-time round-robin queue with Bernoulli input and general arithmetic service time distributions. Acta Informatica, 15(3):251–263, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kemp:1981:LGG

[349] R. Kemp. LR(0) grammars generated by LR(0) parsers. Acta Informatica, 15(3):265–280, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Levin:1981:PTC

[350] Gary Marc Levin and David Gries. A proof technique for communicating sequential processes. Acta Informatica, 15(3):281–302, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anderaa:1981:EHN

[351] Stal O. Anderaa and Egon Börger. The equivalence of Horn and network complexity for Boolean functions. Acta Informatica, 15(3):303–307, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mayr:1981:PVR

[352] Ernst W. Mayr. Persistence of vector replacement systems is decidable. Acta Informatica, 15(3):309–318, June 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Booth:1981:CEA

[353] Kellogg S. Booth and Richard J. Lipton. Computing extremal and approximate distances in graphs having unit cost edges. Acta Informatica, 15(4):319–328, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lipski:1981:EAF

[354] Witold Lipski, Jr. and Franco P. Preparata. Efficient algorithms for finding maximum matchings in convex bipartite graphs and related problems. Acta Informatica, 15(4):329–346, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iglehart:1981:RSR

[355] Donald L. Iglehart and Gerald S. Shedler. Regenerative simulation of response times in networks of queues: Statistical efficiency. Acta Informatica, 15(4):347–363, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cartwright:1981:LA

[356] Robert Cartwright and Derek C. Oppen. The logic of aliasing. Acta Informatica, 15(4):365–384, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

deBruin:1981:GSS

[357] Arie de Bruin. Goto statements: Semantics and deduction systems. Acta Informatica, 15(4):385–424, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wurges:1981:STB

[358] Harald Würges. A specification technique based on predicate transformers. Acta Informatica, 15(4):425–445, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tokuda:1981:EUR

[359] Takehiro Tokuda. Eliminating unit reductions from LR(k) parsers using minimum contexts. Acta Informatica, 15(4):447–470, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Casanova:1981:EGP

[360] Marco A. Casanova and Philip A. Bernstein. Erratum: General Purpose Schedulers for Database System. Acta Informatica, 15(4):471, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [320].

Kedem:1981:CDG

[361] Zvi M. Kedem and Abraham Silberschatz. A characterization of database graphs admitting a simple locking protocol. Acta Informatica, 16(1):1– 13, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Clint:1981:UHV

[362] M. Clint. On the use of history variables. Acta Informatica, 16(1):15– 30, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hamlet:1981:RTP

[363] Richard G. Hamlet. Reliability theory of program testing. Acta Informatica, 16(1):31–43, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Stadel:1981:BVI

[364] Manfred Stadel. Behandlung verschiedener INTEGER-Darstellungen durch optimierende Compiler. (German) [treatment of various INTEGERrepresentation by optimizing compiler]. Acta Informatica, 16(1):45–56, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schassberger:1981:RTD

[365] R. Schassberger. On the response time distribution in a discrete roundrobin queue. Acta Informatica, 16(1):57–62, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Janssens:1981:CCF

[366] Dirk Janssens and Grzegorz Rozenberg. A characterization of contextfree string languages by directed node label controlled graph grammars. Acta Informatica, 16(1):63–85, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pritchard:1981:ALL

[367] Paul Pritchard. Another look at the “longest ascending subsequence” problem. Acta Informatica, 16(1):87–91, August 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Best:1981:FMA

[368] Eike Best and Brian Randell. A formal model of atomicity in asynchronous systems. Acta Informatica, 16(1):93–124, August 1981. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ostrand:1981:PRG

[369] Thomas J. Ostrand, Marvin C. Paull, and Elaine J. Weyuker. Parsing regular grammars with finite lookahead. Acta Informatica, 16(2):125– 138, October 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Claus:1981:BEP

[370] Volker Claus. The (n, k)-bounded emptiness-problem for probabilistic acceptors and related problems. Acta Informatica, 16(2):139–160, October 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Olderog:1981:SCH

[371] Ernst-Rüdiger Olderog. Sound and complete Hoare-like calculi based on copy rules. Acta Informatica, 16(2):161–197, October 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Blikle:1981:CTI

[372] Andrzej Blikle. The clean termination of iterative programs. Acta Informatica, 16(2):199–217, October 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Martin:1981:ADS

[373] Alain J. Martin. An axiomatic definition of synchronization primitives. Acta Informatica, 16(2):219–235, October 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fenner:1981:ATR

[374] Trevor I. Fenner and George Loizou. An analysis of two related loopfree algorithms for generating integer partitions. Acta Informatica, 16 (2):237–252, October 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ramanathan:1981:PAD

[375] Jayashree Ramanathan and Ken Kennedy. Pathlistings applied to data flow analysis. Acta Informatica, 16(3):253–273, November 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1981:FPO

[376] Joost Engelfriet and Gilberto Filè. The formal power of one-visit attribute grammars. Acta Informatica, 16(3):275–302, November 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Goldschlager:1981:PCF

[377] Leslie M. Goldschlager. ε-productions in context-free grammars. Acta Informatica, 16(3):303–308, November 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hennessy:1981:FDR

[378] John L. Hennessy and Richard B. Kieburtz. The formal definition of a real-time language. Acta Informatica, 16(3):309–345, November 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Klingler:1981:GBA

[379] Thomas Klingler and Stefan Reisch. A gap between the actual complexity of permutations and their entropy defined by Stoß. Acta Informatica, 16 (3):347–362, November 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [36].

Markowsky:1981:BHT

[380] George Markowsky. Best Huffman trees. Acta Informatica, 16(3):363– 370, November 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Manna:1981:PFP

[381] Zohar Manna and Richard J. Waldinger. Problematic features of programming languages: a situational-calculus approach. Acta Informatica, 16(4):371–426, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Alblas:1981:CAE

[382] Henk Alblas. A characterization of attribute evaluation in passes. Acta Informatica, 16(4):427–464, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: memorandum 315, Onderafdeling der Informatica, Tech. Hogeschool Twente (September 1980).

Lengauer:1981:BWP

[383] Thomas Lengauer. Black-white pebbles and graph separation. Acta Informatica, 16(4):465–475, December 1981. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gergely:1982:TIP

[384] T. Gergely and L. Úry. A theory of interactive programming. Acta Informatica, 17(1):1–20, April 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Arnold:1982:SBP

[385] A. Arnold. Synchronized behaviours of processes and rational relations. Acta Informatica, 17(1):21–29, April 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Liu:1982:SST

[386] C. L. Liu, Jane W. S. Liu, and Arthur L. Liestman. Scheduling with slack time. Acta Informatica, 17(1):31–41, April 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Shore:1982:ITA

[387] John E. Shore. Information theoretic approximations for M/G/1 und G/G/1 queuing systems. Acta Informatica, 17(1):43–61, April 1982. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Miyano:1982:HTM

[388] Satoru Miyano. A hierarchy theorem for multihead stack-counter automata. Acta Informatica, 17(1):63–67, April 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rozenberg:1982:CEF

[389] Grzegorz Rozenberg and R. Verraedt. Completeness of EOL forms is decidable. Acta Informatica, 17(1):69–87, April 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huynh:1982:RCI

[390] Thiet-Dung Huynh. Remarks on the complexity of an invariant of context-free grammars. Acta Informatica, 17(1):89–99, April 1982. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Martinez:1982:PBP

[391] Michel Martinez. Program behavior prediction and prepaging. Acta Informatica, 17(1):101–120, April 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rabin:1982:CCP

[392] Michael O. Rabin. The choice coordination problem. Acta Informatica, 17(2):121–134, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kessels:1982:ACM

[393] Joep L. W. Kessels. Arbitration without common modifiable variables. Acta Informatica, 17(2):135–141, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vasudevan:1982:ILF

[394] Sridhar Vasudevan. Inner loops in flowgraphs and code optimization. Acta Informatica, 17(2):143–155, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huddleston:1982:NDS

[395] Scott Huddleston and Kurt Mehlhorn. A new data structure for representing sorted lists. Acta Informatica, 17(2):157–184, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Raiha:1982:TAG

[396] Kari-Jouko Räihä and Mikko Saarinen. Testing attribute grammars for circularity. Acta Informatica, 17(2):185–192, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Autebert:1982:FLG

[397] Jean-Michel Autebert, Joffroy Beauquier, and Luc Boasson. Formes de langages et de grammaires. (French) [Grammar and language forms]. Acta Informatica, 17(2):193–213, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Itai:1982:RG

[398] Alon Itai and Michael Rodeh. Representation of graphs. Acta Informatica, 17(2):215–219, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huwig:1982:MPP

[399] Hagen Huwig. Ein Modell des P = NP -Problems mit einer positiven Lösung. (German) [A model of the P = NP problem with a positive solution]. Acta Informatica, 17(2):221–243, June 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Doberkat:1982:DRH

[400] Ernst-Erich Doberkat. Deleting the root of a heap. Acta Informatica, 17 (3):245–265, August 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Overmars:1982:DMD

[401] Mark H. Overmars and Jan van Leeuwen. Dynamic multi-dimensional data structures based on quad- and K-D trees. Acta Informatica, 17 (3):267–285, August 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baccelli:1982:PAE

[402] François Baccelli and Thierry Fleury. On parsing arithmetic expressions in a multiprocessing environment. Acta Informatica, 17(3):287–310, August 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Staphylopatis:1982:PCP

[403] A. Staphylopatis. Performance considerations in the parallel execution of numerical algorithms on two processors. Acta Informatica, 17(3):311– 325, August 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Snyder:1982:RSI

[404] Lawrence Snyder. Recognition and selection of idioms for code optimization. Acta Informatica, 17(3):327–348, August 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Demers:1982:SDP

[405] A. Demers, C. Keleman, and Bernd Reusch. On some decidable properties of finite state translations. Acta Informatica, 17(3):349–364, August 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cristian:1982:RDT

[406] Flaviu Cristian. Robust data types. Acta Informatica, 17(4):365–397, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Leung:1982:EFL

[407] Clement H. C. Leung and Q. H. Choo. The effect of fixed-length record implementation on file system response. Acta Informatica, 17(4):399– 409, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

JaJa:1982:SEA

[408] Joseph JáJá and Janos Simon. Space efficient algorithms for some graph theoretical problems. Acta Informatica, 17(4):411–423, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Blum:1982:PCR

[409] Norbert Blum. On the power of chain rules in context free grammars. Acta Informatica, 17(4):425–433, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Soisalon-Soininen:1982:CRC

[410] Eljas Soisalon-Soininen and Derick Wood. On a covering relation for context-free grammars. Acta Informatica, 17(4):435–449, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Asveld:1982:CTO

[411] Peter R. J. Asveld and J. V. Tucker. Complexity theory and the operational structure of algebraic programming systems. Acta Informatica, 17 (4):451–476, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pritchard:1982:EWS

[412] Paul Pritchard. Explaining the wheel sieve. Acta Informatica, 17(4):477– 485, October 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dewar:1982:SMA

[413] Robert B. K. Dewar, Susan M. Merritt, and Micha Sharir. Some modified algorithms for Dijkstra’s longest upsequence problem. Acta Informatica, 18(1):1–15, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ramirez:1982:ORP

[414] Raúl J. Ramírez, Frank Wm. Tompa, and J. Ian Munro. Optimum reorganization points for arbitrary database costs. Acta Informatica, 18 (1):17–30, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Budd:1982:TNC

[415] Timothy A. Budd and Dana Angluin. Two notions of correctness and their relation to testing. Acta Informatica, 18(1):31–45, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Broy:1982:PAT

[416] Manfred Broy and Martin Wirsing. Partial abstract types. Acta Informatica, 18(1):47–64, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Leguy:1982:LSC

[417] Jeannine Leguy. Langages saturés et cones décroissants langages et cones bifidèles. (French) [Saturated languages and decreasing cones. bifaithful languages and bifaithful cones]. Acta Informatica, 18(1):65–78, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Langmaack:1982:TPF

[418] Hans Langmaack. On termination problems for finitely interpreted ALGOL-like programs. Acta Informatica, 18(1):79–108, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Frougny:1982:HGC

[419] Christiane Frougny, Jacques Sakarovitch, and Erich Valkema. On the Hotz group of a context-free grammar. Acta Informatica, 18(1):109–115, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reisig:1982:DBS

[420] Wolfgang Reisig. Deterministic buffer synchronization of sequential processes. Acta Informatica, 18(2):117–134, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Carter:1982:FAC

[421] Lynn Robert Carter. Further analysis of code generation for a single register machine. Acta Informatica, 18(2):135–147, November 1982. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ernst:1982:VPP

[422] George W. Ernst, Jainendra K. Navlakha, and William F. Ogden. Verification of programs with procedure-type parameters. Acta Informatica, 18 (2):149–169, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nakatsu:1982:LCS

[423] Narao Nakatsu, Yahiko Kambayashi, and Shuzo Yajima. A longest common subsequence algorithm suitable for similar text strings. Acta Informatica, 18(2):171–179, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pettorossi:1982:DVE

[424] Alberto Pettorossi and Rod M. Burstall. Deriving very efficient algorithms for evaluating linear recurrence relations using the program transformation technique. Acta Informatica, 18(2):181–206, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Brown:1982:LBL

[425] Donna J. Brown, Brenda S. Baker, and Howard P. Katseff. Lower bounds for on-line two-dimensional packing algorithms. Acta Informatica, 18(2): 207–225, November 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nicolas:1982:LII

[426] Jean-Marie M. Nicolas. Logic for improving integrity checking in relational data bases. Acta Informatica, 18(3):227–253, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Allen:1982:CON

[427] Brian Allen. On the costs of optimal and near-optimal binary search trees. Acta Informatica, 18(3):255–263, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nielson:1982:DFD

[428] Flemming Nielson. A denotational framework for data flow analysis. Acta Informatica, 18(3):265–287, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anderson:1982:AII

[429] S. O. Anderson and R. C. Backhouse. An alternative implementation of an insertion-only recovery technique. Acta Informatica, 18(3):289–298, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Winklmann:1982:CSPa

[430] Karl Winklmann. On the complexity of some problems concerning the use of procedures. I. Acta Informatica, 18(3):299–318, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Agrawala:1982:ESG

[431] Ashok K. Agrawala and Satish K. Tripathi. On an exponential server with general cyclic arrivals. Acta Informatica, 18(3):319–334, December 1982. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1983:SAV

[432] Karel Culik II, J. Gruska, and Arto Salomaa. Systolic automata for VLSI on balanced trees. Acta Informatica, 18(4):335–344, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vanLeeuwen:1983:SBS

[433] Jan van Leeuwen and Mark H. Overmars. Stratified balanced search trees. Acta Informatica, 18(4):345–359, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Eriksen:1983:CAR

[434] Ole Eriksen and Jørgen Staunstrup. Concurrent algorithms for root searching. Acta Informatica, 18(4):361–376, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Supowit:1983:CDT

[435] Kenneth J. Supowit and Edward M. Reingold. The complexity of drawing trees nicely. Acta Informatica, 18(4):377–392, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See corrigendum [1172].

Slutzki:1983:FSR

[436] Giora Slutzki. Finite state relational programs. Acta Informatica, 18 (4):393–409, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Winklmann:1983:CSP

[437] Karl Winklmann. On the complexity of some problems concerning the use of procedures. II. Acta Informatica, 18(4):411–430, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Banatre:1983:NDW

[438] Jean-Pierre Banâtre, Patrice Frison, and Patrice Quinton. A network for the detection of words in continuous speech. Acta Informatica, 18 (4):431–448, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nain:1983:PTE

[439] Philippe Nain. Partage de tâches entre [deux] processeurs homogènes. (French) [Task scheduling between two homogeneous processors]. Acta Informatica, 18(4):449–466, January 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kung:1983:OTC

[440] H. T. Kung and Christos H. Papadimitriou. An optimality theory of concurrency control for databases. Acta Informatica, 19(1):1–11, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tsuda:1983:TLTa

[441] Takao Tsuda and Takashi Sato. Transposition of large tabular data structures with applications to physical database organization. Part I. transposition of tabular data structures. Acta Informatica, 19(1):13– 33, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kuspert:1983:SUT

[442] Klaus Küspert. Storage utilization in B^{∗}-trees with a generalized overflow technique. Acta Informatica, 19(1):35–55, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Taylor:1983:CAS

[443] Richard N. Taylor. Complexity of analyzing the synchronization structure of concurrent programs. Acta Informatica, 19(1):57–84, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iwama:1983:UPU

[444] Kazuo Iwama. The universe problem for unrestricted flow languages. Acta Informatica, 19(1):85–96, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bergstra:1983:SMS

[445] Jan A. Bergstra and J. Terlouw. Standard model semantics for DSL. A data type specification language. Acta Informatica, 19(1):97–113, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

File:1983:IRA

[446] Gilberto Filé. Interpretation and reduction of attribute grammars. Acta Informatica, 19(2):115–150, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: memorandum 359, Onderafdeling der Informatica, Tech. Hogeschool Twente (1981).

Pfaff:1983:COI

[447] Günther E. Pfaff. The construction of operator interfaces based on logical input devices. Acta Informatica, 19(2):151–166, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tsuda:1983:TLTb

[448] Takao Tsuda, Akira Urano, and Takashi Sato. Transposition of large tabular data structures with applications to physical database organization. Part II. applications to physical database organization. Acta Informatica, 19(2):167–182, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schurfeld:1983:NLB

[449] Ute Schürfeld. New lower bounds on the formula size of Boolean functions. Acta Informatica, 19(2):183–194, April 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Queille:1983:FRP

[450] J.-P. P. Queille and Joseph Sifakis. Fairness and related properties in transition systems: a temporal logic to deal with fairness. Acta Informatica, 19(3):195–220, July 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kearns:1983:IRC

[451] John P. Kearns and Mary Lou Soffa. The implementation of retention in a coroutine environment. Acta Informatica, 19(3):221–233, July 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engels:1983:OSS

[452] Gregor Engels, Udo Pletat, and Hans-Dieter Ehrich. An operational semantics for specifications of abstract data types with error handling. Acta Informatica, 19(3):235–253, July 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nielson:1983:CSW

[453] Hanne Riis Nielson. Computation sequences: a way to characterize classes of attribute grammars. Acta Informatica, 19(3):255–268, July 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: report DAIMI PB-138, Computer Science Department, Aarhus University, Århus, Denmark (1981).

MeyeraufderHeide:1983:EUP

[454] Friedhelm Meyer auf der Heide. Efficiency of universal parallel computers. Acta Informatica, 19(3):269–296, July 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schmitt:1983:NRO

[455] Alfred Schmitt. On the number of relational operators necessary to compute certain functions of real variables. Acta Informatica, 19(3):297–304, July 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vardi:1983:IMD

[456] Moshe Y. Vardi. Inferring multivalued dependencies from functional and join dependencies. Acta Informatica, 19(4):305–324, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Suzuki:1983:TMS

[457] Ichiro Suzuki and Tadao Kasami. Three measures for synchronic dependence in Petri nets. Acta Informatica, 19(4):325–338, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

El-Affendi:1983:MEA

[458] M. A. El-Affendi and Demetres D. Kouvatsos. A maximum entropy analysis of the M/G/1 and G/M/1 queueing systems at equilibrium. Acta Informatica, 19(4):339–355, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ruohonen:1983:SVP

[459] Keijo Ruohonen. On some variants of Post’s correspondence problem. Acta Informatica, 19(4):357–367, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Agrawal:1983:EIL

[460] Rakesh Agrawal and Keith D. Detro. An efficient incremental LR parser for grammars with epsilon productions. Acta Informatica, 19(4):369–376, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hromkovic:1983:OWM

[461] Juraj Hromkovič. One-way multihead deterministic finite automata. Acta Informatica, 19(4):377–384, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Klein:1983:LTB

[462] Peter Klein and Friedhelm Meyer auf der Heide. A lower time bound for the knapsack problem on random access machines. Acta Informatica, 19 (4):385–395, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sommerhalder:1983:PLR

[463] R. Sommerhalder and S. C. van Westrhenen. Parallel language recognition in constant time by cellular automata. Acta Informatica, 19(4): 397–407, September 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wirsing:1983:HAD

[464] Martin Wirsing, Peter Pepper, Helmut Partsch, Walter Dosch, and Manfred Broy. On hierarchies of abstract data types. Acta Informatica, 20(1): 1–33, October 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

He:1983:GPT

[465] Jifeng He. General predicate transformer and the semantics of a programming language with go to statement. Acta Informatica, 20(1):35– 57, October 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Damm:1983:SRH

[466] Werner Damm and Bernhard Josko. A sound and relatively ^{∗}complete Hoare-logic for a language with higher type procedures. Acta Informatica, 20(1):59–101, October 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chandra:1983:SBV

[467] Ashok K. Chandra, Lawrence T. Kou, George Markowsky, and Shmuel Zaks. On sets of Boolean n-vectors with all k-projections surjective. Acta Informatica, 20(1):103–111, October 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kameda:1983:NMQ

[468] Hisao Kameda. A note on multi-queue scheduling of two tasks. Acta Informatica, 20(2):113–120, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Akdag:1983:PAC

[469] Herman Akdag. Performances of an algorithm constructing a nearly optimal binary tree. Acta Informatica, 20(2):121–132, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bucher:1983:TSS

[470] W. Bucher. Two-symbol DOS systems generating regular languages. Acta Informatica, 20(2):133–142, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mannila:1983:RMO

[471] Heikki Mannila and Kari-Jouko Räihä. On the relationship of minimum and optimum covers for a set of functional dependencies. Acta Informatica, 20(2):143–158, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coppo:1983:SP

[472] Mario Coppo. On the semantics of polymorphism. Acta Informatica, 20 (2):159–170, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Goodman:1983:NCP

[473] Nathan Goodman and Oded Shmueli. NP-complete problems simplified on tree schemas. Acta Informatica, 20(2):171–178, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pansiot:1983:HFC

[474] Jean-Jacques Pansiot. Hiérarchie et fermeture de certaines classes de tagsystèmes. (French) [Hierarchy and closure properties of certain classes of tag-systems]. Acta Informatica, 20(2):179–196, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ronse:1983:TSC

[475] Christian Ronse. A three-stage construction for multiconnection networks. Acta Informatica, 20(2):197–206, November 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ben-Ari:1983:TLB

[476] Mordechai Ben-Ari, Amir Pnueli, and Zohar Manna. The temporal logic of branching time. Acta Informatica, 20(3):207–226, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Simon:1983:PMT

[477] Hans-Ulrich U. Simon. Pattern matching in trees and nets. Acta Informatica, 20(3):227–248, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Marque-Pucheu:1983:RST

[478] G. Marque-Pucheu. Rational set of trees and the algebraic semantics of logic programming. Acta Informatica, 20(3):249–260, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bannister:1983:TAF

[479] Joseph A. Bannister and Kishor S. Trivedi. Task allocation in faulttolerant distributed systems. Acta Informatica, 20(3):261–281, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pohlmann:1983:LPA

[480] Werner Pohlmann. LR parsing for affix grammars. Acta Informatica, 20 (4):283–300, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Martin:1983:GPR

[481] Alain J. Martin. A general proof rule for procedures in predicate transformer semantics. Acta Informatica, 20(4):301–313, December 1983. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mili:1983:RAD

[482] Ali Mili. A relational approach to the design of deterministic programs. Acta Informatica, 20(4):315–328, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Francez:1983:PPT

[483] Nissim Francez. Product properties and their direct verification. Acta Informatica, 20(4):329–344, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Flajolet:1983:PEE

[484] Philippe Flajolet. On the performance evaluation of extendible hashing and trie searching. Acta Informatica, 20(4):345–369, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fraenkel:1983:TCP

[485] Aviezri S. Fraenkel, Moshe Mor, and Y. Perl. Is text compression by prefixes and suffixes practical? Acta Informatica, 20(4):371–389, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kleijn:1983:GPR

[486] H. C. M. Kleijn and Grzegorz Rozenberg. On the generative power of regular pattern grammars. Acta Informatica, 20(4):391–411, December 1983. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bayer:1984:DMS

[487] Rudolf Bayer and Peter Schlichtiger. Data management support for database management. Acta Informatica, 21(1):1–28, May 1984. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chapman:1984:LPG

[488] N. P. Chapman. LALR(1, 1) parser generation for regular right part grammars. Acta Informatica, 21(1):29–45, May 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Szwarcfiter:1984:OMS

[489] Jayme Luiz Szwarcfiter. Optimal multiway search trees for variable size keys. Acta Informatica, 21(1):47–60, May 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hennessy:1984:AFD

[490] Matthew Hennessy. Axiomatising finite delay operators. Acta Informatica, 21(1):61–88, May 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Best:1984:FCS

[491] Eike Best and Klaus Voss. Free choice systems have home states. Acta Informatica, 21(1):89–100, May 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tsakalidis:1984:MOG

[492] Athanasios K. Tsakalidis. Maintaining order in a generalized linked list. Acta Informatica, 21(1):101–112, May 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huang:1984:GBS

[493] Shou Hsuan Stephen Huang and C. K. Wong. Generalized binary split trees. Acta Informatica, 21(1):113–123, May 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lubachevsky:1984:AAV

[494] Boris D. Lubachevsky. An approach to automating the verification of compact parallel coordination programs: Part 1. Acta Informatica, 21(2): 125–169, August 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Part II: Technical Report No. 64, February (1983), Courant Institute of Mathematical Sciences, New York University.

Blum:1984:CAO

[495] Norbert Blum and Martin Seysen. Characterization of all optimal networks for a simultaneous computation of AND and NOR. Acta Informatica, 21(2):171–181, August 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gillett:1984:BTE

[496] Will D. Gillett. On binary tree encodements. Acta Informatica, 21(2): 183–192, August 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ibarra:1984:CSB

[497] Oscar H. Ibarra and Sam M. Kim. A characterization of systolic binary tree automata and applications. Acta Informatica, 21(2):193–207, August 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Autebert:1984:BLA

[498] Jean-Michel Autebert, Joffroy Beauquier, Luc Boasson, and Françoise Gire. Bicentres de langages algébriques. (French) [Bicenters of contextfree languages]. Acta Informatica, 21(2):209–227, August 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Devroye:1984:PAH

[499] Luc Devroye. A probabilistic analysis of the height of tries and of the complexity of triesort. Acta Informatica, 21(3):229–237, October 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bird:1984:UCP

[500] R. S. Bird. Using circular programs to eliminate multiple traversals of data. Acta Informatica, 21(3):239–250, October 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). (LA has).

Barringer:1984:LCU

[501] H. Barringer, J. H. Cheng, and Cliff B. Jones. A logic covering undefinedness in program proofs. Acta Informatica, 21(3):251–269, October 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Guting:1984:ODC

[502] Ralf Hartmut Güting. Optimal divide-and-conquer to compute measure and contour for a set of iso-rectangles. Acta Informatica, 21(3):271–291, October 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bergstra:1984:ASP

[503] Jan A. Bergstra and J. V. Tucker. The axiomatic semantics of programs based on Hoare’s logic. Acta Informatica, 21(3):293–320, October 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iglehart:1984:SOA

[504] Donald L. Iglehart and Gerald S. Shedler. Simulation output analysis for local area computer networks. Acta Informatica, 21(4):321–338, November 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mehlhorn:1984:RDS

[505] Kurt Mehlhorn and Uzi Vishkin. Randomized and deterministic simulations of PRAMs by parallel machines with restricted granularity of parallel memories. Acta Informatica, 21(4):339–374, November 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Deransart:1984:SCT

[506] Pierre Deransart, Martin Jourdan, and Bernard Lorho. Speeding up circularity tests for attribute grammars. Acta Informatica, 21(4):375– 391, November 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: rapport RR 211, INRIA, Rocquencourt (May

Choffrut:1984:RTC

[507] Christian Choffrut and Karel Culik II. On real-time cellular automata and trellis automata. Acta Informatica, 21(4):393–407, November 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coffman:1984:PGG

[508] Edward G. Coffman, Jr. and Michael A. Langston. A performance guarantee for the greedy set-partitioning algorithm. Acta Informatica, 21(4): 409–415, November 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Costa:1984:FCC

[509] Gerardo Costa and Colin Stirling. A fair calculus of communicating systems. Acta Informatica, 21(5):417–441, December 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sannella:1984:STS

[510] D. T. Sannella. A set–theoretic semantics for CLEAR. Acta Informatica, 21(5):443–472, December 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bulyonkov:1984:PMC

[511] Mikhail A. Bulyonkov. Polyvariant mixed computation for analyzer programs. Acta Informatica, 21(5):473–484, December 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Leung:1984:PDQ

[512] Clement H. C. Leung and Q. H. Choo. The paging drum queue: a uniform perspective and further results. Acta Informatica, 21(5):485–500, December 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hertel:1984:SSS

[513] Stefan Hertel, Martti Mäntylä, Kurt Mehlhorn, and Jürg Nievergelt. Space sweep solves intersection of convex polyhedra. Acta Informatica, 21 (5):501–519, December 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1984:FCR

[514] Günther Bauer and Friedrich Otto. Finite complete rewriting systems and the complexity of the word problem. Acta Informatica, 21(5):521– 540, December 1984. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wright:1985:SAP

[515] William E. Wright. Some average performance measures for the B-tree. Acta Informatica, 21(6):541–557, March 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fisher:1985:PLB

[516] A. J. Fisher. Practical LL(1)-based parsing of van Wijngaarden grammars. Acta Informatica, 21(6):559–584, March 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chen:1985:DFR

[517] Marina C. Chen and Martin Rem. Deadlock-freedom in resource contentions. Acta Informatica, 21(6):585–598, March 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berry:1985:DSS

[518] Daniel M. Berry. A denotational semantics for shared-memory parallelism and nondeterminism. Acta Informatica, 21(6):599–627, March 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lenfant:1985:PDO

[519] Jacques Lenfant and Serge Tahé. Permuting data with the Omega network. Acta Informatica, 21(6):629–641, March 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Valk:1985:RVS

[520] Rüdiger Valk and Matthias Jantzen. The residue of vector sets with applications to decidability problems in Petri nets. Acta Informatica, 21 (6):643–674, March 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bechtold:1985:TSC

[521] Michael Bechtold, Guy Pujolle, and Otto Spaniol. Throughput of a satellite channel communication. Acta Informatica, 22(1):1–14, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kingston:1985:ATA

[522] Jeffrey H. Kingston. Analysis of tree algorithms for the simulation event list. Acta Informatica, 22(1):15–33, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schneider:1985:TET

[523] Fred B. Schneider, Richard Conway, and Dale Skeen. Thrifty execution of task pipelines. Acta Informatica, 22(1):35–45, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mili:1985:SIF

[524] Ali Mili, Jules Desharnais, and Jean-Raymond Gagné. Strongest invariant functions: Their use in the systematic analysis of while statements. Acta Informatica, 22(1):47–66, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Jacobs:1985:GCU

[525] Dean Jacobs and David Gries. General correctness: a unification of partial and total correctness. Acta Informatica, 22(1):67–83, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ottmann:1985:PTU

[526] Thomas Ottmann, Michael Schrapp, and Derick Wood. Purely top-down updating algorithms for stratified search trees. Acta Informatica, 22(1): 85–100, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Loveland:1985:PBB

[527] Donald W. Loveland. Performance bounds for binary testing with arbitrary weights. Acta Informatica, 22(1):101–114, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Monien:1985:RNA

[528] Burkhard Monien and Ewald Speckenmeyer. Ramsey numbers and an approximation algorithm for the vertex cover problem. Acta Informatica, 22(1):115–123, April 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). NCPY.

Meghini:1985:QFR

[529] Carlo Meghini and Costantino Thanos. Querying fragmented relations in a distributed database. Acta Informatica, 22(2):125–138, June 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bruno:1985:STE

[530] John Bruno. On scheduling tasks with exponential service times and intree precedence constraints. Acta Informatica, 22(2):139–148, June 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cunningham:1985:RSL

[531] R. J. Cunningham and A. J. J. Dick. Rewrite systems on a lattice of types. Acta Informatica, 22(2):149–169, June 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sack:1985:AMH

[532] Jörg-Rüdiger Sack and Thomas Strothotte. An algorithm for merging heaps. Acta Informatica, 22(2):171–186, June 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chiba:1985:DPG

[533] Norishige Chiba, Kazunori Onoguchi, and Takao Nishizeki. Drawing plane graphs nicely. Acta Informatica, 22(2):187–201, June 1985. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kamimura:1985:EGI

[534] Tsutomu Kamimura. An effectively given initial semigroup. Acta Informatica, 22(2):203–227, June 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dunne:1985:LBM

[535] Paul E. S. Dunne. A 2.5n lower bound on the monotone network complexity of T_{3}^{n}. Acta Informatica, 22(2):229–240, June 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rembowski:1985:PQI

[536] Bogdan Rembowski. A priority queue with interruptions of service permitted after a time quantum. Acta Informatica, 22(3):241–251, August 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Krishnamurthy:1985:SPT

[537] Balakrishnan Krishnamurthy. Short proofs for tricky formulas. Acta Informatica, 22(3):253–275, August 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gonczarowski:1985:DPM

[538] Jakob Gonczarowski. Decidable properties of monadic recursive schemas with a depth parameter. Acta Informatica, 22(3):277–310, August 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rudnicki:1985:PPP

[539] Piotr Rudnicki and Włodzimierz Drabent. Proving properties of Pascal programs in MIZAR 2. Acta Informatica, 22(3):311–331, August 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bruno:1985:PBD

[540] John L. Bruno and Peter J. Downey. Probabilistic bounds for dual bin-packing. Acta Informatica, 22(3):333–345, August 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kfoury:1985:NSC

[541] A. J. Kfoury and Paweł Urzyczyn. Necessary and sufficient conditions for the universality of programming formalisms. Acta Informatica, 22 (4):347–377, October 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chung:1985:CPS

[542] Moon-Jung Chung, W. Michael Evangelist, and Ivan Hal Sudborough. Complete problems for space bounded subclasses of NP. Acta Informatica, 22(4):379–395, October 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sonnenschein:1985:GSC

[543] Michael Sonnenschein. Global storage cells for attributes in an attribute grammar. Acta Informatica, 22(4):397–420, October 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huynh:1985:CWP

[544] Dung T. Huynh. Complexity of the word problem for commutative semigroups of fixed dimension. Acta Informatica, 22(4):421–432, October 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Beynon:1985:RCE

[545] Meurig Beynon. Replaceability and computational equivalence for monotone Boolean functions. Acta Informatica, 22(4):433–449, October 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lloyd:1985:WCP

[546] Errol L. Lloyd and Michael C. Loui. On the worst case performance of buddy systems. Acta Informatica, 22(4):451–473, October 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Miyakawa:1985:ODT

[547] Masahiro Miyakawa. Optimum decision trees—an optimal variable theorem and its related applications. Acta Informatica, 22(5):475–498, December 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Heilbrunner:1985:TPC

[548] Stephan Heilbrunner. Truly prefix-correct chain-free LR(1) parsers. Acta Informatica, 22(5):499–536, December 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Moller:1985:ASI

[549] Bernhard Möller. On the algebraic specification of infinite objects— ordered and continuous models of algebraic types. Acta Informatica, 22 (5):537–578, December 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Latteux:1985:FOC

[550] Michel Latteux, B. Leguy, and B. Ratoandromanana. The family of one-counter languages is closed under quotient. Acta Informatica, 22(5): 579–588, December 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hromkovic:1985:FTW

[551] Juraj Hromkovič. Fooling a two-way nondeterministic multihead automaton with reversal number restriction. Acta Informatica, 22(5):589–594, December 1985. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Caspi:1986:FMD

[552] Paul Caspi and Nicolas Halbwachs. A functional model for describing and reasoning about time behaviour of computing systems. Acta Informatica, 22(6):595–627, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nipkow:1986:NDD

[553] Tobias Nipkow. Non-deterministic data types: models and implementations. Acta Informatica, 22(6):629–661, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reusch:1986:MCI

[554] Bernd Reusch and W. Merzenich. Minimal coverings for incompletely specified sequential machines. Acta Informatica, 22(6):663–678, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Diekert:1986:IHG

[555] Volker Diekert. Investigations on Hotz groups for arbitrary grammars. Acta Informatica, 22(6):679–698, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rudnicki:1986:PPP

[556] Piotr Rudnicki and Włodzimierz Drabent. Proving properties of Pascal programs in MIZAR 2. Acta Informatica, 22(6):699–707, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dijkstra:1986:SFA

[557] Edsger W. Dijkstra and A. J. M. van Gasteren. A simple fixpoint argument without the restriction to continuity. Acta Informatica, 23(1): 1–7, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Olderog:1986:SOS

[558] Ernst-Rüdiger R. Olderog and C. A. R. Hoare. Specification-oriented semantics for communicating processes. Acta Informatica, 23(1):9–66, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Duponcheel:1986:AFP

[559] L. Duponcheel and M. Duponcheel. Acceptable functional programming systems. Acta Informatica, 23(1):67–98, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Otto:1986:DWM

[560] Friedrich Otto. On deciding whether a monoid is a free monoid or is a group. Acta Informatica, 23(1):99–110, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mahmoud:1986:AIP

[561] Hosam M. Mahmoud. On the average internal path length of m-ary search trees. Acta Informatica, 23(1):111–117, March 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bengelloun:1986:IPS

[562] S. A. Bengelloun. An incremental primal sieve. Acta Informatica, 23(2): 119–125, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Heckmann:1986:EEP

[563] Reinhold Heckmann. An efficient ELL(1)-parser generator. Acta Informatica, 23(2):127–148, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nakata:1986:GEL

[564] Ikuo Nakata and Masataka Sassa. Generation of efficient LALR parsers for regular right part grammars. Acta Informatica, 23(2):149–162, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Becker:1986:ARP

[565] M. Becker and Kurt Mehlhorn. Algorithms for routing in planar graphs. Acta Informatica, 23(2):163–176, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rosenberg:1986:FAE

[566] Catherine Rosenberg. Files d’attente exponentielles ayant des paramètres non-stationnaires dans le temps. (French) [Exponential queueing systems with non-stationary [time] parameters]. Acta Informatica, 23(2):177– 192, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Soundararajan:1986:TCC

[567] N. Soundararajan. Total correctness of CSP programs. Acta Informatica, 23(2):193–215, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paterson:1986:NOH

[568] Michael S. Paterson and Ingo Wegener. Nearly optimal hierarchies for network and formula size. Acta Informatica, 23(2):217–221, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wu:1986:FAA

[569] Y. F. Wu, Peter Widmayer, and C. K. Wong. A faster approximation algorithm for the Steiner problem in graphs. Acta Informatica, 23(2): 223–229, May 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Makowsky:1986:EPD

[570] Johann A. Makowsky and Moshe Y. Vardi. On the expressive power of data dependencies. Acta Informatica, 23(3):231–244, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bucher:1986:RTD

[571] W. Bucher. A regularity test for dual bordered OS systems. Acta Informatica, 23(3):245–253, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Calzarossa:1986:WMR

[572] Maria Calzarossa, M. Italiani, and Giuseppe Serazzi. A workload model representative of static and dynamic characteristics. Acta Informatica, 23(3):255–266, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tamminen:1986:CMS

[573] Markku Tamminen, W. K. Luk, Paolo Sipala, L. S. Woo, and C. K. Wong. Constructing maximal slicings from geometry. Acta Informatica, 23(3):267–288, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rozenberg:1986:GTC

[574] Grzegorz Rozenberg and Emo Welzl. Graph theoretic closure properties of the family of boundary NLC graph languages. Acta Informatica, 23 (3):289–309, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Krivanek:1986:NHP

[575] Mirko Křivánek and Jaroslav Morávek. NP-hard problems in hierarchical-tree clustering. Acta Informatica, 23(3):311–323, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wagner:1986:CCP

[576] Klaus W. Wagner. The complexity of combinatorial problems with succinct input representation. Acta Informatica, 23(3):325–356, June 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bijlsma:1986:EGM

[577] A. Bijlsma, J. G. Wiltink, and P. A. Matthews. Equivalence of the Gries and Martin proof rules for procedure calls. Acta Informatica, 23(4):357– 360, August 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wyrostek:1986:PTW

[578] Piotr Wyrostek. Precedence technique is not worse than SLR(1). Acta Informatica, 23(4):361–392, August 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Farrow:1986:CSO

[579] Rodney Farrow and Daniel M. Yellin. A comparison of storage optimizations in automatically-generated attribute evaluators. Acta Informatica, 23(4):393–427, August 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: Technical Report, Department of Comp. Sc., Columbia University, New York, NY (January 1985).

Koutny:1986:MRP

[580] Maciej Koutny. The Merlin-Randell problem of train journeys. Acta Informatica, 23(4):429–463, August 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nicola:1986:SSQ

[581] Victor F. Nicola. A single server queue with mixed types of interruptions. Acta Informatica, 23(4):465–486, August 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hehner:1986:PM

[582] Eric C. R. Hehner, Lorene E. Gupta, and Andrew J. Malton. Predicative methodology. Acta Informatica, 23(5):487–505, September 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See erratum [671].

Graf:1986:LSP

[583] Susanne Graf and Joseph Sifakis. A logic for the specification and proof of regular controllable processes of CCS. Acta Informatica, 23(5):507–527, September 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lee:1986:GBT

[584] C. C. Lee, D. T. Lee, and C. K. Wong. Generating binary trees of bounded height. Acta Informatica, 23(5):529–544, September 1986. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kouvatsos:1986:MEG

[585] Demetres D. Kouvatsos. Maximum entropy and the G/G/1/N queue. Acta Informatica, 23(5):545–565, September 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reichardt:1986:DGG

[586] Johannes Reichardt. Deterministic grammars and grammar morphisms. Acta Informatica, 23(5):567–583, September 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Maon:1986:EST

[587] Yael Maon. On the equivalence of some transductions involving letter to letter morphisms on regular languages. Acta Informatica, 23(5):585–596, September 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1986:SDP

[588] Karel Culik II and Juhani Karhumäki. Synchronizable deterministic pushdown automata and the decidability of their equivalence. Acta Informatica, 23(5):597–605, September 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berzins:1986:MSE

[589] Valdis Berzins. On merging software extensions. Acta Informatica, 23 (6):607–619, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Geist:1986:RLC

[590] Robert Geist, Mark Smotherman, Kishor S. Trivedi, and Joanne Bechta Dugan. The reliability of life-critical computer systems. Acta Informatica, 23(6):621–642, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1986:ADC

[591] Erol Gelenbe, David Finkel, and Satish K. Tripathi. Availability of a distributed computer system with failures. Acta Informatica, 23(6):643– 655, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cantor:1986:ITA

[592] J. Cantor, A. Ephremides, and D. Horton. Information theoretic analysis for a general queueing system at equilibrium with application to queues in tandem. Acta Informatica, 23(6):657–678, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Balcazar:1986:SSG

[593] José L. Balcázar and Ronald V. Book. Sets with small generalized Kolmogorov complexity. Acta Informatica, 23(6):679–688, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bublitz:1986:DGM

[594] Siegfried Bublitz. Decomposition of graphs and monotone formula size of homogeneous functions. Acta Informatica, 23(6):689–696, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iliopoulos:1986:MCC

[595] Costas S. Iliopoulos. Monte Carlo circuits for the Abelian permutation group intersection problem. Acta Informatica, 23(6):697–705, November 1986. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cousot:1987:EII

[596] Patrick Cousot and Radhia Cousot. Sometime = always + recursion ≡ always on the equivalence of the intermittent and invariant assertions methods for proving inevitability properties of programs. Acta Informatica, 24(1):1–31, February 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Janicki:1987:FSC

[597] Ryszard Janicki. A formal semantics for concurrent systems with a priority relation. Acta Informatica, 24(1):33–55, February 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Takeichi:1987:PPE

[598] Masato Takeichi. Partial parameterization eliminates multiple traversals of data structures. Acta Informatica, 24(1):57–77, February 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Goldreich:1987:ELR

[599] Oded Goldreich and Liuba Shrira. Electing a leader in a ring with link failures. Acta Informatica, 24(1):79–91, February 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Leung:1987:ASA

[600] Clement H. C. Leung. Analysis of space allocation in a generally fragmented linear store. Acta Informatica, 24(1):93–104, February 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vidyasankar:1987:GTS

[601] K. Vidyasankar. Generalized theory of serializability. Acta Informatica, 24(1):105–119, February 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kunde:1987:LBS

[602] Manfred Kunde. Lower bounds for sorting on mesh-connected architectures. Acta Informatica, 24(2):121–130, April 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Loizou:1987:LPJ

[603] G. Loizou and P. Thanisch. Losslessness and project-join constructibility in relational databases. Acta Informatica, 24(2):131–144, April 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Murtagh:1987:RPN

[604] Thomas P. Murtagh. Redundant proofs of non-interference in Levin-Gries CSP program proofs. Acta Informatica, 24(2):145–156, April 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Heilbrunner:1987:UUM

[605] Stephan Heilbrunner and Steffen Hölldobler. The undecidability of the unification and matching problem for canonical theories. Acta Informatica, 24(2):157–171, April 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Szpankowski:1987:ACR

[606] Wojciech Szpankowski. An analysis of a contention resolution algorithm: Another approach. Acta Informatica, 24(2):173–190, April 1987. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Harris:1987:SLP

[607] Lawrence A. Harris. SLR(1) and LALR(1) parsing for unrestricted grammars. Acta Informatica, 24(2):191–209, April 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

DeNicola:1987:EET

[608] Rocco De Nicola. Extensional equivalences for transition systems. Acta Informatica, 24(2):211–237, April 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mili:1987:RHD

[609] Ali Mili, Jules Desharnais, and Fatma Mili. Relational heuristics for the design of deterministic programs. Acta Informatica, 24(3):239–276, June 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Devroye:1987:BPA

[610] Luc Devroye. Branching processes in the analysis of the heights of trees. Acta Informatica, 24(3):277–298, June 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Alblas:1987:OPT

[611] Henk Alblas. One-pass transformations of attributed program trees. Acta Informatica, 24(3):299–352, June 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: memorandum INF-86-30, Onderafdeling der Informatica, Tech. Hogeschool Twente (October 1986).

Kumar:1987:LSA

[612] S. Kiran Kumar and C. Pandu Rangan. A linear space algorithm for the LCS problem. Acta Informatica, 24(3):353–362, June 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Becker:1987:ETO

[613] Bernd Becker. An easily testable optimal-time VLSI-multiplier. Acta Informatica, 24(4):363–380, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoogewijs:1987:PPL

[614] Albert Hoogewijs. Partial-predicate logic in computer science. Acta Informatica, 24(4):381–393, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kapur:1987:SCR

[615] Deepak Kapur, Paliath Narendran, and Hantào Zhang. On sufficient completeness and related properties of term rewriting systems. Acta Informatica, 24(4):395–415, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Peters:1987:PAS

[616] Joseph G. Peters and Larry Rudolph. Parallel approximation schemes for subset sum and knapsack problems. Acta Informatica, 24(4):417– 432, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schmidt:1987:LUP

[617] Ursula Schmidt. Long unavoidable patterns. Acta Informatica, 24(4): 433–445, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dunne:1987:RVG

[618] Paul E. S. Dunne. A result on k-valent graphs and its application to a graph embedding problem. Acta Informatica, 24(4):447–459, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kempf:1987:TOL

[619] M. Kempf, Rudolf Bayer, and Ulrich Güntzer. Time optimal left to right construction of position trees. Acta Informatica, 24(4):461–474, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bender:1987:OWC

[620] Edward A. Bender, Cheryl E. Praeger, and Nicholas C. Wormald. Optimal worst case trees. Acta Informatica, 24(4):475–489, August 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

deBakker:1987:OMS

[621] J. W. de Bakker and J.-J. Ch. Meyer. Order and metric in the stream semantics of elemental concurrency. Acta Informatica, 24(5):491–511, September 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Blazewicz:1987:MMF

[622] J. Błażewicz, W. Kubiak, H. Röck, and J. Szwarcfiter. Minimizing mean flow-time with parallel processors and resource constraints. Acta Informatica, 24(5):513–524, September 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Duda:1987:PEF

[623] Andrzej Duda and Tadeusz Czachórski. Performance evaluation of fork and join synchronization primitives. Acta Informatica, 24(5):525–553, September 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chowdhury:1987:WCP

[624] Shyamal K. Chowdhury and Pradip K. Srimani. Worst case performance of weighted buddy systems. Acta Informatica, 24(5):555–564, September 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chazelle:1987:STG

[625] Bernard Chazelle. Some techniques for geometric searching with implicit set representations. Acta Informatica, 24(5):565–582, September 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cunto:1987:ITS

[626] Walter Cunto and Jose Luis Gascon. Improving time and space efficiency in generalized binary search trees. Acta Informatica, 24(5):583– 594, September 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huang:1987:DSI

[627] Chua-Huang Huang and Christian Lengauer. The derivation of systolic implementations of programs. Acta Informatica, 24(6):595–632, November 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kastens:1987:LAA

[628] Uwe Kastens. Lifetime analysis for attributes. Acta Informatica, 24(6): 633–651, November 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lao:1987:CBC

[629] Marek J. Lao. Combinator-based compilation of recursive functions with different parameter passing modes. Acta Informatica, 24(6):653–678, November 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Horwitz:1987:EGI

[630] Susan Horwitz, Alan J. Demers, and Tim Teitelbaum. An efficient general iterative algorithm for dataflow analysis. Acta Informatica, 24(6): 679–694, November 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kamin:1987:ETS

[631] Samuel Kamin. The expressive theory of stacks. Acta Informatica, 24 (6):695–709, November 1987. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hehner:1988:TCC

[632] Eric C. R. Hehner and Andrew J. Malton. Termination conventions and comparative semantics. Acta Informatica, 25(1):1–14, January 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Finkel:1988:FNO

[633] Alain Finkel and Annie Choquet. FIFO nets without order deadlock. Acta Informatica, 25(1):15–36, January 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tsakalidis:1988:NCA

[634] Athanasios K. Tsakalidis. The nearest common ancestor in a dynamic tree. Acta Informatica, 25(1):37–54, January 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vianu:1988:DSU

[635] Victor Vianu. Database survivability under dynamic constraints. Acta Informatica, 25(1):55–84, January 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ganapathi:1988:ICG

[636] Mahadevan Ganapathi and Charles N. Fischer. Integrating code generation and peephole optimization. Acta Informatica, 25(1):85–109, January 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1988:CEA

[637] Friedrich L. Bauer and Martin Wirsing. Crypt-equivalent algebraic specifications. Acta Informatica, 25(2):111–153, February 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reps:1988:IEA

[638] Thomas W. Reps. Incremental evaluation for attribute grammars with unrestricted movement between tree modifications. Acta Informatica, 25 (2):155–178, February 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bouge:1988:ESA

[639] Luc Bougé. On the existence of symmetric algorithms to find leaders in networks of communicating sequential processes. Acta Informatica, 25 (2):179–201, February 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1988:RUM

[640] Andrzej Ehrenfeucht, Hendrik Jan Hoogeboom, and Grzegorz Rozenberg. Recording the use of memory in right-boundary grammars and push-down automata. Acta Informatica, 25(2):203–231, February 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sannella:1988:TFD

[641] Donald T. Sannella and Andrzej Tarlecki. Toward formal development of programs from algebraic specifications: implementations revisited. Acta Informatica, 25(3):233–281, April 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Zhang:1988:SOT

[642] Ming-Hua Zhang. A second order theory of data types. Acta Informatica, 25(3):283–303, April 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sobel:1988:PSD

[643] A. E. K. Sobel and N. Soundararajan. A proof system for distributed processes. Acta Informatica, 25(3):305–332, April 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reif:1988:CRD

[644] John H. Reif and Scott A. Smolka. The complexity of reachability in distributed communicating processes. Acta Informatica, 25(3):333–354, April 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Giegerich:1988:CEA

[645] Robert Giegerich. Composition and evaluation of attribute coupled grammars. Acta Informatica, 25(4):355–423, May 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schmerl:1988:RFT

[646] Ulf R. Schmerl. Resolution on formula trees. Acta Informatica, 25(4): 425–438, May 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Biela:1988:PSA

[647] Andrzej Biela. Program-substitution and admissibility of rules in algorithmic logic. Acta Informatica, 25(4):439–473, May 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chan:1988:GDS

[648] Edward P. F. Chan and Héctor J. Hernández. On generating database schemes bounded or constant-time-maintainable by extensibility. Acta Informatica, 25(5):475–496, June 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mitchell:1988:ICR

[649] William P. R. Mitchell. Inductive completion with retracts. Acta Informatica, 25(5):497–514, June 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pudlak:1988:GC

[650] Pavel Pudlák, Vojtěch Rödl, and Petr Savický. Graph complexity. Acta Informatica, 25(5):515–535, June 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1988:AGG

[651] Joost Engelfriet, George Leih, and Grzegorz Rozenberg. Apex graph grammars and attribute grammars. Acta Informatica, 25(5):537–571, June 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Narendran:1988:EFO

[652] Paliath Narendran and Friedrich Otto. Elements of finite order for finite weight-reducing and confluent Thue systems. Acta Informatica, 25(5): 573–591, June 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Back:1988:CRP

[653] R. J. R. Back. A calculus of refinements for program derivations. Acta Informatica, 25(6):593–624, August 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fiadeiro:1988:SVD

[654] José Fiadeiro and Amílcar Sernadas. Specification and verification of database dynamics. Acta Informatica, 25(6):625–661, August 1988. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Shen:1988:CDD

[655] Sheldon Shen. Cooperative distributed dynamic load balancing. Acta Informatica, 25(6):663–676, August 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tripathi:1988:LSD

[656] Satish K. Tripathi, David Finkel, and Erol Gelenbe. Load sharing in distributed systems with failures. Acta Informatica, 25(6):677–689, August 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Levi:1988:CSL

[657] Giorgio Levi and Catuscia Palamidessi. Contributions to the semantics of logic perpetual processes. Acta Informatica, 25(6):691–711, August 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anonymous:1988:AIV

[658] Anonymous. Author index for volumes 1–25 (1971–1988). Acta Informatica, 25(6):715–735, August 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vandeVorst:1988:FDP

[659] Johannes G. G. van de Vorst. The formal development of a parallel program performing LU-decomposition. Acta Informatica, 26(1–2):1– 17, October 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Jacobs:1988:CSS

[660] Dean Jacobs and Martin S. Feather. Corrections to A Synthesis of Several Sorting Algorithms by J. Darlington. Acta Informatica, 26(1–2):19–23, October 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [241, 299].

Bertino:1988:CSI

[661] Elisa Bertino and Daniela Musto. Correctness of semantic integrity checking in database management systems. Acta Informatica, 26(1–2): 25–57, October 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Degano:1988:DOS

[662] Pierpaolo Degano, Rocco De Nicola, and Ugo Montanari. A distributed operational semantics for CCS based on condition/event systems. Acta Informatica, 26(1–2):59–91, October 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sutner:1988:MPA

[663] Klaus Sutner and Wolfgang Maass. Motion planning among time dependent obstacles. Acta Informatica, 26(1–2):93–122, October 1988. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Devroye:1988:ATR

[664] Luc Devroye. Applications of the theory of records in the study of random trees. Acta Informatica, 26(1–2):123–130, October 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1988:HLT

[665] Joost Engelfriet and Heiko Vogler. High level tree transducers and iterated pushdown tree transducers. Acta Informatica, 26(1–2):131–192, October 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cunto:1988:TUM

[666] Walter Cunto and Patricio V. Poblete. Transforming unbalanced multiway trees into a practical external data structure. Acta Informatica, 26 (3):193–211, November 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lipps:1988:ARE

[667] Peter Lipps, Ulrich Möncke, Matthias Olk, and Reinhard Wilhelm. Attribute (re)evaluation in OPTRAN. Acta Informatica, 26(3):213–239, November 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See also: ESPRIT PROSPECTRA Project Report S.1.3 - R.4.0, University des Saarlandes, Saarbrücken (1986).

Kouvatsos:1988:MET

[668] Demetres D. Kouvatsos and John Almond. Maximum entropy twostation cyclic queues with multiple general servers. Acta Informatica, 26 (3):241–267, November 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See errata [683, 699].

Levcopoulos:1988:BST

[669] Christos Levcopoulos and Mark H. Overmars. A balanced search tree with O(1) worst-case update time. Acta Informatica, 26(3):269–277, November 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Szelepcsenyi:1988:MFE

[670] Róbert Szelepcsényi. The method of forced enumeration for nondeterministic automata. Acta Informatica, 26(3):279–284, November 1988. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hehner:1988:EPM

[671] Eric C. R. Hehner, Lorene E. Gupta, and Andrew J. Malton. Erratum: “Predicative methodology” [Acta Inform. 23 (1986), no. 5, 487– 505; MR 87k:68090]. Acta Informatica, 26(3):285, November 1988. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [582].

Morris:1989:LDR

[672] Joseph M. Morris. Laws of data refinement. Acta Informatica, 26(4): 287–308, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hesselink:1989:PTS

[673] Wim H. Hesselink. Predicate-transformer semantics of general recursion. Acta Informatica, 26(4):309–332, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vogler:1989:FSD

[674] Walter Vogler. Failures semantics and deadlocking of modular Petri nets. Acta Informatica, 26(4):333–348, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baeza-Yates:1989:MSF

[675] Ricardo A. Baeza-Yates. Modeling splits in file structures. Acta Informatica, 26(4):349–362, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kobler:1989:CA

[676] Johannes Köbler, Uwe Schöning, and Jacobo Torán. On counting and approximation. Acta Informatica, 26(4):363–379, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Belaga:1989:TMM

[677] Edward G. Belaga. Through the mincing machine with a Boolean layer cake. Nonstandard computations over Boolean circuits in the lowerbounds-to-circuit-size complexity proving. Acta Informatica, 26(4):381– 407, February 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bijlsma:1989:SPR

[678] A. Bijlsma, P. A. Matthews, and J. G. Wiltink. A sharp proof rule for procedures in wp semantics. Acta Informatica, 26(5):409–419, March 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Zhang:1989:UOB

[679] Bin Zhang and Meichun Hsu. Unsafe operations in B-trees. Acta Informatica, 26(5):421–438, March 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baeza-Yates:1989:EBT

[680] Ricardo A. Baeza-Yates. Expected behaviour of B^{+}-trees under random insertions. Acta Informatica, 26(5):439–471, March 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huang:1989:SDK

[681] Bing-Chao Huang and Michael A. Langston. Stable duplicate-key extraction with optimal time and space bounds. Acta Informatica, 26(5): 473–484, March 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Helman:1989:FNC

[682] Paul Helman. A family of NP-complete data aggregation problems. Acta Informatica, 26(5):485–499, March 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kouvatsos:1989:EMEa

[683] Demetres D. Kouvatsos and John Almond. Erratum: “Maximum entropy two-station cyclic queues with multiple general servers”. Acta Informatica, 26(5):501, March 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [668, 699].

Pierre:1989:RIV

[684] Laurent Pierre and Sylviane R. Schwer. Rational index of vector addition systems languages. Acta Informatica, 26(6):503–525, July 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Seidl:1989:FDA

[685] Helmut Seidl. On the finite degree of ambiguity of finite tree automata. Acta Informatica, 26(6):527–542, July 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huang:1989:AMP

[686] Yennun Huang and Pankaj Jalote. Analytic models for the primary site approach to fault-tolerance. Acta Informatica, 26(6):543–557, July 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Akyildiz:1989:CAN

[687] Ian F. Akyildiz and Horst von Brand. Computational algorithms for networks of queues with rejection blocking. Acta Informatica, 26(6):559– 576, July 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ramarao:1989:CDC

[688] K. V. S. Ramarao. Complexity of distributed commit protocols. Acta Informatica, 26(6):577–595, July 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Litwin:1989:CTH

[689] Witold Litwin, Yehoshua Sagiv, and K. Vidyasankar. Concurrency and trie hashing. Acta Informatica, 26(7):597–614, September 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Roth:1989:NVN

[690] Mark A. Roth, Henry F. Korth, and Abraham Silberschatz. Null values in nested relational databases. Acta Informatica, 26(7):615–642, September 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See correction [760] and addendum [761].

Han:1989:TLB

[691] Yijie Han and Yoshihide Igarashi. Time lower bounds for parallel sorting on a mesh-connected processor array. Acta Informatica, 26(7):643–655, September 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Habel:1989:MDP

[692] Annegret Habel, Hans-Jörg Kreowski, and Walter Vogler. Metatheorems for decision problems on hyperedge replacement graph languages. Acta Informatica, 26(7):657–677, September 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Adiri:1989:SMF

[693] Igal Adiri, John Bruno, Esther Frostig, and A. H. G. Rinnooy Kan. Single machine flow-time scheduling with a single breakdown. Acta Informatica, 26(7):679–696, September 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Csirik:1989:LAV

[694] J. Csirik. An on-line algorithm for variable-sized bin packing. Acta Informatica, 26(8):697–709, October 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kemp:1989:EAW

[695] R. Kemp. The expected additive weight of trees. Acta Informatica, 26 (8):711–740, October 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sen:1989:SOF

[696] Arunabha Sen. Supercube: An optimally fault tolerant network architecture. Acta Informatica, 26(8):741–748, October 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Autebert:1989:IGC

[697] Jean-Michel Autebert and Joaquim Gabarró. Iterated GSM’s and co-CFL. Acta Informatica, 26(8):749–769, October 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Simon:1989:CRA

[698] Hans-Ulrich Simon. Continuous reductions among combinatorial optimization problems. Acta Informatica, 26(8):771–785, October 1989. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kouvatsos:1989:EMEb

[699] Demetres D. Kouvatsos and John Almond. Erratum: “Maximum entropy two-station cyclic queues with multiple general servers”. Acta Informatica, 26(8):787, October 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [668, 683].

Alblas:1989:ITP

[700] Henk Alblas. Iteration of transformation passes over attributed program trees. Acta Informatica, 27(1):1–40, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Langenhop:1989:MDB

[701] Carl E. Langenhop and William E. Wright. A model of the dynamic behavior of B-trees. Acta Informatica, 27(1):41–59, November 1989. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ruzicka:1989:ALR

[702] Peter Ružička and Igor Prívara. An almost linear Robinson unification algorithm. Acta Informatica, 27(1):61–71, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lau:1989:NSC

[703] K. K. Lau. A note on synthesis and classification of sorting algorithms. Acta Informatica, 27(1):73–80, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gonczarowski:1989:SVC

[704] Jakob Gonczarowski and Manfred K. Warmuth. Scattered versus context-sensitive rewriting. Acta Informatica, 27(1):81–95, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huang:1989:IMD

[705] Chua-Huang Huang and Christian Lengauer. An incremental mechanical development of systolic solutions to the algebraic path problem. Acta Informatica, 27(2):97–124, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Taubner:1989:SFS

[706] Dirk Taubner and Walter Vogler. Step failures semantics and a complete proof system. Acta Informatica, 27(2):125–156, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rego:1989:SEC

[707] Vernon Rego. Some efficient computational algorithms related to phase models. Acta Informatica, 27(2):157–177, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1989:HMC

[708] Karel Culik II and Juhani Karhumäki. HDTOL matching of computations of multitape automata. Acta Informatica, 27(2):179–191, November 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:1989:MAP

[709] Friedrich L. Bauer. In memoriam: Andrei Petrovich Ershov, 19 April 1931–8 December 1988. Acta Informatica, 27(3):193–194, December 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Salzberg:1989:MSR

[710] Betty Salzberg. Merging sorted runs using large main memory. Acta Informatica, 27(3):195–215, December 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schmidt:1989:DSD

[711] Gunther Schmidt, Rudolf Berghammer, and Hans Zierer. Describing semantic domains with sprouts. Acta Informatica, 27(3):217–245, December 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kouvatsos:1989:MEP

[712] Demetres D. Kouvatsos and Nasreddine Tabet-Aouel. A maximum entropy priority approximation for a stable G/G/1 queue. Acta Informatica, 27(3):247–286, December 1989. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Morris:1990:TPT

[713] Joseph M. Morris. Temporal predicate transformers and fair termination. Acta Informatica, 27(4):287–313, March 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1990:PSSa

[714] Andrzej Ehrenfeucht and Grzegorz Rozenberg. Partial (set) 2-structures. Part I: Basic notions and the representation problem. Acta Informatica, 27(4):315–342, March 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1990:PSSb

[715] Andrzej Ehrenfeucht and Grzegorz Rozenberg. Partial (set) 2-Structures. part II: State spaces of concurrent systems. Acta Informatica, 27(4):343– 368, March 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kou:1990:EIA

[716] Lawrence T. Kou. On efficient implementation of an approximation algorithm for the Steiner tree problem. Acta Informatica, 27(4):369–380, March 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Meyer:1990:HSS

[717] J.-J. Ch. Meyer and Ernst-Rüdiger Olderog. Hiding in stream semantics of uniform concurrency. Acta Informatica, 27(5):381–397, April 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lautemann:1990:CGL

[718] Clemens Lautemann. The complexity of graph languages generated by hyperedge replacement. Acta Informatica, 27(5):399–421, April 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Overmars:1990:MRT

[719] Mark H. Overmars, Michiel H. M. Smid, Mark T. de Berg, and Marc J. van Kreveld. Maintaining range trees in secondary memory. part I: Partitions. Acta Informatica, 27(5):423–452, April 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Smid:1990:MRT

[720] Michiel H. M. Smid and Mark H. Overmars. Maintaining range trees in secondary memory. part II: Lower bounds. Acta Informatica, 27(5):453– 480, April 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Morgan:1990:DRC

[721] Carroll Morgan and P. H. B. Gardiner. Data refinement by calculation. Acta Informatica, 27(6):481–503, May 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sondergaard:1990:RTD

[722] Harald Søndergaard and Peter Sestoft. Referential transparency, definiteness and unfoldability. Acta Informatica, 27(6):505–517, May 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gelenbe:1990:OCA

[723] Erol Gelenbe and Marisela Hernández. Optimum checkpoints with age dependent failures. Acta Informatica, 27(6):519–531, May 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Taubner:1990:RCP

[724] Dirk Taubner. Representing CCS programs by finite predicate/transition nets. Acta Informatica, 27(6):533–565, May 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1990:ASO

[725] Joost Engelfriet and Willem de Jong. Attribute storage optimization by stacks. Acta Informatica, 27(6):567–581, May 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). Preliminary version: rapport 88-30, Vakgroep Informatica, Rijksuniv. te Leiden (December 1988).

Back:1990:DSL

[726] R. J. R. Back and J. von Wright. Duality in specification languages: a lattice–theoretical approach. Acta Informatica, 27(7):583–625, July 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Yu:1990:MTS

[727] Lin Yu and Daniel J. Rosenkrantz. Minimizing time-space cost for database version control. Acta Informatica, 27(7):627–663, July 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berman:1990:PRL

[728] A. Michael Berman, Marvin C. Paull, and Barbara G. Ryder. Proving relative lower bounds for incremental algorithms. Acta Informatica, 27 (7):665–683, July 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Turski:1990:SMC

[729] Wladyslaw M. Turski. On specification of multiprocessor computing. Acta Informatica, 27(8):685–696, September 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gouda:1990:ISS

[730] Mohamed G. Gouda, Rodney R. Howell, and Louis E. Rosier. The instability of self-stabilization. Acta Informatica, 27(8):697–724, September 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cleaveland:1990:TBM

[731] Rance Cleaveland. Tableau-based model checking in the propositional μ-calculus. Acta Informatica, 27(8):725–747, September 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Weber:1990:VFT

[732] Andreas Weber. On the valuedness of finite transducers. Acta Informatica, 27(8):749–780, September 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Meduna:1990:CFD

[733] Alexander Meduna. Context free derivations on word monoids. Acta Informatica, 27(8):781–786, September 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Knuth:1990:NS

[734] Donald E. Knuth. Nested satisfiability. Acta Informatica, 28(1):1–6, November 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lodaya:1990:PTE

[735] K. Lodaya and R. K. Shyamasundar. Proof theory for exception handling in a tasking environment. Acta Informatica, 28(1):7–41, November 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Khoshnevisan:1990:EMT

[736] Hessam Khoshnevisan. Efficient memo-table management strategies. Acta Informatica, 28(1):43–81, November 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1990:CSR

[737] Andrzej Ehrenfeucht and Grzegorz Rozenberg. A characterization of set representable labeled partial 2-structures through decompositions. Acta Informatica, 28(1):83–94, November 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Dick:1990:AKB

[738] Jeremy Dick, John Kalmus, and Ursula Martin. Automating the Knuth Bendix ordering. Acta Informatica, 28(2):95–119, December 1990. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Marlowe:1990:PDF

[739] Thomas J. Marlowe and Barbara G. Ryder. Properties of data flow frameworks. A unified model. Acta Informatica, 28(2):121–163, December 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). An overview of data flow frameworks and their characterizing properties is given. Contains many references to the field of data flow analysis.

Andersson:1990:BST

[740] Arne Andersson, Christian Icking, Rolf Klein, and Thomas Ottmann. Binary search trees of almost optimal height. Acta Informatica, 28(2): 165–178, December 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Latteux:1990:CRE

[741] Michel Latteux and Paavo Turakainen. On characterizations of recursively enumerable languages. Acta Informatica, 28(2):179–186, December 1990. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hennicker:1991:OIA

[742] Rolf Hennicker. Observational implementation of algebraic specifications. Acta Informatica, 28(3):187–230, February 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Best:1991:CBP

[743] Eike Best, Raymond Devillers, Astrid Kiehn, and Lucia Pomello. Concurrent bisimulations in Petri nets. Acta Informatica, 28(3):231–264, February 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tomaras:1991:MHD

[744] Panagiotis J. Tomaras and Demetres D. Kouvatsos. MRE hierarchical decomposition of general queueing network models. Acta Informatica, 28 (3):265–295, February 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anderson:1991:NEG

[745] James H. Anderson and Mohamed G. Gouda. A new explanation of the glitch phenomenon. Acta Informatica, 28(4):297–309, April 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kapur:1991:SCG

[746] Deepak Kapur, Paliath Narendran, Daniel J. Rosenkrantz, and Hantao Zhang. Sufficient–completeness, ground–reducibility and their complexity. Acta Informatica, 28(4):311–350, April 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bozapalidis:1991:ECS

[747] Symeon Bozapalidis. Effective construction of the syntactic algebra of a recognizable series on trees. Acta Informatica, 28(4):351–363, April 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hulsmann:1991:TFH

[748] Klaus Hülsmann and Gunter Saake. Theoretical foundations of handling large substitution sets in temporal integrity monitoring. Acta Informatica, 28(4):365–407, April 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Habermann:1991:AJP

[749] A. Nico Habermann. Alan J. Perlis (1922–1990). Acta Informatica, 28 (5):409–410, May 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Snelting:1991:CCR

[750] Gregor Snelting. The calculus of context relations. Acta Informatica, 28(5):411–445, May 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Critchlow:1991:EPD

[751] Carol Critchlow and Prakash Panangaden. The expressive power of delay operators in SCCS. Acta Informatica, 28(5):447–452, May 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bovet:1991:MDS

[752] Daniel P. Bovet and Pierluigi Crescenzi. Minimum-delay schedules in layered networks. Acta Informatica, 28(5):453–461, May 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kolla:1991:VFP

[753] Reiner Kolla and Bernd Serf. The virtual feedback problem in hierarchical representations of combinational circuits. Acta Informatica, 28(5): 463–476, May 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Otto:1991:DPF

[754] Friedrich Otto and Louxin Zhang. Decision problems for finite special string-rewriting systems that are confluent on some congruence class. Acta Informatica, 28(5):477–510, May 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vandenBos:1991:PCO

[755] Jan van den Bos and Chris Laffra. PROCOL: a concurrent objectoriented language with protocols delegation and constraints. Acta Informatica, 28(6):511–538, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kastens:1991:ADT

[756] Uwe Kastens and W. M. Waite. An abstract data type for name analysis. Acta Informatica, 28(6):539–558, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Harrison:1991:ENL

[757] Peter G. Harrison. On the expansion of non-linear functions. Acta Informatica, 28(6):559–574, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1991:BPP

[758] Joost Engelfriet. Branching processes of Petri nets. Acta Informatica, 28(6):575–591, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Faigle:1991:SOD

[759] Ulrich Faigle and W. Kern. Some order dimension bounds for communication complexity problems. Acta Informatica, 28(6):593–601, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Levene:1991:CNV

[760] Mark Levene and George Loizou. Correction to Null Values in Nested Relational Databases by Mark A. Roth, H. F. Korth, and A. Silberschatz. Acta Informatica, 28(6):603–605, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [690, 761].

Roth:1991:ANV

[761] Mark A. Roth, Henry F. Korth, and Abraham Silberschatz. Addendum to Null values in nested relational databases. Acta Informatica, 28(6):

607–610, July 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See [690, 760].

Pohlmann:1991:FPA

[762] Werner Pohlmann. A fixed point approach to parallel discrete event simulation. Acta Informatica, 28(7):611–629, October 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Qian:1991:EPB

[763] Xiaolei Qian. The expressive power of the bounded-iteration construct. Acta Informatica, 28(7):631–656, October 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bernabeu-Auban:1991:RFS

[764] José M. Bernabéu-Aubán, Mustaque Ahamad, and Mostafa H. Ammar. Resource finding in store-and-forward networks. Acta Informatica, 28 (7):657–680, October 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Yen:1991:PSM

[765] Hsu-Chun Yen. Priority systems with many identical processes. Acta Informatica, 28(7):681–692, October 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cantor:1991:FMP

[766] David G. Cantor and Erich Kaltofen. On fast multiplication of polynomials over arbitrary algebras. Acta Informatica, 28(7):693–701, October 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Whaley:1991:PTE

[767] Thomas P. Whaley. Postorder trees and Eulerian numbers. Acta Informatica, 28(7):703–712, October 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Horwitz:1991:ECP

[768] Susan Horwitz and Thomas W. Reps. Efficient comparison of program slices. Acta Informatica, 28(8):713–732, November 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Pritchard:1991:OAE

[769] Paul Pritchard. Opportunistic algorithms for eliminating supersets. Acta Informatica, 28(8):733–754, November 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gillies:1991:GRS

[770] Donald W. Gillies and Jane W.-S. Liu. Greed in resource scheduling. Acta Informatica, 28(8):755–775, November 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Atzeni:1991:IDS

[771] Paolo Atzeni and Edward P. F. Chan. Independent database schemes under functional and inclusion dependencies. Acta Informatica, 28(8): 777–799, November 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Alfa:1991:AQL

[772] Attahiru Sule Alfa and Mingyuan Chen. Approximating queue lengths in M(t)/G/1 queue using the maximum entropy principle. Acta Informatica, 28(8):801–815, November 1991. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rajasekaran:1992:PIS

[773] Sanguthevar Rajasekaran and Sandeep Sen. On parallel integer sorting. Acta Informatica, 29(1):1–15, February 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Josephs:1992:RPT

[774] Mark B. Josephs. Receptive process theory. Acta Informatica, 29(1):17– 31, February 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hayes:1992:MRZ

[775] Ian J. Hayes. Multi-relations in Z: a cross between multi-sets and binary relations. Acta Informatica, 29(1):33–62, February 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hinz:1992:RTL

[776] Didier Y. Hinz. A run-time load balancing strategy for highly parallel systems. Acta Informatica, 29(1):63–94, February 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Roth:1992:EBP

[777] Peter Roth. Every binary pattern of length six is avoidable on the twoletter alphabet. Acta Informatica, 29(1):95–107, February 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fischer:1992:SFE

[778] Charles N. Fischer and Jon Mauney. A simple, fast, and effective LL(1) error repair algorithm. Acta Informatica, 29(2):109–120, April 1992. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Xu:1992:LMD

[779] J. Xu. On-line multiversion database concurrency control. Acta Informatica, 29(2):121–160, April 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1992:CFH

[780] Joost Engelfriet and Linda Heyker. Context-free hypergraph grammars have the same term-generating power as attribute grammars. Acta Informatica, 29(2):161–210, April 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Harrison:1992:SFI

[781] Peter G. Harrison and Hessam Khoshnevisan. On the synthesis of function inverses. Acta Informatica, 29(3):211–239, June 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

ONeil:1992:STI

[782] Patrick E. O’Neil. The SB-tree. an index-sequential structure for highperformance sequential access. Acta Informatica, 29(3):241–265, June 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Carlsson:1992:PPS

[783] Svante Carlsson and Jingsen Chen. On partitions and presortedness of sequences. Acta Informatica, 29(3):267–280, June 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Du:1992:MFA

[784] M. W. Du and S. C. Chang. A model and a fast algorithm for multiple errors spelling correction. Acta Informatica, 29(3):281–302, June 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baeten:1992:APC

[785] Jos C. M. Baeten and Frits W. Vaandrager. An algebra for process creation. Acta Informatica, 29(4):303–334, July 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ouksel:1992:RES

[786] M. Aris Ouksel and Otto Mayer. A robust and efficient spatial data structure. the nested interpolation-based grid file. Acta Informatica, 29 (4):335–373, July 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lee:1992:SMF

[787] Chung-Yee Lee and Surya Danusaputro Liman. Single machine flowtime scheduling with scheduled maintenance. Acta Informatica, 29(4): 375–382, July 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hinrichs:1992:ARS

[788] Klaus Hinrichs, Jürg Nievergelt, and Peter Schorn. An all-round sweep algorithm for 2-dimensional nearest-neighbor problems. Acta Informatica, 29(4):383–394, July 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Richards:1992:SQD

[789] Dana S. Richards and Jeffrey S. Salowe. Stacks, queues, and deques with order-statistic operations. Acta Informatica, 29(5):395–414, September 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tcha:1992:PST

[790] Dong wan Tcha, Bum-Il Lee, and Young duck Lee. Processors selection and traffic splitting in a parallel processors system. Acta Informatica, 29 (5):415–423, September 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Slobodova:1992:CAM

[791] Anna Slobodová. Communication for alternating machines. Acta Informatica, 29(5):425–441, September 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baeza-Yates:1992:UMT

[792] Ricardo A. Baeza-Yates and Walter Cunto. Unbalanced multiway trees improved by partial expansions. Acta Informatica, 29(5):443–460, September 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fisher:1992:YYP

[793] Anthony J. Fisher. A “yo-yo” parsing algorithm for a large class of van Wijngaarden grammars. Acta Informatica, 29(5):461–481, September 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

deLuca:1992:FRS

[794] Aldo de Luca and Stefano Varricchio. On finitely recognizable semigroups. Acta Informatica, 29(5):483–498, September 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Peng:1992:ACC

[795] Wuxu Peng and S. Purushothaman. Analysis of a class of communicating finite state machines. Acta Informatica, 29(6–7):499–522, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Halbwachs:1992:EPR

[796] Nicolas Halbwachs, Fabienne Lagnier, and Christophe Ratel. An experience in proving regular networks of processes by modular model checking. Acta Informatica, 29(6–7):523–543, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Harju:1992:DSF

[797] Tero Harju, H. C. M. Kleijn, and Michel Latteux. Deterministic sequential functions. Acta Informatica, 29(6–7):545–554, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nielsen:1992:ETS

[798] Mogens Nielsen, Grzegorz Rozenberg, and P. S. Thiagarajan. Elementary transition systems and refinement. Acta Informatica, 29(6–7):555–578, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rego:1992:NAH

[799] Vernon Rego. Naive asymptotics for hitting time bounds in Markov chains. Acta Informatica, 29(6–7):579–594, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Schoett:1992:TIT

[800] Oliver Schoett. Two impossibility theorems on behaviour specification of abstract data types. Acta Informatica, 29(6–7):595–621, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fachini:1992:CST

[801] E. Fachini, Andrea Maggiolo-Schettini, and Davide Sangiorgi. Classes of systolic Y -tree automata and a comparison with systolic trellis automata. Acta Informatica, 29(6–7):623–643, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Marcus:1992:EOE

[802] Leo Marcus and Telis Menas. Expressibility of output equals input. Negative and positive results. Acta Informatica, 29(6–7):645–662, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Weber:1992:LVF

[803] Andreas Weber. On the lengths of values in a finite transducer. Acta Informatica, 29(6–7):663–687, November 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Sannella:1992:TFD

[804] Donald Sannella, Stefan Sokołowski, and Andrzej Tarlecki. Toward formal development of programs from algebraic specifications: Parameterisation revisited. Acta Informatica, 29(8):689–736, December 1992. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

ftp://ftp.dcs.ed.ac.uk/pub/dts/param.ps.

Arun-Kumar:1992:EPP

[805] S. Arun-Kumar and M. Hennessy. An efficiency preorder for processes. Acta Informatica, 29(8):737–760, December 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fachini:1992:LAS

[806] E. Fachini, Angelo Monti, Margherita Napoli, and Domenico Parente. Languages accepted by systolic Y -tree automata: Structural characterizations. Acta Informatica, 29(8):761–778, December 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Atanasiu:1992:CCB

[807] Adrian Atanasiu. A class of coders based on GSM. Acta Informatica, 29 (8):779–791, December 1992. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Thomsen:1993:PCS

[808] Bent Thomsen. Plain CHOCS: a second generation calculus for higher order processes. Acta Informatica, 30(1):1–59, January 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Stewart:1993:LSC

[809] Iain A. Stewart. Logical and schematic characterization of complexity classes. Acta Informatica, 30(1):61–87, January 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Petit:1993:RTL

[810] Antoine Petit. Recognizable trace languages, distributed automata and the distribution problem. Acta Informatica, 30(1):89–101, January 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Marriott:1993:FAI

[811] Kim Marriott. Frameworks for abstract interpretation. Acta Informatica, 30(2):103–129, March 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hac:1993:PRI

[812] Anna Hac. Performance and reliability improvement by using asynchronous algorithms in disk buffer cache memory. Acta Informatica, 30(2):131–146, March 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Navarro:1993:CRS

[813] Marisa Navarro, Fernando Orejas, and Jean-Luc Rémy. Contextual rewriting as a sound and complete proof method for conditional LOGspecifications. Acta Informatica, 30(2):147–180, March 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Nicollin:1993:ATG

[814] Xavier Nicollin, Joseph Sifakis, and Sergio Yovine. From ATP to timed graphs and hybrid systems. Acta Informatica, 30(2):181–202, March 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Amerins:1993:EE

[815] Paul S. Amerins, Ricardo A. Baeza-Yates, and Derick Wood. On efficient entreeings. Acta Informatica, 30(3):203–213, May 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ding:1993:RMM

[816] Yuzheng Ding and Mark Allen Weiss. The relaxed min-max heap. A mergeable double-ended priority queue. Acta Informatica, 30(3):215– 231, May 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Poblete:1993:AHS

[817] Patricio V. Poblete. The analysis of heuristics for search trees. Acta Informatica, 30(3):233–248, May 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anderson:1993:FGS

[818] James H. Anderson. A fine-grained solution to the mutual exclusion problem. Acta Informatica, 30(3):249–265, May 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iwata:1993:RAS

[819] Shigeki Iwata, Takumi Kasai, and Etsuro Moriya. Relations among simultaneous complexity classes of nondeterministic and alternating Turing machines. Acta Informatica, 30(3):267–278, May 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1993:SMR

[820] Karel Culik II and Simant Dube. L-systems and mutually recursive function systems. Acta Informatica, 30(3):279–302, May 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lehmann:1993:OSL

[821] Thomas Lehmann and Jacques Loeckx. OBSCURE: a specification language for abstract data types. Acta Informatica, 30(4):303–350, July 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paun:1993:SPC

[822] Gheorghe Păun. On the synchronization in parallel communicating grammar systems. Acta Informatica, 30(4):351–367, July 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Yellin:1993:SDT

[823] Daniel M. Yellin. Speeding up dynamic transitive closure for bounded degree graphs. Acta Informatica, 30(4):369–384, July 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cassaigne:1993:UBP

[824] Julien Cassaigne. Unavoidable binary patterns. Acta Informatica, 30(4): 385–395, July 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kratochvil:1993:SCN

[825] Jan Kratochvíl and Mirko Křivánek. Satisfiability of co-nested formulas. Acta Informatica, 30(4):397–403, July 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Spuler:1993:OBS

[826] David Spuler. The optimal binary search tree for Andersson’s search algorithm. Acta Informatica, 30(5):405–407, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coffman:1993:SSF

[827] Edward G. Coffman, Jr., Leopold Flatto, and A. Ya. Kreĭnin. Scheduling saves in fault-tolerant computations. Acta Informatica, 30(5):409–423, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Aref:1993:DWM

[828] Walid G. Aref and Hanan Samet. Decomposing a window into maximal quadtree blocks. Acta Informatica, 30(5):425–439, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mateescu:1993:SPS

[829] Alexandru Mateescu and Arto Salomaa. On simplest possible solutions for Post correspondence problems. Acta Informatica, 30(5):441–457, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Devroye:1993:EHF

[830] Luc Devroye. On the expected height of fringe-balanced trees. Acta Informatica, 30(5):459–466, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kumar:1993:IPU

[831] K. Narayan Kumar and Paritosh K. Pandya. Infinitary parallelism without unbounded nondeterminism in CSP. Acta Informatica, 30(5):467– 487, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

VandenBussche:1993:MDC

[832] Jan Van den Bussche. On minimizing the ∀-¬ degree of a connectivefree formula. Acta Informatica, 30(5):489–502, August 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Singh:1993:PRF

[833] Ambuj K. Singh. Program refinement in fair transition systems. Acta Informatica, 30(6):503–535, September 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bevers:1993:PTC

[834] Eddy Bevers and Johan Lewi. Proving termination of (conditional) rewrite systems. A semantic approach. Acta Informatica, 30(6):537–568, September 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Qian:1993:ASH

[835] Zhenyu Qian. An algebraic semantics of higher-order types with subtypes. Acta Informatica, 30(6):569–607, September 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Manna:1993:MR

[836] Zohar Manna and Amir Pnueli. Models for reactivity. Acta Informatica, 30(7):609–678, October 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tuzhilin:1993:QDP

[837] Alexander Tuzhilin. Querying Datalog programs with temporal logic. Acta Informatica, 30(7):679–700, October 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoare:1993:NFA

[838] C. A. R. Hoare, Jifeng He, and A. C. A. Sampaio. Normal form approach to compiler design. Acta Informatica, 30(8):701–739, November 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hernandez:1993:ENR

[839] Héctor J. Hernández. Extended nested relations. Acta Informatica, 30 (8):741–771, November 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Buckle:1993:CMJ

[840] John Buckle. A characterisation of meet and join respecting pre-orders and congruences on finite lattices. Acta Informatica, 30(8):773–785, November 1993. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Scott:1994:NH

[841] Dana Scott. A. Nico Habermann 1932–1993. Acta Informatica, 31(1):1– 3, January 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Costa:1994:OIB

[842] J. F. Costa, Amílcar Sernadas, and Cristina Sernadas. Object inheritance beyond subtyping. Acta Informatica, 31(1):5–26, January 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Su:1994:DPS

[843] Jianwen Su. Dependency preservation in semantic databases. Acta Informatica, 31(1):27–54, January 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

DeFrancesco:1994:PFC

[844] Nicoletta De Francesco and Paola Inverardi. Proving finiteness of CCS processes by non-standard semantics. Acta Informatica, 31(1):55–80, January 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Baier:1994:CBE

[845] Christel Baier and Mila E. Majster-Cederbaum. The connection between an event structure semantics and an operational semantics for TCSP. Acta Informatica, 31(1):81–104, January 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vonWright:1994:LDR

[846] J. von Wright. The lattice of data refinement. Acta Informatica, 31(2): 105–135, ???? 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mongenet:1994:GTM

[847] Catherine Mongenet, Philippe Clauss, and Guy-René Perrin. Geometrical tools to map systems of affine recurrence equations on regular arrays.

Acta Informatica, 31(2):137–160, ???? 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ehrenfeucht:1994:CFT

[848] Andrzej Ehrenfeucht, Paulien ten Pas, and Grzegorz Rozenberg. Context-free text grammars. Acta Informatica, 31(2):161–206, ???? 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). From Joachim Schrod: This paper describes “the basic theory behind text grammars needed for document structure specification”.

Grassi:1994:DEH

[849] Vincenzo Grassi. Dependability evaluation of hierarchical systems. Acta Informatica, 31(3):207–233, April 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bozapalidis:1994:TFF

[850] Symeon Bozapalidis and George Rahonis. On two families of forests. Acta Informatica, 31(3):235–260, April 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lee:1994:BLC

[851] Myung-Joon Lee and Kwang-Moo Choe. Boundedly LR(k)-conflictable grammars. Acta Informatica, 31(3):261–283, April 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wu:1994:CMA

[852] Hongzhong Wu. On n-column 0, 1-matrices with all k-projections surjective. Acta Informatica, 31(3):285–299, April 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Katajainen:1994:SMS

[853] Jyrki Katajainen and Tomi Pasanen. Sorting multisets stably in minimum space. Acta Informatica, 31(4):301–313, June 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ibarra:1994:CBS

[854] Oscar H. Ibarra and Nicholas Q. Trân. On communication-bounded synchronized alternating finite automata. Acta Informatica, 31(4):315–327, June 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Meinke:1994:RSO

[855] Karl Meinke. A recursive second order initial algebra specification of primitive recursion. Acta Informatica, 31(4):329–340, June 1994. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:1994:CFG

[856] Joost Engelfriet, Linda Heyker, and George Leih. Context-free graph languages of bounded degree are generated by apex graph grammars. Acta Informatica, 31(4):341–378, June 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Diekert:1994:DAA

[857] Volker Diekert and Anca Muscholl. Deterministic asynchronous automata for infinite traces. Acta Informatica, 31(4):379–397, June 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). A preliminary version was presented at STACS’93, Lecture Notes in Computer Science, 665 (1993).

Jones:1994:TLP

[858] Cliff B. Jones and C. A. (Kees) Middelburg. A typed logic of partial functions reconstructed classically. Acta Informatica, 31(5):399–430, August 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kuhnemann:1994:SIF

[859] Armin Kühnemann and Heiko Vogler. Synthesized and inherited functions. A new computational model for syntax-directed semantics. Acta Informatica, 31(5):431–477, August 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Farr:1994:PSC

[860] Graham Farr. On problems with short certificates. Acta Informatica, 31 (5):479–502, August 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Woeginger:1994:HPM

[861] Gerhard J. Woeginger. Heuristics for parallel machine scheduling with delivery times. Acta Informatica, 31(6):503–512, September 1994. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hull:1994:DIR

[862] Richard Hull and Jianwen Su. Domain independence and the relational calculus. Acta Informatica, 31(6):513–524, September 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Paun:1994:PTG

[863] Gheorghe Păun and Grzegorz Rozenberg. Prescribed teams of grammars. Acta Informatica, 31(6):525–537, September 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

deLuca:1994:WQO

[864] Aldo de Luca and Stefano Varricchio. Well quasi-orders and regular languages. Acta Informatica, 31(6):539–557, September 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Falin:1994:ITA

[865] G. I. Falin, M. Martìn Dìaz, and J. R. Artalejo. Information theoretic approximations for the M/G/1 retrial queue. Acta Informatica, 31(6): 559–571, September 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Torlone:1994:UOD

[866] Riccardo Torlone. Update operations in deductive databases with functional dependencies. Acta Informatica, 31(6):573–600, September 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kastens:1994:MRA

[867] Uwe Kastens and William M. Waite. Modularity and reusability in attribute grammars. Acta Informatica, 31(7):601–627, October 1994. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ruhl:1994:BRT

[868] Astrid R. Rühl. On bounds of response time performance achievable by multiclass single-server queues. Acta Informatica, 31(7):629–650, October 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bernot:1994:BAA

[869] Gilles Bernot, Michel Bidoit, and Teodor Knapik. Behavioural approaches to algebraic specifications: a comparative study. Acta Informatica, 31(7):651–671, October 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Flajolet:1994:MTA

[870] Philippe Flajolet and Mordecai Golin. Mellin transforms and asymptotics: The mergesort recurrence. Acta Informatica, 31(7):673–696, October 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kiehn:1994:CLC

[871] Astrid Kiehn. Comparing locality and causality based equivalences. Acta Informatica, 31(8):697–718, November 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hauschildt:1994:PNA

[872] Dirk Hauschildt and Matthias Jantzen. Petri net algorithms in the theory of matrix grammars. Acta Informatica, 31(8):719–728, November 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Spuler:1994:OST

[873] David Spuler. Optimal search trees using two-way key comparisons. Acta Informatica, 31(8):729–740, November 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ferdinand:1994:TAC

[874] Christian Ferdinand, Helmut Seidl, and Reinhard Wilhelm. Tree automata for code selection. Acta Informatica, 31(8):741–760, November 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Culik:1994:PSI

[875] Karel Culik II and Jarkko Kari. On the power of L-systems in image generation. Acta Informatica, 31(8):761–773, November 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kirschenhofer:1994:PLR

[876] Peter Kirschenhofer and Helmut Prodinger. The path length of random skip lists. Acta Informatica, 31(8):775–792, November 1994. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Biskup:1995:OSC

[877] Joachim Biskup, Pratul Dublish, and Yehoshua Sagiv. Optimization of a subclass of conjunctive queries. Acta Informatica, 32(1): 1–26, February 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032001/50320001.htm.

Drabent:1995:WFA

[878] Włodzimierz Drabent. What is failure? An approach to constructive negation. Acta Informatica, 32(1):27–59, February 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/5032001/
50320027.htm.

deGuzman:1995:FIB

[879] I. P. de Guzmán, M. Ojeda, and A. Valverde. A formal identification between tuples and lists with an application to listarithmetic categories. Acta Informatica, 32(1):61–78, February 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/5032001/50320061.htm.

Vaidyanathan:1995:PIS

[880] Ramachandran Vaidyanathan, Carlos R. P. Hartmann, and Pramod K. Varshney. Parallel integer sorting using small operations (*). Acta Informatica, 32(1):79–92, February 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/5032001/50320079.htm.

Chin:1995:TMD

[881] Wei-Ngan Chin and Masami Hagiya. A transformation method for dynamic-sized tabulation. Acta Informatica, 32(2):93–115, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/5032002/50320093.htm.

Walz:1995:IAG

[882] Janet A. Walz and Gregory F. Johnson. Inductive attribute grammars: a basis for incremental program execution. Acta Informatica, 32 (2):117–144, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032002/50320117.htm.

Hsu:1995:PTC

[883] W.-J. Hsu and C. V. Page. Parallel tree contraction and prefix computations on a large family of interconnection topologies. Acta Informatica, 32(2):145–153, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/5032002/50320145.htm.

Kirousis:1995:EAC

[884] Lefteris M. Kirousis and Andreas G. Veneris. Efficient algorithms for checking the atomicity of a run of read and write operations. Acta Informatica, 32(2):155–170, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/5032002/50320155.htm.

Eiter:1995:GBE

[885] Thomas Eiter. Generating Boolean μ-expressions. Acta Informatica, 32 (2):171–187, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032002/50320171.htm.

Cherkasova:1995:BSS

[886] Ludmila Cherkasova, Rodney R. Howell, and Louis E. Rosier. Bounded self-stabilizing Petri nets. Acta Informatica, 32(3):189–207, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/5032003/50320189.htm.

LeCharlier:1995:RAI

[887] Baudouin Le Charlier and Pascal Van Hentenryck. Reexecution in abstract interpretation of Prolog. Acta Informatica, 32(3):209– 253, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032003/50320209.htm.

Buchsbaum:1995:LSS

[888] Adam L. Buchsbaum, Rajamani Sundar, and Robert Endre Tarjan. Lazy structure sharing for query optimization. Acta Informatica, 32 (3):255–270, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032003/50320255.htm.

Krishnaswamy:1995:CCC

[889] Vasudha Krishnaswamy and John Bruno. On the complexity of concurrency control using semantic information (*). Acta Informatica, 32 (3):271–284, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032003/50320271.htm.

Meduna:1995:SCS

[890] Alexander Meduna. Syntactic complexity of scattered context grammars. Acta Informatica, 32(3):285–298, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/5032003/50320285.htm.

Kindler:1995:ICS

[891] Ekkart Kindler. Invariants, composition, and substitution (*). Acta Informatica, 32(4):299–312, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/5032004/50320299.htm.

Devillers:1995:IAG

[892] Raymond R. Devillers. S-invariant analysis of general recursive Petri boxes. Acta Informatica, 32(4):313–345, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/5032004/50320313.
htm.

Bernardeschi:1995:PNS

[893] Cinzia Bernardeschi, Nicoletta De Francesco, and Gigliola Vaglini. A Petri nets semantics for data flow networks. Acta Informatica, 32 (4):347–374, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/5032004/50320347.htm.

Hennessy:1995:MLM

[894] M. Hennessy and X. Liu. A modal logic for message passing processes. Acta Informatica, 32(4):375–393, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/5032004/50320375.htm.

Tomita:1995:EEP

[895] Etsuji Tomita and Kazushi Seino. The extended equivalence problem for a class of non-real-time deterministic pushdown automata. Acta Informatica, 32(4):395–413, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/5032004/50320395.htm.

Chen:1995:CRM

[896] X. J. Chen and Carlo Montangero. Compositional refinements in multiple blackboard systems. Acta Informatica, 32(5):415–458, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/5032005/50320415.htm.

Yang:1995:LAP

[897] Wuu Yang. On the look-ahead problem in lexical analysis. Acta Informatica, 32(5):459–476, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://cissun51.cis.nctu.edu.

tw/~wuuyang/lookahead.ps.Z;   http://link.springer-ny.com/link/
service/journals/00236/bibs/5032005/50320459.htm.

Neraud:1995:DMI

[898] Jean Néraud. Detecting morphic images of a word on the rank of a pattern. Acta Informatica, 32(5):477–489, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/5032005/50320477.
htm.

Jurgensen:1995:LHD

[899] H. Jürgensen and Ludwig Staiger. Local Hausdorff dimension. Acta Informatica, 32(5):491–507, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/5032005/50320491.htm.

Hennessy:1995:CTP

[900] Matthew Hennessy. Concurrent testing of processes. Acta Informatica, 32(6):509–543, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Montanari:1995:CN

[901] Ugo Montanari and Francesca Rossi. Contextual nets. Acta Informatica, 32(6):545–596, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/5032006/50320545.htm.

Sur:1995:IGN

[902] Sumit Sur and Pradip K. Srimani. IEH graphs. A novel generalization of hypercube graphs. Acta Informatica, 32(6):597–609, ???? 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/5032006/50320597.htm.

Peterson:1995:VSR

[903] W. Wesley Peterson. Variance of storage requirements for B+-trees. Acta Informatica, 32(7):611–625, October 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/5032007/50320611.htm.

Gold:1995:CDS

[904] Robert Gold. A compositional dataflow semantics for Petri nets. Acta Informatica, 32(7):627–645, October 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/5032007/50320627.htm.

Badouel:1995:TNP

[905] Eric Badouel and Philippe Darondeau. Trace nets and process automata. Acta Informatica, 32(7):647–679, October 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/5032007/50320647.htm.

Culik:1995:IWF

[906] Karel Culik II and Peter Rajčáni. Iterative weighted finite transductions. Acta Informatica, 32(7):681–703, October 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/5032007/50320681.htm.

Leavens:1995:SVO

[907] Gary T. Leavens and William E. Weihl. Specification and verification of object-oriented programs using supertype abstraction. Acta Informatica, 32(8):705–778, November 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/5032008/50320705.htm.  An

earlier version appeared as Iowa State University technical report TR#92-28d, under the title “Subtyping, Modular Specification, and Modular Verification for Applicative Object-Oriented Programs,” 1992.

Rhee:1995:FMM

[908] Chong Kye Rhee and Y. Daniel Liang. Finding a maximum matching in a permutation graph. Acta Informatica, 32(8):779–792, November 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/5032008/50320779.htm.

Anonymous:1995:AR

[909] Anonymous. Acknowledgement to referees. Acta Informatica, 32(8): 793–??, November 1995. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Taubenfeld:1996:PIR

[910] Gadi Taubenfeld and Shlomo Moran. Possibility and impossibility results in a shared memory environment. Acta Informatica, 33 (1):1–20, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033001/60330001.htm.

Duggan:1996:KTI

[911] Dominic Duggan, Gordon V. Cormack, and John Ophel. Kinded type inference for parametric overloading. Acta Informatica, 33(1): 21–68, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033001/60330021.htm.

Sangiorgi:1996:TBC

[912] Davide Sangiorgi. A theory of bisimulation for the π-calculus. Acta Informatica, 33(1):69–97, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/6033001/60330069.htm. Earlier

version published as Report ECS-LFCS-93-270, University of Edinburgh. An extended abstract appeared in the Proceedings of CONCUR ’93, LNCS 715.

Matsumoto:1996:AAP

[913] Aki Matsumoto, D. S. Han, and Takao Tsuda. Alias analysis of pointers in Pascal and Fortran 90: Dependence analysis between pointer references. Acta Informatica, 33(2):99–130, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/6033002/60330099.htm.

Lee:1996:MGL

[914] John Lee and Alan Fekete. Multi-granularity locking for nested transactions: a proof using a possibilities mapping. Acta Informatica, 33 (2):131–152, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033002/60330131.htm.

Cau:1996:PCA

[915] A. Cau and P. Collette. Parallel composition of assumption-commitment specifications: a unifying approach for shared variable and distributed message passing concurrency. Acta Informatica, 33(2):153– 176, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033002/60330153.htm.

Haldar:1996:SEW

[916] S. Haldar and K. Vidyasankar. Simple extensions of 1-writer atomic variable constructions to multiwriter ones. Acta Informatica, 33(2): 177–202, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033002/60330177.htm.

Tuzhilin:1996:MDI

[917] Alexander Tuzhilin and Zvi M. Kedem. Modeling data-intensive reactive systems with relational transition systems. Acta Informatica, 33 (3):203–231, May 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033003/60330203.htm.

Hesselink:1996:BDF

[918] Wim H. Hesselink. Bounded delay for a free address. Acta Informatica, 33(3):233–254, May 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/6033003/60330233.htm.

Kumar:1996:OCH

[919] Akhil Kumar and Kavindra Malik. Optimizing the costs of hierarchical quorum consensus. Acta Informatica, 33(3):255–275, May 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/6033003/60330255.htm.

Staalmarck:1996:SRP

[920] Gunnar Stålmarck. Short resolution proofs for a sequence of tricky formulas. Acta Informatica, 33(3):277–280, May 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/6033003/60330277.htm.

Senizergues:1996:RSF

[921] Géraud Sénizergues. On the rational subsets of the free group. Acta Informatica, 33(3):281–296, May 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/6033003/60330281.htm.

Desel:1996:SPP

[922] Jörg Desel and Wolfgang Reisig. The synthesis problem of Petri nets. Acta Informatica, 33(4):297–315, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/6033004/60330297.htm.

Aceto:1996:TCP

[923] Luca Aceto and David Murphy. Timing and causality in process algebra. Acta Informatica, 33(4):317–350, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/6033004/60330317.htm.

ONeil:1996:LSM

[924] Patrick E. O’Neil, Edward Cheng, Dieter Gawlick, and Elizabeth J. O’Neil. The log-structured merge-tree (LSM-tree). Acta Informatica, 33(4):351–385, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/6033004/60330351.htm.

Diaz:1996:PAS

[925] J. Díaz, M. J. Serna, and Jacobo Torán. Parallel Approximation Schemes for problems on planar graphs. Acta Informatica, 33(4): 387–408, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033004/60330387.htm.

Jancar:1996:FAC

[926] Petr Jančar, František Mráz, and Martin Plátek. Forgetting automata and context-free languages. Acta Informatica, 33(5):409– 420, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033005/60330409.htm.

Harman:1996:AMM

[927] N. A. Harman and J. V. Tucker. Algebraic models of microprocessors: Architecture and organisation. Acta Informatica, 33(5):

421–456, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033005/60330421.htm.

Meduna:1996:SCC

[928] Alexander Meduna. Syntactic complexity of context-free grammars over word monoids. Acta Informatica, 33(5):457–462, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/6033005/60330457.htm.

Wanke:1996:URU

[929] Egon Wanke. Undecidability of restricted uniform recurrence equations. Acta Informatica, 33(5):463–475, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/6033005/60330463.htm.

Diaconescu:1996:CBM

[930] Răzvan Diaconescu. Category-based modularisation for equational logic programming. Acta Informatica, 33(5):477–510, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/6033005/
60330477.htm.

Thorup:1996:DGE

[931] Mikkel Thorup. Disambiguating grammars by exclusion of sub-parse trees. Acta Informatica, 33(6):511–522, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/6033006/60330511.
htm.

Maggiolo-Schettini:1996:KLP

[932] Andrea Maggiolo-Schettini and Józef Winkowski. A kernel language for programmed rewriting of (hyper)graphs. Acta Informatica, 33(6): 523–546, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033006/60330523.htm.

Nurmi:1996:CBS

[933] Otto Nurmi and Eljas Soisalon-Soininen. Chromatic binary search trees: a structure for concurrent rebalancing. Acta Informatica, 33 (6):547–557, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033006/60330547.htm.

Deineko:1996:RPS

[934] Vladimir Deineko, Rüdiger Rudolf, and Gerhard J. Woeginger. On the recognition of permuted Supnick and incomplete Monge matrices. Acta Informatica, 33(6):559–569, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/6033006/60330559.htm.

Ehrenfeucht:1996:LLE

[935] Andrzej Ehrenfeucht, Gheorghe Păun, and Grzegorz Rozenberg. The linear landscape of external contextual languages. Acta Informatica, 33 (6):571–593, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/6033006/60330571.htm.

KrishnaRao:1996:RCI

[936] M. R. K. Krishna Rao. Relating confluence, innermost-confluence and outermost-confluence properties of term rewriting systems. Acta Informatica, 33(6):595–606, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/6033006/60330595.htm.

Saxena:1996:PIS

[937] Sanjeev Saxena. Parallel integer sorting and simulation amongst CRCW models. Acta Informatica, 33(7):607–619, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/6033007/60330607.
htm.

Ameur:1996:TEN

[938] Foued Ameur, Paul Fischer, Klaus-Uwe Höffgen, and Friedhelm Meyer auf der Heide. Trial and Error: a new approach to space-bounded learning. Acta Informatica, 33(7):621–630, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/6033007/60330621.htm.

Bertsch:1996:OSR

[939] Eberhard Bertsch. An observation on suffix redundancy in LL(1) error repair. Acta Informatica, 33(7):631–639, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/6033007/60330631.
htm.

Degano:1996:AAN

[940] Pierpaolo Degano, José Meseguer, and Ugo Montanari. Axiomatizing the algebra of net computations and processes. Acta Informatica, 33(7):641–667, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/6033007/60330641.htm.

Bause:1996:APN

[941] Falko Bause. On the analysis of Petri nets with static priorities. Acta Informatica, 33(7):669–685, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/6033007/60330669.htm.

Sloan:1996:RRT

[942] Robert H. Sloan and Ugo A. Buy. Reduction rules for time Petri nets. Acta Informatica, 33(7):687–706, ???? 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/6033007/60330687.htm.

Milner:1996:CI

[943] Robin Milner. Calculi for interaction. Acta Informatica, 33(8):707–737, November 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL ftp://ftp.cl.cam.ac.uk/users/rm135/ac9.ps;
http://link.springer-ny.com/link/service/journals/00236/bibs/
6033008/60330707.htm.

Reps:1996:SNI

[944] Thomas W. Reps. On the sequential nature of interprocedural program-analysis problems. Acta Informatica, 33(8):739–757, November 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/6033008/60330739.htm.

Biela:1996:RSL

[945] Andrzej Biela and Jakub Borowczyk. RETRPROV, a system that looks for axioms. Acta Informatica, 33(8):759–780, November 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/6033008/60330759.htm.

Heilbrunner:1996:DCC

[946] Stephan Heilbrunner. A direct complement construction for LR(1) grammars. Acta Informatica, 33(8):781–797, November 1996. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/6033008/
60330781.htm.

Wang:1997:WID

[947] Ke Wang, Weining Zhang, and Siu-Cheung Chau. Weakly independent database schemes. Acta Informatica, 34(1):1–22, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034001/70340001.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/7034001/70340001.
pdf;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034001/70340001.ps.gz.

Keesmaat:1997:NBC

[948] N. W. Keesmaat and H. C. M. Kleijn. Net-based control versus rational control. The relation between ITNC vector languages and rational relations. Acta Informatica, 34(1):23–57, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/7034001/70340023.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
7034001/70340023.pdf;   http://link.springer-ny.com/link/service/
journals/00236/papers/7034001/70340023.ps.gz.

Fulop:1997:MER

[949] Zoltán Fülöp and Sándor Vágvölgyi. Minimal equational representations of recognizable tree languages. Acta Informatica, 34(1): 59–84, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034001/70340059.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034001/
70340059.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034001/70340059.ps.gz;   http://link.springer-
ny.com/link/service/journals/00236/tocs/mailto:helpdesk@link.
springer.de.

Esparza:1997:DMC

[950] Javier Esparza. Decidability of model checking for infinite-state concurrent systems. Acta Informatica, 34(2):85–107, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/7034002/
70340085.htm;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034002/70340085.pdf;   http://link.springer-ny.

com/link/service/journals/00236/papers/7034002/70340085.ps.
gz.

Eiter:1997:DMP

[951] Thomas Eiter and Heikki Mannila. Distance measures for point sets and their computation. Acta Informatica, 34(2):109–133, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034002/70340109.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/7034002/70340109.
pdf;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034002/70340109.ps.gz.

Levene:1997:APF

[952] Mark Levene and George Loizou. The additivity problem for functional dependencies in incomplete relations. Acta Informatica, 34(2): 135–149, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034002/70340135.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034002/
70340135.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034002/70340135.ps.gz.

Culik:1997:CFG

[953] Karel Culik II and Jarkko Kari. Computational fractal geometry with WFA. Acta Informatica, 34(2):151–166, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/7034002/70340151.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034002/70340151.pdf;   http://link.springer-ny.com/
link/service/journals/00236/papers/7034002/70340151.ps.gz;
http://link.springer-ny.com/link/service/journals/00236/tocs/
mailto:helpdesk@link.springer.de.

Orman:1997:RDC

[954] Levent V. Orman. Relational database constraints as counterexamples. Acta Informatica, 34(3):167–189, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/7034003/70340167.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
7034003/70340167.pdf;   http://link.springer-ny.com/link/service/
journals/00236/papers/7034003/70340167.ps.gz.

Rus:1997:UGC

[955] Teodor Rus and Sriram V. Pemmaraju. Using graph coloring in an algebraic compiler. Acta Informatica, 34(3):191–209, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034003/70340191.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/7034003/70340191.
pdf;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034003/70340191.ps.gz.

Shapiro:1997:GDM

[956] Alexander Shapiro. A generalized distribution model for random recursive trees. Acta Informatica, 34(3):211–216, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/7034003/
70340211.htm;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034003/70340211.pdf;   http://link.springer-ny.
com/link/service/journals/00236/papers/7034003/70340211.ps.
gz.

Hakala:1997:SWE

[957] Ismo Hakala and Juha Kortelainen. On the system of word equations x_{1}^{i}x_{2}^{i} ⋯ x_{m}^{i} = y_{1}^{i}y_{2}^{i} ⋯ y_{n}^{i} (i = 1, 2, …) in a free monoid. Acta Informatica, 34(3):217–230, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034003/70340217.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/7034003/
70340217.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034003/70340217.ps.gz.

Djidjev:1997:RCS

[958] Hristo N. Djidjev and Shankar M. Venkatesan. Reduced constants for simple cycle graph separation. Acta Informatica, 34(3):231– 243, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034003/70340231.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034003/
70340231.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034003/70340231.ps.gz;   http://link.springer-
ny.com/link/service/journals/00236/tocs/mailto:helpdesk@link.
springer.de.

Savicky:1997:EAT

[959] Petr Savický and Ingo Wegener. Efficient algorithms for the transformation between different types of binary decision diagrams. Acta Informatica, 34(4):245–256, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034004/70340245.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/7034004/
70340245.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034004/70340245.ps.gz.

Mitrana:1997:IBS

[960] Victor Mitrana. On the interdependence between shuffle and crossing-over operations. Acta Informatica, 34(4):257–266, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034004/70340257.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/7034004/70340257.
pdf;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034004/70340257.ps.gz.

Russmann:1997:DLP

[961] Arnd Rußmann. Dynamic LL(k) parsing. Acta Informatica, 34(4): 267–289, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034004/70340267.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034004/
70340267.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034004/70340267.ps.gz.

Corradini:1997:LBS

[962] Flavio Corradini and Rocco De Nicola. Locality based semantics for process algebras. Acta Informatica, 34(4):291–324, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034004/70340291.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/7034004/70340291.
pdf;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034004/70340291.ps.gz;   http://link.springer-ny.com/
link/service/journals/00236/tocs/mailto:helpdesk@link.springer.
de.

Yamazaki:1997:HCA

[963] Koichi Yamazaki. A hierarchy of the class of apex NLC graph languages by bounds on the number of nonterminal nodes in productions. Acta Informatica, 34(5):325–335, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034005/70340325.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/7034005/
70340325.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034005/70340325.ps.gz.

Liang:1997:MFV

[964] Y. Daniel Liang and Maw-Shang Chang. Minimum feedback vertex sets in cocomparability graphs and convex bipartite graphs. Acta Informatica, 34(5):337–346, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034005/70340337.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034005/
70340337.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034005/70340337.ps.gz.

Culik:1997:IDW

[965] Karel Culik II and Simant Dube. Implementing Daubechies wavelet transform with weighted finite automata. Acta Informatica, 34(5): 347–366, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034005/70340347.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034005/
70340347.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034005/70340347.ps.gz.

Janicki:1997:FMC

[966] Ryszard Janicki and Maciej Koutny. Fundamentals of modelling concurrency using discrete relational structures. Acta Informatica, 34 (5):367–388, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034005/70340367.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034005/
70340367.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034005/70340367.ps.gz.

Morita:1997:HUP

[967] Kenichi Morita, Noritaka Nishihara, Yasunori Yamamoto, and Zhiguo Zhang. A hierarchy of uniquely parsable grammar classes and deterministic acceptors. Acta Informatica, 34(5):389–410, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034005/70340389.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/7034005/70340389.
pdf;   http://link.springer-ny.com/link/service/journals/00236/
papers/7034005/70340389.ps.gz;   http://link.springer-ny.com/
link/service/journals/00236/tocs/mailto:helpdesk@link.springer.
de.

Trogemann:1997:PAP

[968] Georg Trogemann and Matthias Gente. Performance analysis of parallel programs based on directed acyclic graphs. Acta Informatica, 34 (6):411–428, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034006/70340411.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034006/
70340411.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034006/70340411.ps.gz.

Efe:1997:OEV

[969] Kemal Efe and Nancy Eleser. An optimal emulator and VLSI layout for complete binary trees. Acta Informatica, 34(6):429– 447, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034006/70340429.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034006/
70340429.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034006/70340429.ps.gz.

Barbuti:1997:ACM

[970] Roberto Barbuti, Nicoletta Di Francesco, and Antonella Santone. Algebraic computational models of OR-parallel execution of Prolog. Acta Informatica, 34(6):449–489, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034006/70340449.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/7034006/
70340449.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/7034006/70340449.ps.gz;   http://link.springer-
ny.com/link/service/journals/00236/tocs/mailto:helpdesk@link.
springer.de.

Stephens:1997:SSP

[971] Robert Stephens. A survey of stream processing. Acta Informatica, 34(7):491–541, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034007/70340491.htm.

Rangarajan:1997:SMT

[972] Sampath Rangarajan, Yennun Huang, and Satish K. Tripathi. On the scalability and mean-time to failure of k resilient protocols. Acta Informatica, 34(7):543–556, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034007/70340543.htm.

Brisaboa:1997:TBC

[973] Nieves R. Brisaboa and Héctor J. Hernández. Testing bag-containment of conjunctive queries. Acta Informatica, 34(7):557–578, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer-ny.com/link/service/journals/00236/bibs/
7034007/70340557.htm;   http://link.springer-ny.com/link/service/
journals/00236/tocs/mailto:helpdesk@link.springer.de.

Hui:1997:MCC

[974] Chi-Chung Hui and Samuel T. Chanson. Minimal communication cost software construction in the Internet environment. Acta Informatica, 34 (8):579–595, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034008/70340579.htm.

Nymeyer:1997:CGB

[975] Albert Nymeyer and Joost-Pieter Katoen. Code generation based on formal BURS theory and heuristic search. Acta Informatica, 34(8): 597–635, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034008/70340597.htm.

Aczel:1997:NFS

[976] János Aczél and Wolfgang Ertel. A new formula for speedup and its characterization. Acta Informatica, 34(8):637–652, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034008/70340637.htm;   http://link.springer-
ny.com/link/service/journals/00236/tocs/mailto:helpdesk@link.
springer.de.

Hsieh:1997:FGD

[977] C. Samuel Hsieh. A fine-grained data-flow analysis framework. Acta Informatica, 34(9):653–665, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/7034009/70340653.htm.

Garg:1997:UCD

[978] Vijay K. Garg and Alexander I. Tomlinson. Using the causal domain to specify and verify distributed programs. Acta Informatica, 34 (9):667–686, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034009/70340667.htm.

Burnetas:1997:AIE

[979] Apostolos Burnetas, Daniel Solow, and Rishi Agarwal. An analysis and implementation of an efficient in-place bucket sort. Acta Informatica, 34(9):687–700, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034009/70340687.htm.

Baier:1997:MSP

[980] Christel Baier and Mila E. Majster-Cederbaum. Metric semantics from partial order semantics. Acta Informatica, 34(9):701– 735, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034009/70340701.htm.

Poetzsch-Heffter:1997:PRP

[981] Arnd Poetzsch-Heffter. Prototyping realistic programming languages based on formal specifications. Acta Informatica, 34(10):737–772, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/7034010/70340737.htm.

A tool supporting the generation of language-specific software from specifications is presented. Static semantics is defined by an attribution technique (e.g. for the specification of flow graphs). The dynamic semantics is defined by ASMs. As an example, an object-oriented programming language with parallelism is specified. This work is partly based upon work described in the author’s book Compiler Construction, pp. 233–247, 1994, Springer-Verlag.

Engelfriet:1997:CFG

[982] Joost Engelfriet and Jan Joris Vereijken. Context-free graph grammars and concatenation of graphs. Acta Informatica, 34(10):773– 803, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034010/70340773.htm.

Corradini:1997:PPC

[983] Flavio Corradini, Roberto Gorrieri, and Marco Roccetti. Performance preorder and competitive equivalence. Acta Informatica, 34 (11):805–835, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034011/70340805.htm.

Fernau:1997:UTR

[984] Henning Fernau. Unconditional transfer in regulated rewriting. Acta Informatica, 34(11):837–857, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/7034011/70340837.htm.

Hemaspaandra:1997:ESH

[985] Lane A. Hemaspaandra, Jörg Rothe, and Gerd Wechsung. Easy sets and hard certificate schemes. Acta Informatica, 34(11):859– 879, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034011/70340859.htm.

Bruno:1997:OFT

[986] John Bruno and Edward G. Coffman, Jr. Optimal fault-tolerant computing on multiprocessor systems. Acta Informatica, 34(12): 881–904, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034012/70340881.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/7034012/
70340881.pdf.

Laurent:1997:UDT

[987] Dominique Laurent, V. Phan Luong, and Nicolas Spyratos. The use of deleted tuples in database, querying and updating. Acta Informatica, 34 (12):905–925, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/7034012/70340905.htm;   http://

link.springer-ny.com/link/service/journals/00236/papers/7034012/

70340905.pdf. See erratum [1007].

terHofstede:1997:ACF

[988] A. H. M. ter Hofstede, E. Lippe, and Th. P. van der Weide. Applications of a categorical framework for conceptual data modeling. Acta Informatica, 34(12):927–963, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/7034012/70340927.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/7034012/
70340927.pdf.

Anonymous:1997:AR

[989] Anonymous. Acknowledgement to referees. Acta Informatica, 34(12): 965–??, ???? 1997. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Blondel:1998:SNP

[990] Vincent D. Blondel. Structured numbers: Properties of a hierarchy of operations on binary trees. Acta Informatica, 35(1):1– 15, January 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035001/80350001.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035001/
80350001.pdf.

Kemp:1998:GWL

[991] Rainer Kemp. Generating words lexicographically: an average-case analysis. Acta Informatica, 35(1):17–89, January 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/8035001/80350017.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/8035001/80350017.pdf.

Sanders:1998:DRM

[992] Beverly A. Sanders. Data refinement of mixed specifications. A generalization of UNITY. Acta Informatica, 35(2):91–129, February 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035002/80350091.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035002/80350091.
pdf.

Back:1998:RFA

[993] Ralph J. R. Back and Qiwen Xu. Refinement of fair action systems. Acta Informatica, 35(2):131–165, February 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/8035002/80350131.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
8035002/80350131.pdf.

Andrei:1998:ACC

[994] Ştefan Andrei and Cristian Masalagiu. About the Collatz conjecture. Acta Informatica, 35(2):167–179, February 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/8035002/80350167.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/8035002/80350167.pdf.

Locuratolo:1998:CCS

[995] Elvira Locuratolo and Fausto Rabitti. Conceptual classes and system classes in object databases. Acta Informatica, 35(3):181– 210, March 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035003/80350181.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035003/
80350181.pdf.

FernandezdelaVega:1998:ACC

[996] Wenceslas Fernandez de la Vega, Vangelis Th. Paschos, and Andreas N. Stafylopatis. Average-case complexity for the execution of recursive definitions on relational databases. Acta Informatica, 35(3): 211–243, March 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035003/80350211.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035003/
80350211.pdf.

Cheng:1998:CFL

[997] Edward Y. C. Cheng and Michael Kaminski. Context-free languages over infinite alphabets. Acta Informatica, 35(3):245–267, March 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035003/80350245.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035003/80350245.
pdf.

Kourie:1998:LML

[998] Derrick G. Kourie and G. Deon Oosthuizen. Lattices in machine learning: Complexity issues. Acta Informatica, 35(4):269– 292, April 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035004/80350269.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035004/
80350269.pdf.

Rabinovich:1998:MEN

[999] Alexander Rabinovich. Modularity and expressibility for nets of relations. Acta Informatica, 35(4):293–327, April 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/8035004/80350293.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
8035004/80350293.pdf.

Buchholz:1998:TCF

[1000] Thomas Buchholz and Martin Kutrib. On time computability of functions in one-way cellular automata. Acta Informatica, 35(4): 329–352, April 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035004/80350329.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035004/
80350329.pdf.

Boreale:1998:FAS

[1001] Michele Boreale and Davide Sangiorgi. A fully abstract semantics for causality in the π-calculus. Acta Informatica, 35(5):353– 400, May 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035005/80350353.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035005/
80350353.pdf.

Kari:1998:DCS

[1002] Lila Kari, Gheorghe Păun, Grzegorz Rozenberg, Arto Salomaa, and Sheng Yu. DNA computing, sticker systems, and universality. Acta Informatica, 35(5):401–420, May 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/8035005/80350401.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/8035005/
80350401.pdf.

BarYehuda:1998:PSF

[1003] Reuven Bar Yehuda and Sergio Fogel. Partitioning a sequence into few monotone subsequences. Acta Informatica, 35(5):421– 440, May 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035005/80350421.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035005/
80350421.pdf.

Baldoni:1998:CRA

[1004] Roberto Baldoni, Jean-Michel Helary, and Michel Raynal. Consistent records in asynchronous computations. Acta Informatica, 35(6): 441–455, June 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035006/80350441.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035006/
80350441.pdf.

Sagiv:1998:LBA

[1005] Mooly Sagiv, Nissim Francez, Michael Rodeh, and Reinhard Wilhelm. A logic-based approach to program flow analysis. Acta Informatica, 35 (6):457–504, June 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035006/80350457.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035006/
80350457.pdf.

Esik:1998:NAE

[1006] Zoltán Ésik and Michael Bertol. Nonfinite axiomatizability of the equational theory of shuffle. Acta Informatica, 35(6):505– 539, June 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035006/80350505.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035006/
80350505.pdf.

Laurent:1998:EUD

[1007] D. Laurent, V. Phan Luong, and N. Spyratos. Erratum: “The use of deleted tuples in database querying and updating”. Acta Informatica, 35(6):541, June 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035006/80350541.htm;   http://

link.springer-ny.com/link/service/journals/00236/papers/8035006/
80350541.pdf;   http://link.springer-ny.com/link/service/journals/
00236/papers/8035006/80350541.ps.gz. See [987].

Sibeyn:1998:LRM

[1008] Jop F. Sibeyn. List ranking on meshes. Acta Informatica, 35(7): 543–566, July 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035007/80350543.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035007/
80350543.pdf.

Diekert:1998:AT

[1009] Volker Diekert and Paul Gastin. Approximating traces. Acta Informatica, 35(7):567–593, July 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035007/80350567.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035007/
80350567.pdf.

Leung:1998:FAL

[1010] Hing Leung. On finite automata with limited nondeterminism. Acta Informatica, 35(7):595–624, July 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/8035007/80350595.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/8035007/
80350595.pdf.

Honkala:1998:DPC

[1011] Juha Honkala. Decision problems concerning thinness and slenderness of formal languages. Acta Informatica, 35(7):625–636, July 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035007/80350625.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035007/80350625.
pdf.

VandenBussche:1998:CUF

[1012] Jan Van den Bussche and Luca Cabibbo. Converting untyped formulas to typed ones. Acta Informatica, 35(8):637–643, August 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035008/80350637.htm;   http://link.springer-

ny.com/link/service/journals/00236/papers/8035008/80350637.
pdf.

Beauquier:1998:PMC

[1013] Danièle Beauquier and Anatol Slissenko. Polytime model checking for timed probabilistic computation tree logic. Acta Informatica, 35(8): 645–664, August 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035008/80350645.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035008/
80350645.pdf.

Caporaso:1998:RBU

[1014] Salvatore Caporaso and Michele Zito. On a relation between uniform coding and problems of the form DTIMEF(F) =?DSPACEF(F). Acta Informatica, 35(8):665–672, August 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/8035008/80350665.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/8035008/
80350665.pdf.

Yen:1998:PCF

[1015] Hsu-Chun Yen. Priority conflict-free Petri nets. Acta Informatica, 35 (8):673–688, August 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035008/80350673.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035008/
80350673.pdf.

Shyr:1998:BCS

[1016] H. J. Shyr and S. S. Yu. Bi-catenation and shuffle product of languages. Acta Informatica, 35(8):689–707, August 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/8035008/80350689.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
8035008/80350689.pdf.

Fan:1998:WL

[1017] Chen-Ming Fan, H. J. Shyr, and S. S. Yu. d-words and d-languages. Acta Informatica, 35(8):709–727, August 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/8035008/80350709.htm;

http://link.springer-ny.com/link/service/journals/00236/papers/
8035008/80350709.pdf.

Sernadas:1998:DSO

[1018] Amílcar Sernadas, Cristina Sernadas, and Carlos Caleiro. Denotational semantics of object specification. Acta Informatica, 35(9):729– 773, September 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035009/80350729.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035009/
80350729.pdf.

Moffat:1998:TBM

[1019] Alistair Moffat, Ola Petersson, and Nicholas C. Wormald. A tree-based Mergesort. Acta Informatica, 35(9):775–793, September 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035009/80350775.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035009/80350775.
pdf.

Sanlaville:1998:MSA

[1020] Eric Sanlaville and Günter Schmidt. Machine scheduling with availability constraints. Acta Informatica, 35(9):795–811, September 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035009/80350795.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035009/80350795.
pdf.

Best:1998:NAH

[1021] Eike Best, Wojciech Fraçzak, Richard P. Hopkins, Hanna Klaudel, and Elisabeth Pelz. M-nets: An algebra of high-level Petri nets, with an application to the semantics of concurrent programming languages. Acta Informatica, 35(10):813–857, October 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035010/80350813.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035010/80350813.
pdf.

Larsen:1998:ACR

[1022] Kim S. Larsen. Amortized constant relaxed rebalancing using standard rotations. Acta Informatica, 35(10):859–874, October 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035010/80350859.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035010/80350859.
pdf.

Wedler:1998:LLR

[1023] Christoph Wedler and Christian Lengauer. On linear list recursion in parallel. Acta Informatica, 35(10):875–909, October 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/8035010/80350875.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/8035010/80350875.
pdf.

Hwang:1998:AEM

[1024] Hsien-Kuei Hwang. Asymptotic expansions of the mergesort recurrences. Acta Informatica, 35(11):911–919, November 1998. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/8035011/80350911.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/8035011/80350911.pdf.

Back:1998:FSE

[1025] Ralph-Johan Back and Michael Butler. Fusion and simultaneous execution in the refinement calculus. Acta Informatica, 35(11):921– 949, November 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035011/80350921.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035011/
80350921.pdf.

Bidoit:1998:MCP

[1026] Michel Bidoit and Rolf Hennicker. Modular correctness proofs of behavioural implementations. Acta Informatica, 35(11):951–1005, November 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035011/80350951.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035011/
80350951.pdf.

Bijlsma:1998:DSP

[1027] Lex Bijlsma and Rob Nederpelt. Dijkstra-Scholten predicate calculus: concepts and misconceptions. Acta Informatica, 35(12):1007– 1036, December 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035012/80351007.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035012/
80351007.pdf.

DeFrancesco:1998:TSC

[1028] Nicoletta De Francesco and Antonella Santone. A transformation system for concurrent processes. Acta Informatica, 35(12):1037– 1073, December 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035012/80351037.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035012/
80351037.pdf.

Engelfriet:1998:AGG

[1029] Joost Engelfriet and Tjalling Gelsema. Axioms for generalized graphs, illustrated by a Cantor-Bernstein proposition. Acta Informatica, 35(12): 1075–1096, December 1998. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/8035012/80351075.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/8035012/
80351075.pdf.

Schenke:1999:TDRa

[1030] Michael Schenke and Ernst-Rüdiger Olderog. Transformational design of real-time systems. Part I: From requirements to program specifications. Acta Informatica, 36(1):1–65, January 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/9036001/90360001.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
9036001/90360001.pdf.

Schenke:1999:TDRb

[1031] Michael Schenke and Ernst-Rüdiger Olderog. Transformational design of real-time systems. Part II: From program specifications to programs. Acta Informatica, 36(1):67–96, January 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/9036001/90360067.htm;

 http://link.springer-ny.com/link/service/journals/00236/papers/
9036001/90360067.pdf.

Schewe:1999:TTC

[1032] Klaus-Dieter Schewe and Bernhard Thalheim. Towards a theory of consistency enforcement. Acta Informatica, 36(2):97–141, February 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/9036002/90360097.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/9036002/90360097.
pdf.

Yen:1999:OAS

[1033] William C. K. Yen and C. Y. Tang. An optimal algorithm for solving the searchlight guarding problem on weighted two-terminal series-parallel graphs. Acta Informatica, 36(2):143–172, February 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/9036002/90360143.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/9036002/90360143.
pdf.

Vincent:1999:SFR

[1034] Millist W. Vincent. Semantic foundations of 4NF in relational database design. Acta Informatica, 36(3):173–213, March 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/9036003/90360173.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/9036003/90360173.pdf.

Han:1999:PPC

[1035] Yijie Han and Yoshihide Igarashi. Parallel PROFIT/COST algorithms through fast derandomization. Acta Informatica, 36(3):215– 232, March 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036003/90360215.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036003/
90360215.pdf.

Cerna:1999:CEN

[1036] Ivana Černá, Mojmír Křetínský, and Antonín Kučera. Comparing expressibility of normed BPA and normed BPP processes. Acta Informatica, 36(3):233–256, March 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/9036003/90360233.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/9036003/
90360233.pdf.

Proietti:1999:OAD

[1037] Guido Proietti. An optimal algorithm for decomposing a window into maximal quadtree blocks. Acta Informatica, 36(4):257– 266, July 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036004/90360257.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036004/
90360257.pdf.

Khardon:1999:REP

[1038] Roni Khardon, Heikki Mannila, and Dan Roth. Reasoning with examples: propositional formulae and database dependencies. Acta Informatica, 36(4):267–286, July 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036004/90360267.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036004/
90360267.pdf.

Fiat:1999:LSS

[1039] Amos Fiat and Gerhard J. Woeginger. On-line scheduling on a single machine: minimizing the total completion time. Acta Informatica, 36 (4):287–293, July 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036004/90360287.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036004/
90360287.pdf.

Back:1999:RAA

[1040] R. J. R. Back and J. von Wright. Reasoning algebraically about loops. Acta Informatica, 36(4):295–334, July 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/9036004/90360295.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
9036004/90360295.pdf.

Degano:1999:CDM

[1041] Pierpaolo Degano, Corrado Priami, Lone Leth, and Bent Thomsen. Causality for debugging mobile agents. Acta Informatica, 36(5):335– 374, September 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036005/90360335.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036005/
90360335.pdf.

Konikowska:1999:RFO

[1042] Beata Konikowska and Marcin Bialasik. Reasoning with first order nondeterministic specifications. Acta Informatica, 36(5):375– 403, September 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036005/90360375.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036005/
90360375.pdf.

Shen:1999:FMV

[1043] Hong Shen. Finding the k most vital edges with respect to minimum spanning tree. Acta Informatica, 36(5):405–424, September 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/9036005/90360405.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/9036005/90360405.
pdf.

Lundberg:1999:OBG

[1044] Lars Lundberg and Håkan Lennerstad. Optimal bounds on the gain of permitting dynamic allocation of communication channels in distributed computing. Acta Informatica, 36(6):425–446, October 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/9036006/90360425.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/9036006/90360425.
pdf.

Dolev:1999:MRS

[1045] Shlomi Dolev, Mohamed G. Gouda, and Marco Schneider. Memory requirements for silent stabilization. Acta Informatica, 36(6): 447–462, October 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036006/90360447.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036006/
90360447.pdf.

Abolhassan:1999:CEP

[1046] Ferri Abolhassan, Jörg Keller, and Wolfgang J. Paul. On the costeffectiveness of PRAMs. Acta Informatica, 36(6):463–487, October 1999. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/9036006/90360463.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/9036006/90360463.
pdf.

Ravelo:1999:TGA

[1047] Jesús N. Ravelo. Two graph algorithms derived. Acta Informatica, 36 (6):489–510, October 1999. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/9036006/90360489.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/9036006/
90360489.pdf.

Bonner:2000:QSD

[1048] Anthony J. Bonner and Giansalvatore Mecca. Querying sequence databases with transducers. Acta Informatica, 36(7):511–544, January 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/0036007/00360511.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/0036007/00360511.
pdf.

Schmidt:2000:HCS

[1049] Karsten Schmidt. How to calculate symmetries of Petri nets. Acta Informatica, 36(7):545–590, January 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/0036007/00360545.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/0036007/
00360545.pdf.

Ehrich:2000:SCD

[1050] Hans-Dieter Ehrich and Carlos Caleiro. Specifying communication in distributed information systems. Acta Informatica, 36(8):591– 616, March 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0036008/00360591.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0036008/
00360591.pdf.

Leavens:2000:CAC

[1051] Gary T. Leavens and Don Pigozzi. A complete algebraic characterization of behavioral subtyping. Acta Informatica, 36(8):617– 663, March 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0036008/00360617.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0036008/
00360617.pdf.

Ilie:2000:ESS

[1052] Lucian Ilie and Arto Salomaa. On the expressiveness of subset-sum representations. Acta Informatica, 36(8):665–672, March 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/0036008/
00360665.htm;   http://link.springer-ny.com/link/service/journals/
00236/papers/0036008/00360665.pdf.

Cheng:2000:SMS

[1053] T. C. E. Cheng and Q. Ding. Single machine scheduling with deadlines and increasing rates of processing times. Acta Informatica, 36 (9/10):673–692, April 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0036009/00360673.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0036009/
00360673.pdf.

Kinoshita:2000:DRA

[1054] Yoshiki Kinoshita and John Power. Data refinement and algebraic structure. Acta Informatica, 36(9/10):693–719, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/0036009/00360693.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/0036009/00360693.pdf.

Forst:2000:MHT

[1055] Gunnar Forst and Anders Thorup. Minimal Huffman trees. Acta Informatica, 36(9/10):721–734, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/0036009/00360721.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/0036009/
00360721.pdf.

Mahmoud:2000:AVB

[1056] Hosam Mahmoud, Philippe Flajolet, Philippe Jacquet, and Mireille Régnier. Analytic variations on bucket selection and sorting. Acta Informatica, 36(9/10):735–760, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/0036009/00360735.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/0036009/
00360735.pdf.

Gorlatch:2000:APD

[1057] Sergei Gorlatch and Christian Lengauer. Abstraction and performance in the design of parallel programs: an overview of the SAT approach. Acta Informatica, 36(9/10):761–803, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/0036009/00360761.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
0036009/00360761.pdf.

Honkala:2000:SLB

[1058] Juha Honkala. On slender 0L languages over the binary alphabet. Acta Informatica, 36(9/10):805–815, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/0036009/00360805.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
0036009/00360805.pdf.

Intrigila:2000:GHK

[1059] Benedetto Intrigila and Stefano Varricchio. On the generalization of Higman and Kruskal’s theorems to regular languages and rational trees. Acta Informatica, 36(9/10):817–835, April 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/0036009/00360817.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
0036009/00360817.pdf.

Kesten:2000:VCH

[1060] Yonit Kesten, Zohar Manna, and Amir Pnueli. Verification of clocked and hybrid systems. Acta Informatica, 36(11):837–912, May 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/0036011/00360837.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/0036011/00360837.
pdf.

Csuhaj-Varju:2000:ESL

[1061] Erzsébet Csuhaj-Varjú and Victor Mitrana. Evolutionary systems: a language generating device inspired by evolving communities of cells. Acta Informatica, 36(11):913–926, May 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/0036011/00360913.htm;
http://link.springer-ny.com/link/service/journals/00236/papers/
0036011/00360913.pdf.

Tip:2000:CHS

[1062] Frank Tip and Peter F. Sweeney. Class hierarchy specialization. Acta Informatica, 36(12):927–982, July 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/0036012/00360927.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/0036012/
00360927.pdf.

Carpi:2000:SFP

[1063] Arturo Carpi and Aldo de Luca. Special factors, periodicity, and an application to Sturmian words. Acta Informatica, 36(12):983– 1006, July 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0036012/00360983.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0036012/
00360983.pdf.

Middeldorp:2000:TIE

[1064] Aart Middeldorp and Hitoshi Ohsaki. Type introduction for equational rewriting. Acta Informatica, 36(12):1007–1029, July 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/0036012/
00361007.htm;   http://link.springer-ny.com/link/service/journals/
00236/papers/0036012/00361007.pdf.

Anonymous:2000:AR

[1065] Anonymous. Acknowledgement to referees. Acta Informatica, 36(12): 1031–1032, July 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0036012/00361031.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0036012/
00361031.pdf.

Kobuchi:2000:SAT

[1066] Youichi Kobuchi, Takashi Saito, and Hidenobu Nunome. Semantics analysis through elementary meanings. Acta Informatica, 37 (1):1–19, August 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0037001/00370001.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0037001/
00370001.pdf.

Ranjan:2000:DSO

[1067] Desh Ranjan, Enrico Pontelli, and Gopal Gupta. Data structures for order-sensitive predicates in parallel nondeterministic systems. Acta Informatica, 37(1):21–43, August 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/0037001/00370021.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/0037001/
00370021.pdf.

Berard:2000:AZW

[1068] Béatrice Bérard and Claudine Picaronny. Accepting Zeno words: a way toward timed refinements. Acta Informatica, 37(1):45– 81, August 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0037001/00370045.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0037001/
00370045.pdf.

Ben-Amram:2000:CCP

[1069] Amir M. Ben-Amram and Neil D. Jones. Computational complexity via programming languages: constant factors do matter. Acta Informatica, 37(2):83–120, September 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/0037002/00370083.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/0037002/
00370083.pdf.

Dube:2000:EBP

[1070] Danny Dubé and Marc Feeley. Efficiently building a parse tree from a regular expression. Acta Informatica, 37(2):121–144, September 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/0037002/00370121.htm;   http://link.springer-

ny.com/link/service/journals/00236/papers/0037002/00370121.
pdf.

Andrei:2000:SRC

[1071] Stefan Andrei, Manfred Kudlek, and Radu Stefan Niculescu. Some results on the Collatz problem. Acta Informatica, 37(2):145–160, September 2000. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/0037002/00370145.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/0037002/
00370145.pdf.

Chaudhuri:2000:OWA

[1072] Soma Chaudhuri, Martha J. Kosa, and Jennifer L. Welch. One-write algorithms for multivalued regular and atomic registers. Acta Informatica, 37(3):161–192, November 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Yang:2000:CSM

[1073] Boting Yang and Paul Gillard. The class Steiner minimal tree problem: a lower bound and test problem generation. Acta Informatica, 37(3): 193–211, November 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hoogeboom:2000:FSL

[1074] Hendrik Jan Hoogeboom and Nikè van Vugt. Fair sticker languages. Acta Informatica, 37(3):213–225, November 2000. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Broy:2001:ELE

[1075] Manfred Broy. Editorial: Letter from the Editor. Acta Informatica, 37 (4/5):227–228, January 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

vanGlabbeek:2001:RAE

[1076] Rob van Glabbeek and Ursula Goltz. Refinement of actions and equivalence notions for concurrent systems. Acta Informatica, 37(4/5):229–327, January 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

McIver:2001:DAU

[1077] A. K. McIver and C. Morgan. Demonic, angelic and unbounded probabilistic choices in sequential programs. Acta Informatica, 37(4/5):329– 354, January 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chanzy:2001:ARS

[1078] Philippe Chanzy, Luc Devroye, and Carlos Zamora-Cura. Analysis of range search for random k-d trees. Acta Informatica, 37(4/5):355–383, January 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hayes:2001:SRT

[1079] Ian J. Hayes and Mark Utting. A sequential real-time refinement calculus. Acta Informatica, 37(6):385–448, February 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Tiplea:2001:SES

[1080] Ferucio Laurentiu Tiplea, Erkki Mäkinen, and Corina Apachite. Synchronized extension systems. Acta Informatica, 37(6):449–465, February 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Corradini:2001:CIP

[1081] Flavio Corradini and Marco Pistore. ‘Closed Interval Process Algebra’ versus ‘Interval Process Algebra’. Acta Informatica, 37(7):467–509, March 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fernau:2001:PCG

[1082] Henning Fernau. Parallel communicating grammar systems with terminal transmission. Acta Informatica, 37(7):511–540, March 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Morris:2001:TB

[1083] Joseph M. Morris and Alexander Bunkenburg. A theory of bunches. Acta Informatica, 37(8):541–561, May 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/1037008/10370541.htm;
http://link.springer-ny.com/link/service/journals/00236/papers/
1037008/10370541.pdf.

Moreau:2001:CDR

[1084] Luc Moreau and Jean Duprat. A construction of distributed reference counting. Acta Informatica, 37(8):563–595, May 2001. CODEN AINFA2.

ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/1037008/10370563.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/1037008/10370563.pdf.

Carpi:2001:PLW

[1085] Arturo Carpi and Aldo de Luca. Periodic-like words, periodicity, and boxes. Acta Informatica, 37(8):597–618, May 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer-ny.com/link/service/journals/00236/bibs/1037008/10370597.
htm;   http://link.springer-ny.com/link/service/journals/00236/
papers/1037008/10370597.pdf.

Kim:2001:ERA

[1086] Changwook Kim. Efficient recognition algorithms for boundary and linear eNCE graph languages. Acta Informatica, 37(9):619– 632, June 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1037009/10370619.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1037009/
10370619.pdf.

Aycock:2001:EFG

[1087] John Aycock, Nigel Horspool, Jan Janousek, and Borivoj Melichar. Even faster generalized LR parsing. Acta Informatica, 37(9):633– 651, June 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1037009/10370633.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1037009/
10370633.pdf.

Alonso:2001:TIP

[1088] Laurent Alonso and René Schott. On the tree inclusion problem. Acta Informatica, 37(9):653–670, June 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/1037009/10370653.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/1037009/
10370653.pdf.

Morimoto:2001:YAG

[1089] Shin ichi Morimoto and Masataka Sassa. Yet another generation of LALR parsers for regular right part grammars. Acta Informatica, 37 (9):671–697, June 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1037009/10370671.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1037009/
10370671.pdf.

Greco:2001:ESD

[1090] Sergio Greco, Domenico Saccà, and Carlo Zaniolo. Extending stratified datalog to capture complexity classes ranging from P to QH. Acta Informatica, 37(10):699–725, July 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/1037010/10370699.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/1037010/
10370699.pdf.

vanStee:2001:RJC

[1091] Rob van Stee and Han La Poutré. Running a job on a collection of partly available machines, with on-line restarts. Acta Informatica, 37 (10):727–742, July 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1037010/10370727.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1037010/
10370727.pdf.

Larsen:2001:RBS

[1092] Kim S. Larsen, Thomas Ottmann, and Eljas Soisalon-Soininen. Relaxed balance for search trees with local rebalancing. Acta Informatica, 37 (10):743–763, July 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1037010/10370743.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1037010/
10370743.pdf.

Ramon:2001:PTC

[1093] Jan Ramon and Maurice Bruynooghe. A polynomial time computable metric between point sets. Acta Informatica, 37(10):765– 780, July 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1037010/10370765.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1037010/
10370765.pdf.

Best:2001:RPN

[1094] Eike Best, Raymond Devillers, and Maciej Koutny. Recursion and Petri nets. Acta Informatica, 37(11-12):781–829, August 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer-ny.com/link/service/journals/00236/bibs/1037011/
10370781.htm;   http://link.springer-ny.com/link/service/journals/
00236/papers/1037011/10370781.pdf.

Astesiano:2001:LTL

[1095] E. Astesiano and G. Reggio. Labelled transition logic: an outline. Acta Informatica, 37(11-12):831–879, August 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/1037011/10370831.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
1037011/10370831.pdf.

Chakka:2001:MMM

[1096] Ram Chakka and Peter G. Harrison. A Markov modulated multi-server queue with negative customers — the MM CPP/GE/c/L G-queue. Acta Informatica, 37(11-12):881–919, August 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/1037011/10370881.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
1037011/10370881.pdf.

Colomb:2001:CTF

[1097] Robert M. Colomb, C. N. G. Dampney, and Michael Johnson. Categorytheoretic fibration as an abstraction mechanism in information systems. Acta Informatica, 38(1):1–44, September 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-

ny.com/link/service/journals/00236/bibs/1038001/10380001.htm;
 http://link.springer-ny.com/link/service/journals/00236/papers/
1038001/10380001.pdf.

Meinke:2001:CDS

[1098] K. Meinke and L. J. Steggles. Correctness of dataflow and systolic algorithms using algebras of streams. Acta Informatica, 38(1): 45–88, September 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1038001/10380045.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1038001/
10380045.pdf.

LaTorre:2001:TTA

[1099] Salvatore La Torre and Margherita Napoli. Timed tree automata with an application to temporal logic. Acta Informatica, 38(2):89– 116, November 2001. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer-ny.com/link/
service/journals/00236/bibs/1038002/10380089.htm;   http://
link.springer-ny.com/link/service/journals/00236/papers/1038002/
10380089.pdf.

Ito:2001:GWF

[1100] Masami Ito, Carlos Martín-Vide, and Victor Mitrana. Group weighted finite transducers. Acta Informatica, 38(2):117–129, November 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/1038002/10380117.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/1038002/10380117.
pdf.

Dassow:2001:TSM

[1101] Jürgen Dassow, Gheorghe Paun, Gabriel Thierrin, and Sheng Yu. Tree-systems of morphisms. Acta Informatica, 38(2):131–153, November 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/1038002/10380131.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/1038002/10380131.
pdf.

Rensink:2001:PAA

[1102] Arend Rensink and Heike Wehrheim. Process algebra with action dependencies. Acta Informatica, 38(3):155–234, December 2001. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/1038003/10380155.htm;   http://link.springer-
ny.com/link/service/journals/00236/papers/1038003/10380155.
pdf.

Makinen:2002:MAT

[1103] Erkki Mäkinen and Tarja Systä. Minimally adequate teacher synthesizes statechart diagrams. Acta Informatica, 38(4):235–259, January 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL http://link.springer-ny.com/link/service/
journals/00236/bibs/2038004/20380235.htm;   http://link.springer-

ny.com/link/service/journals/00236/papers/2038004/20380235.
pdf.

Drmota:2002:VHD

[1104] Michael Drmota. The variance of the height of digital search trees. Acta Informatica, 38(4):261–276, January 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.

com/link/service/journals/00236/bibs/2038004/20380261.htm;
http://link.springer-ny.com/link/service/journals/00236/papers/
2038004/20380261.pdf.

Lin:2002:ATA

[1105] Huimin Lin and Wang Yi. Axiomatising timed automata. Acta Informatica, 38(4):277–305, January 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer-ny.com/

link/service/journals/00236/bibs/2038004/20380277.htm;   http:
//link.springer-ny.com/link/service/journals/00236/papers/2038004/
20380277.pdf.

Simovici:2002:IMD

[1106] Dan A. Simovici, Dana Cristofor, and Laurentiu Cristofor. Impurity measures in databases. Acta Informatica, 38(5):307–324, March 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Liu:2002:CDP

[1107] Jixue Liu and Millist Vincent. Containment and disjointedness in partitioned normal form relations. Acta Informatica, 38(5):325–342, March 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hesselink:2002:ACA

[1108] Wim H. Hesselink. An assertional criterion for atomicity. Acta Informatica, 38(5):343–366, March 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Duggan:2002:OTCb

[1109] Dominic Duggan. Object type constructors. Acta Informatica, 38(6): 367–408, April 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Carpi:2002:WUF

[1110] Arturo Carpi, Aldo de Luca, and Stefano Varricchio. Words, univalent factors, and boxes. Acta Informatica, 38(6):409–436, April 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Frederiks:2002:DPI

[1111] P. J. M. Frederiks and T. P. van der Weide. Deriving and paraphrasing information grammars using object-oriented analysis models. Acta Informatica, 38(7):437–488, June 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Penabad:2002:GPC

[1112] Miguel R. Penabad, Nieves R. Brisaboa, Héctor J. Hernández, et al. A general procedure to check conjunctive query containment. Acta Informatica, 38(7):489–529, June 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Santone:2002:AVC

[1113] Antonella Santone. Automatic verification of concurrent systems using a formula-based compositional approach. Acta Informatica, 38(8):531– 564, July 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Larsen:2002:RRB

[1114] Kim S. Larsen. Relaxed red-black trees with group updates. Acta Informatica, 38(8):565–586, July 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chakaravarthy:2002:NAP

[1115] Venkatesan T. Chakaravarthy and Susan Horwitz. On the nonapproximability of points-to analysis. Acta Informatica, 38(8):587–598, July 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Cordero:2002:BCS

[1116] P. Cordero, M. Enciso, and I. P. de Guzmán. Bases for closed sets of implicants and implicates in temporal logic. Acta Informatica, 38(9):599– 619, August 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Giannella:2002:APC

[1117] Chris Giannella and Dirk Van Gucht. Adding a path connectedness operator to FO + poly (linear). Acta Informatica, 38(9):621–648, August 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Berstel:2002:FPX

[1118] Jean Berstel and Luc Boasson. Formal properties of XML grammars and languages. Acta Informatica, 38(9):649–671, August 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coffman:2002:PRS

[1119] E. G. Coffman, Jr., Peter J. Downey, and Peter Winkler. Packing rectangles in a strip. Acta Informatica, 38(10):673–693, September 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bottoni:2002:MSP

[1120] Paolo Bottoni, Carlos Martín-Vide, Gheorghe Păun, et al. Membrane systems with promoters/inhibitors. Acta Informatica, 38(10):695–720, September 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Madhu:2002:GNF

[1121] Mutyam Madhu and Kamala Krithivasan. Generalized normal form for rewriting P systems. Acta Informatica, 38(10):721–734, September 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Corradini:2002:CWC

[1122] Flavio Corradini, Walter Vogler, and Lars Jenner. Comparing the worstcase efficiency of asynchronous systems with PAFAS. Acta Informatica, 38(11/12):735–792, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Vajnovszki:2002:GVM

[1123] Vincent Vajnovszki. Gray visiting Motzkins. Acta Informatica, 38(11/ 12):793–811, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mahmoud:2002:SRB

[1124] Hosam M. Mahmoud. The size of random bucket trees via urn models. Acta Informatica, 38(11/12):813–838, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). See erratum [1179].

Honkala:2002:CIP

[1125] Iiro Honkala and Antoine Lobstein. On the complexity of the identification problem in Hamming spaces. Acta Informatica, 38(11/12):839–845, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Iwamoto:2002:QST

[1126] Chuzo Iwamoto, Katsuyuki Tateishi, Kenichi Morita, et al. A quadratic speedup theorem for iterative arrays. Acta Informatica, 38(11/12):847– 858, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anonymous:2002:AR

[1127] Anonymous. Acknowledgement to referees. Acta Informatica, 38(11/12): 859–860, November 2002. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coulondre:2003:TPP

[1128] Stéphane Coulondre. A top-down proof procedure for generalized data dependencies. Acta Informatica, 39(1):1–29, January 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Penna:2003:SRE

[1129] Giuseppe Della Penna, Benedetto Intrigila, Enrico Tronci, et al. Synchronized regular expressions. Acta Informatica, 39(1):31–70, January 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Burrieza:2003:FAT

[1130] Alfredo Burrieza and Inma P. de Guzmán. A functional approach for temporal × modal logics. Acta Informatica, 39(2):71–96, February 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Epstein:2003:SR

[1131] Leah Epstein. Bin stretching revisited. Acta Informatica, 39(2):97–117, February 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gawo:2003:SCD

[1132] Ján Gawo and Martin Nehéz. Stochastic cooperative distributed grammar systems and random graphs. Acta Informatica, 39(2):119–140, February 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Bauer:2003:EWD

[1133] Friedrich L. Bauer and Manfred Broy. Edsger W. Dijkstra — Acta Informatica and Marktoberdorf. Acta Informatica, 39(3):141–142, March 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kiepuszewski:2003:FCF

[1134] B. Kiepuszewski, A. H. M. ter Hofstede, and W. M. P. van der Aalst. Fundamentals of control flow in workflows. Acta Informatica, 39(3):143– 209, March 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Hesselink:2003:PRF

[1135] Wim H. Hesselink. Preference rankings in the face of uncertainty. Acta Informatica, 39(3):211–231, March 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Wabenhorst:2003:SDF

[1136] Axel Wabenhorst. Stepwise development of fair distributed systems. Acta Informatica, 39(4):233–271, April 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Reisig:2003:GTS

[1137] W. Reisig. On Gurevich’s theorem on sequential algorithms. Acta Informatica, 39(4):273–305, April 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Meduna:2003:CES

[1138] A. Meduna. Coincidental extension of scattered context languages. Acta Informatica, 39(5):307–314, May 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ying:2003:RAP

[1139] M. Ying. Reasoning about probabilistic sequential programs in a probabilistic logic. Acta Informatica, 39(5):315–389, May 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Biskup:2003:AID

[1140] Joachim Biskup and Torsten Polle. Adding inclusion dependencies to an object-oriented data model with uniqueness constraints. Acta Informatica, 39(6/7):391–449, June 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Currie:2003:WLW

[1141] James D. Currie and Erica Moodie. A word on 7 letters which is nonrepetitive up to mod5. Acta Informatica, 39(6/7):451–468, June 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Srba:2003:SBS

[1142] Jiří Srba. Strong bisimilarity of simple process algebras: complexity lower bounds. Acta Informatica, 39(6/7):469–499, June 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Fokkink:2003:SOS

[1143] Wan Fokkink and Thuy Duong Vu. Structural operational semantics and bounded nondeterminism. Acta Informatica, 39(6/7):501–516, June 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Castellanos:2003:NEP

[1144] Juan Castellanos, Carlos Martín-Vide, Victor Mitrana, et al. Networks of evolutionary processors. Acta Informatica, 39(6/7):517–529, June 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Majster-Cederbaum:2003:TAR

[1145] Mila Majster-Cederbaum and Jinzhao Wu. Towards action refinement for true concurrent real time. Acta Informatica, 39(8):531–577, July 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Silva:2003:NPP

[1146] Pedro V. Silva. A note on pure and p-pure languages. Acta Informatica, 39(8):579–595, July 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Coffman:2003:IPS

[1147] E. G. Coffman, J. Sethuraman, and V. G. Timkovsky. Ideal preemptive schedules on two processors. Acta Informatica, 39(8):597–612, July 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Engelfriet:2003:CPT

[1148] Joost Engelfriet and Sebastian Maneth. A comparison of pebble tree transducers with macro tree transducers. Acta Informatica, 39(9):613– 698, August 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Meduna:2003:ECE

[1149] Alexander Meduna. Erratum: Coincidental extension of scattered context languages. Acta Informatica, 39(9):699, August 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anonymous:2003:AR

[1150] Anonymous. Acknowledgement to referees. Acta Informatica, 39(9): 701, August 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anonymous:2003:MPP

[1151] Anonymous. Message from the Publisher: Publication in the Online First directory. Acta Informatica, 40(1):1, September 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Boyar:2003:EAF

[1152] Joan Boyar, Lene M. Favrholdt, Kim S. Larsen, et al. Extending the accommodating function. Acta Informatica, 40(1):3–35, September 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Doberkat:2003:PMS

[1153] Ernst-Erich Doberkat. Pipelines: Modelling a software architecture through relations. Acta Informatica, 40(1):37–79, September 2003. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Anonymous:2003:MP

[1154] Anonymous. Message from the Publisher. Acta Informatica, 40(1):??, September 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ben-Amram:2003:EDO

[1155] Amir M. Ben-Amram, Omer Berkman, and Holger Petersen. Element distinctness on one-tape Turing machines: a complete solution. Acta Informatica, 40(2):81–94, October 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Khomenko:2003:CPP

[1156] Victor Khomenko, Maciej Koutny, and Walter Vogler. Canonical prefixes of Petri net unfoldings. Acta Informatica, 40(2):95–118, October 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Kari:2003:SFO

[1157] Lila Kari, Stavros Konstantinidis, Elena Losseva, et al. Sticky-free and overhang-free DNA languages. Acta Informatica, 40(2):119–157, October 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lesley:2003:PVS

[1158] N. Lesley and A. Fekete. Providing view synchrony for group communication services. Acta Informatica, 40(3):159–210, November 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Layuan:2003:DQA

[1159] Li Layuan and Li Chunlin. A distributed QoS-aware multicast routing protocol. Acta Informatica, 40(3):211–233, November 2003. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Walther:2004:PRE

[1160] Andrea Walther. Program reversals for evolutions with non-uniform step costs. Acta Informatica, 40(4):235–263, February 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Charpentier:2004:STP

[1161] Michel Charpentier and K. Mani Chandy. Specification transformers: a predicate transformer approach to composition. Acta Informatica, 40 (4):265–301, February 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Lee:2004:BCS

[1162] Wen-Chiung Lee, Chin-Chia Wu, and Hua-Jung Sung. A bi-criterion single-machine scheduling problem with learning considerations. Acta Informatica, 40(4):303–315, February 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Barbuti:2004:TAU

[1163] Roberto Barbuti and Luca Tesei. Timed automata with urgent transitions. Acta Informatica, 40(5):317–347, March 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=40&issue=5&spage=317.

Andrei:2004:SEC

[1164] Stefan Andrei, Wei-Ngan Chin, and Salvador Valerio Cavadini. Selfembedded context-free grammars with regular counterparts. Acta Informatica, 40(5):349–365, March 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=40&issue=5&
spage=349.

He:2004:OAS

[1165] Yong He and Yiwei Jiang. Optimal algorithms for semi-online preemptive scheduling problems on two uniform machines. Acta Informatica, 40(5): 367–383, March 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=40&issue=5&spage=367.

Engelfriet:2004:NNS

[1166] Joost Engelfriet and Tjalling Gelsema. A new natural structural congruence in the pi-calculus with replication. Acta Informatica, 40(6–7): 385–430, May 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=40&issue=6&spage=385.

Markey:2004:PFC

[1167] Nicolas Markey. Past is for free: on the complexity of verifying linear temporal properties with past. Acta Informatica, 40(6–7):431– 458, May 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=40&issue=6&spage=431.

Scott:2004:RND

[1168] Elizabeth Scott and Adrian Johnstone. Reducing non-determinism in right nulled GLR parsers. Acta Informatica, 40(6–7):459–489, May 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=40&issue=6&spage=459.

Domaratzki:2004:TBC

[1169] Michael Domaratzki. Trajectory-based codes. Acta Informatica, 40(6–7): 491–527, May 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=40&issue=6&spage=491.

Grumbach:2004:ERA

[1170] Stéphane Grumbach, Maurizio Rafanelli, and Leonardo Tininini. On the equivalence and rewriting of aggregate queries. Acta Informatica, 40(8): 529–584, July 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=40&issue=8&spage=529.

Bacchelli:2004:EGC

[1171] Silvia Bacchelli, Elena Barcucci, Elisabetta Grazzini, and Elisa Pergola. Exhaustive generation of combinatorial objects by ECO. Acta Informatica, 40(8):585–602, July 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=40&issue=8&
spage=585.

Akkerman:2004:CDT

[1172] Thorsten Akkerman, Christoph Buchheim, Michael Jünger, and Daniel Teske. On the complexity of drawing trees nicely: corrigendum. Acta Informatica, 40(8):603–607, July 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=40&issue=8&

spage=603. See [435].

Dolev:2004:SSG

[1173] Shlomi Dolev and Elad Schiller. Self-stabilizing group communication in directed networks. Acta Informatica, 40(9):609–636, September 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=40&issue=9&spage=609.

Delvaux:2004:BTA

[1174] Steven Delvaux and Leon Horsten. On best transitive approximations to simple graphs. Acta Informatica, 40(9):637–655, September 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=40&issue=9&spage=637.

Epstein:2004:ASM

[1175] Leah Epstein and Tamir Tassa. Approximation schemes for the Min-Max Starting Time Problem. Acta Informatica, 40(9):657–674, September 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=40&issue=9&spage=657.

Anonymous:2004:AR

[1176] Anonymous. Acknowledgement to referees. Acta Informatica, 40(9): 675, September 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=40&issue=9&spage=675.

Mahmoud:2004:RSI

[1177] Hosam M. Mahmoud. Random sprouts as Internet models, and Pólya processes. Acta Informatica, 41(1):1–18, November 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=1&spage=1.

Bozapalidis:2004:AG

[1178] Symeon Bozapalidis and Antonios Kalampakas. An axiomatization of graphs. Acta Informatica, 41(1):19–61, November 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=1&spage=19.

Mahmoud:2004:ESR

[1179] Hosam M. Mahmoud. Erratum: The size of random bucket trees via urn models. Acta Informatica, 41(1):63, November 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=1&spage=63. See [1124].

Kim:2004:MMT

[1180] SungSuk Kim, Sun Ok Yang, and SangKeun Lee. Maintaining mobile transactional consistency in hybrid broadcast environments. Acta Informatica, 41(2–3):65–81, December 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=2&spage=65.

Grigoriev:2004:PSI

[1181] Alexander Grigoriev and Gerhard J. Woeginger. Project scheduling with irregular costs: complexity, approximability, and algorithms. Acta Informatica, 41(2–3):83–97, December 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=2&spage=83.

Mahmoud:2004:LLT

[1182] Hosam Mahmoud and Tatsuie Tsukiji. Limit laws for terminal nodes in random circuits with restricted fan-out: a family of graphs generalizing binary search trees. Acta Informatica, 41(2–3):99–110, December 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=41&issue=2&spage=99.

Alhazov:2004:TPL

[1183] Artiom Alhazov, Linqiang Pan, and Gheorghe Păun. Trading polarizations for labels in P systems with active membranes. Acta Informatica, 41(2–3):111–144, December 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=41&issue=2&
spage=111.

Frisco:2004:SSA

[1184] Pierluigi Frisco and Hendrik Jan Hoogeboom. P systems with symport/ antiport simulating counter automata. Acta Informatica, 41(2–3):145– 170, December 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=41&issue=2&spage=145.

Li:2004:TEC

[1185] Zheng-Zhu Li and Y. S. Tsai. Three-element codes with one d-primitive word. Acta Informatica, 41(2–3):171–180, December 2004. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=2&spage=171.

Duggan:2005:TBH

[1186] Dominic Duggan. Type-based hot swapping of running modules. Acta Informatica, 41(4–5):181–220, March 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=4&spage=181.

Lamoureux:2005:DSL

[1187] Michael G. Lamoureux and Bradford G. Nickerson. A deterministic skip list for k-dimensional range search. Acta Informatica, 41(4–5):221– 255, March 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=41&issue=4&spage=221.

Csuhaj-Varju:2005:HNE

[1188] Erzsébet Csuhaj-Varjú, Carlos Martín-Vide, and Victor Mitrana. Hybrid networks of evolutionary processors are computationally complete. Acta Informatica, 41(4–5):257–272, March 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=4&spage=257.

Brodal:2005:FAD

[1189] Gerth Stølting Brodal, Erik D. Demaine, and J. Ian Munro. Fast allocation and deallocation with an improved buddy system. Acta Informatica, 41(4–5):273–291, March 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=41&issue=4&
spage=273.

Morvan:2005:FAC

[1190] Christophe Morvan and Chloé Rispal. Families of automata characterizing context-sensitive languages. Acta Informatica, 41(4–5):293– 314, March 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=41&issue=4&spage=293.

Jiang:2005:POA

[1191] Yiwei Jiang and Yong He. Preemptive online algorithms for scheduling with machine cost. Acta Informatica, 41(6):315–340, May 2005. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=41&issue=6&spage=315.

Klonowska:2005:ORS

[1192] Kamilla Klonowska, Håkan Lennerstad, Lars Lundberg, and Charlie Svahnberg. Optimal recovery schemes in fault tolerant distributed computing. Acta Informatica, 41(6):341–365, May 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=6&spage=341.

Vansummeren:2005:CDT

[1193] Stijn Vansummeren. On the complexity of deciding typability in the relational algebra. Acta Informatica, 41(6):367–381, May 2005. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=41&issue=6&spage=367.

Chen:2005:WS

[1194] Yifeng Chen and J. W. Sanders. The weakest specifunction. Acta Informatica, 41(7–8):383–414, June 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=41&issue=7&
spage=383.

Kucera:2005:SPR

[1195] Antonín Kucera and Jan Strejcek. The stuttering principle revisited. Acta Informatica, 41(7–8):415–434, June 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=7&spage=415.

Zuliani:2005:CQP

[1196] Paolo Zuliani. Compiling quantum programs. Acta Informatica, 41(7–8): 435–474, June 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=41&issue=7&spage=435.

Schmitt:2005:CDS

[1197] Ingo Schmitt and Gunter Saake. A comprehensive database schema integration method based on the theory of formal concepts. Acta Informatica, 41(7–8):475–524, June 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=41&issue=7&
spage=475.

Ying:2005:CNC

[1198] Mingsheng Ying. π-calculus with noisy channels. Acta Informatica, 41(9):525–593, October 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=41&issue=9&
spage=525.

Epstein:2005:OSC

[1199] Leah Epstein and Rob van Stee. Online square and cube packing. Acta Informatica, 41(9):595–606, October 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=41&issue=9&spage=595.

Anonymous:2005:AR

[1200] Anonymous. Acknowledgement to referees. Acta Informatica, 41(9): 607, October 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=41&issue=9&spage=607.

Cieslik:2005:LCC

[1201] Iwona Cieślik. On-line coloring and cliques covering for K_{s,t}-free graphs. Acta Informatica, 42(1):1–20, October 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=1&spage=1.

Buttner:2005:EPC

[1202] Markus Büttner. Enhanced prefetching and caching strategies for single- and multi-disk systems. Acta Informatica, 42(1):21–42, October 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=42&issue=1&spage=21.

Geerts:2005:DVD

[1203] Floris Geerts, Lieven Smits, and Jan Van den Bussche. N-dimensional versus (N − 1)-dimensional connectivity testing of first-order queries to semi-algebraic sets. Acta Informatica, 42(1):43–56, October 2005. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=42&issue=1&spage=43.

Jacobsen:2005:EDN

[1204] Lars Jacobsen and Kim Skak Larsen. Exponentially decreasing number of operations in balanced trees. Acta Informatica, 42(1):57–78, October 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=42&issue=1&spage=57.

DeNicola:2005:TC

[1205] Rocco De Nicola and Davide Sangiorgi. Types in concurrency. Acta Informatica, 42(2–3):79–81, November 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=42&issue=
2&spage=79.

Berger:2005:GC

[1206] Martin Berger, Kohei Honda, and Nobuko Yoshida. Genericity and the π-calculus. Acta Informatica, 42(2–3):83–141, November 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=42&issue=2&spage=83.

Bettini:2005:MCM

[1207] Lorenzo Bettini, Betti Venneri, and Viviana Bono. MOMI: a calculus for mobile mixins. Acta Informatica, 42(2–3):143–190, November 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=42&issue=2&spage=143.

Gay:2005:SST

[1208] Simon Gay and Malcolm Hole. Subtyping for session types in the pi calculus. Acta Informatica, 42(2–3):191–225, November 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=2&spage=191.

Hennessy:2005:SLC

[1209] Matthew Hennessy, Julian Rathke, and Nobuko Yoshida. SAFEDPI: a language for controlling mobile code. Acta Informatica, 42(4–5):227– 290, December 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=42&issue=4&spage=227.

Kobayashi:2005:TBI

[1210] Naoki Kobayashi. Type-based information flow analysis for the πcalculus. Acta Informatica, 42(4–5):291–347, December 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=4&spage=291.

Konig:2005:GFT

[1211] Barbara König. A general framework for types in graph rewriting. Acta Informatica, 42(4–5):349–388, December 2005. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=4&spage=349.

Majster-Cederbaum:2006:RAR

[1212] Mila Majster-Cederbaum, Jinzhao Wu, and Houguang Yue. Refinement of actions for real-time concurrent systems with causal ambiguity. Acta Informatica, 42(6–7):389–418, March 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=6&spage=389.

Ehrenfeucht:2006:ELO

[1213] Andrzej Ehrenfeucht, Tero Harju, and Grzegorz Rozenberg. Embedding linear orders in grids. Acta Informatica, 42(6–7):419–428, March 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=42&issue=6&spage=419.

Levi:2006:TEB

[1214] Francesca Levi. A typed encoding of boxed into safe ambients. Acta Informatica, 42(6–7):429–500, March 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=6&spage=429.

Epstein:2006:VAS

[1215] Leah Epstein and Tamir Tassa. Vector assignment schemes for asymmetric settings. Acta Informatica, 42(6–7):501–514, March 2006. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=42&issue=6&spage=501.

Trombetta:2006:EOE

[1216] Alberto Trombetta and Danilo Montesi. Equivalences and optimizations in an expressive XSLT subset. Acta Informatica, 42(6–7):515– 539, March 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=42&issue=6&spage=515.

Meduna:2006:DPA

[1217] Alexander Meduna. Deep pushdown automata. Acta Informatica, 42(8– 9):541–552, April 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=42&issue=8&spage=541.

Bozapalidis:2006:RGP

[1218] Symeon Bozapalidis and Antonios Kalampakas. Recognizability of graph and pattern languages. Acta Informatica, 42(8–9):553–581, April 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=42&issue=8&spage=553.

Hesselink:2006:SFS

[1219] Wim H. Hesselink. Splitting forward simulations to cope with liveness. Acta Informatica, 42(8–9):583–602, April 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=8&spage=583.

Brlek:2006:NUR

[1220] Srecko Brlek, Elisa Pergola, and Olivier Roques. Non uniform random generation of generalized Motzkin paths. Acta Informatica, 42(8–9): 603–616, April 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=42&issue=8&spage=603.

Tatti:2006:SPB

[1221] Nikolaj Tatti. Safe projections of binary data sets. Acta Informatica, 42(8–9):617–638, April 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=42&issue=8&
spage=617.

Tiplea:2006:ADT

[1222] Ferucio Laurentiu Tiplea and Constantin Enea. Abstractions of data types. Acta Informatica, 42(8–9):639–671, April 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=42&issue=8&spage=639.

Indermark:2006:ACP

[1223] Klaus Indermark and Thomas Noll. Algebraic correctness proofs for compiling recursive function definitions with strictness information. Acta Informatica, 43(1):1–43, June 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=43&issue=1&
spage=1.

Dingel:2006:CAC

[1224] Juergen Dingel. Compositional analysis of C/C++ programs with VeriSoft. Acta Informatica, 43(1):45–71, June 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=1&spage=45.

Corradini:2006:FAS

[1225] F. Corradini, M. R. Di Berardini, and W. Vogler. Fairness of actions in system computations. Acta Informatica, 43(2):73–130, August 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=43&issue=2&spage=73.

Pan:2006:SHS

[1226] Linqiang Pan and Artiom Alhazov. Solving HPP and SAT by P systems with active membranes and separation rules. Acta Informatica, 43(2): 131–145, August 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=2&spage=131.

Arora:2006:TAW

[1227] Amrinder Arora, Fanchun Jin, Gokhan Sahin, Hosam Mahmoud, and Hyeong-Ah Choi. Throughput analysis in wireless networks with multiple users and multiple channels. Acta Informatica, 43(3):147–164, October 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=43&issue=3&spage=147.

Harju:2006:PEW

[1228] Tero Harju and Dirk Nowotka. Periods in extensions of words. Acta Informatica, 43(3):165–171, October 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=43&issue=3&
spage=165.

Li:2006:CDL

[1229] Zheng-Zhu Li, H. J. Shyr, and Y. S. Tsai. Classifications of dense languages. Acta Informatica, 43(3):173–194, October 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=3&spage=173.

Hesselink:2006:RVL

[1230] Wim H. Hesselink. Refinement verification of the lazy caching algorithm. Acta Informatica, 43(3):195–222, October 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=3&spage=195.

Bordihn:2006:PGT

[1231] Henning Bordihn and Markus Holzer. Programmed grammars and their relation to the LBA problem. Acta Informatica, 43(4):223–242, November 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=4&spage=223.

Aguech:2006:DRD

[1232] Rafik Aguech, Nabil Lasmar, and Hosam Mahmoud. Distances in random digital search trees. Acta Informatica, 43(4):243–264, November 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=43&issue=4&spage=243.

Carayol:2006:LBI

[1233] Arnaud Carayol and Antoine Meyer. Linearly bounded infinite graphs. Acta Informatica, 43(4):265–292, November 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=4&spage=265.

Amigo:2006:RIP

[1234] José María Amigó. Representing the integers with powers of 2 and 3. Acta Informatica, 43(5):293–306, December 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=5&spage=293.

Khomenko:2006:MPN

[1235] Victor Khomenko, Alex Kondratyev, Maciej Koutny, and Walter Vogler. Merged processes: a new condensed representation of Petri net behaviour. Acta Informatica, 43(5):307–330, December 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=5&spage=307.

Alhazov:2006:NNU

[1236] Artiom Alhazov, Carlos Martín-Vide, and Yurii Rogozhin. On the number of nodes in universal networks of evolutionary processors. Acta Informatica, 43(5):331–339, December 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=5&spage=331.

Parama:2006:SAO

[1237] José R. Paramá, Nieves R. Brisaboa, Miguel R. Penabad, and Ángeles S. Places. A semantic approach to optimize linear datalog programs. Acta Informatica, 43(5):341–370, December 2006. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=5&spage=341.

Janssen:2007:TTP

[1238] Wim Janssen, Alexandr Korlyukov, and Jan Van den Bussche. On the tree-transformation power of XSLT. Acta Informatica, 43(6):371– 393, January 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=6&spage=371.

Konstantinidis:2007:RUA

[1239] Stavros Konstantinidis, Nicolae Santean, and Sheng Yu. Representation and uniformization of algebraic transductions. Acta Informatica, 43(6): 395–417, January 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=6&spage=395.

Honkala:2007:NBD

[1240] Juha Honkala. A new bound for the D0L sequence equivalence problem. Acta Informatica, 43(6):419–429, January 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=6&spage=419.

Gravell:2007:VCC

[1241] Andrew M. Gravell. Verification conditions are code. Acta Informatica, 43(6):431–447, January 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=43&issue=6&
spage=431.

Olderog:2007:EHS

[1242] Ernst-Rüdiger Olderog and Anders P. Ravn. Editorial: Hybrid systems. Acta Informatica, 43(7):449–450, February 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=7&spage=449.

Asarin:2007:HMA

[1243] Eugene Asarin, Thao Dang, and Antoine Girard. Hybridization methods for the analysis of nonlinear systems. Acta Informatica, 43(7):451– 476, February 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=7&spage=451.

Tabuada:2007:SMC

[1244] Paulo Tabuada. Symbolic models for control systems. Acta Informatica, 43(7):477–500, February 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=43&issue=7&
spage=477.

Wisniewski:2007:GAN

[1245] Rafael Wisniewski and Martin Raussen. Geometric analysis of nondeterminacy in dynamical systems. Acta Informatica, 43(7):501–519, February 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=7&spage=501.

Currie:2007:AAB

[1246] James D. Currie and Terry I. Visentin. On Abelian 2-avoidable binary patterns. Acta Informatica, 43(8):521–533, March 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=43&issue=8&spage=521.

Fu:2007:FA

[1247] Yuxi Fu. Fair ambients. Acta Informatica, 43(8):535–594, March 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=43&issue=8&spage=535.

Anonymous:2007:ARa

[1248] Anonymous. Acknowledgement to referees. Acta Informatica, 43(8): 595, March 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=43&issue=8&spage=595.

Vagner:2007:PLPa

[1249] Ladislav Vagner and Borivoj Melichar. Parallel LL parsing. Acta Informatica, 44(1):1–21, April 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=44&issue=1&
spage=1.

Dai:2007:EST

[1250] Tian-Shyr Dai and Yuh-Dauh Lyuu. An exact subexponential-time lattice algorithm for Asian options. Acta Informatica, 44(1):23–39, April 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=44&issue=1&spage=23.

Derrick:2007:UDA

[1251] John Derrick and Heike Wehrheim. On using data abstractions for model checking refinements. Acta Informatica, 44(1):41–71, April 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=44&issue=1&spage=41.

Vagner:2007:PLPb

[1252] Ladislav Vagner and Borivoj Melichar. Parallel LL parsing. Acta Informatica, 44(1):73, April 2007. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0001-5903&volume=44&issue=1&spage=73.

Bergstra:2007:DPP

[1253] Jan A. Bergstra, Inge Bethke, and Alban Ponse. Decision problems for pushdown threads. Acta Informatica, 44(2):75–90, May 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=44&issue=2&spage=75.

Kahrs:2007:IRM

[1254] Stefan Kahrs. Infinitary rewriting: meta-theory and convergence. Acta Informatica, 44(2):91–121, May 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=44&issue=2&
spage=91.

Hesselink:2007:CAR

[1255] Wim H. Hesselink. A criterion for atomicity revisited. Acta Informatica, 44(2):123–151, May 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=44&issue=2&
spage=123.

Kari:2007:SMH

[1256] Lila Kari, Kalpana Mahalingam, and Gabriel Thierrin. The syntactic monoid of hairpin-free languages. Acta Informatica, 44(3–4):153– 166, July 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=3&spage=153.

Okhotin:2007:RDP

[1257] Alexander Okhotin. Recursive descent parsing for Boolean grammars. Acta Informatica, 44(3–4):167–189, July 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=44&issue=3&spage=167.

Halava:2007:IMP

[1258] Vesa Halava and Mika Hirvensalo. Improved matrix pair undecidability results. Acta Informatica, 44(3–4):191–205, July 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=44&issue=3&spage=191.

Vincent:2007:EBF

[1259] Millist W. Vincent, Jixue Liu, and Mukesh Mohania. On the equivalence between FDs in XML and FDs in relations. Acta Informatica, 44(3–4): 207–247, July 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=3&spage=207.

Geeraerts:2007:WSL

[1260] Gilles Geeraerts, Jean-François Raskin, and Laurent Van Begin. Wellstructured languages. Acta Informatica, 44(3–4):249–288, July 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=44&issue=3&spage=249.

Afrati:2007:VSR

[1261] Foto Afrati, Rada Chirkova, Manolis Gergatsoulis, and Vassia Pavlaki. View selection for real conjunctive queries. Acta Informatica, 44(5):289– 321, September 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=5&spage=289.

Morris:2007:DUN

[1262] Joseph M. Morris and Malcolm Tyrrell. Dual unbounded nondeterminacy, recursion, and fixpoints. Acta Informatica, 44(5):323–344, September 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=5&spage=323.

Iwamoto:2007:TLD

[1263] Chuzo Iwamoto, Naoki Hatayama, Yoshiaki Nakashiba, Kenichi Morita, and Katsunobu Imai. Translational lemmas for DLOGTIME-uniform circuits, alternating TMs, and PRAMs. Acta Informatica, 44(5):345– 359, September 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=5&spage=345.

Bernini:2007:GEG

[1264] Antonio Bernini, Elisabetta Grazzini, Elisa Pergola, and Renzo Pinzani. A general exhaustive generation algorithm for Gray structures. Acta Informatica, 44(5):361–376, September 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=44&issue=5&spage=361.

Fuzzati:2007:DCR

[1265] Rachele Fuzzati, Massimo Merro, and Uwe Nestmann. Distributed consensus, revisited. Acta Informatica, 44(6):377–425, October 2007. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=44&issue=6&spage=377.

Scott:2007:BCT

[1266] Elizabeth Scott, Adrian Johnstone, and Rob Economopoulos. BRNGLR: a cubic Tomita-style GLR parsing algorithm. Acta Informatica, 44(6):

427–461, October 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=6&spage=427.

Haddad:2007:RPN

[1267] Serge Haddad and Denis Poitrenaud. Recursive Petri nets. Acta Informatica, 44(7–8):463–508, December 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=44&issue=
7&spage=463.

Nishimura:2007:SSU

[1268] Naomi Nishimura, Prabhakar Ragde, and Stefan Szeider. Solving #sat using vertex covers. Acta Informatica, 44(7–8):509–523, December 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=44&issue=7&spage=509.

Bergstra:2007:SCE

[1269] J. A. Bergstra and C. A. Middelburg. Synchronous cooperation for explicit multi-threading. Acta Informatica, 44(7–8):525–569, December 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=44&issue=7&spage=525.

Jiang:2007:OSO

[1270] Yiwei Jiang and Yong He. Optimal semi-online algorithms for preemptive scheduling problems with inexact partial information. Acta Informatica, 44(7–8):571–590, December 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=44&issue=7&
spage=571.

Calders:2007:CSC

[1271] Toon Calders. The complexity of satisfying constraints on databases of transactions. Acta Informatica, 44(7–8):591–624, December 2007. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=44&issue=7&spage=591.

Anonymous:2007:ARb

[1272] Anonymous. Acknowledgement to referees. Acta Informatica, 44(7–8): 625, December 2007. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=44&issue=7&spage=625.

Bozapalidis:2008:PD

[1273] Symeon Bozapalidis. Picture deformation. Acta Informatica, 45(1):1– 31, February 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=45&issue=1&spage=1.

Elmasry:2008:ASI

[1274] Amr Elmasry and Michael L. Fredman. Adaptive sorting: an information theoretic perspective. Acta Informatica, 45(1):33–42, February 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=45&issue=1&spage=33.

Duan:2008:DPP

[1275] Zhenhua Duan, Cong Tian, and Li Zhang. A decision procedure for propositional projection temporal logic with infinite models. Acta Informatica, 45(1):43–78, February 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=45&issue=1&
spage=43.

Cieslik:2008:LGC

[1276] Iwona Cieślik. On-line graph coloring of P_{5}-free graphs. Acta Informatica, 45(2):79–91, April 2008. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0001-5903&volume=45&issue=2&spage=79.

Magnani:2008:MIP

[1277] Matteo Magnani and Danilo Montesi. Management of interval probabilistic data. Acta Informatica, 45(2):93–130, April 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=45&issue=2&spage=93.

Brazdil:2008:DPB

[1278] Tomás Brázdil, Antonín Kucera, and Oldrich Strazovský. Deciding probabilistic bisimilarity over infinite-state probabilistic systems. Acta Informatica, 45(2):131–154, April 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=45&issue=2&
spage=131.

Epstein:2008:TDP

[1279] Leah Epstein, Asaf Levin, and Rob van Stee. Two-dimensional packing with conflicts. Acta Informatica, 45(3):155–175, May 2008. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=45&issue=3&spage=155.

Kutrib:2008:BCL

[1280] Martin Kutrib, Andreas Malcher, and Detlef Wotschke. The Boolean closure of linear context-free languages. Acta Informatica, 45(3):177– 191, May 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=45&issue=3&spage=177.

Elmasry:2008:TTR

[1281] Amr Elmasry, Claus Jensen, and Jyrki Katajainen. Two-tier relaxed heaps. Acta Informatica, 45(3):193–210, May 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=45&issue=3&spage=193.

Berghammer:2008:ARA

[1282] Rudolf Berghammer. Applying relation algebra and RelView to solve problems on orders and lattices. Acta Informatica, 45(3):211–236, May 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=45&issue=3&spage=211.

Broutin:2008:WHR

[1283] N. Broutin, L. Devroye, and E. McLeish. Weighted height of random trees. Acta Informatica, 45(4):237–277, June 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=45&issue=4&spage=237.

Janicki:2008:RSM

[1284] Ryszard Janicki. Relational structures model of concurrency. Acta Informatica, 45(4):279–320, June 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=45&issue=4&
spage=279.

Meinicke:2008:ARP

[1285] Larissa Meinicke and Ian J. Hayes. Algebraic reasoning for probabilistic action systems and while-loops. Acta Informatica, 45(5):321– 382, July 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=45&issue=5&spage=321.

Brijder:2008:FRR

[1286] Robert Brijder and Hendrik Jan Hoogeboom. The fibers and range of reduction graphs in ciliates. Acta Informatica, 45(5):383–402, July 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=45&issue=5&spage=383.

Godlin:2008:IRP

[1287] Benny Godlin and Ofer Strichman. Inference rules for proving the equivalence of recursive procedures. Acta Informatica, 45(6):403–439, August 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=45&issue=6&spage=403.

Morris:2008:MHO

[1288] Joseph M. Morris and Malcolm Tyrrell. Modelling higher-order dual nondeterminacy. Acta Informatica, 45(6):441–465, August 2008. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=45&issue=6&spage=441.

Fan:2008:RAL

[1289] Chen-Ming Fan, C. C. Huang, and H. J. Shyr. Regular autodense languages. Acta Informatica, 45(7–8):467–477, December 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=45&issue=7&spage=467.

Damiani:2008:TSS

[1290] Ferruccio Damiani, Elena Giachino, Paola Giannini, and Sophia Drossopoulou. A type safe state abstraction for coordination in Java-like languages. Acta Informatica, 45(7–8):479–536, December 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=45&issue=7&spage=479.

Klaudel:2008:MNS

[1291] Hanna Klaudel and Franck Pommereau. M-nets: a survey. Acta Informatica, 45(7–8):537–564, December 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=45&issue=
7&spage=537.

Link:2008:CCF

[1292] Sebastian Link. Charting the completeness frontier of inference systems for multivalued dependencies. Acta Informatica, 45(7–8):565–591, December 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=45&issue=7&spage=565.

Meduna:2008:SCG

[1293] Alexander Meduna and Jiří Techet. Scattered context grammars that erase nonterminals in a generalized k-limited way. Acta Informatica, 45(7–8):593–608, December 2008. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=45&issue=7&
spage=593.

Bozzelli:2009:DLM

[1294] Laura Bozzelli, Mojmír Křetínsky, Vojtěch Řehák, and Jan Strejček. On decidability of LTL model checking for process rewrite systems. Acta Informatica, 46(1):1–28, February 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=46&issue=1&
spage=1.

Varany:2009:SST

[1295] Vince Várány. Semi-synchronous transductions. Acta Informatica, 46 (1):29–42, February 2009. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0001-5903&volume=46&issue=1&spage=29.

Bashirov:2009:ECP

[1296] Rza Bashirov, Fabrice Kordon, and Hüseyin Lort. Exploiting colored Petri nets to decide on permutation admissibility. Acta Informatica, 46(1):43–55, February 2009. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0001-5903&volume=46&issue=1&spage=43.

Ben-Amram:2009:CTR

[1297] Amir M. Ben-Amram. A complexity tradeoff in ranking-function termination proofs. Acta Informatica, 46(1):57–72, February 2009. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=46&issue=1&spage=57.

Aravind:2009:QBM

[1298] Alex A. Aravind and Wim H. Hesselink. A queue based mutual exclusion algorithm. Acta Informatica, 46(1):73–86, February 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=1&spage=73.

Meyer:2009:TSS

[1299] Roland Meyer. A theory of structural stationarity in the πcalculus. Acta Informatica, 46(2):87–137, April 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=46&issue=2&spage=87.

Engelfriet:2009:TCT

[1300] Joost Engelfriet. The time complexity of typechecking tree-walking tree transducers. Acta Informatica, 46(2):139–154, April 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=46&issue=2&spage=139.

Subramani:2009:RWS

[1301] K. Subramani, Hong-Jian Lai, and Xiaofeng Gu. Random walks for selected boolean implication and equivalence problems. Acta Informatica, 46(2):155–168, April 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=46&issue=2&
spage=155.

Sawa:2009:HEC

[1302] Zdeněk Sawa and Petr Jančar. Hardness of equivalence checking for composed finite-state systems. Acta Informatica, 46(3):169–191, May 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=3&spage=169.

Campeanu:2009:CPE

[1303] Cezar Câmpeanu and Nicolae Santean. On the closure of pattern expressions languages under intersection with regular languages. Acta Informatica, 46(3):193–207, May 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=46&issue=3&
spage=193.

Corradini:2009:LMA

[1304] Flavio Corradini, Maria Rita Di Berardini, and Walter Vogler. Liveness of a mutex algorithm in a fair process algebra. Acta Informatica, 46(3): 209–235, May 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=46&issue=3&spage=209.

Best:2009:DTF

[1305] Eike Best and Philippe Darondeau. A decomposition theorem for finite persistent transition systems. Acta Informatica, 46(3):237–254, May 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=3&spage=237.

Brucker:2009:SCA

[1306] Achim D. Brucker and Burkhart Wolff. Semantics, calculi, and analysis for object-oriented specifications. Acta Informatica, 46(4):255– 284, July 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=46&issue=4&spage=255.

Klonowska:2009:MGI

[1307] Kamilla Klonowska, Lars Lundberg, and Håkan Lennerstad. The maximum gain of increasing the number of preemptions in multiprocessor scheduling. Acta Informatica, 46(4):285–295, July 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=46&issue=4&spage=285.

Armendariz-Inigo:2009:CPD

[1308] José Enrique Armendáriz-Iñigo, José Ramón González de Mendívil, José Ramón Garitagoitia, and Francesc D. Muñoz-Esco. Correctness proof of a database replication protocol under the perspective of the I/O automaton model. Acta Informatica, 46(4):297–330, July 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=4&spage=297.

Bresolin:2009:TUP

[1309] Davide Bresolin, Angelo Montanari, and Gabriele Puppis. A theory of ultimately periodic languages and automata with an application to time granularity. Acta Informatica, 46(5):331–360, August 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=5&spage=331.

Ciobanu:2009:FSB

[1310] Gabriel Ciobanu and Sergiu Rudeanu. Final and sequential behaviours of M-automata. Acta Informatica, 46(5):361–374, August 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=5&spage=361.

Bergstra:2009:MSO

[1311] J. A. Bergstra and C. A. Middelburg. Machine structure oriented control code logic. Acta Informatica, 46(5):375–401, August 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=5&spage=375.

Zuliani:2009:RAF

[1312] Paolo Zuliani. Reasoning about faulty quantum programs. Acta Informatica, 46(6):403–432, October 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=46&issue=6&
spage=403.

Khomenko:2009:SDS

[1313] Victor Khomenko, Mark Schaefer, Walter Vogler, and Ralf Wollowski. STG decomposition strategies in combination with unfolding. Acta Informatica, 46(6):433–474, October 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=46&issue=6&
spage=433.

Shahriar:2009:PKX

[1314] Md. Sumon Shahriar and Jixue Liu. Preserving key in XML data transformation. Acta Informatica, 46(7):475–507, November 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=46&issue=7&spage=475.

Quesada:2009:PDP

[1315] Argimiro Arratia Quesada and Iain A. Stewart. On the power of deep pushdown stacks. Acta Informatica, 46(7):509–531, November 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=7&spage=509.

Janousek:2009:RTL

[1316] Jan Janoušek and Bořivoj Melichar. On regular tree languages and deterministic pushdown automata. Acta Informatica, 46(7):533–547, November 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=46&issue=7&spage=533.

Aycock:2009:EAE

[1317] John Aycock and Angelo Borsotti. Early action in an Earley parser. Acta Informatica, 46(8):549–559, December 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=46&issue=8&spage=549.

Engelfriet:2009:EMB

[1318] Joost Engelfriet, Eric Lilin, and Andreas Maletti. Extended multi bottom-up tree transducers: Composition and decomposition. Acta Informatica, 46(8):561–590, December 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=46&issue=
8&spage=561.

Carpi:2009:STA

[1319] Arturo Carpi and Flavio D’Alessandro. Strongly transitive automata and the Černý conjecture. Acta Informatica, 46(8):591–607, December 2009. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=46&issue=8&spage=591.

Habermehl:2010:ABV

[1320] Peter Habermehl, Radu Iosif, and Tom Vojnar. Automata-based verification of programs with tree updates. Acta Informatica, 47(1):1– 31, February 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=1&spage=1.

Jaghoori:2010:SPO

[1321] Mohammad Mahdi Jaghoori, Marjan Sirjani, Mohammad Reza Mousavi, Ehsan Khamespanah, and Ali Movaghar. Symmetry and partial order reduction techniques in model checking Rebeca. Acta Informatica, 47 (1):33–66, February 2010. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://www.springerlink.com/openurl.
asp?genre=article&issn=0001-5903&volume=47&issue=1&spage=33.

Do:2010:MMR

[1322] Tien Van Do. M/M/1 retrial queue with working vacations. Acta Informatica, 47(1):67–75, February 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=47&issue=1&
spage=67.

Berghammer:2010:EMS

[1323] Rudolf Berghammer and Michael Winter. Embedding mappings and splittings with applications. Acta Informatica, 47(2):77–110, March 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=47&issue=2&spage=77.

Merro:2010:OTC

[1324] Massimo Merro. On the observational theory of the CPS-calculus. Acta Informatica, 47(2):111–132, March 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=47&issue=
2&spage=111.

Loos:2010:SUA

[1325] Remco Loos, Florin Manea, and Victor Mitrana. Small universal accepting hybrid networks of evolutionary processors. Acta Informatica, 47(2):

133–146, March 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=2&spage=133.

Aceto:2010:LNF

[1326] Luca Aceto, Wan Fokkink, Anna Ingolfsdottir, and MohammadReza Mousavi. Lifting non-finite axiomatizability results to extensions of process algebras. Acta Informatica, 47(3):147–177, May 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=47&issue=3&spage=147.

Kim:2010:LER

[1327] Ik-Soon Kim and Kwangkeun Yi. LR error repair using the A* algorithm. Acta Informatica, 47(3):179–207, May 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=47&issue=3&spage=179.

Fan:2010:NAR

[1328] Chen-Ming Fan, C. C. Huang, H. J. Shyr, and Kuo-Hsiang Chen. A note on autodense related languages. Acta Informatica, 47(4):209– 219, June 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=4&spage=209.

Ying:2010:QLP

[1329] Mingsheng Ying and Yuan Feng. Quantum loop programs. Acta Informatica, 47(4):221–250, June 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=47&issue=4&
spage=221.

Dax:2010:RTL

[1330] Christian Dax, Felix Klaedtke, and Martin Lange. On regular temporal logics with past. Acta Informatica, 47(4):251–277, June 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=47&issue=4&spage=251.

Lanotte:2010:RRT

[1331] Ruggero Lanotte, Andrea Maggiolo-Schettini, and Angelo Troina. Reachability results for timed automata with unbounded data structures. Acta Informatica, 47(5–6):279–311, September 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=47&issue=5&spage=279.

Dolev:2010:RAS

[1332] Shlomi Dolev and Nir Tzachar. Randomization adaptive selfstabilization. Acta Informatica, 47(5–6):313–323, September 2010. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=47&issue=5&spage=313.

Hochbaum:2010:HAR

[1333] Dorit S. Hochbaum and Asaf Levin. How to allocate review tasks for robust ranking. Acta Informatica, 47(5–6):325–345, September 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=47&issue=5&spage=325.

Huang:2010:NPC

[1334] C. C. Huang. A note on pure codes. Acta Informatica, 47(5–6):347– 357, September 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=5&spage=347.

Boyar:2010:TCL

[1335] Joan Boyar, Martin R. Ehmsen, Jens S. Kohrt, and Kim S. Larsen. A theoretical comparison of LRU and LRU-K. Acta Informatica, 47(7–8):359– 374, December 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=7&spage=359.

Epstein:2010:ETP

[1336] Leah Epstein. Equilibria for two parallel links: the strong price of anarchy versus the price of anarchy. Acta Informatica, 47(7–8):375– 389, December 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=7&spage=375.

Kutrib:2010:SDR

[1337] Martin Kutrib, Hartmut Messerschmidt, and Friedrich Otto. On stateless deterministic restarting automata. Acta Informatica, 47(7–8):391– 412, December 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=47&issue=7&spage=391.

Fan:2010:NPP

[1338] Chen-Ming Fan and C. C. Huang. A note on prefix primitive words. Acta Informatica, 47(7–8):413–423, December 2010. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=47&issue=7&spage=413.

Li:2011:SPD

[1339] Zheng-Zhu Li and Y. S. Tsai. Some properties of the disjunctive languages contained in Q. Acta Informatica, 48(1):1–18, February 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=48&issue=1&spage=1.

Honkala:2011:CRD

[1340] Juha Honkala. A characterization of rational D0L power series. Acta Informatica, 48(1):19–24, February 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=1&spage=19.

Blanchet-Sadri:2011:ABP

[1341] F. Blanchet-Sadri, Robert Mercaş, Sean Simmons, and Eric Weissenstein. Avoidable binary patterns in partial words. Acta Informatica, 48(1):25– 41, February 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=48&issue=1&spage=25. See

erratum [1361].

Mitrana:2011:NBQ

[1342] Victor Mitrana and Cristina Tîrnăucă. New bounds for the query complexity of an algorithm that learns DFAs with correction and equivalence queries. Acta Informatica, 48(1):43–50, February 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=1&spage=43.

Kabir:2011:ESC

[1343] Md. Enamul Kabir, Hua Wang, and Elisa Bertino. Efficient systematic clustering method for k-anonymization. Acta Informatica, 48(1):

51–66, February 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=48&issue=1&spage=51.

Aravind:2011:NDB

[1344] Alex A. Aravind and Wim H. Hesselink. Nonatomic dual bakery algorithm with bounded tokens. Acta Informatica, 48(2):67–96, April 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=48&issue=2&spage=67.

Cavalcanti:2011:TRC

[1345] Ana Cavalcanti and Marie-Claude Gaudel. Testing for refinement in Circus. Acta Informatica, 48(2):97–147, April 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=2&spage=97.

Meduna:2011:OSR

[1346] Alexander Meduna and Petr Zemek. One-sided random context grammars. Acta Informatica, 48(3):149–163, May 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=3&spage=149.

Drewes:2011:MLT

[1347] Frank Drewes, Johanna Högberg, and Andreas Maletti. MAT learners for tree series: an abstract data type and two realizations. Acta Informatica, 48(3):165–189, May 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=48&issue=3&
spage=165.

Seidel:2011:RTL

[1348] Daniel Seidel and Janis Voigtländer. Refined typing to localize the impact of forced strictness on free theorems. Acta Informatica, 48(3): 191–211, May 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=48&issue=3&spage=191.

Polyvyanyy:2011:CWN

[1349] Artem Polyvyanyy, Matthias Weidlich, and Mathias Weske. Connectivity of workflow nets: the foundations of stepwise verification. Acta Informatica, 48(4):213–242, July 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=48&issue=4&
spage=213.

Do:2011:MQV

[1350] Tien Van Do, Ram Chakka, Nam H. Do, and László Pap. A Markovian queue with varying number of servers and applications to the performance comparison of HSDPA user equipment. Acta Informatica, 48(4): 243–269, July 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=48&issue=4&spage=243.

Qiu:2011:MLQ

[1351] Daowen Qiu, Lvzhou Li, Xiangfu Zou, Paulo Mateus, and Jozef Gruska. Multi-letter quantum finite automata: decidability of the equivalence and minimization of states. Acta Informatica, 48(5–6):271–290, August 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=48&issue=5&spage=271.

Rabe:2011:FOC

[1352] Markus N. Rabe and Sven Schewe. Finite optimal control for time-bounded reachability in CTMDPs and continuous-time Markov games. Acta Informatica, 48(5–6):291–315, August 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=5&spage=291.

Vagvolgyi:2011:CRC

[1353] Sándor Vágvölgyi. CHAP and rewrite components. Acta Informatica, 48(5–6):317–361, August 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=48&issue=5&
spage=317.

Perna:2011:CHS

[1354] Juan Perna, Jim Woodcock, Augusto Sampaio, and Juliano Iyoda. Correct hardware synthesis: an algebraic approach. Acta Informatica, 48 (7–8):363–396, December 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=48&issue=7&
spage=363.

Domosi:2011:PA

[1355] Pál Dömösi and György Maróti. On α_{2}-ν_{2}-products of automata. Acta Informatica, 48(7–8):397–408, December 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.

springerlink.com/openurl.asp?genre=article&issn=0001-5903&volume=
48&issue=7&spage=397.

Aman:2011:SWN

[1356] Bogdan Aman and Gabriel Ciobanu. Solving a weak NP-complete problem in polynomial time by using mutual mobile membrane systems. Acta Informatica, 48(7–8):409–415, December 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=7&spage=409.

Chakrabarti:2011:QCE

[1357] Amit Chakrabarti, Venkatesan Guruswami, Andrew Wirth, and Anthony Wirth. The query complexity of estimating weighted averages. Acta Informatica, 48(7–8):417–426, December 2011. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=48&issue=7&spage=417.

Coffman:2012:EAF

[1358] Edward G. Coffman, Dariusz Dereniowski, and Wieslaw Kubiak. An efficient algorithm for finding ideal schedules. Acta Informatica, 49(1): 1–14, February 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=49&issue=1&spage=1.

Wang:2012:UTS

[1359] Yunhe Wang and Li Jiao. Using transition set sequences to partition behaviors of Petri nets. Acta Informatica, 49(1):15–28, February 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=49&issue=1&spage=15.

Bozapalidis:2012:EWT

[1360] Symeon Bozapalidis, Zoltán Fülöp, and George Rahonis. Equational weighted tree transformations. Acta Informatica, 49(1):29–52, February 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=49&issue=1&spage=29.

Blanchet-Sadri:2012:EAB

[1361] F. Blanchet-Sadri, Robert Mercas, Sean Simmons, and Eric Weissenstein. Erratum to: “Avoidable binary patterns in partial words”. Acta Informatica, 49(1):53–54, February 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=49&issue=1&

spage=53. See [1341].

Meduna:2012:NCO

[1362] Alexander Meduna and Petr Zemek. Nonterminal complexity of onesided random context grammars. Acta Informatica, 49(2):55–68, March 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=49&issue=2&spage=55.

Stahl:2012:TBS

[1363] Christian Stahl and Walter Vogler. A trace-based service semantics guaranteeing deadlock freedom. Acta Informatica, 49(2):69–103, March 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=49&issue=2&spage=69.

Lorencs:2012:IPF

[1364] A. Lorencs. The identity problem of finitely generated bi-ideals. Acta Informatica, 49(2):105–115, March 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=49&issue=
2&spage=105.

Lescanne:2012:SCN

[1365] Pierre Lescanne and Matthieu Perrinel. ‘Backward’ coinduction, Nash equilibrium and the rationality of escalation. Acta Informatica, 49(3): 117–137, May 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=49&issue=3&spage=117.

Bergstra:2012:ISP

[1366] J. A. Bergstra and C. A. Middelburg. Instruction sequence processing operators. Acta Informatica, 49(3):139–172, May 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=49&issue=3&spage=139.

Wang:2012:LCC

[1367] Junhu Wang, Jeffrey Xu Yu, Chaoyi Pang, and Chengfei Liu. Least common container of tree pattern queries and its applications. Acta Informatica, 49(3):173–202, May 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=49&issue=3&
spage=173.

Baier:2012:SGL

[1368] Christel Baier, Tomás Brázdil, Marcus Größer, and Antonín Kucera. Stochastic game logic. Acta Informatica, 49(4):203–224, June 2012. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://www.springerlink.com/openurl.asp?genre=article&issn=
0001-5903&volume=49&issue=4&spage=203.

Choffrut:2012:FOL

[1369] Christian Choffrut, Andreas Malcher, Carlo Mereghetti, and Beatrice Palano. First-order logics: some characterizations and closure properties. Acta Informatica, 49(4):225–248, June 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=49&issue=4&spage=225.

Bozzelli:2012:TAS

[1370] Laura Bozzelli, Axel Legay, and Sophie Pinchinat. On timed alternating simulation for concurrent timed games. Acta Informatica, 49(4): 249–279, June 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=49&issue=4&spage=249.

Fan:2012:PPA

[1371] Chen-Ming Fan, Cheng-Chih Huang, Christine Chifen Tseng, and Jen-Tse Wang. Prefix-primitive annihilators of languages under some operations. Acta Informatica, 49(5):281–293, August 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=49&issue=5&spage=281.

Katz:2012:ASS

[1372] Bastian Katz, Ignaz Rutter, and Gerhard Woeginger. An algorithmic study of switch graphs. Acta Informatica, 49(5):295–312, August 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://www.springerlink.com/openurl.asp?genre=
article&issn=0001-5903&volume=49&issue=5&spage=295.

Santone:2012:ARD

[1373] Antonella Santone and Gigliola Vaglini. Abstract reduction in directed model checking CCS processes. Acta Informatica, 49(5):313– 341, August 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=49&issue=5&spage=313.

Guttmann:2012:AII

[1374] Walter Guttmann. Algebras for iteration and infinite computations. Acta Informatica, 49(5):343–359, August 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=49&issue=5&spage=343.

Tan:2012:IEM

[1375] Zhiyi Tan, Long Wan, Qi Zhang, and Wei Ren. Inefficiency of equilibria for the machine covering game on uniform machines. Acta Informatica, 49(6):361–379, September 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.com/

openurl.asp?genre=article&issn=0001-5903&volume=49&issue=6&
spage=361.

Golovach:2012:FVS

[1376] Petr A. Golovach, Bernard Lidický, Barnaby Martin, and Daniël Paulusma. Finding vertex-surjective graph homomorphisms. Acta Informatica, 49(6):381–394, September 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://www.springerlink.

com/openurl.asp?genre=article&issn=0001-5903&volume=49&issue=
6&spage=381.

Vasile:2012:PEN

[1377] Cristian Ioan Vasile, Ana Brândusa Pavel, Ioan Dumitrache, and Gheorghe Paun. On the power of enzymatic numerical P systems. Acta Informatica, 49(6):395–412, September 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http:

//www.springerlink.com/openurl.asp?genre=article&issn=0001-
5903&volume=49&issue=6&spage=395.

Caron:2012:MTB

[1378] Pascal Caron, Jean-Marc Champarnaud, and Ludovic Mignot. Multitilde-bar expressions and their automata. Acta Informatica, 49(6):413– 436, September 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL http://www.springerlink.com/openurl.asp?
genre=article&issn=0001-5903&volume=49&issue=6&spage=413.

Chunhua:2012:DLR

[1379] Cao Chunhua, Yang Di, and Liu Yin. Disjunctive languages related to p-primitive words. Acta Informatica, 49(7–8):437–444, November 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-012-0165-z.

Xu:2012:DRH

[1380] Xian Xu. Distinguishing and relating higher-order and first-order processes by expressiveness. Acta Informatica, 49(7–8):445–484, November 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
012-0168-9.

Brand:2012:DIA

[1381] Michael Brand. Does indirect addressing matter? Acta Informatica, 49(7–8):485–491, November 2012. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-012-0171-1.

Arroyo:2013:ASS

[1382] Fernando Arroyo, Juan Castellanos, Jürgen Dassow, and Victor Mitrana . . .. Accepting splicing systems with permitting and forbidding words. Acta Informatica, 50(1):1–14, February 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-012-0169-8.

Kleijn:2013:SSB

[1383] Jetty Kleijn, Maciej Koutny, and Marta Pietkiewicz-Koutny . . .. Step semantics of boolean nets. Acta Informatica, 50(1):15–39, February 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-012-0170-2.

Dassow:2013:NEP

[1384] Jürgen Dassow, Florin Manea, and Bianca Truthe. Networks of evolutionary processors: the power of subregular filters. Acta Informatica, 50(1):41–75, February 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-012-0172-0.

Bettini:2013:CTC

[1385] Lorenzo Bettini, Ferruccio Damiani, and Ina Schaefer. Compositional type checking of delta-oriented software product lines. Acta Informatica, 50(2):77–122, March 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-012-0173-z.

Kahrs:2013:IRC

[1386] Stefan Kahrs. Infinitary rewriting: closure operators, equivalences and models. Acta Informatica, 50(2):123–156, March 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer.com/article/10.1007/s00236-012-0174-y.

Bilotta:2013:ACB

[1387] Stefano Bilotta, Elisabetta Grazzini, Elisa Pergola, and Renzo Pinzani. Avoiding cross-bifix-free binary words. Acta Informatica, 50(3):157–173, May 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
013-0176-4.

Aizikowitz:2013:CGA

[1388] Tamar Aizikowitz and Michael Kaminski. Conjunctive grammars and alternating pushdown automata. Acta Informatica, 50(3):175–197, May 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
013-0177-3.

Hesselink:2013:VSM

[1389] Wim H. Hesselink. Verifying a simplification of mutual exclusion by Lycklama–Hadzilacos. Acta Informatica, 50(3):199–228, May 2013. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-013-0178-2.

Nagy:2013:DPC

[1390] Benedek Nagy and Friedrich Otto. Deterministic pushdown-CD-systems of stateless deterministic R(1)-automata. Acta Informatica, 50(4): 229–255, June 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-012-0175-x.

Chen:2013:QOI

[1391] Dongfeng Chen, Rada Chirkova, Fereidoon Sadri, and Tiia J. Salo. Query optimization in information integration. Acta Informatica, 50 (4):257–287, June 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-013-0179-1.

Elmasry:2013:HDS

[1392] Amr Elmasry, Arash Farzan, and John Iacono. On the hierarchy of distribution-sensitive properties for data structures. Acta Informatica, 50(4):289–295, June 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-013-0180-8.

Hesselink:2013:DRA

[1393] Wim H. Hesselink. A distributed resource allocation algorithm for many processes. Acta Informatica, 50(5–6):297–329, September 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-013-0181-7.

Halava:2013:NPU

[1394] Vesa Halava and Tero Harju. New proof for the undecidability of the circular PCP. Acta Informatica, 50(5–6):331–341, September 2013. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-013-0183-5.

Christ:2013:OMC

[1395] Marie G. Christ, Lene M. Favrholdt, and Kim S. Larsen. Online multi-coloring on the path revisited. Acta Informatica, 50(5–6):343– 357, September 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-013-0184-4.

Seo:2013:JSS

[1396] Jeongbong Seo and Sungwoo Park. Judgmental subtyping systems with intersection types and modal types. Acta Informatica, 50(7–8): 359–380, December 2013. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-013-0186-2.

DeFelice:2013:NFC

[1397] Clelia De Felice. A note on the factorization conjecture. Acta Informatica, 50(7–8):381–402, December 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-013-0187-1.

Rosa-Velardo:2013:PNN

[1398] Fernando Rosa-Velardo. Petri nets with name creation for transient secure association. Acta Informatica, 50(7–8):403–436, December 2013. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-013-0188-0.

Li:2014:TNQ

[1399] Yangjia Li, Nengkun Yu, and Mingsheng Ying. Termination of nondeterministic quantum programs. Acta Informatica, 51(1):1–24, February 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
013-0185-3.

Bozzelli:2014:VRE

[1400] Laura Bozzelli and César Sánchez. Visibly rational expressions. Acta Informatica, 51(1):25–49, February 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-013-0190-6.

Li:2014:DEE

[1401] Suo Ping Li, Yong Qiang Zhou, and Yong Zhou. Delay and energy efficiency analysis of multicast cooperative ARQ over wireless networks. Acta Informatica, 51(1):51–60, February 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-013-0192-4.

Mandrali:2014:WFO

[1402] Eleni Mandrali and George Rahonis. On weighted first-order logics with discounting. Acta Informatica, 51(2):61–106, March 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-013-0193-3.

Ranzato:2014:ESA

[1403] Francesco Ranzato. An efficient simulation algorithm on Kripke structures. Acta Informatica, 51(2):107–125, March 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.

springer.com/article/10.1007/s00236-014-0195-9.

Peled:2014:ESI

[1404] Doron Peled and Sven Schewe. Editorial: special issue on synthesis. Acta Informatica, 51(3–4):127–128, June 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-014-0198-6;
 http://link.springer.com/content/pdf/10.1007/s00236-014-0198-
6.pdf.

Chatterjee:2014:SSM

[1405] Krishnendu Chatterjee, Mickael Randour, and Jean-François Raskin. Strategy synthesis for multi-dimensional quantitative objectives. Acta Informatica, 51(3–4):129–163, June 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-013-0182-6.

Bulychev:2014:ECS

[1406] Peter Bulychev, Alexandre David, Kim G. Larsen, and Guangyuan Li. Efficient controller synthesis for a fragment of MTL_{0,∞}. Acta Informatica, 51(3–4):165–192, June 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-013-0189-z.

Bloem:2014:SRS

[1407] Roderick Bloem, Krishnendu Chatterjee, and Karin Greimel. Synthesizing robust systems. Acta Informatica, 51(3–4):193–220, June 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-013-0191-5.

Fridman:2014:DSR

[1408] Wladimir Fridman and Bernd Puchala. Distributed synthesis for regular and contextfree specifications. Acta Informatica, 51(3–4):221–260, June 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
014-0194-x.

Fahrenberg:2014:GQS

[1409] Uli Fahrenberg and Axel Legay. General quantitative specification theories with modal transition systems. Acta Informatica, 51(5):261–295, August 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
014-0196-8.

Fulop:2014:FBA

[1410] Zoltán Fülöp and Heiko Vogler. Forward and backward application of symbolic tree transducers. Acta Informatica, 51(5):297–325, August 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
014-0197-7.

Meduna:2014:CFA

[1411] Alexander Meduna and Petr Zemek. Controlled finite automata. Acta Informatica, 51(5):327–337, August 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-014-0199-5.

Chunhua:2014:SKP

[1412] Cao Chunhua, Yang Shuang, and Yang Di. Some kinds of primitive and non-primitive words. Acta Informatica, 51(6):339–346, September 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-014-0200-3.

Bertran:2014:FCE

[1413] Miquel Bertran, Francesc Babot, and August Climent. Formal communication elimination and sequentialization equivalence proofs for distributed system models. Acta Informatica, 51(6):347–418, September 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
014-0203-0.

Cazzola:2014:IGS

[1414] Walter Cazzola and Edoardo Vacchi. On the incremental growth and shrinkage of LR goto-graphs. Acta Informatica, 51(7):419–447, October 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
014-0201-2.

Sorge:2014:EHM

[1415] Manuel Sorge, Hannes Moser, Rolf Niedermeier, and Mathias Weller. Exploiting a hypergraph model for finding Golomb rulers. Acta Informatica, 51(7):449–471, October 2014. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-014-0202-1.

Belmonte:2014:PCT

[1416] Rémy Belmonte, Petr A. Golovach, Pim van ’t Hof, and Daniël Paulusma. Parameterized complexity of three edge contraction problems with degree constraints. Acta Informatica, 51(7):473–497, October 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-014-0204-z.

Vogler:2014:TFB

[1417] Walter Vogler, Christian Stahl, and Richard Müller. Trace- and failure-based semantics for responsiveness. Acta Informatica, 51(8): 499–552, November 2014. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-014-0205-y.

Kutrib:2014:HSH

[1418] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Head and state hierarchies for unary multi-head finite automata. Acta Informatica, 51(8):553–569, November 2014. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-014-0206-x.

Olderog:2015:LME

[1419] Ernst-Rüdiger Olderog. Letter from the Managing Editor. Acta Informatica, 52(1):1–2, February 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-014-0212-z;   http://link.springer.com/
content/pdf/10.1007/s00236-014-0212-z.pdf.

vanGlabbeek:2015:SICa

[1420] Rob J. van Glabbeek, Ursula Goltz, and Ernst-Rüdiger Olderog. Special issue on “Combining Compositionality and Concurrency”: part 1. Acta Informatica, 52(1):3–4, February 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-014-0213-y;
 http://link.springer.com/content/pdf/10.1007/s00236-014-0213-
y.pdf.

Bruni:2015:RCC

[1421] Roberto Bruni, Ugo Montanari, and Matteo Sammartino. Revisiting causality, coalgebraically. Acta Informatica, 52(1):5–33, February 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-014-0207-9.

Best:2015:SRP

[1422] Eike Best and Raymond Devillers. Synthesis and reengineering of persistent systems. Acta Informatica, 52(1):35–60, February 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-014-0209-7.

Bernardo:2015:RBM

[1423] Marco Bernardo, Rocco De Nicola, and Michele Loreti. Revisiting bisimilarity and its modal logic for nondeterministic and probabilistic processes. Acta Informatica, 52(1):61–106, February 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-014-0210-1.

Luttgen:2015:SIC

[1424] Gerald Lüttgen and Flavio Corradini. Special issue on “Comprehending asynchrony in specification and analysis” dedicated to Walter Vogler on the occasion of his 60th birthday. Acta Informatica, 52(2–3):107–108, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s00236-015-0214-5;   http://link.springer.com/article/
10.1007/s00236-015-0214-5.

Dang:2015:MAP

[1425] Han-Hing Dang and Bernhard Möller. Modal algebra and Petri nets. Acta Informatica, 52(2–3):109–132, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-015-0216-3.

Best:2015:SSA

[1426] Eike Best and Raymond Devillers. State space axioms for T -systems. Acta Informatica, 52(2–3):133–152, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-015-0219-0.

Desel:2015:OLP

[1427] Jörg Desel and Görkem Kilinç. Observable liveness of Petri nets. Acta Informatica, 52(2–3):153–174, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-015-0218-1.

vanGlabbeek:2015:CF

[1428] Rob J. van Glabbeek and Peter Höfner. CCS: It’s not fair! Acta Informatica, 52(2–3):175–205, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-015-0221-6.

Valmari:2015:CUL

[1429] Antti Valmari. On constructibility and unconstructibility of LTS operators from other LTS operators. Acta Informatica, 52(2–3):207–234, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
015-0217-2.

Hennicker:2015:MIT

[1430] Rolf Hennicker and Alexander Knapp. Moving from interface theories to assembly theories. Acta Informatica, 52(2–3):235–268, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-015-0220-7.

Benes:2015:RCP

[1431] Nikola Benes, Jan Kretínský, Kim G. Larsen, Mikael H. Møller, Salomon Sickert, and Jirí Srba. Refinement checking on parametric modal transition systems. Acta Informatica, 52(2–3):269–297, April 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-015-0215-4.

Broy:2015:LDI

[1432] Manfred Broy. A life dedicated to informatics: an obituary for Prof. Friedrich L. Bauer. Acta Informatica, 52(4–5):299–301, June 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
015-0241-2;   http://link.springer.com/content/pdf/10.1007/
s00236-015-0241-2.pdf.

vanGlabbeek:2015:SICb

[1433] Rob J. van Glabbeek, Ursula Goltz, and Ernst-Rüdiger Olderog. Special issue on “Combining Compositionality and Concurrency”: part 2. Acta Informatica, 52(4–5):303–304, June 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-015-0240-3;
 http://link.springer.com/content/pdf/10.1007/s00236-015-0240-
3.pdf.

Luttgen:2015:RIA

[1434] Gerald Lüttgen, Walter Vogler, and Sascha Fendrich. Richer interface automata with optimistic and pessimistic compatibility. Acta Informatica, 52(4–5):305–336, June 2015. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-014-0211-0.

Garavel:2015:CVA

[1435] Hubert Garavel, Frédéric Lang, and Radu Mateescu. Compositional verification of asynchronous concurrent systems using CADP. Acta Informatica, 52(4–5):337–392, June 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-015-0226-1.

Aguado:2015:DFP

[1436] Joaquín Aguado, Michael Mendler, Reinhard von Hanxleden, and Insa Fuhrmann. Denotational fixed-point semantics for constructive scheduling of synchronous concurrency. Acta Informatica, 52(4–5):393–442, June 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
015-0238-x.

Klein:2015:CCM

[1437] Joachim Klein, Christel Baier, and Sascha Klüppelholz. Compositional construction of most general controllers. Acta Informatica, 52 (4–5):443–482, June 2015. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-015-0239-9.

Fan:2015:SPI

[1438] Chen-Ming Fan, Jen-Tse Wang, and Cheng-Chih Huang. Some properties of involution binary relations. Acta Informatica, 52(6):483– 495, September 2015. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-014-0208-8.

Drewes:2015:CHR

[1439] Frank Drewes and Berthold Hoffmann. Contextual hyperedge replacement. Acta Informatica, 52(6):497–524, September 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-015-0223-4.

Sanchez:2015:PII

[1440] Alejandro Sánchez and César Sánchez. Parametrized invariance for infinite state processes. Acta Informatica, 52(6):525–557, September 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-015-0222-5.

Engelfriet:2015:TWP

[1441] Joost Engelfriet. Two-way pebble transducers for partial functions and their composition. Acta Informatica, 52(7–8):559–571, November 2015. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-

015-0224-3;   http://link.springer.com/content/pdf/10.1007/
s00236-015-0224-3.pdf.

Bernini:2015:GCO

[1442] A. Bernini, S. Bilotta, R. Pinzani, A. Sabri, and V. Vajnovszki. Gray code orders for q-ary words avoiding a given factor. Acta Informatica, 52(7– 8):573–592, November 2015. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-015-0225-2.

Otto:2015:DOR

[1443] Friedrich Otto and Frantisek Mráz. Deterministic ordered restarting automata for picture languages. Acta Informatica, 52(7–8):593– 623, November 2015. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-015-0230-5.

Li:2016:ADH

[1444] Suoping Li, Yongqiang Zhou, Duo Peng, Zufang Dou, and Yong Zhou. Analysis of dual-hop and multiple relays cooperative truncated ARQ with relay selection in WSNs. Acta Informatica, 53(1):1–22, February 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-015-0232-3.

Cao:2016:CKK

[1445] Chunhua Cao, Haiyan Liu, and Di Yang. Characterizations of k-comma codes and k-comma intercodes. Acta Informatica, 53(1):23–33, February 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
015-0233-2.

Janicki:2016:ST

[1446] Ryszard Janicki, Jetty Kleijn, Maciej Koutny, and Łukasz Mikulski. Step traces. Acta Informatica, 53(1):35–65, February 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-015-0244-z.

Han:2016:SCD

[1447] Yo-Sub Han, Sang-Ki Ko, and Kai Salomaa. State complexity of deletion and bipolar deletion. Acta Informatica, 53(1):67–85, February 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-015-0245-y.

Sanchez:2016:SIT

[1448] Cesar Sanchez, K. Brent Venable, and Esteban Zimanyi. Special issue on temporal representation and reasoning (TIME’13). Acta Informatica, 53(2):87–88, March 2016. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-016-0260-7;   http://link.springer.com/content/
pdf/10.1007/s00236-016-0260-7.pdf.

Hunsberger:2016:EED

[1449] Luke Hunsberger. Efficient execution of dynamically controllable simple temporal networks with uncertainty. Acta Informatica, 53(2):89–147, March 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
015-0227-0.

Condotta:2016:OTQ

[1450] Jean-François Condotta, Souhila Kaci, and Yakoub Salhi. Optimization in temporal qualitative constraint networks. Acta Informatica, 53 (2):149–170, March 2016. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-015-0228-z.

Bersani:2016:TDS

[1451] Marcello M. Bersani, Matteo Rossi, and Pierluigi San Pietro. A tool for deciding the satisfiability of continuous-time metric temporal logic. Acta Informatica, 53(2):171–206, March 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-015-0229-y.

Aceto:2016:CCE

[1452] Luca Aceto, Dario Della Monica, Valentin Goranko, Anna Ingólfsdóttir, Angelo Montanari, and Guido Sciavicco. A complete classification of the expressiveness of interval logics of Allen’s relations: the general and the dense cases. Acta Informatica, 53(3):207–246, April 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-015-0231-4.

Schuppan:2016:EUC

[1453] Viktor Schuppan. Extracting unsatisfiable cores for LTL via temporal resolution. Acta Informatica, 53(3):247–299, April 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://

link.springer.com/article/10.1007/s00236-015-0242-1.

Reynolds:2016:MTL

[1454] Mark Reynolds. Metric temporal logic revisited. Acta Informatica, 53 (3):301–324, April 2016. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-015-0243-0.

Finkbeiner:2016:SIR

[1455] Bernd Finkbeiner and Cesar Sanchez. Special issue on rich models, EU-COST action IC0901 rich-model toolkit. Acta Informatica, 53 (4):325–326, June 2016. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-016-0259-0;   http://link.springer.com/content/
pdf/10.1007/s00236-016-0259-0.pdf.

Peleg:2016:SAR

[1456] Hila Peleg, Sharon Shoham, Eran Yahav, and Hongseok Yang. Symbolic automata for representing big code. Acta Informatica, 53(4):327–356, June 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
015-0234-1.

Abdulla:2016:VHM

[1457] Parosh Aziz Abdulla, Lukás Holík, Bengt Jonsson, Ondrej Lengál, Cong Quy Trinh, and Tomás Vojnar. Verification of heap manipulating programs with ordered data by extended forest automata. Acta Informatica, 53(4):357–385, June 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.com/

article/10.1007/s00236-015-0235-0.

Leroux:2016:GCI

[1458] Jérôme Leroux, Philipp Rümmer, and Pavle Subotić. Guiding Craig interpolation with domain-specific abstractions. Acta Informatica, 53 (4):387–424, June 2016. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL http://link.springer.com/article/10.
1007/s00236-015-0236-z.

vonEssen:2016:SES

[1459] Christian von Essen, Barbara Jobstmann, David Parker, and Rahul Varshneya. Synthesizing efficient systems in probabilistic environments. Acta Informatica, 53(4):425–457, June 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/article/10.1007/s00236-015-0237-y.

Peng:2016:OCR

[1460] Xiaoning Peng and Zhijun Xiao. Optimal covers in the relational database model. Acta Informatica, 53(5):459–468, August 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s00236-015-0247-9;   http://link.springer.com/article/
10.1007/s00236-015-0247-9.

Borger:2016:CAS

[1461] Egon Börger and Klaus-Dieter Schewe. Concurrent abstract state machines. Acta Informatica, 53(5):469–492, August 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-015-0249-7;
 http://link.springer.com/content/pdf/10.1007/s00236-015-0249-
7.pdf.

Lin:2016:LTA

[1462] Anthony W. Lin and Sanming Zhou. A linear-time algorithm for the orbit problem over cyclic groups. Acta Informatica, 53(5):493–508, August 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s00236-015-0251-0;   http://link.springer.com/article/
10.1007/s00236-015-0251-0.

Axelsen:2016:RTM

[1463] Holger Bock Axelsen and Robert Glück. On reversible Turing machines and their function universality. Acta Informatica, 53(5):509–543, August 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s00236-015-0253-y;   http://link.springer.com/article/
10.1007/s00236-015-0253-y.

Bresolin:2016:SIS

[1464] Davide Bresolin and Guido Sciavicco. Special issue: selected papers from the 21st International Symposium on Temporal Representations and Reasoning (TIME-2014). Acta Informatica, 53(6–8):545–546, October 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
016-0262-5;   http://link.springer.com/content/pdf/10.1007/
s00236-016-0262-5.pdf.

Combi:2016:MAI

[1465] Carlo Combi and Pietro Sala. Mining approximate interval-based temporal dependencies. Acta Informatica, 53(6–8):547–585, October 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/accesspage/article/10.
1007/s00236-015-0246-x;   http://link.springer.com/article/
10.1007/s00236-015-0246-x.

Molinari:2016:CIP

[1466] Alberto Molinari, Angelo Montanari, Aniello Murano, Giuseppe Perelli, and Adriano Peron. Checking interval properties of computations. Acta Informatica, 53(6–8):587–619, October 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00236-015-0250-1;   http://
link.springer.com/article/10.1007/s00236-015-0250-1.

Montanari:2016:MPN

[1467] Angelo Montanari, Marco Pazzaglia, and Pietro Sala. Metric propositional neighborhood logic with an equivalence relation. Acta Informatica, 53(6–8):621–648, October 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00236-016-0256-3;   http://
link.springer.com/article/10.1007/s00236-016-0256-3.

Mayer:2016:PEF

[1468] Marta Cialdea Mayer, Andrea Orlandini, and Alessandro Umbrico. Planning and execution with flexible timelines: a formal account. Acta Informatica, 53(6–8):649–680, October 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00236-015-0252-z;   http://
link.springer.com/article/10.1007/s00236-015-0252-z.

Cimatti:2016:DCT

[1469] Alessandro Cimatti, Luke Hunsberger, Andrea Micheli, Roberto Posenato, and Marco Roveri. Dynamic controllability via Timed Game Automata. Acta Informatica, 53(6–8):681–722, October 2016. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/accesspage/article/10.1007/s00236-
016-0257-2;   http://link.springer.com/article/10.1007/s00236-
016-0257-2.

Nilsson:2016:EPS

[1470] Mikael Nilsson, Jonas Kvarnström, and Patrick Doherty. Efficient processing of simple temporal networks with uncertainty: algorithms for dynamic controllability verification. Acta Informatica, 53(6–8):723–752, October 2016. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s00236-015-0248-8;   http://
link.springer.com/article/10.1007/s00236-015-0248-8.

Aceto:2017:SIS

[1471] Luca Aceto and David de Frutos Escrig. Special issue: Selected papers from the 26th International Conference on Concurrency Theory (CONCUR 2015). Acta Informatica, 54(1):1–2, February 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
017-0292-7;   http://link.springer.com/content/pdf/10.1007/
s00236-017-0292-7.pdf.

Hunter:2017:RSR

[1472] Paul Hunter, Guillermo A. Pérez, and Jean-François Raskin. Reactive synthesis without regret. Acta Informatica, 54(1):3–39, February 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
016-0268-z;   http://link.springer.com/content/pdf/10.1007/
s00236-016-0268-z.pdf. See [1634].

Brenguier:2017:AAS

[1473] Romain Brenguier, Jean-François Raskin, and Ocan Sankur. Assumeadmissible synthesis. Acta Informatica, 54(1):41–83, February 2017. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

http://link.springer.com/article/10.1007/s00236-016-0273-2;
 http://link.springer.com/content/pdf/10.1007/s00236-016-0273-
2.pdf.

Brihaye:2017:PIA

[1474] Thomas Brihaye, Gilles Geeraerts, Axel Haddad, and Benjamin Monmege. Pseudopolynomial iterative algorithm to solve total-payoff games and min-cost reachability games. Acta Informatica, 54(1):85–125, February 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
016-0276-z;   http://link.springer.com/content/pdf/10.1007/
s00236-016-0276-z.pdf.

Bonchi:2017:GAC

[1475] Filippo Bonchi, Daniela Petrişan, Damien Pous, and Jurriaan Rot. A general account of coinduction up-to. Acta Informatica, 54(2):127–190, March 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
016-0271-4;   http://link.springer.com/content/pdf/10.1007/
s00236-016-0271-4.pdf.

Esparza:2017:VPP

[1476] Javier Esparza, Pierre Ganty, Jérôme Leroux, and Rupak Majumdar. Verification of population protocols. Acta Informatica, 54(2):191–215, March 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
016-0272-3;   http://link.springer.com/content/pdf/10.1007/
s00236-016-0272-3.pdf.

Soudjani:2017:DBN

[1477] Sadegh Esmaeil Zadeh Soudjani, Alessandro Abate, and Rupak Majumdar. Dynamic Bayesian networks for formal verification of structured stochastic processes. Acta Informatica, 54(2):217–242, March 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL http://link.springer.com/article/10.1007/s00236-
016-0287-9;   http://link.springer.com/content/pdf/10.1007/
s00236-016-0287-9.pdf.

Carbone:2017:MST

[1478] Marco Carbone, Fabrizio Montesi, Carsten Schürmann, and Nobuko Yoshida. Multiparty session types as coherence proofs. Acta Informatica, 54(3):243–269, May 2017. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s00236-016-0285-y.pdf.

Kouzapas:2017:CBH

[1479] Dimitrios Kouzapas, Jorge A. Pérez, and Nobuko Yoshida. Characteristic bisimulation for higher-order session processes. Acta Informatica, 54(3):271–341, May 2017. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s00236-016-0289-7.pdf.

Jaghoori:2017:CSA

[1480] Mohammad Mahdi Jaghoori, Frank de Boer, Delphine Longuet, Tom Chothia, and Marjan Sirjani. Compositional schedulability analysis of real-time actor-based systems. Acta Informatica, 54(4):343–378, June 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s00236-015-0254-x.pdf.

Kari:2017:DOP

[1481] Lila Kari and Manasi S. Kulkarni. Disjunctivity and other properties of sets of pseudo-bordered words. Acta Informatica, 54(4):379–398, June 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Degano:2017:RCF

[1482] Pierpaolo Degano, Gian-Luigi Ferrari, and Gianluca Mezzetti. Regular and context-free nominal traces. Acta Informatica, 54(4):399–433, June 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Aman:2017:ESP

[1483] Bogdan Aman and Gabriel Ciobanu. Efficiently solving the Bin Packing problem through bio-inspired mobility. Acta Informatica, 54(4):435–445, June 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Marin:2017:RBM

[1484] A. Marin and S. Rossi. On the relations between Markov chain lumpability and reversibility. Acta Informatica, 54(5):447–485, August 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chatterjee:2017:RWC

[1485] Kingshuk Chatterjee and Kumar Sankar Ray. Reversible Watson–Crick automata. Acta Informatica, 54(5):487–499, August 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Mostefaoui:2017:SFA

[1486] Achour Mostéfaoui and Michel Raynal. Signature-free asynchronous Byzantine systems: from multivalued to binary consensus with t < n/3, O(n^{2}) messages, and constant time. Acta Informatica, 54(5):501–520, August 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Gu:2017:ESA

[1487] Ke Gu, Weijia Jia, Guojun Wang, and Sheng Wen. Efficient and secure attribute-based signature for monotone predicates. Acta Informatica, 54 (5):521–541, August 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Chatterjee:2017:SIS

[1488] Krishnendu Chatterjee and Rüdiger Ehlers. Special issue: Synthesis and SYNT 2014. Acta Informatica, 54(6):543–544, September 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://

link.springer.com/content/pdf/10.1007/s00236-017-0299-0.pdf.

Bohy:2017:SAM

[1489] Aaron Bohy, Véronique Bruyère, Jean-François Raskin, and Nathalie Bertrand. Symblicit algorithms for mean-payoff and shortest path in monotonic Markov decision processes. Acta Informatica, 54(6):545–587, September 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Ceska:2017:PPS

[1490] Milan Ceska, Frits Dannenberg, Nicola Paoletti, Marta Kwiatkowska, and Lubos Brim. Precise parameter synthesis for stochastic biochemical systems. Acta Informatica, 54(6):589–623, September 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Huang:2017:MCI

[1491] Chung-Hao Huang, Sven Schewe, and Farn Wang. Model-checking iterated games. Acta Informatica, 54(7):625–654, November 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-016-0277-y.

Tomita:2017:SLS

[1492] Takashi Tomita, Atsushi Ueno, Masaya Shimakawa, Shigeki Hagihara, and Naoki Yonezaki. Safraless LTL synthesis considering maximal realizability. Acta Informatica, 54(7):655–692, November 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-016-0280-3;   https://link.springer.com/content/pdf/
10.1007/s00236-016-0280-3.pdf.

Jha:2017:TFS

[1493] Susmit Jha and Sanjit A. Seshia. A theory of formal synthesis via inductive learning. Acta Informatica, 54(7):693–726, November 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-017-0294-5.

Baier:2017:SII

[1494] Christel Baier and Cesare Tinelli. Special issue of the 21st International Conference on Tools and Algorithms for the Construction and Analysis of Systems (TACAS 2015). Acta Informatica, 54(8):727–728, December 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-017-0298-1;   https://link.springer.com/content/pdf/
10.1007/s00236-017-0298-1.pdf.

Chistikov:2017:ACS

[1495] Dmitry Chistikov, Rayna Dimitrova, and Rupak Majumdar. Approximate counting in SMT and value estimation for probabilistic programs. Acta Informatica, 54(8):729–764, December 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

URL https://link.springer.com/article/10.1007/s00236-017-
0297-2;   https://link.springer.com/content/pdf/10.1007/s00236-
017-0297-2.pdf.

Giacobbe:2017:MCE

[1496] Mirco Giacobbe, Calin C. Guet, Ashutosh Gupta, Thomas A. Henzinger, Tiago Paixão, and Tatjana Petrov. Model checking the evolution of gene regulatory networks. Acta Informatica, 54(8):765–787, December 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-016-0278-x;   https://link.springer.com/content/pdf/
10.1007/s00236-016-0278-x.pdf.

Abdulla:2017:SMC

[1497] Parosh Aziz Abdulla, Stavros Aronis, Mohamed Faouzi Atig, Bengt Jonsson, Carl Leonardsson, and Konstantinos Sagonas. Stateless model checking for TSO and PSO. Acta Informatica, 54(8):789–818, December 2017. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-016-0275-0;   https://link.springer.com/content/pdf/
10.1007/s00236-016-0275-0.pdf.

Bonnet:2018:SSA

[1498] Édouard Bonnet and Vangelis Th. Paschos. Sparsification and subexponential approximation. Acta Informatica, 55(1):1–15, February 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-016-0281-2.

Bjorklund:2018:CQC

[1499] Henrik Björklund, Wim Martens, and Thomas Schwentick. Conjunctive query containment over trees using schema information. Acta Informatica, 55(1):17–56, February 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-016-0282-1.

Favrholdt:2018:OEC

[1500] Lene M. Favrholdt and Jesper W. Mikkelsen. Online edge coloring of paths and trees with a fixed number of colors. Acta Informatica, 55 (1):57–80, February 2018. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-016-0283-0.

Honkala:2018:NBD

[1501] Juha Honkala. A new bound for the D0L language equivalence problem. Acta Informatica, 55(1):81–88, February 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-016-0286-x.

Esparza:2018:PSI

[1502] Javier Esparza and Enrico Tronci. Preface for the special issue GandALF 2015. Acta Informatica, 55(2):89–90, March 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

URL https://link.springer.com/article/10.1007/s00236-018-
0315-z;   https://link.springer.com/content/pdf/10.1007/s00236-
018-0315-z.pdf.

Bouyer:2018:AEG

[1503] Patricia Bouyer, Nicolas Markey, Mickael Randour, Kim G. Larsen, and Simon Laursen. Average-energy games. Acta Informatica, 55 (2):91–127, March 2018. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-016-0274-1.

Zimmermann:2018:PLT

[1504] Martin Zimmermann. Parameterized linear temporal logics meet costs: still not costlier than LTL. Acta Informatica, 55(2):129–152, March 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-016-0279-9.

Dimitrova:2018:RAR

[1505] Rayna Dimitrova and Rupak Majumdar. Reachability analysis of reversal-bounded automata on series–parallel graphs. Acta Informatica, 55(2):153–189, March 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.com/

article/10.1007/s00236-016-0290-1;   https://link.springer.
com/content/pdf/10.1007/s00236-016-0290-1.pdf.

Torfah:2018:CCM

[1506] Hazem Torfah and Martin Zimmermann. The complexity of counting models of linear-time temporal logic. Acta Informatica, 55(3):191– 212, May 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-016-0284-z.

Daltrophe:2018:BDI

[1507] Hadassa Daltrophe, Shlomi Dolev, and Zvi Lotker. Big data interpolation using functional representation. Acta Informatica, 55(3):213– 225, May 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-016-0288-8.

Barbuti:2018:GCR

[1508] Roberto Barbuti, Roberta Gori, Francesca Levi, and Paolo Milazzo. Generalized contexts for reaction systems: definition and study of dynamic causalities. Acta Informatica, 55(3):227–267, May 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-017-0296-3.

Damiani:2018:CCD

[1509] Ferruccio Damiani, Luca Padovani, Ina Schaefer, and Christoph Seidl. A core calculus for dynamic delta-oriented programming. Acta Informatica, 55(4):269–307, June 2018. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-017-0293-6.

Basin:2018:AMR

[1510] David Basin, Felix Klaedtke, and Eugen Zalinescu. Algorithms for monitoring real-time properties. Acta Informatica, 55(4):309–338, June 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-017-0295-4.

Devillers:2018:FTS

[1511] Raymond Devillers. Factorisation of transition systems. Acta Informatica, 55(4):339–362, June 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.com/

article/10.1007/s00236-017-0300-y.

Fares:2018:EAT

[1512] Elie Fares, Jean-Paul Bodeveix, and Mamoun Filali. Event algebra for transition systems composition application to timed automata. Acta Informatica, 55(5):363–400, August 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-017-0302-9.

Cranen:2018:PGR

[1513] Sjoerd Cranen, Jeroen J. A. Keiren, and Tim A. C. Willemse. Parity game reductions. Acta Informatica, 55(5):401–444, August 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-017-0301-x;   https://link.springer.com/content/pdf/
10.1007/s00236-017-0301-x.pdf.

Chunhua:2018:FSC

[1514] Cao Chunhua, Lu Qing, and Yang Di. A first step in characterizing threeelement codes. Acta Informatica, 55(5):445–457, August 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-017-0309-2.

deBoer:2018:PSI

[1515] Frank de Boer and Nikolaj Bjorner. Preface for the special issue “FM15”. Acta Informatica, 55(6):459–460, September 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

URL https://link.springer.com/article/10.1007/s00236-018-
0323-z;   https://link.springer.com/content/pdf/10.1007/s00236-
018-0323-z.pdf.

Zhang:2018:PBR

[1516] Lijun Zhang, Pengfei Yang, Lei Song, Holger Hermanns, Christian Eisentraut, David N. Jansen, and Jens Chr. Godskesen. Probabilistic bisimulation for realistic schedulers. Acta Informatica, 55(6):461– 488, September 2018. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0313-1.

Debois:2018:RRA

[1517] Søren Debois, Thomas T. Hildebrandt, and Tijs Slaats. Replication, refinement & reachability: complexity in dynamic condition-response graphs. Acta Informatica, 55(6):489–520, September 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-017-0303-8.

Amato:2018:DCN

[1518] Gianluca Amato, Simone Di Nardo Di Maio, Maria Chiara Meo, and Francesca Scozzari. Descending chains and narrowing on template abstract domains. Acta Informatica, 55(6):521–545, September 2018. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-016-0291-0.

Borsotti:2018:FDP

[1519] Angelo Borsotti, Luca Breveglieri, Stefano Crespi Reghizzi, and Angelo Morzenti. Fast deterministic parsers for transition networks. Acta Informatica, 55(7):547–574, November 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-017-0308-3.

Best:2018:BCF

[1520] Eike Best, Raymond Devillers, and Uli Schlachter. Bounded choicefree Petri net synthesis: algorithmic issues. Acta Informatica, 55(7): 575–611, November 2018. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-017-0310-9.

Zhang:2018:AMS

[1521] Hongbo Zhang. An analysis of the M^{X} /M/1 queue with multiple working vacations by GI/M/1 type Markov process. Acta Informatica, 55(7): 613–624, November 2018. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0316-y.

Finkbeiner:2018:PSI

[1522] Bernd Finkbeiner, Geguang Pu, and Lijun Zhang. Preface for the special issue for ATVA 2015. Acta Informatica, 55(8):625–626, December 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-018-0330-0;   https://link.springer.com/content/pdf/
10.1007/s00236-018-0330-0.pdf.

Hunter:2018:LMP

[1523] Paul Hunter, Guillermo A. Pérez, and Jean-François Raskin. Looking at mean payoff through foggy windows. Acta Informatica, 55(8):627–647, December 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-017-0304-7;   https://link.springer.com/content/pdf/
10.1007/s00236-017-0304-7.pdf.

Wang:2018:TTL

[1524] Chao Wang, Yi Lv, and Peng Wu. TSO-to-TSO linearizability is undecidable. Acta Informatica, 55(8):649–668, December 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-017-0305-6;   https://link.springer.com/content/pdf/
10.1007/s00236-017-0305-6.pdf.

Berwanger:2018:HIS

[1525] Dietmar Berwanger, Anup Basil Mathew, and Marie van den Bogaard. Hierarchical information and the synthesis of distributed strategies. Acta Informatica, 55(8):669–701, December 2018. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

URL https://link.springer.com/article/10.1007/s00236-017-
0306-5;   https://link.springer.com/content/pdf/10.1007/s00236-
017-0306-5.pdf.

Faran:2018:SSS

[1526] Rachel Faran and Orna Kupferman. Spanning the spectrum from safety to liveness. Acta Informatica, 55(8):703–732, December 2018. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-017-0307-4;   https://link.springer.com/content/pdf/
10.1007/s00236-017-0307-4.pdf.

Ebrahimi:2019:SCF

[1527] Masoud Ebrahimi, Gholamreza Sotudeh, and Ali Movaghar. Symbolic checking of Fuzzy CTL on Fuzzy Program Graph. Acta Informatica, 56(1):1–33, February 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.com/

article/10.1007/s00236-018-0311-3;   https://link.springer.
com/content/pdf/10.1007/s00236-018-0311-3.pdf.

Fernau:2019:PCI

[1528] Henning Fernau, Lakshmanan Kuppusamy, and Indhumathi Raman. On path-controlled insertion-deletion systems. Acta Informatica, 56 (1):35–59, February 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0312-2.

Baldan:2019:PND

[1529] Paolo Baldan and Fabio Gadducci. Petri nets are dioids: a new algebraic foundation for non-deterministic net theory. Acta Informatica, 56 (1):61–92, February 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0314-0.

Desel:2019:NCP

[1530] Jörg Desel, Javier Esparza, and Philipp Hoffmann. Negotiation as concurrency primitive. Acta Informatica, 56(2):93–159, March 2019. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-018-0318-9.

Abbasi:2019:VAS

[1531] Rosa Abbasi, Fatemeh Ghassemi, and Ramtin Khosravi. Verification of asynchronous systems with an unspecified component. Acta Informatica, 56(2):161–203, March 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0317-x.

Fiedor:2019:NAW

[1532] Tomás Fiedor, Lukás Holík, Ondrej Lengál, and Tomás Vojnar. Nested antichains for WS1S. Acta Informatica, 56(3):205–228, April 2019. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-018-0331-z.

Reisig:2019:ACC

[1533] Wolfgang Reisig. Associative composition of components with doublesided interfaces. Acta Informatica, 56(3):229–253, April 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-018-0328-7.

Kaminski:2019:HAP

[1534] Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Christoph Matheja. On the hardness of analyzing probabilistic programs. Acta Informatica, 56(3):255–285, April 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-018-0321-1.

Fendrich:2019:GTI

[1535] Sascha Fendrich and Gerald Lüttgen. A generalised theory of interface automata, component compatibility and error. Acta Informatica, 56(4):287–319, June 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0319-8.

Bandyopadhyay:2019:ECP

[1536] Soumyadip Bandyopadhyay, Dipankar Sarkar, and Chittaranjan Mandal. Equivalence checking of Petri net models of programs using static and dynamic cut-points. Acta Informatica, 56(4):321–383, June 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-018-0320-2.

Monniaux:2019:DEP

[1537] David Monniaux. On the decidability of the existence of polyhedral invariants in transition systems. Acta Informatica, 56(4):385–389, June 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-018-0324-y.

Elmasry:2019:RBT

[1538] Amr Elmasry, Mostafa Kahla, Fady Ahdy, and Mahmoud Hashem. Red-black trees with constant update time. Acta Informatica, 56 (5):391–404, July 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-019-00335-9.

Cenciarelli:2019:DCD

[1539] Pietro Cenciarelli, Daniele Gorla, and Ivano Salvo. Depletable channels: dynamics, behaviour, and efficiency in network design. Acta Informatica, 56(5):405–431, July 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0329-6.

Srinivasan:2019:PBA

[1540] Sathyanarayanan Srinivasan and Ramesh Kandukoori. A Paxos based algorithm to minimize the overhead of process recovery in consensus. Acta Informatica, 56(5):433–446, July 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-019-00334-w.

Fulop:2019:WIL

[1541] Zoltán Fülöp and Heiko Vogler. Weighted iterated linear control. Acta Informatica, 56(5):447–469, July 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-018-0325-x.

Chen:2019:IBS

[1542] Huiyan Chen and Chenchen Zhang. Identity-based signatures in standard model. Acta Informatica, 56(6):471–486, September 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-018-0322-0.

Krebs:2019:MCF

[1543] Andreas Krebs, Arne Meier, and Martin Mundhenk. The model checking fingerprints of CTL operators. Acta Informatica, 56(6):487–519, September 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL https://link.springer.com/article/10.
1007/s00236-018-0326-9.

Atanasiu:2019:PMP

[1544] Adrian Atanasiu, Ghajendran Poovanandran, and Wen Chean Teh. Parikh matrices for powers of words. Acta Informatica, 56(6):521– 535, September 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-018-0327-8.

Jancar:2019:SLP

[1545] Petr Jancar and David Purser. Structural liveness of Petri nets is ExpSpace-hard and decidable. Acta Informatica, 56(6):537–552, September 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00338-6.

Castellani:2019:RSF

[1546] Ilaria Castellani, Mariangiola Dezani-Ciancaglini, and Paola Giannini. Reversible sessions with flexible choices. Acta Informatica, 56(7–8): 553–583, November 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-019-00332-y.

Rot:2019:DLM

[1547] Jurriaan Rot. Distributive laws for monotone specifications. Acta Informatica, 56(7–8):585–617, November 2019. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-019-00333-x;   https://
link.springer.com/content/pdf/10.1007/s00236-019-00333-x.pdf.

Bonsma:2019:UCS

[1548] Paul Bonsma and Daniël Paulusma. Using contracted solution graphs for solving reconfiguration problems. Acta Informatica, 56(7–8):619– 648, November 2019. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-019-00336-8.

Bloem:2020:PS

[1549] Roderick Bloem and Paulo Tabuada. Preface for the SYNT. Acta Informatica, 57(1–2):1, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-019-00354-6;   https://
link.springer.com/content/pdf/10.1007/s00236-019-00354-6.pdf.

Luttenberger:2020:PSR

[1550] Michael Luttenberger, Philipp J. Meyer, and Salomon Sickert. Practical synthesis of reactive systems from LTL specifications via parity games. Acta Informatica, 57(1–2):3–36, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-019-00349-3.

Firman:2020:PHG

[1551] Elizabeth Firman, Shahar Maoz, and Jan Oliver Ringert. Performance heuristics for GR(1) synthesis and related algorithms. Acta Informatica, 57(1–2):37–79, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-019-00351-9;   https://
link.springer.com/content/pdf/10.1007/s00236-019-00351-9.pdf.

Jacobs:2020:SAL

[1552] Swen Jacobs and Mouhammad Sakr. A symbolic algorithm for lazy synthesis of eager strategies. Acta Informatica, 57(1–2):81–106, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00344-8.

Dimitrova:2020:RSM

[1553] Rayna Dimitrova, Mahsa Ghasemi, and Ufuk Topcu. Reactive synthesis with maximum realizability of linear temporal logic specifications. Acta Informatica, 57(1–2):107–135, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-019-00348-4.

Finkbeiner:2020:SH

[1554] Bernd Finkbeiner, Christopher Hahn, Philip Lukert, and Marvin Stenger. Synthesis from hyperproperties. Acta Informatica, 57(1–2):137– 163, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00358-2;   https://link.springer.com/content/pdf/
10.1007/s00236-019-00358-2.pdf.

Peleg:2020:PPF

[1555] Hila Peleg, Shachar Itzhaky, Sharon Shoham, and Eran Yahav. Programming by predicates: a formal model for interactive synthesis. Acta Informatica, 57(1–2):165–193, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-019-00340-y.

Neider:2020:SOR

[1556] Daniel Neider, Alexander Weinert, and Martin Zimmermann. Synthesizing optimally resilient controllers. Acta Informatica, 57(1–2):195– 221, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00345-7;   https://link.springer.com/content/pdf/
10.1007/s00236-019-00345-7.pdf.

Abate:2020:AFS

[1557] Alessandro Abate, Iury Bessa, Lucas Cordeiro, and Cristina David. Automated formal synthesis of provably safe digital controllers for continuous plants. Acta Informatica, 57(1–2):223–244, April 2020. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-019-00359-
1;   https://link.springer.com/content/pdf/10.1007/s00236-019-
00359-1.pdf.

Girard:2020:SSI

[1558] Antoine Girard and Gregor Gössler. Safety synthesis for incrementally stable switched systems using discretization-free multi-resolution abstractions. Acta Informatica, 57(1–2):245–269, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-019-00341-x.

Mirzaie:2020:PSS

[1559] Nahal Mirzaie, Fathiyeh Faghih, Swen Jacobs, and Borzoo Bonakdarpour. Parameterized synthesis of self-stabilizing protocols in symmetric networks. Acta Informatica, 57(1–2):271–304, April 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-019-00361-7.

Hofner:2020:P

[1560] Peter Höfner, Carroll Morgan, and Vaughan Pratt. Preface. Acta Informatica, 57(3–5):305–311, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-020-00382-7;   https://
link.springer.com/content/pdf/10.1007/s00236-020-00382-7.pdf.

Goltz:2020:SAC

[1561] Ursula Goltz and Jens-W. Schicke-Uffmann. Synchronous and asynchronous communication(s) between three parties. Acta Informatica, 57 (3–5):313–320, October 2020. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00381-8.

Pratt:2020:MTR

[1562] Vaughan Pratt. My time with Rob. Acta Informatica, 57(3–5): 321–322, October 2020. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00380-9.

Fehnker:2020:CR

[1563] Ansgar Fehnker. Out for coffee: with Rob. Acta Informatica, 57(3–5):323–327, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-020-00378-3;   https://
link.springer.com/content/pdf/10.1007/s00236-020-00378-3.pdf.

Gazda:2020:COP

[1564] Maciej Gazda, Wan Fokkink, and Vittorio Massaro. Congruence from the operator’s point of view. Acta Informatica, 57(3–5):329–351, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

URL https://link.springer.com/article/10.1007/s00236-019-
00355-5;   https://link.springer.com/content/pdf/10.1007/s00236-
019-00355-5.pdf.

Valmari:2020:ACB

[1565] Antti Valmari. All congruences below stability-preserving fair testing or CFFD. Acta Informatica, 57(3–5):353–383, October 2020. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-019-00364-
4;   https://link.springer.com/content/pdf/10.1007/s00236-019-
00364-4.pdf.

Bergstra:2020:ADT

[1566] Jan A. Bergstra and Alban Ponse. Arithmetical datatypes with true fractions. Acta Informatica, 57(3–5):385–402, October 2020. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-019-00352-
8;   https://link.springer.com/content/pdf/10.1007/s00236-019-
00352-8.pdf.

Mestel:2020:TBM

[1567] David Mestel and A. W. Roscoe. Translating between models of concurrency. Acta Informatica, 57(3–5):403–438, October 2020. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-020-00372-
9;   https://link.springer.com/content/pdf/10.1007/s00236-020-
00372-9.pdf.

Bisping:2020:CSF

[1568] Benjamin Bisping, Uwe Nestmann, and Kirstin Peters. Coupled similarity: the first 32 years. Acta Informatica, 57(3–5):439–463, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00356-4.

Baier:2020:PBS

[1569] Christel Baier, Pedro R. D’Argenio, and Holger Hermanns. On the probabilistic bisimulation spectrum with silent moves. Acta Informatica, 57 (3–5):465–512, October 2020. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00379-2.

Vogler:2020:LTB

[1570] Walter Vogler and Gerald Lüttgen. A linear-time branching-time perspective on interface automata. Acta Informatica, 57(3–5):513–550, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/

s00236-020-00369-4. See correction [1610].

Bouwman:2020:SAA

[1571] Mark Bouwman, Bas Luttik, and Tim Willemse. Off-the-shelf automated analysis of liveness properties for just paths. Acta Informatica, 57(3–5):551–590, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-020-00371-w;   https://
link.springer.com/content/pdf/10.1007/s00236-020-00371-w.pdf.

Gieseking:2020:SHL

[1572] Manuel Gieseking, Ernst-Rüdiger Olderog, and Nick Würdemann. Solving high-level Petri games. Acta Informatica, 57(3–5):591–626, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/

s00236-020-00368-5. See correction [1609].

Qin:2020:SBG

[1573] Xudong Qin, Simon Bliudze, Eric Madelaine, Zechen Hou, and Yuxin Deng. SMT-based generation of symbolic automata. Acta Informatica, 57 (3–5):627–656, October 2020. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00367-6.

Zhang:2020:MCD

[1574] Chenyi Zhang. Minimal consistent DFA from sample strings. Acta Informatica, 57(3–5):657–670, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-020-00365-8.

Jasper:2020:CIH

[1575] Marc Jasper, Maximilian Schlüter, and Bernhard Steffen. Characteristic invariants in Hennessy–Milner logic. Acta Informatica, 57(3–5):671–687, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-020-00376-5;   https://
link.springer.com/content/pdf/10.1007/s00236-020-00376-5.pdf.

Jensen:2020:CAW

[1576] Mathias Claus Jensen and Kim Guldstrand Larsen. A complete axiomatization of weighted branching bisimulation. Acta Informatica, 57(3– 5):689–725, October 2020. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00375-6.

Endrullis:2020:TDA

[1577] Jörg Endrullis, Jan Willem Klop, and Rena Bakhshi. Transducer degrees: atoms, infima and suprema. Acta Informatica, 57(3–5):727–758, October 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00353-7.

Peters:2020:PSI

[1578] Kirstin Peters and Simone Tini. Preface to special issue: EXPRESS/SOS 2016 + 2017. Acta Informatica, 57(6):759–760, December 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-019-00342-w.

Baranov:2020:ECB

[1579] Eduard Baranov and Simon Bliudze. Expressiveness of component-based frameworks: a study of the expressiveness of BIP. Acta Informatica, 57 (6):761–800, December 2020. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-019-00337-7. See correction [1581].

Huttel:2020:UST

[1580] Hans Hüttel. Using session types for reasoning about boundedness in the π-calculus. Acta Informatica, 57(6):801–827, December 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-019-00339-5.

Baranov:2020:CEC

[1581] Eduard Baranov and Simon Bliudze. Correction to: Expressiveness of component-based frameworks: a study of the expressiveness of BIP. Acta Informatica, 57(6):829, December 2020. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-019-00343-9. See [1579].

Lanese:2021:SVD

[1582] Ivan Lanese, Doriana Medić, and Claudio Antares Mezzina. Static versus dynamic reversibility in CCS. Acta Informatica, 58(1–2):1–34, April 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00346-6.

Engels:2021:SVS

[1583] Steven Engels, Tony Tan, and Jan Van den Bussche. Subsequence versus substring constraints in sequence pattern languages. Acta Informatica, 58(1–2):35–56, April 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-019-00347-5.

Bhattacharyya:2021:FMA

[1584] Arnab Bhattacharyya, Ashutosh Gupta, and Mukund Thattai. A formal methods approach to predicting new features of the eukaryotic vesicle traffic system. Acta Informatica, 58(1–2):57–93, April 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-019-00357-3.

Engelfriet:2021:LBC

[1585] Joost Engelfriet, Kazuhiro Inaba, and Sebastian Maneth. Linearbounded composition of tree-walking tree transducers: linear size increase and complexity. Acta Informatica, 58(1–2):95–152, April 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-019-00360-
8.

Nagy:2021:DSW

[1586] Benedek Nagy and Shaghayegh Parchami. On deterministic sensing 5′ → 3′ Watson–Crick finite automata: a full hierarchy in 2detLIN. Acta Informatica, 58(3):153–175, June 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-019-00362-6.

Bjorklund:2021:ABM

[1587] Johanna Björklund and Loek Cleophas. Aggregation-based minimization of finite state automata. Acta Informatica, 58(3):177–194, June 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-019-00363-5.

Borsotti:2021:DPA

[1588] Angelo Borsotti, Luca Breveglieri, and Angelo Morzenti. A deterministic parsing algorithm for ambiguous regular expressions. Acta Informatica, 58(3):195–229, June 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00366-7.

Ganguly:2021:WCQ

[1589] Debayan Ganguly, Kingshuk Chatterjee, and Kumar Sankar Ray. Watson–Crick quantum finite automata. Acta Informatica, 58(3):231– 240, June 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-020-00370-x.

Fernau:2021:PMK

[1590] Henning Fernau, Andreas Malcher, and Giovanni Pighizzini. Preface to Martin Kutrib Festschrift. Acta Informatica, 58(4):241–242, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00401-1.

Ostovar:2021:NLG

[1591] Ahmad Ostovar, Suna Bensch, and Thomas Hellström. Natural language guided object retrieval in images. Acta Informatica, 58(4): 243–261, August 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00400-2.

Bordihn:2021:RPC

[1592] Henning Bordihn and György Vaszil. Reversible parallel communicating finite automata systems. Acta Informatica, 58(4):263–279, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00396-9.

Dassow:2021:OCR

[1593] Jürgen Dassow. Operational complexity and right linear grammars. Acta Informatica, 58(4):281–299, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-020-00386-3.

Bordihn:2021:NAS

[1594] Henning Bordihn and Markus Holzer. On the number of active states in finite automata. Acta Informatica, 58(4):301–318, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-021-00397-8.

Kamilya:2021:NPP

[1595] Supreeti Kamilya and Jarkko Kari. Nilpotency and periodic points in non-uniform cellular automata. Acta Informatica, 58(4):319–333, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-020-00390-7.

Jakobi:2021:DPQ

[1596] Sebastian Jakobi, Katja Meckel, and Beatrice Palano. The descriptional power of queue automata of constant length. Acta Informatica, 58 (4):335–356, August 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00398-7.

Konstantinidis:2021:SPD

[1597] Stavros Konstantinidis, António Machiavelo, and Rogério Reis. On the size of partial derivatives and the word membership problem. Acta Informatica, 58(4):357–375, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-021-00399-6.

Morita:2021:ISR

[1598] Kenichi Morita. An instruction set for reversible Turing machines. Acta Informatica, 58(4):377–396, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-020-00388-1.

Otto:2021:RSO

[1599] Friedrich Otto and Matthias Wendlandt. Reversibility for stateless ordered RRWW-automata. Acta Informatica, 58(4):397–425, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-020-00389-0.

Umeo:2021:NCS

[1600] Hiroshi Umeo, Naoki Kamikawa, and Gen Fujita. A new class of the smallest FSSP partial solutions for 1D rings of length n = 2^{k} − 1. Acta Informatica, 58(4):427–450, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-020-00391-6.

Worsch:2021:FAB

[1601] Thomas Worsch. A faster algorithm for the Birthday Song Singers Synchronization Problem (FSSP) in one-dimensional CA with multiple speeds. Acta Informatica, 58(4):451–462, August 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-020-00383-6.

Geffert:2021:CTW

[1602] Viliam Geffert, Christos A. Kapoutsis, and Mohammad Zakzok. Complement for two-way alternating automata. Acta Informatica, 58(5): 463–495, October 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00373-8.

Chen:2021:IDP

[1603] Mingshuai Chen, Martin Fränzle, and Naijun Zhan. Indecision and delays are the parents of failure-taming them algorithmically by synthesizing delay-resilient control. Acta Informatica, 58(5):497–528, October 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-020-00374-7.

Gorrieri:2021:TBA

[1604] Roberto Gorrieri. Team bisimilarity, and its associated modal logic, for BPP nets. Acta Informatica, 58(5):529–569, October 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-020-00377-4.

Das:2021:BDA

[1605] Litan Kumar Das and Kumar Sankar Ray. Bitopological duality for algebras of Fitting’s logic and natural duality extension. Acta Informatica, 58 (5):571–584, October 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00384-5.

Gutierrez:2021:EGC

[1606] Julian Gutierrez, Aniello Murano, and Michael Wooldridge. Equilibria for games with combined qualitative and quantitative objectives. Acta Informatica, 58(6):585–610, December 2021. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-020-00385-4.

Mazhar:2021:DEA

[1607] Rabia Mazhar and Muddassar Azam Sindhu. DKL: an efficient algorithm for learning deterministic Kripke structures. Acta Informatica, 58(6): 611–651, December 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00387-2.

Xu:2021:MCR

[1608] Ming Xu, Cheng-Chao Huang, and Yuan Feng. Measuring the constrained reachability in quantum Markov chains. Acta Informatica, 58 (6):653–674, December 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-020-00392-5.

Gieseking:2021:CSH

[1609] Manuel Gieseking, Ernst-Rüdiger Olderog, and Nick Würdemann. Correction to: Solving high-level Petri games. Acta Informatica, 58(6): 675–676, December 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00405-x. See [1572].

Vogler:2021:CLT

[1610] Walter Vogler and Gerald Lüttgen. Correction to: A linear-time branching-time perspective on interface automata. Acta Informatica, 58(6):677, December 2021. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00406-w. See [1570].

Toth:2022:CVT

[1611] Tamás Tóth and István Majzik. Configurable verification of timed automata with discrete variables. Acta Informatica, 59(1):1–35, February 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-020-00393-4.

Cano:2022:SBC

[1612] Mauricio Cano, Hugo A. López, and Camilo Rueda. Session-based concurrency, declaratively. Acta Informatica, 59(1):1–87, February 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00395-w.

Baxter:2022:SRE

[1613] James Baxter, Pedro Ribeiro, and Ana Cavalcanti. Sound reasoning in tock-CSP. Acta Informatica, 59(1):125–162, February 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:

//link.springer.com/article/10.1007/s00236-020-00394-3. See

[1618].

Kutrib:2022:FAU

[1614] Martin Kutrib, Andreas Malcher, and Christian Schneider. Finite automata with undirected state graphs. Acta Informatica, 59(1):163–181, February 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00402-0.

Ergurtuna:2022:ASR

[1615] Mert Ergurtuna, Beyazit Yalcinkaya, and Ebru Aydin Gol. An automated system repair framework with signal temporal logic. Acta Informatica, 59(2–3):183–209, June 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-021-00403-z.

Marin:2022:PLP

[1616] Andrea Marin, Carla Piazza, and Sabina Rossi. Proportional lumpability and proportional bisimilarity. Acta Informatica, 59(2–3):211– 244, June 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00404-y.

Elmasry:2022:RNS

[1617] Amr Elmasry and Jyrki Katajainen. Regular numeral systems for data structures. Acta Informatica, 59(2–3):245–281, June 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-021-00407-9.

Baxter:2022:CSR

[1618] James Baxter, Pedro Ribeiro, and Ana Cavalcanti. Correction to: Sound reasoning in tock-CSP. Acta Informatica, 59(2–3):283, June 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00409-7. See [1613].

Fernau:2022:PKJ

[1619] Henning Fernau, Markus Holzer, and Petra Wolf. Preface to Klaus-Jörn Lange Festschrift. Acta Informatica, 59(4):285–287, August 2022. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-022-00426-
0.

Allender:2022:DFS

[1620] Eric Allender, Archit Chauhan, and Samir Datta. Depth-first search in directed planar graphs, revisited. Acta Informatica, 59(4):289–319, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00425-1.

Cadilhac:2022:RLW

[1621] Michaël Cadilhac and Charles Paperman. The regular languages of wire linear AC^{0}. Acta Informatica, 59(4):321–336, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-022-00432-2.

Dassow:2022:OCP

[1622] Jürgen Dassow and Ismaël Jecker. Operational complexity and pumping lemmas. Acta Informatica, 59(4):337–355, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-022-00431-3.

Diekert:2022:PGS

[1623] Volker Diekert, Henning Fernau, and Petra Wolf. Properties of graphs specified by a regular language. Acta Informatica, 59(4):357–385, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00427-z.

Erlebach:2022:EEK

[1624] Thomas Erlebach and Jakob T. Spooner. Exploration of k-edge-deficient temporal graphs. Acta Informatica, 59(4):387–407, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00421-5.

Fujioka:2022:MRC

[1625] Kaoru Fujioka, Fumiya Okubo, and Takashi Yokomori. L-reduction computation revisited. Acta Informatica, 59(4):409–426, August 2022. CO-DEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL

https://link.springer.com/article/10.1007/s00236-022-00418-
0.

Bockenhauer:2022:RPP

[1626] Hans-Joachim Böckenhauer, Elisabet Burjons, and Peter Rossmanith. Reoptimization of parameterized problems. Acta Informatica, 59(4): 427–450, August 2022. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-022-00428-y.

Jain:2022:LGA

[1627] Sanjay Jain, Birzhan Moldagaliyev, and Tien Dat Tran. Lamplighter groups and automata. Acta Informatica, 59(4):451–478, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00423-3.

Shabana:2022:CSP

[1628] Hanan Shabana and M. V. Volkov. Careful synchronization of partial deterministic finite automata. Acta Informatica, 59(4):479–504, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00433-1.

Wolf:2022:DFP

[1629] Petra Wolf. On the decidability of finding a positive ILP-instance in a regular set of ILP-instances. Acta Informatica, 59(4):505–519, August 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00429-x.

Schinko:2022:IAS

[1630] Ayleen Schinko, Walter Vogler, and Gerald Lüttgen. Interface automata for shared memory. Acta Informatica, 59(5):521–556, October 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00408-8.

Kocman:2022:JWC

[1631] Radim Kocman, Zbynek Krivka, and Benedek Nagy. A jumping 5′ → 3′ Watson–Crick finite automata model. Acta Informatica, 59 (5):557–584, October 2022. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00413-x.

Kretinsky:2022:IAR

[1632] Jan Kretínský, Tobias Meggendorfer, and Maximilian Weininger. Index appearance record with preorders. Acta Informatica, 59(5):585– 618, October 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL https://link.springer.com/article/10.
1007/s00236-021-00412-y.

Geffert:2022:ICT

[1633] Viliam Geffert, Christos A. Kapoutsis, and Mohammad Zakzok. Improved complement for two-way alternating automata. Acta Informatica, 59(5):619–669, October 2022. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00414-w.

Hunter:2022:CRS

[1634] Paul Hunter, Guillermo A. Pérez, and Jean-François Raskin. Correction to: Reactive synthesis without regret. Acta Informatica, 59 (5):671, October 2022. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-021-00410-0. See [1472].

Michalak:2022:HHB

[1635] Marcin Michalak. Hierarchical heuristics for Boolean-reasoning-based binary bicluster induction. Acta Informatica, 59(6):673–685, December 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-021-00415-9.

Chrobak:2022:HWA

[1636] Marek Chrobak, Mordecai Golin, J. Ian Munro, and Neal E. Young. On Huang and Wong’s algorithm for generalized binary split trees. Acta Informatica, 59(6):687–708, December 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-021-00411-z.

Bille:2022:REM

[1637] Philip Bille and Inge Li Gørtz. From regular expression matching to parsing. Acta Informatica, 59(6):709–724, December 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-022-00420-6.

Bandyopadhyay:2022:TVC

[1638] Soumyadip Bandyopadhyay, Dipankar Sarkar, Chittaranjan Mandal, and Holger Giese. Translation validation of coloured Petri net models of programs on integers. Acta Informatica, 59(6):725–759, December 2022. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-022-00419-z.

Guo:2023:MST

[1639] Haiyan Guo and Bo Zhou. Minimum status of trees with a given degree sequence. Acta Informatica, 60(1):1–10, March 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-022-00416-2.

vanGlabbeek:2023:RBS

[1640] Rob van Glabbeek. Reactive bisimulation semantics for a process algebra with timeouts. Acta Informatica, 60(1):11–57, March 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-022-00417-1.

Dolev:2023:SSR

[1641] Shlomi Dolev and Yin Li. Secret-shared RAM indefinite private and secure RAM execution of perfectly unrevealed programs. Acta Informatica, 60(1):59–78, March 2023. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-022-00422-4.

Kuske:2023:ACC

[1642] Dietrich Kuske and Christian Schwarz. Alternating complexity of counting first-order logic for the subword order. Acta Informatica, 60(1): 79–100, March 2023. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-022-00424-2.

Levin:2023:WTS

[1643] Asaf Levin and Tal Shusterman. Weighted throughput in a single machine preemptive scheduling with continuous controllable processing times. Acta Informatica, 60(2):101–122, June 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-022-00430-4.

Pighizzini:2023:PAC

[1644] Giovanni Pighizzini and Luca Prigioniero. Pushdown automata and constant height: decidability and bounds. Acta Informatica, 60(2): 123–144, June 2023. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-022-00434-0.

Samadi:2023:DRV

[1645] Mahboubeh Samadi, Fatemeh Ghassemi, and Ramtin Khosravi. Decentralized runtime verification of message sequences in message-based systems. Acta Informatica, 60(2):145–178, June 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-022-00435-z.

Lodato:2023:PIR

[1646] Ivano Lodato, Snehal M. Shekatkar, and Tian An Wong. On partial information retrieval: the unconstrained 100 prisoner problem. Acta Informatica, 60(2):179–208, June 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-022-00436-y.

Khaireddine:2023:TTP

[1647] Besma Khaireddine, Aleksandr Zakharchenko, Matias Martinez, and Ali Mili. Toward a theory of program repair. Acta Informatica, 60(3): 209–255, September 2023. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00438-4.

Xie:2023:UDC

[1648] Jingnan Xie and Harry B. Hunt III. On the undecidability and descriptional complexity of synchronized regular expressions. Acta Informatica, 60(3):257–278, September 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-023-00439-3.

Kochdumper:2023:CPZ

[1649] Niklas Kochdumper and Matthias Althoff. Constrained polynomial zonotopes. Acta Informatica, 60(3):279–316, September 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https:/

/link.springer.com/article/10.1007/s00236-023-00437-5.

Domosi:2023:SCA

[1650] Pál Dömösi, Géza Horváth, and Norbert Tihanyi. Simple chain automaton random number generator for IoT devices. Acta Informatica, 60(3): 317–329, September 2023. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00440-w.

Fernau:2023:ECI

[1651] Henning Fernau. Editorial 2023: changes and invariants. Acta Informatica, 60(4):331–333, December 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-023-00447-3.

Sheikhi:2023:DDS

[1652] Farnaz Sheikhi, Behnam Zeraatkar, and Sama Hanaie. Dot to dot, simple or sophisticated: a survey on shape reconstruction algorithms.

Acta Informatica, 60(4):335–359, December 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-023-00443-7.

Lassaigne:2023:TMT

[1653] Richard Lassaigne and Michel de Rougemont. Testing membership for timed automata. Acta Informatica, 60(4):361–384, December 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-023-00442-8.

Aceto:2023:FOR

[1654] Luca Aceto, Ian Cassar, Adrian Francalanza, and Anna Ingólfsdóttir. On first-order runtime enforcement of branching-time properties. Acta Informatica, 60(4):385–451, December 2023. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-023-00441-9.

Cao:2023:SSC

[1655] Chunhua Cao, Jiao Xu, Lei Liao, Di Yang, Guichuan Jia, and Qian Du. The second step in characterizing a three-word code. Acta Informatica, 60 (4):453–465, December 2023. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00444-6.

Tapia-Flores:2024:DWN

[1656] Tonatiuh Tapia-Flores and Ernesto López-Mellado. Discovering workflow nets of concurrent iterative processes. Acta Informatica, 61(1): 1–21, March 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-

0525 (electronic). URL https://link.springer.com/article/10.
1007/s00236-023-00445-5.

Wu:2024:DPS

[1657] Hao Wu, Yu-Fang Chen, Zhilin Wu, Bican Xia, and Naijun Zhan. A decision procedure for string constraints with string/integer conversion and flat regular constraints. Acta Informatica, 61(1):23–52, March 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-023-00446-4.

Gao:2024:BAS

[1658] Shuyang Gao, Leen Hatem, and Hosam Mahmoud. Balancing m-ary search trees with compressions on the fringe. Acta Informatica, 61 (1):53–66, March 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00448-2.

Liu:2024:EPC

[1659] Haiyan Liu, Rongdong Cui, and Tianjie Zhang. n-PS-codes, 2-infix-outfix codes and some related classes of codes. Acta Informatica, 61 (1):67–81, March 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00449-1.

Dolev:2024:NMR

[1660] Shlomi Dolev, Sayaka Kamei, Yoshiaki Katayama, Fukuhito Ooshita, and Koichi Wada. Neighborhood mutual remainder: self-stabilizing distributed implementation and applications. Acta Informatica, 61 (1):83–100, March 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00450-8.

Alonistiotis:2024:ASS

[1661] Giannis Alonistiotis, Antonis Antonopoulos, Nikolaos Melissinos, Aris Pagourtzis, Stavros Petsalakis, and Manolis Vasilakis. Approximating subset sum ratio via partition computations. Acta Informatica, 61 (2):101–113, June 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00451-7.

Dede:2024:NFL

[1662] Cahit Dede. New families of Laplacian borderenergetic graphs. Acta Informatica, 61(2):115–129, June 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-024-00454-y.

Li:2024:EDQ

[1663] Hao Li, Daowen Qiu, Le Luo, and Paulo Mateus. Exact distributed quantum algorithm for generalized Simon’s problem. Acta Informatica, 61(2):131–159, June 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-024-00455-x.

Bagossy:2024:ELC

[1664] Attila Bagossy and Péter Battyányi. An encoding of the λ-calculus in the String MultiSet Rewriting calculus. Acta Informatica, 61(2):

161–181, June 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-024-00456-w.

Yang:2024:DEM

[1665] Gang Yang, Jiannan Zhou, Changxiang He, and Yaping Mao. Distanceedge-monitoring sets of networks. Acta Informatica, 61(2):183–198, June 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-024-00453-z.

Ghiasi-Shirazi:2024:RRB

[1666] Kamaledin Ghiasi-Shirazi, Taraneh Ghandi, Ali Taghizadeh, and Ali Rahimi-Baigi. Revisiting 2-3 red–black trees with a pedagogically sound yet efficient deletion algorithm: parity-seeking. Acta Informatica, 61(3): 199–229, September 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-023-00452-6.

Chen:2024:RAL

[1667] Shiping Chen and Xinyu Ge. Reachability analysis of linear systems. Acta Informatica, 61(3):231–260, September 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-024-00458-8.

Ghardallou:2024:IRA

[1668] Wided Ghardallou, Hessamaldin Mohammadi, Richard C. Linger, Mark Pleszkoch, JiMeng Loh, and Ali Mili. Invariant relations for affine loops. Acta Informatica, 61(3):261–314, September 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-024-00457-9.

Lai:2024:LLD

[1669] Wenfeng Lai, Adiesha Liyanage, Binhai Zhu, and Peng Zou. The longest letter-duplicated subsequence and related problems. Acta Informatica, 61 (3):315–329, September 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-024-00459-7.

Fernau:2024:EMF

[1670] Henning Fernau. Editorial 2024: moving forwards in the electronic age. Acta Informatica, 61(4):331–332, December 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-024-00466-8.

Sucu:2024:CEB

[1671] Burkay Sucu and Ebru Aydin Gol. Cycle encoding-based parameter synthesis for timed automata safety. Acta Informatica, 61(4):333– 356, December 2024. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-024-00460-0.

Kumar:2024:CLH

[1672] R. Mahendra Kumar and N. Sadagopan. A closer look at Hamiltonicity and domination through the lens of diameter and convexity. Acta Informatica, 61(4):357–382, December 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-024-00461-z.

Fleischmann:2024:WRG

[1673] Pamela Fleischmann, Lukas Haschke, Tim Löck, and Dirk Nowotka. Word-representable graphs from a word’s perspective. Acta Informatica, 61(4):383–400, December 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-024-00462-y.

De:2024:PAD

[1674] Koustav De, Harshil Mittal, Palash Dey, and Neeldhara Misra. Parameterized aspects of distinct Kemeny rank aggregation. Acta Informatica, 61(4):401–414, December 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.springer.

com/article/10.1007/s00236-024-00463-x.

Jargalsaikhan:2024:SPA

[1675] Davaajav Jargalsaikhan, Diptarama Hendrian, Yohei Ueki, Ryo Yoshinaka, and Ayumi Shinohara. Serial and parallel algorithms for orderpreserving pattern matching based on the duel-and-sweep paradigm. Acta Informatica, 61(4):415–444, December 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic). URL https://link.

springer.com/article/10.1007/s00236-024-00464-w.

Inenaga:2024:LSS

[1676] Shunsuke Inenaga. Linear-size suffix tries and linear-size CDAWGs simplified and improved. Acta Informatica, 61(4):445–468, December 2024. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525

(electronic). URL https://link.springer.com/article/10.1007/
s00236-024-00465-9.

Ramos:2025:CGS

[1677] Lucas P. Ramos, Felipe A. Louza, and Guilherme P. Telles. Comparative genomics with succinct colored de Bruijn graphs. Acta Informatica, 62(1):??, March 2025. CODEN AINFA2. ISSN 0001-5903 (print),

1432-0525 (electronic). URL https://link.springer.com/article/
10.1007/s00236-024-00467-7.